A Complete Bibliography of Publications in the Reviews in Fish Biology and Fisheries

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

15 May 2024 Version 1.16

Title word cross-reference

$25 [Ste05]. 3 [JJMD13, OBS08]. _{1} [BVMF13]. _{2} [BVMF13]. δ^{13} [CNH22]. δ^{15} [CNH22]. × [IJ01].

-D [JJMD13]. -phosphate [OBS08].

0 [Coc05]. 0-632-06389-0 [Coc05]. 0-691-11545-1 [MAJ05].

1 [Ste05]. 1-55963-324-7 [Ste05]. 120-year [HETS23]. 15-year [VCD12]. 16S [AVL07]. 19 [BPA^{+}23, LSDRS23, OSB^{+}23]. 1950s [CBH15]. 1991 [Ano94a]. 1997 [Pit98b]. 19th [BGHC23]. 19th-century [BGHC23].

20-year [YWS^{+}24]. 2001 [Sul04]. 2004 [Zim05, Wil07]. 2006 [JBX07]. 2008 [KGW10, KGW11]. 2030 [MMM22b, PAN22]. 21st [Utt94]. 26th [Hus04]. 2nd [Zim05].

342pp. [Els04]. 3R [GS23]. 3R-Guidelines [GS23].

5S [MPO11, RZV12]. 5th [Coc05, Gil93a, Lor93].

6-species-fish [BHS19].

7 [Ste05].

8th [GB23].

A. [Els04]. abalone [HTH00, HSC^{+}24]. Abandoned [SEH24]. Abiotic [MÁMS05, VMB07, PMP21, TJLLC10]. Aboriginal [SE16]. abrupt [ZL23]. absence [Gau01]. absolute [HKN18]. abundance [BGOW23, CCR^{+}24, CBB19, EBS20, HWA14, JNW14, OV24, PWM23, PBS14, STP22]. Acanthocybium [ZFT13]. access [Mar06]. accessory [WOR09]. acclimation [APLL07, Sän93]. according [GS23]. account [Wil07]. accumulation [EDP10, Wie96]. accuracy [Vil18]. achievable [NAP22a, NAP22b]. achieving [HFM19]. Achilles [ANP^{+}23]. acid [GGZ10, GSD18]. acidification [SRD93]. acids [ZJDZ10]. Acipenser [WOR09]. Acipenseridae [CCO20, PVJ07]. Acipenseriformes [SVC21]. Acoustic [PM19, SVM^{+}24a, SVM^{+}24b, AVB^{+}23, KCF14, KM20, LMH21, MKL^{+}23, MBM06]. Acoustics [Bur07, Ano92m, Jua02, Mis97, Bar95]. acrosome [AHL12]. across [BHS19, Gau01, MWP^{+}23, PDN^{+}24, dSSHK21, SJH22]. Action [WBW09, HP14, LSH15, MVM99]. activation [CBP02, SZG10]. activities [Koc01, OGR17, ZL23]. activity [AE22, BMN21, DBT15, OK12, Ree02]. actors [APP^{+}23]. adaptability [SDS15, ZLH^{+}23]. adaptation [AS95, FCP19, FHvP14, HYW13, HP14, HCvP16, JLC21, PCW17, TFP22]. adapting [MTAP22]. Adaption [WG19]. Adaptive [DPV^{+}23, HFP14, SVM^{+}24a, SVM^{+}24b, HII16, MBH14]. address [OGR17]. addressing [KM20]. adipose [WLB13]. adjacent [HSC20, How94, PCG19, Rai94, SRGS04]. Adriatic [CUT07, CCR^{+}24]. Adult [dSSHK21]. advance [BEC11, MTC19]. Advancements [DD13a, DD13b, AS08, GOP^{+}23]. Advances [Cad00, DW11, LW17, Smi98]. Advancing [ANB19, CWN11, GMP12, SVRS19, SK23]. advantages [HKTS^{+}23]. advice [GOP^{+}23, PWB99]. aerobic [HKN18]. Aeromonas [SZG10]. Affairs [WBW09]. affect [BCA^{+}23, Ros00, YH18]. affected [BGOW23, DD15]. affecting [DCS20, TØH08]. affects [PBS14]. affinis [Pyk05]. Aflatoxins [SCC08]. Africa [JCL07, MC16, MR18, OSB07, PGJ15, Tur95]. African [MS07, Rei99, BOD21, CBR98, Cra92, HLD07, Mar93, PB98, Whi99]. AFS [Har98]. after [CBF^{+}24, CGRCM19, LQW19, SW10]. against [Jon07b]. Agassiz [dRRGC12]. Age [AKGB11, STS07, Vil18, LE24, MRMJJ17, WMA20]. ageing [Vil18, WDP19]. agenda [CWN11]. Agents [RT05, PDGBR07]. aggregating [SOB20]. aggregation [CSSO02, DT04]. aggregations [MP07b]. aggression [MGB15]. Agnathans [For95]. Agulhas [OSB07]. Ahl [Pau10]. al [DCS12, RPA13, TBK11]. alalunga [NMF17]. Alarm [Smi92]. Alaska [AAH98, OWW04, PDA12, WHS04]. albacares [PZC17]. albacore [NMF17]. albatrosses [Koc01, PWM23]. Alboran [BCA^{+}23]. Alburnus [LMCB^{+}23]. albus [Shi05]. alert [PBC12]. algorithm [MGRRJ21]. Alien [DAN02, BCA^{+}23, McD06]. alligator [MAC02]. Allis [MBA15]. allocation [Kam08]. Allodontichthys [Web02]. Alloparental [Wis99]. allopatric [WM02]. Allozymes [PBB00]. alone [ENF^{+}23]. along [HMN17, MMG18, PTOS^{+}24, SHHK21]. Alopias [FCCS15, MMF18]. Alosa [MBA15]. Alps [GLC16]. alterations [IPT10]. alternate [NAP22a, NAP22b]. Alternative [Smi91, Bas93, Cad91, CSSO02, LWS17, LSF14, PVL21, SA12]. Amacuzac [TJLLC10]. Amazon [CPN11, CSA11, CSA12, CBB19, HKTS^{+}23, MCN12, PBG04]. Amazonia [MJSOF16]. Ameiurus [CTB16, RBK13]. America [Nel92, Ano94j, ABK14, CBL17, DCS20, DAN02, JNW14, JZ00, MGTD18, MKS12, Nor03, Nor06, QSBV18, REA^{+}23]. American [Soi99, WGL14, Bag11, BP08, CDC09, CD01, CM92, DC05, DML^{+}24, Gre93b, Knu01, MS02, PVJ07, QSBV18, SBZR17, VEK10]. Americans [Gil93b]. americanus [Bag11]. Americas [AGMB^{+}23, BSM17]. ammocoetes [MJM15]. among [APP^{+}23, BOV09a, BOV09b, DBT15, DD13a, DD13b, LBS19, MPO11, MHvH16, SRBS21, Smi93, ZLF21]. Amphibious [SD91]. amphidromous [AWC17, McD09b, McD10]. amphidromy [WIT14]. amplified [AHW04]. ampullary [WT08]. Anadromous [LPC^{+}24, BGBE^{+}23, Cha95, CPM14, KC14, MC04, PQS14, QGM15, SDS15, SAC^{+}23]. Anadromy [QM04, CBA10]. Anaesthetic [Mul10]. anagenesis [WM02]. analyses [MFL21, WGL17, Duc19]. analysis [AMV13, Ano92s, BSL18, BSV11, CNH22, CWN11, ENP18, GSD18, HW24, IRPG21, JEL10, JLC21, LMCB^{+}23, MPO11, McD97, Mok93, PASF13, PH97, dSSHK21, SAN^{+}23, SBG17, SJB^{+}23, SW14, TFF09, TRB13, VCZ19, VLD20, Woo93, WPF16, YOC15]. analytical [EBS20, ROP^{+}23]. analyzing [SMN08]. Anarhichas [FIØ04]. anatomical [GH17]. Anchovy [Csi99]. Andaman [STS07]. Andean [VE05]. Andrew [Tur99a]. anecdote [FHvP14]. anesthesia [CWB11]. angiru [PBM14]. Angler [Ano95a, BGM^{+}24, Ano94b, AAJ21, GAD10]. anglers [BPA^{+}23]. Anglia [Ano94a]. Angling [BB05, SEA22]. Anguilla [Ara14, BLD19, Jel22a, Jel22b, LF03, vGM05]. anguillid [WJP23]. animal [Ano95g, CWB16, GS23, MKL^{+}23, RR11]. animals [Ano94p, Avi00, DZZZ22, DCS11, DCS12, KBV11, MHW07, SYL^{+}23, TBK11, Mul10]. Anne [Hus04, RHP99]. Annotated [Abl06, RCCASG02]. Announcement [Ano96a, Bra99]. Annual [Hus04, NRCD^{+}23]. Antarctic [BOV09a, BOV09b, JLD^{+}24, KdSD18, MWP^{+}23]. Antarctica [Fui99]. antarcticum [BOV09a, BOV09b]. anterior [MBS17]. anterior-gear [MBS17]. Anthology [Woo99b]. Anthony [Nor96, Lac02]. Anthropocene [BCP22]. Anthropogenic [WGR20, TRJ24, Utt00, dJFS20]. Antibacterial [RT05]. Antibiotic [Mul11]. antimicrobial [PDGBR07]. Antioxidant [MÁMS05]. Anvironmental [KVH98]. apama [HFG07]. Aphanopus [BMF09]. Aphia [LAI05]. apparatus [RL05]. apparent [WPD12]. appearance [Col10]. Application [Ano94h, ROL14, Ano92c, Bag11, Cal22, HJC09, Jua02, SMS12, TFF09, FQSJ23]. Applications [Vil03, Ano95a, BC03, GVB94, Lin94, Mag95, RSGS^{+}23a, RSGS^{+}23b, Tel09, VCZ19]. applied [AJM22, PY97, UIA11]. apply [Cad99]. Applying [CKN21, GC23, JBX07]. Appraisal [WG94]. approach [ACST17, Ano92g, BOs12, CCA^{+}23, CUT07, CCP13, CAC15, ELC^{+}24, GJA17, HJC09, IWG17, McL94, OWW04, PBM14, Pat92, PMVA19, PH97, RNJ16, Sch18, SSBCL11, dSKV16]. Approaches [AAJ21, SPS07, ANL12, BdST16, CRPI^{+}22, GZT13, JZ00, LAWD06, NRCD^{+}23]. appropriate [CHFTV22]. approval [BTW15]. Approximate [PB98]. Aquaculture [ZLH^{+}23, AY10, BRI00, BO12, Cas93, DGV11, DW09, DZZZ22, FGL10, KMF13, LW17, PTP14, RPD16, RNJ16, ROL14, RBC16a, RBC16b, RM01, SYL^{+}23, TRB13, UE02, Vil03, WMDS18]. aquacultured [UIA11]. aquaria [BCK18]. Aquatic [EN00, Ano93e, Ano95d, Avi00, CDS16, DCS11, DCS12, DDD16, DAN02, KCF14, LSP11, MKL^{+}23, Mul10, MT18, SCC08, SX16, SYL^{+}23, TBK11, ZWW20, VAJ10]. Arabian [Moo12, BHD23]. Arabian/Persian [BHD23]. arapaima [CSA11, CSA12]. archaeological [JNW14]. archival [LR11]. Arctic [BD20, Nor95]. Arctica [RR11]. Arcto [SP98]. Arcto-boreal [SP98]. Area [ACST17, GLG12, Jon02, MLC17, PP92, PAK^{+}23, SW12, TV15]. area-based [PAK^{+}23]. areas [AFB15, BBW09, HCK21, HSC20, Jon07b, Kel23, LET^{+}23, Ros00, Sot02, TRJ24, Whi17]. Argentina [ACB08, HCMT24a, HCMT24b]. Arginine [Kul95]. Arguments [Jon07b]. arid [Mag13, WWS17]. arid-land [Mag13, WWS17]. arsenal [Har24]. art [BSM17]. Arthur [MBJ12]. Article [DD13a, SDJ13]. Artificial [Ano93a, GCSR09, Mar93, PCG19]. Artisanal [ASS^{+}23, DC05, APP^{+}23, BCA^{+}23, BHD23, MCFC21, Sur23]. Arve [Gla00]. aseptic [CMC11]. Asian [Wil96, ENF^{+}23, RPD16, YWS^{+}24]. aspects [AY10, BGTA19, DBR15, FCH16a, FCH16b, IJ01, Moo12, PY97, WPD12, Har99b]. aspirations [CDS16, NBF22]. assemblage [AGO04, CGKSP13]. assemblages [dPAGGB16, ACB08, AF04, BP08, EBRLGB02, JLD^{+}24, LQW19, LLSDTJ09, MSR14, TJLLC10, Whi99, WA03]. assembly [JJC22]. assess [CCA^{+}23, GMS17]. assessed [CPM14]. Assessing [CDH^{+}23a, LE24, MMJ22, PMVA19, SAN^{+}23, TBA20, YKS14, BHP11, HWA14, dMV20]. Assessment [ALJ08, CCR21a, AMVV13, AVA19, Ano92j, Ano92v, AS96, CUT07, CDH^{+}23b, FCP19, GBR22, GS23, KHW09, LWS17, LS07b, LCC16, LFMP21, MLSP^{+}23, MGW93, OSB^{+}23, PLW17, PM14, PWM23, PH97, ROW^{+}23, SA12, SBZR17, TRJ24, VEK10, WM96, CCR21b, Gla00].

Assessments [Har99c, WCP97]. associated [AVA19, EBRLGB02, GSD18, PBMF12, PWC06]. associations [HW24, PMP21]. associative [FD00]. Assumption [CS05]. assumptions [Pau96]. Astacidae [GCSR09]. Astacus [HH04]. Astroblepidae [VE05]. Astroblepus [VE05]. Astyanax [CMA15, CCA17, PBMF12]. asymmetry [All11]. at-vessel [GH17]. Atlantic [Qui12, Ste05, dC98, AS95, Ano94f, ABK14, BGR21, BOs12, Ber93, BD20, BAA18, CS05, Car92, CGRCM19, CMA15, DHG18, FCCS15, Fle96, Gib93, GBR22, HK14, IJ01, IJ03, Jue95, LFdSRM16, LFMP21, MCL08, Man94, MT14, MP16, McQ97, MFS18, MAP21, MMF18, PQS14, dSSHK21, SFO14b, SFO14a, SAK14, SW14, TØH08, VVU22, VGA11, WGL14]. atlanticus [Jon07a]. atoll [PDN^{+}24]. Atractosteus [CFGG13, MAC02]. attract [Duc19]. attributes [AGO04, AMVV13, CSS20]. audiometry [LF13]. Auditory [LF13, CC22]. aurata [SZC11]. Australasia [CTM92]. Australia [BDS05, CJV13, FHvP14, GYH10, HFG07, HMV17, Ken95, LY07, Lig16, NWG07, RBC16a, RBC16b, SE16, SQR09, SLH11, ZFT13]. Australian [ABS15, BCK18, FCP19, Gau01, HFG07, HSC^{+}24, IWG17, NWG21, OSB^{+}23, PMV18, RTT12, YMR12]. australis [HP07, MP07b, SM07]. Australoheros [PBM14, PNC11]. Author [Ano92a, Ano99a, Ano00a, Ano02a, Ano02b, Ano04a, Ano05a, Ano05b]. authors [Ano01]. autopolyploidy [SVC21]. autotrophic [DBT15]. autumn [KZ07]. availability [vPKW18]. Avila [WBW09]. avoidance [RCP19]. away [OSC20]. axiom [Cad91]. ayraudi [MS13]. ayu [WIT14].

B [Bla98, JBJ06, Mul10, Nor02, Rei99, Web09, dC98, CCA17, NVA12, WZB^{+}24]. B. [Qui12]. back [DP12, Pau97]. bad [Fro99]. Bag [DC05]. Bahamas [How94]. bait [GH17, GCS20, LSF14]. baited [WFH17]. baits [LSF14]. Baja [RVRCB11, RCCASG02]. balance [WCP97]. Balanced [ZKvZ19]. balancing [AAJ21]. balao [TdSL15]. Balon [Woo99b]. Baltic [Soi99, Mad07, OK10, OK12]. balticus [OK10]. ban [CKN21, MTL21]. bandwagons [Dav96]. Bang [Hus04]. Bangladesh [IH04]. Bank [OSB07]. Barbed [Dav96]. barbel [KKW12]. barbs [SN00]. Barbus [KKW12]. barcode [LFdSRM16, SL07]. barcoding [DHB17, SKW22]. Barents [MMG18, DD15]. barnacle [PTOS^{+}24]. Barrier [CFB14, STP22, SBP07, BC03, PQS14, ZRZ20]. barriers [KZ20, MSM20]. basal [LG97, dSPPM12]. Based [Han05, AVL07, Asw05, BBS10, CCR^{+}24, CBR98, CCP13, CWN11, CGR11, EG03, ETWE12, FB05, GMS17, GOP^{+}23, HCB11, IH04, LFdSRM16, LM19, MMB13, MKS12, MLSP^{+}23, MSLN02, MMB22, MG18, OGR17, OK12, PGA^{+}24, PAK^{+}23, RSGS^{+}23a, RSGS^{+}23b, SBSŠ21, Sch18, TBB^{+}24, YWS^{+}24, vDBV02]. baseline [PE08]. basic [CZF22, PY97]. Basin [LQW19, VRRCAG02, CPN11, EGMM02, MKS12, BO12, SGD^{+}23]. basins [PBM14]. basis [MAC02]. bass [PV19, TP14, TRB13]. Bathygobius [LFdSRM16]. batoid [GLHD^{+}24a, GLHD^{+}24b]. Batoidea [VVU22, WSC09]. batoids [VVU22]. Bayesian [PH97]. be [CHFTV22, Gau01, Jel22a, Jel22b, McD10]. beak [SMNK^{+}23]. Bear [MLK13]. beavers [CG00]. become [GJ07]. beds [Whi17]. been [Ann96]. Behavior [Mol92, MJM15, RHP99, CGCG18, MS02, RSSS23, dSKV16, Fos08]. Behavioral [WLB^{+}23, AS08, ZLH^{+}23]. Behaviour [Est05, Gla00, ADC15, Ano92x, BCP22, BHK00, Bas93, Bre93, CHS18, DW93, FD00, HFP14, HWA14, HII16, JR97, Jue95, KC92, LSF14, McL94, MBK12, ODS10, Rob92, SHS14, Bar94]. behavioural [CCO20, JN18, MFS18, MHvH16, Woo98a]. behaviours [MFV19]. behind [BHB22, Pau96]. Belgian [LFM13]. Beller [WBW09]. belonging [CGK11]. below [SCC09]. Belt [AMVC20]. benefit [CFS^{+}23]. benefits [BO12, PQS14, WMDS18]. Benguela [Ano95b, YKS14]. Benin [TFE07]. benthic [CD01, SLH11]. Bergen [Hus04]. Bering [KZ07]. Beringian [CBF^{+}24]. Berit [Hus04]. Bermuda [Abl06]. best [CKN21, SJH22]. better [LCC19]. between [ABK14, BCP22, BCM20, BBW09, BSKB22, BZM12, CE96, HETS23, HII16, LKK10, LSH15, MFS18, MWS22, NvPV22, NVA12, OCB12, PBM14, PCG19, PQS14, SBN04, TRJ24, WIT14, WGL14, ZZ04, dMV20]. Beverton [SHH95, GLC98, Pau98, Pit98b, WK99, vPKW18]. Beyond [Pau98, CAC15, PAN22]. biased [Kva98, PDDDE21]. biases [Duc19]. Bibinagar [PR05]. bibliography [Wil07]. biennial [BP08]. Big [HFM19, Hus04, FN02, WFH17]. bigeye [FCCS15, MMF18]. bigger [MWS22]. billfish [KWM^{+}22]. bimaculatus [PBMF12]. biochemical [KdSD18]. biochronology [SG24]. Biodiversidad [WBW05]. Biodiversity [LQW19, CBEGR02, DMK^{+}24, GLC16, LM93, McD06, RPA13, RUL95, SPL12, SDJ13, WMTL22, ZWW20, JBJ06]. bioenergetics [OWW04, Woo95]. bioengineering [CDC09, Sol93]. biofouling [SM07]. biogeochemical [RSGS^{+}23a, RSGS^{+}23b]. biogeographical [BVM13, MKS12, PBM14]. Biogeography [CBF^{+}24, KDF13, QSBV18]. Biological [Ano92f, Fui99, Paw99, SCC09, Zim05, Ano93g, Ano95f, BOs12, PDB16, SPB00, SQR09, SMM94, WMM92]. Biological-physical [Ano92f]. biologically [MFV19]. biologists [ABK14]. Biology [Ano99b, ALJ08, BL00, BDS05, CM92, JBJ06, KGW10, KGW11, LAI05, Mar93, MJSOF16, Mil99, MCH95, Pyk05, QSBV18, Ram08, Sei08, Woo99b, AJM22, AY10, ARL12, AHL12, Ano92e, Ano95j, AAJ21, BP93, BGBE^{+}23, Bur10, CWO19, CGCG18, CJV13, FCCS15, FCH16a, FCH16b, GK18a, Han96, Hun21, JZ00, KMF13, LMCB^{+}23, Lis10, LPC^{+}24, MDF14, MBC21, MBA15, MLC17, Nel99, NMF17, Nor95, OSB07, PJ08, PZC17, Per93, PVJ07, PHH10, RPD16, RMN11, RM01, RTT12, SOB20, Tho92, TKR04, WPD12, WMM92, WSC09, YC20, ZFT13, JBS02, Zim05, Paw99, Rei99]. Biomanipulation [SQR09]. biomarkers [BCL21]. biomass [HMN17]. biomineralized [RSGS^{+}23a, RSGS^{+}23b]. biophysical [MVGW18]. bioregion [GP15]. Biotic [MÁMS05, MSLN02, TJLLC10]. Birstein [BL02]. Blache [MR18]. Black [CCP13, Paw99, BMF09, CTB16, RMK24, RBK13, Bar93]. Blackwell [Coc05, Els04]. bladder [dMlTB19]. bleak [LMCB^{+}23]. Bleeker [GLM20]. blennies [AS95]. blood [GAD10, Nor09]. Blue [AFBB23, BNC22, HK14, KWI07, MG18, NBF22]. bluefin [PAT15, VGA11]. bluefish [JBS02]. Bo [KCR17]. Bob [HNSS02]. Bockmann [dBdSA12]. bodies [WJP23, Bil02]. body [BSL18, FN02, MKK10, Nor14, dSSHK21, VSM12, ZJDZ10]. bolster [MKL^{+}23]. bonitos [JJMD13]. bony [FN02]. Book [Ano92b, Ano93b, Ano94a, Ano96b, Ano96c, Ano97b, Ano97a, Ano97c, Ano97d, Ano98a, Ano98b, Bar94, Bil02, Bla98, BFD91, Bro00b, Bur07, Coc05, Csi99, Dri05, Els04, FQSJ23, Fos08, Gil93a, Gla00, GCO03, Har99a, HC04, Har98, Har99b, HES95, Hil99, How94, Hus04, IBJ04, JBJ06, JCR95, KCR17, Lor93, Mil99, Mul09, Mul10, MAJ05, Nor02, Pav05, Paw99, Pis00, Qui12, Ram08, Rei99, RHP99, SHM96, Sei08, SMC95, Soi99, Ste05, Sul04, Tur99a, VAJ10, Web09, Wil07, WBW05, WBW09, Woo98a, Woo99b, Zim05, dC98, RSGS^{+}23a, RSGS^{+}23b]. Books [Ano93c, Ano95c, Ano96d, Ano96e, Ano98c, Ano00b, Ano92p]. border [IOHM23]. boreal [MLM19, MT18, SP98]. Boreman [dC98]. Born [BOV09a, BOV09b]. both [BCA^{+}23, EDP10]. bottom [MIA^{+}23b, MIA^{+}23a, VdSD21]. bottom-up [MIA^{+}23a]. bound [MTAP22]. boundaries [CAC15, Gau01]. boundary [BSWA14, HSS21, VBC^{+}23]. box [RMK24]. Brachyplatystoma [PBG04]. Brains [KVH98, Bas93]. branch [RSB16]. Branson [Mul09]. brasiliensis [TdSL15]. Bravo [CBEGR02]. Brazil [AMV13, AVB09, BO12, CMA15, ELC^{+}24, PBM14, PNC11, TdSL15]. Brazilian [FTH15, HKTS^{+}23, dSPPM12]. breadth [BSL18, SHHK21]. bream [IPT10]. breathing [PP92, PKF92, RT91]. Breeding [RBK10, CCMS12, KGF10, YWS^{+}24]. breviceps [Bea07]. bridging [BBW09]. brief [BTW15, YCT11]. bright [JLC21]. British [Ano93e]. broadcast [HTH00]. broadnose [SRBS21]. Broadtail [CUT07]. brood [SW95b]. broodstock [CB00, Tan03]. Brook [MT18, CBA10, DAN02]. Browman [Hus04]. Brown [Fos08, CTM92, GLC16, LPC^{+}24, Qui12, RBK13]. browsing [PFW16, PFW17]. Bruce [MBJ12]. Brünner [RMN11]. BRUVver [WFH17]. building [Ano95f, BNC22]. bulletin [Ano93e]. bullhead [CTB16, RBK13]. Bureau [Web09]. Burrow [AVB^{+}23, SA12]. Bycatch [AH96, Dia04, AA10, BCO21, Bro00a, GCS19, GCK21, GHMG22, GC23, GCB^{+}24, HGC17, HB19, JLD^{+}24, Ken95, KB21, LB23, MCL08, MTM22, MC12]. Bycatches [GK18b, Hal96]. Byczkowska [SW12].

C [Har99a, Soi99, VAJ10, WBW09, CNH22]. Ca [KKW12]. Cabo [MTM22]. Caddy [LFJ08]. Cadmium [SW10]. cage [Jue95]. cage-rearing [Jue95]. calamary [HP07]. calcium [PGC19]. California [RVRCB11, RCCASG02, BBS10, MVGW18, MC04, QGM15]. call [PAT15, TTC11]. camera [MBM06]. campechanus [EBS20]. camtschaticus [DD18]. can [CCA17, Dia04, GMS17, Jel22a, Jel22b, KLC18, KW08, NKF21]. Canada [Ber93, Gil93b, HBC13, MP16, MLK13, OAJ14, Gil93a]. Canadian [Pis00, BD20, Ano93e]. Canal [CGRCM19, MA08]. CanFishPass [HBC13]. Cannibalism [IK10, SR91, Smi93, SW98, DPS22, Kva98, NPS17, PAW17, SW95b]. capability [SX16]. capacity [MBH14]. Captive [How93, SAK14, CCMS12, ODS10]. captivity [HHCN11]. Capture [MBD07, Cal22, DCS11, DCS12, Kel23, TBK11]. Carangidae [AGG03]. Carassius [PEP12]. carbo [BMF09]. carbonate [PGC19]. Carcharhinus [DHG18, MG18, YC20]. carcharias [BKB17]. Carcharodon [BKB17]. care [AS95, CWB16, SW95a, Wis99]. career [NRJ^{+}23]. Caretta [AA10, MTM22]. Caribbean [MOF11]. Carl [MAJ05]. Carnation [Ano93e]. Carp [PR05, APLL07, CM92, NMS13, Vil18]. carp-SFF [NMS13]. carpio [Vil18]. carrying [RZV12]. cartilaginous [CC22]. cascade [CLX15]. cascaded [ZDJ19]. cascades [Bas93]. Case [Asw05, ABS15, AS95, Ano94d, ASS^{+}23, BCA^{+}23, BCK18, CDC09, CS04, CNT21, CTM92, ENF^{+}23, EBRLGB02, Gau01, KL23, MGB15, MSV^{+}23, SPB00, SLH11, VaR92, Whi99, WM02, ZLH^{+}23]. Casellas [WBW09]. Caspian [ADS11, GBOK23, RSSS23]. Castor [CG00]. catadromous [AWC17]. Catalog [Gre93a]. Catch [Ano92c, BB05, Bis06, GH17]. Catch-and-Release [BB05]. Catchability [AS96, War08]. catches [CKN21, LC02]. catecholamines [PKF92, RT91]. Categorising [GKC19]. categorization [VE05]. category [KL23]. catfish [AKGB11, CHS18, PBG04, RPD16, VE05, dBdSA12]. caudal [Mun95]. Caught [Gou16, MIA^{+}23a, FTH15, MIA^{+}23b]. cause [Kva98]. Causes [WGC18, HAC18]. caution [PAT15]. caveats [GC23]. CBE [SHH95]. cell [BEBFM14, MGL^{+}23, Vil03]. cells [JB95, Kot91]. cellular [SZG10]. Center [RBK10]. Central [CCR21a, CCR21b, CCR^{+}24, CUT07, EBRLGB02, MSLN02, RSB16, GP15, MCN12, MC04]. Century [STW^{+}24, BGHC23, BH17, Cad99, OSC20, Smi98, Utt94]. cepedianus [SRBS21]. Cephalopod [CR98, NWG07, JBX07, PGG07a, PGG07b, SBSŠ21, SPS07, ZHP07]. Cephalopoda [Bol07, HLD07, OJB07, PDGBR07, SBP07, SL07, YVR07]. cephalopods [IK10, IRPG21, KZ07, LY07, MHW07, SMNK^{+}23, YVR07, dlCFP19]. cephalus [WPD12]. Cerame [WBW09]. Cerame-Vivas [WBW09]. Ceresole [KGW10, KGW11]. CERF [BP08]. cernua [GH16]. cetacean [CCR21a, CCR21b]. cetaceans [Ano95e, BHB22]. chain [APP^{+}23, SRB^{+}23]. challenges [COGFPV23, FD22, IRPG21, IPT10, KWM^{+}22, LE24, NAP22a, NAP22b, NRJ^{+}23, PBR19, ROL14, SK15, SDJ13, SBG17, STW^{+}24, TKB18, dMlTB19, GOP^{+}23]. chalumnae [Tho92]. Chamelea [CCR^{+}24].

Champotón [LLSDTJ09]. Change [CR98, Csi99, Shi05, ASS^{+}23, BGR21, BTZ^{+}23, Bis06, CBL17, FMH07, FCP19, Fri04, FHvP14, GYH10, HP14, HCvP16, IDG16, JLC21, Jon07a, JPC14, KPN21, KSK17, LLJC24, MBH14, MGL^{+}23, MLP17, OAJ14, Pau10, PLW17, PJ08, PHB14, PGJ15, PCW17, RWH04, Sot02, VD22, WWS17, ZCW19]. Changes [EBRLGB02, Woo99a, ADC15, ACB08, ACST17, CBEGR02, Cra92, Cus94, GAD10, MSR14, MTPR15, OV24, PBF15, WH07, YKS14, ZLH^{+}23]. changing [BSWA14, CTL17, HJC20, PTOS^{+}24, PBS14, SPL12, TFP22]. Channel [CBO19]. char [BD20]. characid [MPO11, PCG19]. Characidae [CMA15, CCA17, PBMF12, dSPPM12]. Characiformes [BLF10, MBdBC13, PBMF12, dRRGC12]. characterisation [dSPPM12]. Characteristics [WLB13, ARL12, LSF14, MS13, NWO16, Nor12, RBK13, SN00, WIT14]. Characterization [BLF10, dPAGGB16, BEBFM14, CPN11, OBS08, SZC11]. Characterizing [JLD^{+}24, LM19, LZC22]. Charles [Lac02]. charr [CBA10, MBK12]. chart [ATdLCBR02]. Check [Car92]. Check-list [Car92]. Checklist [Abl06, DPC12, RCCASG02]. Cheilinus [SKD03]. chemical [GGZ10, Har94, Kot91]. chemistry [DMD21, EG03, MDM20]. chemoreception [Dou93]. chemosensory [Kot91]. Chile [APLM14]. Chilean [LC02]. Chimaeriformes [LG97]. chimaeroid [Lis10]. China [Ara15, TCS20, HSC20, KPN21, LQW19, LZC22, LLJC24, MTL21, XSC15, ZDJ19, ZWW20]. chinensis [STS07]. Chinese [CDH^{+}23a, FHK04, KLC18, WH24]. Chinook [BCK16, SAK14]. Chionoecetes [MDR14, MMG18]. chloride [JB95]. Choice [Ano92v, RHP99, ANB19, DW93, EP00]. chokka [MS07, OSB07]. chondrichthyan [TRJ24]. chondrichthyans [FCOCNC24, LBS19, SBG17]. Chondrichthyes [Lis10, LSB22, LG97]. Choosing [MBS17]. chromatic [Bur10]. chromatin [CZF22]. Chromosomal [BSV11, CPN11, MdAVA11, MCN12, PNC11, SGF13, BVMF13, MPO11, PASF13]. chromosome [BLF10, BVMF13, BZM12, CCA17, MBdBC13, MdAVA11, PGA^{+}24, RZV12]. chromosomes [Ano92o, BZM12, CPN11, NVA12]. chronicles [Ter01, Ter11, Ter18]. chum [OWW04]. cichlid [WMG00]. Cichlidae [PGA^{+}24, PBM14, Pau10, PNC11, RL05, RTT12]. cichlids [PGA^{+}24, SGF13]. Cichliformes [PGA^{+}24]. Cichlinae [PGA^{+}24]. circadian [Ree02]. circulating [PKF92, PSK12]. circulation [Ran92]. Circumpolar [MWP^{+}23]. City [TFE07, Zim05]. clam [CCR^{+}24, ROW^{+}23]. Clarifying [SSBCL11]. classes [HFG07, PASF13]. classical [Kim93, Kim93]. classification [BCO21, PGA^{+}24]. Cleaner [WSGP22, Pou93]. cleaning [NVH21, Pou93]. Climate [CBL17, HYW13, KPN21, MHP12, BGR21, BTZ^{+}23, CTL17, Cus94, DBT15, FMH07, FCP19, FHvP14, GYH10, HCvP16, JLC21, JPC14, KSK17, LLJC24, MTAP22, MBH14, MLP17, OAJ14, PLW17, PJ08, PHB14, PTOS^{+}24, PGJ15, PCW17, RWH04, SHHK21, Sot02, VD22, WWS17, ZCW19]. climate-driven [MTAP22]. climate-productivity [SHHK21]. climatic [Cus94, KMK16, RSSS23]. clines [TV15]. CLOFETA [Car92]. cloning [TRB13]. closed [JCL07, Ros00]. closely [KL23]. closures [AGJ14]. clues [Col10]. clupeid [Mar93]. Clupeidae [CPM14]. Clupeiformes [RG99]. Clupeoid [CM98]. Co [CBHOH19, DC05, CD01]. Co-evolution [CBHOH19]. Co-management [DC05, CD01]. coast [BOs12, MS07]. Coastal [ADC15, Bla98, BP08, CM98, LFJ08, ABS15, Ano93e, AFBB23, BNC22, CJV13, DCS20, FHK04, HCvP16, IH04, MTL21, MTPR15, PGJ15, RSSS23, RVRCB11, SE16, Bla98]. coastal-dependent [ABS15]. coasts [PTOS^{+}24]. Coco [WBW05]. cod [IJ03, Mad07, MFS18, SQR09, WM96]. coding [DZZZ22]. Coelacanth [Rei93]. coelacanths [Tho92]. coelom [CMC11]. coexistence [NKF21]. Cognition [Fos08]. cognitive [DW93]. Coho [SBN04]. COI [MPO11]. coindetii [CUT07]. cold [IPT10]. cold-induced [IPT10]. coldwater [BTW15]. Collaboration [ABK14, CDS16]. Collaborative [DDD16]. collapse [MCFC21, WM96, dC98]. Collected [Ano95h]. Collection [Woo99a, DML^{+}24, ENP18, MBD07, PRP16]. collective [RFH15]. Collett [Jon07a]. Colloques [Woo99a]. colonial [RR11]. colonization [PQS14]. Color [RHP99, Ste05]. Colorado [VRRCAG02]. colour [ODS10]. Columbia [Ano93e]. column [BBS17]. comb [OJB07]. combined [SSBCL11]. Combining [MN98]. comment [Dav96, Kva98]. comments [CML19, MS02]. Commercial [DD15, BMN21, Cal22, CBHOH19, DMK^{+}24, HFM19, HB19, HP07, MGTD18, MCFC21, MML18, MG18, NWG07, ROW^{+}23, WP96]. commercially [DSU08, NWG21]. committees [CWB16]. common [APP^{+}23, CWD11, CBH15, Kot91, OV24, OBS08, Vil18, Gil93a]. Commonwealth [CBH15]. communities [BSWA14, Cra92, DROCM23, EGMM02, HCvP16, JCL07, MTL21, Nor03, SAC^{+}23]. community [AMVC20, BHS19, JJC22, KHW09]. Comparative [BSL18, HK14, Nor95, TMD^{+}24, TFF09, MBdBC13, UIA11]. Comparing [CTL17, LM19]. Comparison [GKR10, WT08, AKGB11, EAW22, MWS22, MBM06, RBK13, VGMAGB02, WIT14]. Comparisons [SBN04, Nor03, WGL14]. compatible [AFB15]. compensate [CCA17]. compensatory [MKK10]. competing [OCB12]. Competition [BHS19]. competitive [RN04]. Complementary [Smi98, MKL^{+}23]. complete [TTC11, Tur95]. Complex [GOW13, HII16, Bis06, CMA15, CCA17, GK18a, HDI15, MCD09a, WPD12]. Complexities [ELC^{+}24]. complexity [DROCM23, Sto00]. Compliance [ENF^{+}23]. components [CDH^{+}23b, Nor09]. Composition [SE16, Wie96, CGKSP13, CRZHZ02, Cra92, GGZ10, KKW12, LY07, STP22, ZJDZ10]. compounds [HP93]. computational [RNJ16]. concentration [KKM10, LS07a]. Concept [Ano99b, CH94, CH99, Nel99, ZKvZ19]. Concepts [Lar96, MSH14, Pau10, dP99]. conceptual [YCT11]. concerning [GCSR09, PDA12, ZRZ20]. concerns [SL07]. Conchos [EGMM02]. conclusions [Sny05]. condensation [CZF22]. condition [FC07, KKW12]. conditioners [HCB11]. conditions [GKR10, HJC09, KGF10, RSSS23]. Conducting [BLC19]. Conference [Ano96f, Ano97e, Ano98d, BJ00, CM97, DLP95, Ano94e, Hus04]. conflicting [CBR98]. conflicts [GCS19, Kel23]. Confronting [Mil99, MBH14]. Congo [CCA^{+}23]. Congress [GB23]. Connecting [KEP22]. connections [MP16]. Connectivity [BBW09, MWP^{+}23, PDM20, SVRS19]. consequence [JPC14]. consequences [HAC18, WGC18]. Conservation [EGMM02, EN00, JBJ06, Ram08, SBZR17, AGMB^{+}23, Ano95f, Ara15, AVB09, Bea07, BJM20, BL00, CDC09, CSA11, CSA12, CDH^{+}23a, CGCG18, GK18a, GMP12, GCUR22, GLM20, HPdL02, KHW09, Kel23, LCC16, LBS19, MV13, MCFC21, MT14, MMF18, MBA15, OK14, PAK^{+}23, PBR19, QGM15, RPD16, RSSS23, SPL12, SJH22, Tay99, Utt04, VRRCAG02, WGL17, WMTL22, YC20, ZWW20]. conservatism [Cor02]. conserve [GLC16]. consideration [HJC20]. Considerations [GdMD13, BPA^{+}23, BLC19, CWD11, GLM20, JN18, MHW07, SL07]. considering [DROCM23]. constant [Cad91]. constructions [ZDJ19]. consumer [HETS23]. consumption [Nor14, OWW04]. contemporary [SKU20]. content [IRPG21]. Contents [Ano99c, Ano00c, Ano04b, Ano05f, Ano02c, Ano02d, Ano05e, GSD18, YWS^{+}24]. context [FSB14, MSH14]. Contextualising [Sur23]. contextualizing [LSB22]. Continental [Bla98, BMF09, Gil93b, HMN17, MMG18]. continents [Nor03]. continuous [SMN08]. continuum [AWC17, MP07b]. Contrasting [NWG21, PDN^{+}24]. contrasts [RCP19]. contribute [Kos09]. contributed [Ara14]. contributing [MDR14]. Contribution [PAK^{+}23, MBA15]. Contributions [GCO03, REA^{+}23]. control [ARL12, AHL12, CM92, GLG12, GVB94, MSM20, MZ00, PKF92, RT91, RTT12, SMS12, SQR09, SMM94, Vil03]. controversial [KLC18, TdSL15]. controversy [Ter11]. conventional [BBW09, Jon07b]. cool [BTW15, JCL07, McD06]. cool- [BTW15]. cool-temperate [JCL07, McD06]. Cooper [Zim05]. coordinated [KGF10]. copepods [AYN11]. copper [SW10]. copulatory [HII16]. Coral [FCH16a, FCH16b, MGW93, SKD03, AGJ14, Ano92h, Ara15, BCM20, CWP14, CBH15, CBHOH19, PDN^{+}24, PBF15, PCW17, PFW16, PFW17, STP22, SHB21, TSB22, TBB^{+}24, SBP07, YMR12]. coral-dwelling [SHB21]. coral-reef [CBH15, CBHOH19]. Correction [CCR21b, DPS22, GLHD^{+}24a, HCMT24a, Jel22a, MIA^{+}23b, MMM22a, NAP22a, RSGS^{+}23a, SVM^{+}24a]. correlation [BZM12]. Correlations [Mye98, CPN11]. corridor [RSB16]. Corrigendum [Ano92d, Ano93d, Ano98e, Ano05c]. Cortisol [MVM99]. cosmopolitan [WPD12]. Costa [WBW05]. costs [SW95a]. countergradient [TV15]. counting [SA12, WPF16]. countries [LBS19, Rai94]. Covelo [Ste05]. Cover [LFJ08, Pit98a]. COVID [BPA^{+}23, LSDRS23, OSB^{+}23]. COVID-19 [BPA^{+}23, OSB^{+}23]. Coward [Pis00]. Cowx [Hil99]. CPE [APLL07]. crab [AAH98, DD15, DD18, MDR14, MMG18, SBZR17]. crabs [GGL14]. Crafting [Bro05]. Craig [Dri05]. crayfish [GCSR09, HH04, HH06, KZ20, OSC20, ST94, WR16]. credit [Sur23]. Creek [Ano93e]. Creel [Ano94b]. Crenicichla [PNC11]. Crenicichlina [PGA^{+}24]. Crevice [MS02]. Crisis [DC05, McD06]. critical [BSKB22, CRD00, CJV13, CHN18, Fro99, PHK20, SDS15]. croaker [LLC11, QSBV18]. cross [GCS19, IOHM23, VBC^{+}23]. cross-border [IOHM23]. cross-boundary [VBC^{+}23]. cross-taxa [GCS19]. crossed [CGRCM19]. Crowder [JBJ06]. Crustacean [AAH98, CSY12]. crustaceans [SCC09]. Crying [McD06]. cryopreservation [AHL12, LLC11]. Cryptic [DMK^{+}24]. crystal [PGC19]. Crystalline [SBN04]. Ctenopharyngodon [CM92]. Cuban [CFGG13]. cucumber [Gou16]. cue [BLD19]. cultural [LBS^{+}23, MP16, RBC16a, RBC16b]. Culture [KGW10, KGW11, COGFPV23, FIØ04, IPT10, IFW05, MAC02, MKK10, NMS13, Sin93, Vil03]. cultured [DPS22, Fer94b, KKF10, NPS17, SZC11, Utt04, WPF16]. Culum [Fos08]. cumulative [BTZ^{+}23]. cuneata [HHCM10, HHCN11]. Current [PTP14, YKS14, CHFTV22, DT04, Fer94a, FCP19, FD22, HPdL02, HSS21, IRPG21, KB14, KHW09, LE24, LF03, LSDH12, MHP12, MC12, MBK12, OGR17, PZC17, PVJ07, PHH10, PDB16, TP14]. currents [BSWA14]. curves [MF99]. Customary [Asw05, Lig16]. cutthroat [DAN02, LPC^{+}24]. cuttle [AVL07]. cuttlefish [HFG07]. Cuvier [APLL07, OBS08]. CWT [SMN08]. cycle [LET^{+}23, MS07, SLK16]. cylindrospermopsin [SDG15]. Cyprinella [May02, WM02]. Cyprinid [Ano92e]. Cyprinidae [CM92, MS02, May02, TV15, WM02]. cyprinids [Per93]. Cyprinodon [VRRCAG02]. Cyprinodontiformes [Web02]. cyprinodontoid [Nor06, Nor14]. Cyprinus [Vil18]. Cyrino [Web09]. Cytogenetic [dSPPM12, LFdSRM16]. Cytonuclear [Avi00, SPB00].

D [Har98, MAJ05, Paw99, Soi99, Web09, WBW09, Woo99b, JJMD13]. D.C [Dri05]. dagaa [IOHM23]. Dahlgren [Dri05]. daily [MS07]. d’Alba [KGW10, KGW11]. Dale [Soi99]. Dam [QGM15, AGO04, CLX15, Kel23, ZDJ19, LQW19]. damage [GAD10, VdSD21, ZRZ20]. dams [AVA19]. damselfishes [SHB21]. Daniel [Bil02, Ste05]. Danube [JEL10]. Data [PB98, AVL07, Bag11, BBJ12, BSV11, BM09, Bis06, BTW15, CCA^{+}23, CS04, CPM14, CDH^{+}23b, ENP18, GLHD^{+}24a, GLHD^{+}24b, JNW14, Jon07a, LFdSRM16, MBC21, MTC19, MHvH16, OK14, PRP16, SW12, SMN08, WNB12, WMA20, Mil99]. data-less [CCA^{+}23]. data-limited [CS04, CDH^{+}23b, WNB12]. Data-sparse [PB98]. database [GPS18, HBC13, HSS21, SL07]. datasets [WMA20]. David [Bro00b, Bur07, BOJ16]. dead [KM20]. dealing [CWB16]. Death [Cad91]. d’eau [Bil02]. debris [dMV20]. Decade [NAP22a, NAP22b, CHN18, PAN22]. decades [MC16, Utt04]. decapodiform [YVR07]. decision [Ano95f, BJM20, GHMG22, OSB^{+}23, PH97]. Decline [HTH00, AAH98, CBH15, LQW19, MDR14, PDA12, PWM23]. declining [CLB^{+}22, CBHOH19]. deconstructing [CDH^{+}23b]. Dedication [HNSS02]. Deep [MGRRJ21, NBF22, AVB^{+}23, CDH^{+}23a, NWO16, SBG17]. deep-sea [SBG17]. deep-water [AVB^{+}23, NWO16]. deeper [CDS16]. deepwater [BGOW23, WNB12]. defence [SZG10]. defences [FM94]. Defenses [MÁMS05]. deficiency [HAC18]. Defining [Knu01, MSH14]. definitions [DP12]. deformities [MKK10]. dehydrogenase [OBS08]. del [WBW05]. delineation [MKS12]. delivery [MZ00]. demand [HETS23]. demersal [Ken95, KB21, KCE12, MTL21, NWO16, WNB12, YKS14]. demographic [DHG18, LMH21]. demography [TMD^{+}24]. density [CRD00, VSM12]. density-dependent [CRD00, VSM12]. Dentex [MDF14, RM01]. dependent [ABS15, Bis06, CRD00, HCvP16, HSC^{+}24, OCB12, VSM12, WGC18]. depleted [CA04]. depletion [AAH98, PBG04]. depredation [MCL08, MML18, MDV^{+}23, TBA20]. depth [AMVC20, BSL18, STP22]. DeSalle [BL02]. describe [SKU20]. described [Wil07]. desert [VRRCAG02]. Design [Dri05, WMW18]. designation [GGZ10]. Designing [PLW17, FAN22, MVGW18]. detection [KCF14, MGRRJ21]. Detective [Mil99]. determinants [LM19]. Determination [DJ05, CLB^{+}22, LE24, SN10, Vil18]. determined [LY07]. determines [CSS20]. Determining [MS07, IWG17]. deterrents [LB23, PM19]. detoxificant [EDP10]. Developing [Bla98, NAP22b, ASS^{+}23, BEBFM14, BPA^{+}23, PWB99, WNB12, NAP22a]. Development [DML^{+}24, FHK04, FM94, HFM19, ARL12, Bro00a, BC03, CFGG13, CCO20, Gou16, GPS18, HHCM10, LZC22, LSF14, LBS19, MP07a, MAC02, PM09, RMK24, SX16, TS96, WOR09, WGR20, ZWW20, NAP22a, NAP22b]. developmental [BOV09a, BOV09b, LKH16]. Developments [PM94, Woo99b]. device [DT04]. devices [SMR11, SOB20]. devil [AUdS18]. diadromy [McD97]. diagnosis [Nee01]. diamondback [DMK^{+}24]. Diana [Zim05]. diapause [WAS12]. diaspros [RMN11]. Diazinon [MPK10]. Dicentrarchus [TRB13]. dichroma [MS02]. Dicologoglossa [HHCM10, HHCN11]. dictionary [Ano92b]. Diel [AE22, Ree02]. diet [Bea07, Jon07a, KKW12, KBV11, PBF15, PWM23, TRG19, WMM92]. Dietary [SZG10, dSKV16, BSL18, CNH22, KTW12, ROL14, SHPH^{+}22, ZJDZ10]. diets [LY07, MFV19]. differences [LBS19, PCG19, SYL^{+}23]. different [CMC11, KGF10, WLB13, ZJDZ10]. Differential [APLM14, ZCT02, AGJ14, WMG00]. differentiate [GP15, TMKC24]. Differentiated [NLFM12, SZC11]. Differentiation [MBdBC13, BVMF13, DD13a, DD13b, KKF10, MPO11, MGL^{+}23, NLFM12, SN10, ZDJ19]. digestion [GP12]. digestive [CFGG13, GP12, HHCM10, Web09, WOR09]. Diggles [TBK11]. dimensions [HCNH24]. dimorphism [EAW22]. Dionda [MS02]. dipeptide [KTW12]. diploid [BZM12]. direct [Koc01, LS07b]. directions [DT04, GLHD^{+}24a, GLHD^{+}24b, GOP^{+}23, HB19, IRPG21, MHP12, MDV^{+}23, SBG17, YC20]. directly [ZLH^{+}23]. discard [CSS20, CKN21]. discarded [SEH24]. discarding [BHP15]. discards [MMJ22, RCP19]. discrimination [BB96, CNH22]. Disease [Gau01, EEDP18, VaR92, Vil03, YWS^{+}24]. Disease-translocation [Gau01]. diseases [Mun93, NVH21, Lor93]. Disentangling [TdSL15, HKTS^{+}23]. Disintegration [LS07a]. disorders [RHM11]. disparity [ZL23]. dispersal [BD20, FHW^{+}24, McD10, PDDDE21, Sto00]. dispersion [PBMF12]. displayed [BCK18]. disputes [CH99]. Disruption [Bro00b, SRB^{+}23]. disruptors [GELL16]. distinct [MPO11]. Distribution [ACB08, BKB17, EN00, Gla00, KZ07, PWM23, VRRCAG02, VRR12, Bol07, CZF22, Duc19, FHW^{+}24, FCCS15, HMN17, HSS21, JPC14, LY07, LLSDTJ09, MWP^{+}23, OV24, PMP21, PBS14, RCP19, RTT12, SAT11, TV15, TJLLC10, YKS14, ZHP07]. distributional [RCCASG02, SEA22]. distributions [BGOW23, GPS18, MBP14]. ditropis [KWI07]. divergences [Jon02]. diverse [CNT21, Joy23, MFV19, SJH22, Smi93]. diversification [BVM13, LFdSRM16]. Diversity [Ara15, BCK16, CMA15, CZF22, HSC20, Ano95f, AVB09, BO12, CRZHZ02, CCH18, FSSM17, Har94, JJC22, LSP11, LCM22, MMB13, MFS18, MCN12, NWG07, PGC19, STP22, SBP07, TFE07, ZCT02]. dividing [PBM14]. diving [HMV17]. DNA [AHW04, CGL22, DHB17, LFdSRM16, OW97, PASF13, SKU20, SGF13, SKW22, TdSL15]. Do [GCUR22, Mye98, ST94, YH18, Dav94, KW08, MV96, Ros00, SD91]. documented [MLP17]. Does [Asw05, HKL10, CCMS12, MWS22, PM09, WHS04, ZLF21]. dogfish [DBR15]. dolphin [Hal98]. Dolphins [Pie93]. Dombeck [Har99a]. domestication [TFF09]. dominance [RN04]. dominant [NKF21, PEP12]. Domingo [WBW05]. don’t [MC04]. dosage [APLL07]. Dosidicus [SKU20]. Doukakis [BL02]. down [MIA^{+}23b, MIA^{+}23a]. downstream [AVA19, McD09b]. Dr [Ano94c]. Dr. [HNSS02]. drag [MV96]. drift [LKH16]. drifting [WH24]. Drive [Pit98a, ZDJ19]. driven [MTAP22]. drivers [AGM19, EBS20, HSC20, SMNK^{+}23, WJP23]. drives [GSC17]. drought [LCC19]. drought-stricken [LCC19]. Drug [RT05, BTW15]. Dublin [Har99b]. Duffy [WBW09]. Dumeril [AY10]. duplication [ZCT02]. duration [HMN17]. during [BGR21, CMC11, JPC14, LNT^{+}23, Mul11, NAP22a, NAP22b, WOR09, vOR15]. dusky [CGCG18]. duvauceli [STS07]. dwelling [SHB21, VSM12, VCD12]. dynamic [CAC15, Mar08, WCP97]. Dynamics [Ano92f, BDS05, MMG18, Ano92s, Ano92v, Ano94k, BdST16, CSA11, CSA12, CBHOH19, GHF98, GLM20, HJC09, JBX07, LFM13, Mag13, McD10, MVM99, MRMJJ17, Ros00, SKU20, SG24, VSM12, ZMG17, Gla00].

E. [Mul09]. Eagle [CCMS12]. Early [Ano92g, GCO03, McD09b, BHP11, FC07, LKH16, NRJ^{+}23, PHB14, SRD93, VMB07, MBJ12, Soi99]. early-career [NRJ^{+}23]. East [Ano94a, ZZ04, CJV13, FHvP14, TCS20]. Eastern [CGRCM19, RMN11, BP08, Car92, CGKSP13, GYH10, LY07, MJSOF16, Nor03, Nor06, ZFT13, LKK10, MMG18, SKU20]. eat [MWS22]. EC [MPK10]. eco [ZLF21]. eco-physiology [ZLF21]. Ecological [BCM20, BAA18, CR98, GYH10, Mil99, MVGW18, SFO14b, SFO14a, Ano94o, AFB15, CSS20, DT04, DIR20, EBS20, FCOCNC24, Hal98, HSC20, IDG16, JN18, Kam02, Kos09, MCD09a, MAP21, MHvH16, MSH14, SMNK^{+}23, SCC09, SBG17, SLH11, SRB^{+}23, TCS20, VEK10, WA03, dlCFP19]. Ecology [AGMB^{+}23, CHS18, DCS11, DCS12, JZ00, MAJ05, OK10, PVJ07, Rei99, RHP99, SMR11, Smi93, TBK11, Woo98a, Woo98b, Zim05, Ano92h, BKB17, BBV12a, BBV12b, BOJ16, CG00, CGCG18, CGP^{+}24, CJV13, DMD21, FQSJ23, Fle96, GBOK23, Gre93b, GLM20, HMQ16, HK14, Hun21, IH04, JBX07, JBS02, KMF13, LBS^{+}23, LPC^{+}24, MTPR15, Moo12, MBA15, NMS13, NWO16, NTC12, PJ08, PZC17, PHH10, QSBV18, RTT12, SRGS04, SHPH^{+}22, VE05, VBC^{+}23, WPD12, Qui12]. Economic [Har99b, BO12, DBR15, HPY18, HMV17, LBS19, MSR03, TCS20]. Economics [BDS05]. economies [DPV^{+}23]. economy [AFBB23, BNC22, NBF22]. Ecopath [CAC15]. Ecosystem [OGR17, ANL12, Ano93e, BBS10, BAA18, CCR^{+}24, CCP13, ETWE12, GYH10, Lar96, LM19, PMVA19, RSSS23, RSGS^{+}23a, RSGS^{+}23b, SBtIWG12, WZB^{+}24, YMR12, ZCW19, ZMG17, dlCFP19, TCS20, YKS14]. ecosystem-based [BBS10, CCR^{+}24, CCP13, ETWE12, LM19, OGR17, RSGS^{+}23a, RSGS^{+}23b]. Ecosystems [Ste05, Ano92f, BBS10, CAC15, CG00, DDD16, DAN02, GOW13, Jia09, KPN21, LSP11, LM19, OAJ14, SLH11, TBA20, VAJ10, WCP97, WMTL22]. ectotherms [HKN18]. Ed [Sei08, Bil02, Mul09, Sul04, VAJ10, Woo98a]. Edited [Csi99, Har99a, Hil99, Paw99, Pis00, Soi99, Tur99a, Woo99b, Coc05]. Edition [Mul10, Sul04, ALJ08, Coc05, Zim05, Bar94, Gil93a, How94, Lor93]. Editor [Woo99b, Ben06, BDD02, TKR06]. Editorial [Ano91a, Ano92i, Ano96g, Ano96h, CP94, LAS17, Nel99, Pit98b]. edn [Ano93b]. eDNA [JLD^{+}24]. eds [Bla98, Dri05, Els04, GCO03, Har98, Hus04, JBJ06, KCR17, Ram08, Web09, dC98]. Edward [Csi99]. Eel [MR18, Shi05, Sin93, Ara14, CLB^{+}22, HWA14, MJSOF16, SAT11, VEK10, WJP23, vGM05, Coc05]. Eels [DJ05, BLD19, Jel22a, Jel22b, PHK20]. Effect [AA20, GCS20, KKM10, KTW12, SW95b, VdSD21]. effective [AH96, RP91]. Effects [Ano92j, BHK00, DHG18, Els04, GCM18, Gla00, GELL16, HMN17, HP07, IE06, JR97, KKW12, KdSD18, KSK17, OAJ14, PCW17, RWH04, ZJDZ10, ZCW19, ZMG17, vOR15, dPAGGB16, AGO04, Ano92o, AFB15, BTZ^{+}23, BHP11, CNH22, CBB19, CLX15, CG00, CWN11, Cus94, EP00, EBS20, GH17, GCSR09, GGL14, GYH10, HPP15, Kim95, MHP12, MLP17, MC04, PSK12, ROP^{+}23, SRD93, Sny05, SDG15, SM07, VLD20, dJFS20, Els04]. Efficacy [GZT13]. efficiency [CML19]. efficient [Jue95, LSF14]. effort [Ano92c, Bis06, Duc19]. Egg [BTS97, CBP02, Kam05, KKM10, NRCD^{+}23, SM07]. eggs [GCSR09].

Egypt [SK15]. EIFAC [Har99b]. Eigenmann [GdMD13]. eight [Ann96]. eighty [LFM13]. elasmobranch [WGR20]. Elasmobranchs [BDF09, Moo12, FN02, GKC19, PDDDE21, WAS12, WLB^{+}23]. elasticity [WGL17]. electric [MJSOF16, MCN12]. electricus [MJSOF16]. Electrocommunication [Mol92]. electrofishing [Sny05]. electronic [BEC11, CMC11, CWN11, CWD11, ENP18, HL11, LR11, Mul11, OCE11, WCD11]. Electrophorus [MJSOF16]. Elemental [ZHP07]. elements [KKW12]. eleotrid [Nor12]. Elliott [JBJ06]. Elsevier [Ano92b]. elusive [BBJ12]. embryo [SM07, VMB07]. Embryonic [WAS12]. embryos [MPK10]. emergence [AVB^{+}23]. Emerging [dMlTB19, HJC20]. emotion [AH96]. Emphasis [HL98b, CD01, Koc01, RTT12, TØH08, WPD12]. Empirical [War08, Avi00, BHS19, Pat92, SJB^{+}23]. employed [LR11]. Empowering [FMM22]. Endangered [BP16, ABK14, CPM14, GK18b, PDM20]. endangerment [VE05]. endemic [May02, RVRCB11]. endocrine [CE96, GELL16, Bro00b]. endogenous [WOR09]. endurance [VW91]. Energetics [Kam05, Ano92g]. engagement [KEP22, SEA22]. England [Ber93, JNW14, KSK17]. English [Ano92b, GBR22]. enhance [ABK14, TAD14]. enhancement [ADS11, Ara14, DGV11, GJA17, LNT^{+}23, LSL03, MLN05a, MLN05b, Mus03]. Enhancing [VBC^{+}23]. enigma [Jel22a, Jel22b]. enigmatic [May02]. entanglement [HB19]. Entosphenus [CWO19, JMJ19]. entrainment [CCO20]. entry [CMC11]. Environment [CM98, Mye98, Ano92r, IJ01, KDF13, MGRRJ21, MVGW18, Nor12, ZLH^{+}23]. Environmental [FHW^{+}24, GGL14, KW08, Nor09, Per93, RT05, Woo99b, ADC15, ASS^{+}23, CLB^{+}22, DIR20, EG03, EBS20, GLG12, HW24, IDG16, KCR17, LMCB^{+}23, MBS17, PMP21, SHS14, SGD^{+}23, WHS04]. Environmentalist [FB05]. environments [GKC19, Nor06, PCG19, WG19]. epigenetic [MGL^{+}23]. Epinephelus [CGCG18]. epipelagic [AGM19, AE22, KZ07]. Epizootic [KB14]. equilibrium [GLG12]. Equity [AFO22]. era [CBF^{+}24, Knu01]. Erratum [Ano05d, BBV12a, CSA12, FCH16a, KGW11, PFW17, RBC16a, SFO14b]. errors [AMV13, CWD11]. Erythrinidae [BLF10, MBdBC13, dRRGC12]. Erythrinus [MBdBC13]. escape [MV96, Rob92]. Eschscholtz [HH04]. essay [Han96]. essays [Ano95h]. Essential [FD22]. establishing [KGF10]. establishment [BD20]. estimates [BCO21, HWA14, War08]. Estimating [Bag11, BSR19, HPY18, LMH21, MF99, PB98, WMA20, CKN21, LH18]. estimation [CBH15, GLG12, GIT09]. estuaries [BP08, GCUR22, HWA14, JCL07, Nor03, Nor06, WGL14, Whi99, Whi17, Rei99]. Estuarine [BP08, CRZHZ02, CGP^{+}24, GCUR22, HK14, Nor03, RWH04, WTC06]. Estuary [SVM^{+}24a, SVM^{+}24b, HW24, MMB13, WGC18, ZCW19]. estuary-dependent [WGC18]. Ethical [MHW07, BEC11]. Ethics [Pis00]. ethology [DW93]. Eugene [Woo99b]. Eugerres [APLL07]. Eurasian [TV15]. Europe [DGV11, HMQ16]. European [AKGB11, BOs12, CHFTV22, CTB16, CHS18, GP15, LAI05, MBA15, Nor95, Per93, RSB16, SW12, ST94, TRB13, vGM05]. Euryhaline [RT05, CGP^{+}24]. EUS [KB14]. eutrophic [SQR09]. Eutrophication [Rey93]. evacuation [Bro94]. evaluate [BCO21]. Evaluating [ACST17, MTM22, CWB16]. Evaluation [CMC11, GLC16, SAT11, Ano92j, Ano93g, CS04, GMS17]. even [Gau01]. Evidence [Ano92k, CR98, HFG07, RT91, RN04, SRBS21, BHS19, BSWA14, GCS20, GC23, GOP^{+}23, PBMF12, PDDDE21, WMW18, ZKvZ19]. evidence-based [GOP^{+}23]. evoked [LF13]. Evolution [KVH98, Tay99, vGM05, AGMB^{+}23, Ano92k, Ano95d, BZM12, CZF22, CBHOH19, EJD18, FQSJ23, Fle96, GdMD13, LC02, LSH15, MTPR15, MS02, McD97, McL94, PGA^{+}24, PRN95, SGF13, SVRS19, Smi93, Tho92, Utt04, Woo93, dBdSA12]. evolutionarily [Joy23]. Evolutionary [LFdSRM16, PBM14, dRRGC12, BHK00, BLF10, CMA15, MBdBC13, NLFM12, SBSŠ21, SFO14b, SFO14a]. examination [Mok93, WMG00]. Examine [BDS05]. example [BBS10, MV13, WW06]. Examples [PB98, Avi00, CCA^{+}23, MSH14]. exaptation [AS95]. exclusive [Gil93b]. excretion [HP93]. excretory [PDGBR07]. existence [MdAVA11]. exogenous [All11]. exotic [IE06]. expanding [CCA17, CBHOH19, RSGS^{+}23a, RSGS^{+}23b]. expatrial [McD10]. Expectations [Smi98]. Experience [Arn05, HL98a, CA04, LC02, VCD12]. experiences [CBR98, NRJ^{+}23, SAK14]. Experimental [GVB94, BHP11, CD01, GCSR09, GS23, MHW07, VMB07]. experiments [Bro94, DW93, Mol92]. explain [LBS19]. Explaining [KDF13]. explanation [RFH15]. exploitation [Ano92e, Ano92u, Ano94f, HWA14, HTH00, IH04, LBS19, OCB12, WW06, Sei08]. Exploited [Ano93f, Ano94k, CBB19, JPC14, ROW^{+}23, Ros00, RUL95, WCP97]. Exploration [CRZHZ02]. explore [MBC21]. Exploring [HKN18, SHPH^{+}22]. exposure [SW10]. Expression [SZC11, DW11, ITV98, KTW12, RBPPS16, TRB13]. extant [WSC09]. External [GAD10, Col10, HFP14]. extinct [CBAVGR02, WSC09]. Extinction [Pit98a, TRJ24, EP00, FCOCNC24, GKC19, MMB22, WMG00]. extirpation [PJ14]. extract [APLL07]. extreme [PNC11, WG19]. eye [PDB16].

F. [RVRCB11]. face [JLC21]. facetus [PNC11]. facilitate [NKF21]. facilitating [MCD09a]. Fact [Bev98]. Factor [Ano95j]. factorial [HKN18]. Factors [Kos09, MÁMS05, TØH08, CNH22, DPV^{+}23, DCS20, GSD18, HW24, MDR14, PMP21, SHS14, Tan03, TRJ24, TJLLC10, WK99, WMG00]. Facts [VGA11]. FAD [DT04]. failure [Ann96]. faith [CGR11]. faith-based [CGR11]. falciformes [MG18]. falciformis [DHG18]. families [Nor14]. Family [PB98, CGK11, MFV19, YWS^{+}24, dBdSA12]. family-based [YWS^{+}24]. Fantasy [Bev98]. Far [ZZ04]. Farmed [Paw99, Col10, PBC12]. farmers [ZRZ20]. Farming [RT05]. farms [AGT08]. Fast [MRMJJ17, HCvP16]. fast-warming [HCvP16]. fasted [RBPPS16]. fate [MGL^{+}23]. fatty [GGZ10, GSD18, ZJDZ10]. fauna [AGT08, BGHC23, Bol07, ZDJ19]. faunal [BP08]. faunistic [MKS12]. Fausch [Woo99b]. Features [DSK11]. February [JBX07]. fed [RBPPS16]. Federation [BP08]. feed [FSB14, GDC13]. feedbacks [LSH15, NvPV22]. Feeding [BBV12a, BBV12b, KWI07, SK02, BSL18, Bro94, Kam02, Kam08, LSF14, MFV19, NMS13, NKF21, OCB12, RL05, SHPH^{+}22, WOR09, dSKV16, Web09]. female [CB00, GAD10, Kva98]. female-biased [Kva98]. females [NVA12]. fertilization [CBP02]. FIBiol [SHH95]. field [AVB09, BHS19, CWB16, GMS17, MBK12, Nel92]. field-based [GMS17]. Fifty [Hol98]. figures [Ste05]. filamentosum [PBG04]. find [RMK24]. findings [MC16]. Finfish [RT05, PTP14]. Fingerling [ADS11]. Finnish [MH20]. fins [SKW22]. firmer [RFH15]. First [BOs12, Ann96, Ano93b, BOV09a, BOV09b, BH17, Cad99, CGRCM19, SZC11]. Fish [AJM22, Ano92k, Ano92l, Ano93f, Ano94a, Ano94e, Ano94h, Ano99b, Bev98, BOD21, CBEGR02, DT04, Dou93, DW93, Els04, Gla00, Han05, Hus04, IDG16, JCL07, KHW09, KK12, KVH98, Kul99, LAI05, LWS17, MÁMS05, Nor03, Pis00, RT05, RG99, RBK10, SKD03, TFE07, VaR92, VW91, VCZ19, WMW18, Woo95, Woo98b, dPAGGB16, AGO04, AMVC20, ADC15, ACB08, AYN11, AHW04, AF04, AMVV13, AVA19, AA20, Ano92b, Ano92g, Ano92o, Ano92r, Ano92s, Ano92u, Ano92x, Ano94k, Ano94i, Ano94o, Ano95i, Ano95j, Ano95h, AS96, AVB09, BBJ12, BGM^{+}24, BHK00, BEBFM14, BJM20, BP93, BLF10, BHS19, BTW15, BO12, Bro94, BTS97, BEC11, BLC19, BC03, BHP11, COGFPV23, CWB11, CGKSP13, CBB19, CRZHZ02, CSSO02, Cha95, CDH^{+}23a, CLX15]. fish [CAB19, CMC11, CG00, CBAVGR02, CZF22, CWN11, CWD11, CWB16, CCO20, CFS^{+}23, CBP02, Cra92, CTM92, CBHOH19, CE96, DSU08, Dav94, DW09, DBT15, DLP95, DIR20, DPS22, DD15, DD13a, DD13b, EB93, EGMM02, EG03, EEDP18, EBRLGB02, FSSM17, FC07, FD00, Fri04, GLG12, GP12, GDC13, GMS17, GHF98, GSD18, GP15, GS23, GPS18, HP93, Han96, Har94, HL11, HW24, HJC09, HSC20, HMN17, HDI15, HSS21, HCMT24a, HCMT24b, IE06, ITV98, JLD^{+}24, JJC22, Joy23, Jua02, KB14, Kam02, Kam08, KDF13, KL23, KB21, KC92, Kim93, Kim95, Kul95, LF13, LSP11, LNT^{+}23, LKH16, LMM11, LCC19, LSDH12, LQW19, LSF14, LCM22, LLSDTJ09, MDM20, MV13, MGTD18, MC16, MTL21, MSR14, Mar95, MPO11, MBC21, McL94, MSLN02, MIA^{+}23b, MIA^{+}23a, MBM06, Mul11]. fish [MV96, Mun93, MZ00, NMS13, NVH21, NPS17, Nel99, NWO16, OK12, PP92, Pen93, PV19, PBMF12, PMP21, PBC12, PDB16, PM19, QSBV18, RSB16, RMK24, RT91, Ran92, RSSS23, RFH15, RK10, RBPPS16, RHM11, ROP^{+}23, RCCASG02, SHHK21, SHPH^{+}22, SN10, Sän93, SCC09, SPL12, SDJ13, SD91, SRD93, SJH22, She94, SN00, Sin97, SMM94, SLH11, SR91, SG24, SOB20, Sny05, SDG15, SK02, SGD^{+}23, SJB^{+}23, TFF09, Tel09, TTC11, TKR04, TJLLC10, Tur99b, TR94, VBC^{+}23, Vil03, WCD11, WGL17, WG19, WLB13, WGC18, Wie96, WMM92, WA03, WRSG21, WPF16, XSC15, YH18, YKS14, ZDJ19, ZRZ20, ZLH^{+}23, ZCT02, dSPPM12, dSKV16, dJFS20, dMV20, vDBV02, Bil02, Fos08, Mul09, Sul04, Bro00b, Hil99, Nor02, Paw99, Lor93, Woo99b]. fish-based [MSLN02]. fish-estuary [HW24]. fish-eye [PDB16]. Fish-marking [Ano92l]. fished [GCUR22, GGL14, NWG21]. Fisher [REA^{+}23, DCS20, ENF^{+}23, MMJ22, PRP16, RSSS23]. fisher-led [MMJ22]. Fisheries [Ano92m, Arn05, ALJ08, Bar95, Bro05, CR98, Cad99, CS05, DSU08, De 98, FB05, GB23, Gla00, Har98, Har99b, HL98a, KLC18, LAI05, Pat92, Pit98a, PB98, PH97, Smi98, Soi99, Sol93, Ste05, Wil96, APP^{+}23, ANL12, AH23, ATdLCBR02, AMV13, Ano92c, Ano92v, Ano93a, Ano93b, Ano93e, Ano93g, Ano94b, Ano94n, Ano95a, Ano95j, Ano95h, ANP^{+}23, AAH98, Arn96, AAJ21, BGR21, BBS10, BRI00, BHP15, BHD23, BCO21, Bis06, BDF09, BBW09, BH17, BPA^{+}23, BSKB22, BCL21, BSM17, Cal22, CDC09, CH94, CCA^{+}23, CR08, CMF^{+}24, CBSG24, CBO19, CS04, Chr96a, CKN21, CCH18, CCP13, CBL17, CVD21, CNT21, CFS^{+}23, CDH^{+}23b, CJV13, Cus94, CHN18, DPV^{+}23, Dav96, DBR15, DCS20, DCS11, DLP95, DD18, ELC^{+}24, ENF^{+}23, ENP18, EBRLGB02, ETWE12, Fer94a]. fisheries [FMH07, FCP19, FSB14, FCH16a, FCH16b, GCS19, GHMG22, GC23, GCB^{+}24, GOP^{+}23, GK18b, GCUR22, GZT13, GLM20, GSC17, HGC17, HKTS^{+}23, Han96, Har11, HJC20, HCB15, HPY18, HCNH24, Hun21, IH04, IWG17, JLC21, Jia09, Jon07b, Jua02, KWM^{+}22, KHW09, Kel23, Ken95, KCE12, KW08, Knu01, LAWD06, LMH21, LFM13, LW17, Lin94, LM19, LBS^{+}23, LSF14, LCC16, LM93, LFMP21, MTPR15, MDF14, MTM22, MCD09a, MP16, MGW93, Mil12, Mis97, MML18, MLN05a, MLN05b, MRMJJ17, MSV^{+}23, NWO16, NMF17, NWG07, NRJ^{+}23, OGR17, OSB^{+}23, PM94, PGG07a, PGG07b, PLW17, Pau97, PJ08, PZC17, PWB99, PBG04, PAK^{+}23, PMV18, REA^{+}23, RSGS^{+}23a, RSGS^{+}23b, RCP19, RP91, RWH04, ROW^{+}23, Ros97, SBSŠ21, SK15, SE16, SK23, SEH24, SAC^{+}23, TCS20, TKB18, TBA20, Utt94]. fisheries [WP96, WG94, WPD12, WNB12, WMA20, YMR12, ZCW19, ZWW20, Bur07, GCO03, MAJ05, DCS12, TBK11, Pis00]. fishers [ASS^{+}23, BMN21, CCH18, ENF^{+}23, HFP14, MCFC21]. Fishery [Asw05, DC05, Har99c, Lac02, LFJ08, RPD16, SVM^{+}24a, SVM^{+}24b, Tur99a, dC98, AGJ14, ABS15, AUdS18, Ano92b, BCA^{+}23, Bag11, BBS10, Bis06, BMF09, CHFTV22, Cad00, CCR21a, CCR21b, CCR^{+}24, CBR98, CBH15, CBHOH19, Dia04, GCK21, GIT09, HFP14, Koc01, LC02, Mad07, MCL08, Mar08, MLSP^{+}23, MBH14, MMJ22, MLK13, MG18, Neh96, OMAGL02, PRP16, TBA20, VGA11, Csi99]. fishery-dependent [Bis06]. fishery-related [Koc01]. Fishes [Abl06, Fui99, HL98b, Kam05, McD03, Par92, Rei99, Smi91, Soi99, WTC06, Woo99b, AGJ14, AGT08, All11, AGMB^{+}23, AGM19, Ano92e, Ano92h, Ano94l, Ano94m, Ara15, AE22, AWC17, Avi00, BCP22, BTZ^{+}23, Bas93, Ber93, BTW15, BBV12a, BBV12b, BOJ16, CBF^{+}24, CNH22, Car92, CGK11, CC22, CWP14, CRD00, CJV13, DPC12, DSK11, EM97, Fer94b, FN02, FM94, GVB94, Gil93a, Gil93b, GJA17, Gre93a, Gre93b, HAC18, Har24, HPP15, HRC09, HF20, How93, JR97, JJ93, JPC14, LE24, LY07, LI03, LET^{+}23, Lis10, Mag13, MFV19, MA08, McD97, McD02, McD06, MMB22, Mol92, MHvH16, MCH95, MLP17, NCMT23, Nel92, Nor06, Nor09, Nor12, Nor14, NTC12, OW97, OCB12, PLW17, PBF15, PAW17, PRN95, PKF92, PQS14, PY97, PBR19, PHK20, PWC06, PBS14, PGJ15]. fishes [PFW16, PFW17, Rai94, RNJ16, Ree02, Rob92, RWH04, RUL95, SMS12, SL15, STP22, SPB00, SMR11, Smi92, SW95a, SW95b, Tay99, TBB^{+}24, TRJ24, TAS10, TS14, Tur95, UIA11, Utt04, VGMAGB02, WGR20, Whi17, WWS17, Wil07, WDP19, Wis99, YCT11, WBW05, Win06, Zim05, MBJ12, Woo98a, Bar94, Web09, How94]. Fishing [BGHC23, DROCM23, Han05, HPL05, Hup05, Knu01, MFL21, Neh96, AGJ14, dPAGGB16, Ano92t, Ano92x, AFBB23, BGOW23, BMN21, BO12, BSM17, CVD21, GYH10, HFM19, HB19, HP07, HR00, Koc01, Mad07, MG18, PVL21, PMVA19, PKC^{+}24, RSSS23, SEH24, Sur23, WH07, War08, ZCW19, ZL23, vOR15]. fishways [HBC13]. fit [ZLF21]. fitness [EP00]. FitzGerald [Ano94c]. Fitzsimons [Soi99]. five [Ano92b, OV24, PASF13]. Flatfish [Bur10]. flatfishes [DGV11, LR11, Sei08]. flathead [WPD12]. fleet [FTH15, LFM13, WH07]. fleetwide [GCB^{+}24]. flexible [SHB21]. floating [CSSO02]. Flood [AGO04, Mag13]. Flooding [CBB19]. Floodplain [MCN12, Mag13, MJSOF16, REA^{+}23]. floodplains [CBB19, Lig16]. flounder [Ter01, Ter11, Ter18]. flow [YCT11]. flows [BSKB22]. fluctuates [BHS19]. Fluctuating [All11]. fluctuation [FHW^{+}24]. Fluctuations [CCR^{+}24]. fly [Dav94]. flying [Dav94]. focus [JNW14]. focusing [TRB13]. Fold [AMVC20]. following [MTL21]. Food [Ano94d, FAN22, ABS15, CCR21a, CCR21b, FSB14, MTC19, PBF15, Whi17, vPKW18]. forage [BBS10]. foraging [BHB22, NTC12]. forces [YH18]. Forecasted [WWS17]. forecasting [OK10]. Forest [CMA15]. forestry [Ano93e]. forests [Whi17]. Foreword [Nie02]. form [Ran92]. formation [RFH15]. former [AGT08]. Formosa [Jua10]. forthcoming [Ano91b, Ano91c]. Forum [Wil96]. forward [Kel23, KMK16, RNJ16]. fossil [For95, Rei93]. foul [McD06]. found [Nor06]. foundation [LCC16]. four [GBOK23, KSK17, PJ14]. Fourth [Ano94a, Ano94e, Wil96]. fragility [PDM20]. fragmented [FSSM17]. Framework [MA03, Bro00a, PWB99, SP98]. frameworks [GCB^{+}24]. France [CBSG24, PBB00]. Fraser [RMN11]. fraud [LSDRS23]. Fréeon [Gla00]. French [Ano92b, GLC16]. frequencies [NVA12]. fresh [Gib93, Rey93]. Freshwater [CBAVGR02, DJ05, RPA13, SPL12, SRD93, Woo98b, AMVV13, AMV13, Ano92c, Ano93a, BGM^{+}24, BGHC23, BLD19, CBSG24, CDH^{+}23a, CMC11, DBT15, DPC12, DPS22, FSSM17, FMH07, GP15, Gre93b, HKTS^{+}23, HH04, HSC20, Jel22a, Jel22b, KDF13, LSP11, LCC19, Lig16, LBS19, LSB22, MV13, MBC21, McD02, NMS13, NPS17, Nel92, PP92, PV19, RVRCB11, RCCASG02, SPB00, SJH22, SFO14b, SFO14a, Tay99, TFF09, Tur95, VCZ19, WIT14, WR16, WA03, XSC15, dSPPM12, SDJ13]. freshwaters [PEP12, SL15]. Friedrich [Coc05]. Friedrich-Wilhelm [Coc05]. frisii [ADS11]. fronds [SLH11]. Frontier [Bro05]. FRS [SHH95]. Fuiman [GCO03]. function [JB95, LSDH12, RL05, SOB20]. functional [AMVC20, ELC^{+}24, GBOK23, HW24, HKL10, JJC22, SZC11, TSB22]. functionality [SZC11]. functioning [Ano95b, TBB^{+}24]. Functions [She94, LSDH12, Mun95, Web09]. fundamental [AJM22]. Fundulidae [RVRCB11]. Fundulus [RVRCB11]. fusion [RZV12]. Future [MMM22a, AFO22, Ano94f, CGL22, DT04, FAN22, Fer94a, FMM22, FD22, GLHD^{+}24a, GLHD^{+}24b, GOP^{+}23, GLM20, HMH22, HB19, HJC20, IRPG21, LE24, LAWD06, LM93, MHP12, MSM20, MDV^{+}23, MLK13, OGR17, PTP14, PHH10, SBtIWG12, SBG17, STW^{+}24, TSB22, WB94, YC20, MMM22b, PAN22]. futures [NAP22a, NAP22b].

G [Bla98, GCO03, Hil99, Mul10, Web09, Woo98a]. G. [Pyk05]. Gabes [RMN11]. Galaxiid [McD03]. galaxiids [WIT14]. galaxioid [McD06]. gallina [CCR^{+}24]. Gambusia [Pyk05]. Gamete [CBP02]. Gan [RBK10]. Ganga [SPL12]. gap [BBW09, CWN11]. gaps [CCR^{+}24, CCO20, CJV13, DT04, HPP15, YCT11]. gar [CFGG13, MAC02]. Garavello [MdAVA11]. Garcia [WBW09]. Garrison [WBW05]. Gary [Sul04]. gastric [Bro94]. gastropod [LC02]. Gaza [AUdS18]. Gear [CR08, HB19, MBS17, MFL21, SEH24, WH07, War08]. gears [dPAGGB16, Mad07, MF99, PVL21]. gender [Ano95h]. gene [KTW12, RBPPS16, TRB13]. genera [GdMD13, Gre93a, Utt00]. General [EAW22, CSSO02, CG00]. generalisation [FD00]. generalized [CTL17]. Genes [MGL^{+}23]. Genetic [CPM14, DGV11, FSSM17, KKF10, SBP07, Tan03, UE02, Avi00, BM09, CHFTV22, DHG18, Fer94b, GLC16, GJA17, MWP^{+}23, MOF11, MAP21, OK14, PM94, PDDDE21, PDM20, SRBS21, VGA11, WG94, WB94]. genetically [BOs12]. Genetics [Ano95d, Col10, CH94, CAB19, Fer94a, IJ03, Lin94, MMF18, SVRS19, Utt94, Utt04, WPD12]. genome [GKR10, ZCT02]. genomic [PDN^{+}24]. genomics [AJM22, BJM20, CBF^{+}24, CAB19, UIA11, WG19]. genotoxic [Ano92o]. genotype [IJ01]. genotyping [LW17]. genotyping-by-sequencing [LW17]. genus [BLF10, BZM12, CPN11, CM08, Jel22a, Jel22b, LF03, Pau10, PBF15, Web02, WM02, YVR07]. Geographic [MBP14, Gau01]. geometric [LH18]. geophysical [HR00]. Georgia [MBP14, SBN04]. Geotriidae [MBG21]. Gerdeaux [Bil02]. German [Ano92b, Coc05]. Gerry [Ano94c]. gestation [GELL16]. Gestion [Bil02]. Getting [AH23]. Ghana [AFBB23]. Giannoni [WBW09]. Giant [SKD03, GJ07, HFG07]. Giants [HL98a]. gibelio [PEP12]. Gibson [Sei08]. gigas [SKU20]. Gilabert [WBW09]. Gill [PP92, LH18, RHM11, SW12]. Gillnets [Ano95e, HKTS^{+}23]. gills [JB95]. gilthead [IPT10, SZC11]. Ginger [WBW05]. glands [WOR09]. glanis [AKGB11]. glass [HWA14]. glauca [KWI07, MG18]. Global [Ano95f, BPA^{+}23, CR98, CM08, Csi99, MLP17, SBtIWG12, BGR21, BOs12, BSKB22, CA04, CAB19, CKN21, CCH18, DBT15, EEDP18, FMH07, FSB14, GLHD^{+}24a, GLHD^{+}24b, GCS20, GLM20, GSC17, HH06, HW24, HP14, HJC20, KMK16, LCM22, MBS17, PLW17, PAK^{+}23, PE08, PKC^{+}24, RWH04, SAB^{+}23, Sot02, SAC^{+}23, TRJ24, VAJ10, VEK10, VCZ19, WR16, WPD12, WDP19, WRSG21, YOC15, ZWW20, dMlTB19, AH23, Har98, Woo99a]. global-scale [YOC15]. globally [PDM20]. globe [CNT21, LPC^{+}24]. Glossary [Ano94g]. Glyceraldehyde [OBS08]. Glyceraldehyde- [OBS08]. glycine [KTW12]. Gmelin [PTOS^{+}24]. GnRHa [MZ00]. GnRHa-delivery [MZ00]. Goal [HFM19]. Goals [NRJ^{+}23]. gobies [GBOK23, McD09b]. Goby [LAI05]. Godin [Woo98a]. Golden [BP16, GGZ10, RBK10]. Gonadal [HL98b, Fri04]. González [WBW09]. good [BTS97, Fro99]. Goodeidae [Web02]. gorbuscha [RN04]. Gordon [Soi99]. Gorges [LQW19]. governance [AH23, HMH22]. Governing [CHN18]. gradient [SHHK21]. Gradients [VGMAGB02, dPAGGB16]. grading [BHP15]. Graeme [Ram08]. Graham [Wil07]. Grande [CBEGR02, EGMM02, Mag13]. Grande/Rio [CBEGR02]. grands [Bil02]. grass [CM92]. grayling [Nor95]. graylings [EAW22]. Great [CBL17, Com93, Cra92, HMQ16, MSM20, Soi99, GK18a, MLK13, SK23, CFB14, HJC09, MLK13, STP22, SBP07]. Greater [ACST17, AAH98]. Greece [PEP12]. grey [DHB17]. ground [Ara14, Jon02]. Groundfish [CR98, dC98, KSK17]. grounds [BSM17, FD22, Knu01, MLC17]. Group [Zim05, PBMF12, dSPPM12]. grouper [CGCG18, FCH16b, PCW17, FCH16a]. groupers [BGOW23]. groups [GKR10, KGF10]. growing [MRMJJ17]. Growth [AGG03, HL98b, HLD07, AKGB11, CSY12, CTB16, CRD00, DW09, FC07, GBCCO12, GCSR09, HDI15, IJ01, IJ03, Joy23, KTW12, LE24, MHP12, MV96, MRMJJ17, MKK10, Nis00, dSSHK21, Sch18, SG24, TV15, TS96, UIA11, VMB07, VSM12, VLD20, WR16, YWS^{+}24, ZJDZ10]. guardians [FMM22]. guests [OV24]. guidance [CCO20]. Guide [Ano01, Ber93, Dri05, CTL17, Nel92, PHK20, Tur95]. guidelines [CA04, MVGW18, GS23]. guiding [CDH^{+}23b]. Guinea [MR18]. guineensis [AY10]. Gulf [AMVC20, AAH98, EBS20, GPS18, HFG07, BBS10, BHD23, CCR21a, CCR21b, CGR11, Dia04, MR18, Moo12, MVGW18, OGR17, RMN11]. Guppies [RHP99]. gustation [Har94]. gut [TRG19, ZRZ20]. Gymnocephalus [GH16]. Gymnotidae [MCN12, MJSOF16]. Gymnotiformes [CPN11, MJSOF16]. Gymnotus [MCN12].

H [Bla98, BOJ16]. H. [TdSL15]. Habitat [ABS15, BS03, CFB14, LFJ08, BBS17, CWP14, HJC09, MLM19, Qui12, ROW^{+}23, RBC16a, RBC16b, SHS14, SHB21, VRRCAG02]. Habitat-specific [ABS15]. habitats [Ano93a, CG00, MAC02, MT18]. habits [BSL18, KWI07, SK02]. haemoglobins [PRN95, RG99]. Haemulidae [NLFM12]. Haemulon [PBF15]. hake [SLK16]. half [Jon07b]. halfbeak [TdSL15]. halibut [CLH21, IJ01, LR11]. Hamilton [BP16]. Hammerhead [PDM20, GK18a, PJ14]. hammerheads [GK18a]. Handbook [Ano92o, KCR17]. handling [DML^{+}24, HCB11]. handling-related [HCB11]. haplotype [CRRCdL08]. Haq [Bla98]. harbour [MV96, YH18]. hard [WDP19]. harmful [Sny05]. Harold [Pis00].

Harp [Pie92]. Harris [Ram08]. Harttia [BVMF13]. harvest [HH04, SE16, ST94, ZKvZ19]. harvested [GCUR22]. Harvesting [FSB14, AFB15, Lig16, SEA22]. hatch [McD09b]. Hatchery [Sul04, Kos09, MT14, OWW04, SBN04, WHS04, ZZ04]. hatching [KK12, KKM10]. haul [KFS21]. hay [AGT08]. headwaters [VCD12]. Healing [MCD09a]. health [NvPV22, PBC12, ZRZ20]. Heaphy [SHH95]. hearts [PBC12]. Heat [ITV98, DW11]. heavy [EDP10]. heel [ANP^{+}23]. held [BP08, Har99b]. helminth [Mar95]. help [GMS17]. helps [MDM20, SBtIWG12]. Hemingway [Els04]. Hemiramphus [TdSL15]. hemisphere [Nor12, MBG21, WMW18]. hemorrhagic [EEDP18]. hepatic [SZC11]. hepatocytes [SZC11]. Heptapteridae [GdMD13, dBdSA12]. her [FMM22]. herbivorous [PFW16, PFW17, TBB^{+}24]. hermaphroditic [Fri04]. Hermaphroditism [CCA17, PMP21]. Hermetia [TRG19]. Hernández [WBW09]. herring [Cor02, FD22, McQ97, PDA12, dSSHK21, WW06]. herrings [Hun21]. heterogeneity [TR94]. heuristic [DP12]. hidden [IOHM23]. hierarchy [GC23]. High [ANP^{+}23, BHP15, DLP95, MWP^{+}23, RNJ16, SMR11, WH24]. High-grading [BHP15]. high-latitude [SMR11]. high-throughput [RNJ16]. highland [GGZ10]. highly [CBR98, DLP95, SZC11, SOB20]. highways [HK14]. Hilborn [Mil99]. Hippoglossus [IJ01, LR11]. Hiroya [Woo99b]. Histological [PM14, HHCM10]. Historic [Jon07a]. Historical [PJ14, BGHC23, BAA18, DD18, Gre93b, JNW14, JZ00, SKU20, SRBS21, WH07, War08]. histories [Ano92n, FQSJ23, HDI15, MT18, Woo93, Qui12]. History [Abl06, Smi91, AS08, Ano92g, Ano94f, AWC17, BP16, BCK16, BHP11, CGP^{+}24, CTB16, DHG18, EJD18, FC07, FHvP14, GP15, GH16, HPdL02, JJMD13, JBS02, KB14, KDF13, LFMP21, MBJ12, MR18, NWO16, Nor12, PQS14, TP14, VEK10, VGMAGB02, Vil18, VVU22, VMB07, WK99, WIT14, KCR17]. holbrooki [Pyk05]. Holding [OCE11, DML^{+}24, PWC06]. holistic [BBJ12, MT18]. Holocephali [Lis10]. Holothuria [Gou16]. Holt [Pit98b, GLC98, Pau98, WK99, vPKW18]. Homarus [Bag11]. home [OSC20]. Homing [KC14]. homogeneous [BOs12]. Honeyfield [Soi99]. Hong [CS04, MTL21]. Honour [Woo99b]. hook [GH17, GCM18]. Hooking [CML19, GH17]. hooks [MF99]. Hoplerythrinus [dRRGC12]. Hoplias [BLF10]. Hoplostethus [Jon07a]. horizon [HJC20]. Horizons [Pau98]. horizontal [LY07]. hormonal [Bas93]. Hormone [HL98b, BLD19, DW09, GVB94, Kul02, PSK12]. horse [AGG03]. horseshoe [SBZR17]. hosts [PWM23, SMM94]. hotspot [FHvP14, MBP14]. hotspots [HP14, PHB14]. Houde [RHP99]. Howard [Hus04]. Human [AFB15, Cra92, HCNH24, MTPR15, Moo12, NvPV22, TØH08, WGC18]. Human-induced [Cra92]. Humboldt [SKU20]. humoral [SZG10]. hump [GGZ10]. Humphead [SKD03]. Hundred [LFM13, KCE12]. Huppert [Har98]. Hutchings [FQSJ23]. hybrid [HSC^{+}24]. hybridisation [MHvH16]. Hybridization [EN00, SPB00, Avi00, EP00, PBMF12]. hybrids [BRI00]. hydrokinetic [SMR11]. hydrologic [ACST17]. Hydrological [AMVV13]. hydrophila [SZG10]. Hydropower [YCT11, AVA19, BTZ^{+}23, PDB16, WMW18]. Hydropower-related [YCT11]. hydroturbine [TAD14]. hyperspectral [TMKC24]. hypogean [Wil07]. Hypostomus [BZM12, BVM13, PASF13]. Hypotheses [PDA12, APLM14]. hypothesis [CSSO02, FD00, Kul95, Kul02, YH18]. hypoxia [SYL^{+}23].

Iberia [PTOS^{+}24]. Iberian [MV13, SAC^{+}23]. ice [BOV09a, BOV09b, MBM06, Pie92]. Iceland [Arn96, Arn05]. ICES [Bar95]. ichthyofauna [LMM11]. Ichthyofaunal [Whi99, MMB13]. Ichthyological [Rei99]. Ichthyology [Rei99]. Ichthyoplankton [STW^{+}24]. Idaho [SAK14]. idella [CM92]. Identification [Abl06, HP14, PDGBR07, BM09, Cad00, DOP18, DD13a, DD13b, Gau01, Tel09]. identified [HBC13]. Identifying [APP^{+}23, BBJ12, CJV13]. identity [TdSL15]. Iguanodectinae [dSPPM12]. Iguassu [PBM14]. II [AHL12, Ter11]. III [Ter18, WM02]. Ikpoba [TFE07]. illegal [MMJ22]. Illex [CUT07]. illucens [TRG19]. Illusion [CS05]. Imagery [CM98]. imaging [TMKC24]. immune [TRB13]. immunity [RMK24]. Impact [Ano95j, PR05, AJM22, CTM92, DGV11, Koc01, LMM11, TBA20, ZRZ20]. impacted [LWS17]. impacting [GBR22, MIA^{+}23b, MIA^{+}23a]. Impacts [Ano93e, HR00, KMK16, AFB15, DAN02, FHK04, FMH07, Hal98, HYW13, KPN21, MLM19, MBS17, MBH14, OSB^{+}23, PJ08, PHB14, PV19, PMVA19, PWC06, PGJ15, SMR11, SOB20, Sot02, TØH08, WGC18, WA03, YCT11]. Imparfinis [GdMD13]. Imperative [Bla98]. imperiled [HMQ16]. impingement [CCO20]. implantation [BEC11, CMC11, CWN11, CWD11, HL11, LR11, Mul11, OCE11, TTC11, WCD11]. Implications [AMVC20, BB05, CLB^{+}22, RT05, SJB^{+}23, VGA11, Bea07, BGBE^{+}23, BdST16, CVD21, ENP18, HWA14, HSC^{+}24, IWG17, Jia09, LSP11, MMF18, MRMJJ17, MH20, Nis00, OSB^{+}23, OWW04, SMR11, SLK16, SRB^{+}23, VSM12, WGL17, ZMG17, ZWW20, ZKvZ19, vOR15, Sch18]. Importance [CWP14, LFJ08, BSL18, BHP11, KBV11, LBS^{+}23, MHvH16, MLC17, SDJ13]. Important [RT05, DSU08, Gil93b, QSBV18]. impounded [KFS21]. impoundment [LQW19]. improve [CJV13, PBC12]. improved [Moo12]. improvement [Kat97]. Improving [TAD14, YWS^{+}24, Col10, GS23, NRCD^{+}23]. INBio [WBW05]. inbreeding [WHU02]. incidence [PMP21]. including [Bol07]. Incorporating [BdST16, SBSŠ21]. increase [MV96]. increasing [STP22]. increasingly [LCC19]. incubation [GCSR09]. Index [Ano92p, Ano00a, Ano04a, Ano05a, Ano05b, Cha95, Ano92a, Ano92w, Ano99a, Ano02a, Ano02b, MSLN02, MMB22]. India [SPL12, Kel23, LSP11, PR05, RPA13, Rai94, SDJ13, SJB^{+}23, Sur23, Ano93b]. Indian [KWM^{+}22, CBO19, MMF18, TKB18, VRR12]. indicator [TFE07]. indicators [SCC09, SBtIWG12, SLH11, WMM92]. Indigenous [CDS16, DDD16, Lig16, RSSS23, SX16, KZ20, MP16, RSB16]. indirectly [ZLH^{+}23]. IndiSeas [SBtIWG12]. Individual [DOP18, GCB^{+}24, Gra96, TR94, Bag11, SHHK21]. individuals [MFS18]. Indo [Ano94e, Hun21, NWO16, PFW16, PFW17]. Indo-Pacific [Ano94e, NWO16, PFW16, PFW17]. Indo-West [Hun21]. Indonesia [BDS05]. Indonesian [BDF09, STW^{+}24]. induce [APLL07]. induced [Cra92, IPT10, WGC18, WLB^{+}23]. inducing [MP07a]. industrialisation [AFBB23]. industry [Gou16, HFM19, Pie93]. Inertia [MV96]. Infectious [EEDP18, Mun93, RHM11]. infer [JNW14]. inference [SPB00]. inferences [BSV11, CMA15]. inferred [KKF10, PWM23]. infers [Jon07a, RBPPS16]. Influence [PTOS^{+}24, SHS14, CG00, HHCN11, Koc01, LET^{+}23, MDM20, PFL10, TJLLC10, WGR20]. Influences [MGB15, Nor14, EG03, FHW^{+}24, Nor09, ODS10, VMB07]. influencing [VEK10]. inform [CFS^{+}23, DDD16]. Information [HPP15, Jon07a, OK14, SMN08, WP96]. inhabiting [EGMM02]. Initial [MTL21, OSB^{+}23, EP00]. injuries [AVA19]. injury [HCB11]. Inland [Kel23, Rai94, ZWW20, ATdLCBR02, BOD21, DAN02, IE06, JLC21, LCC16, MLP17, PLW17, REA^{+}23, SEH24, CNT21]. Innovating [SJH22]. Innovation [LSDRS23]. innovations [CVD21]. insect [TRG19]. inshore [HCK21, PJ08]. Insights [JJC22, CCR^{+}24, GLHD^{+}24a, GLHD^{+}24b, HJC09, HF20, MFS18, MBA15, PM14, SEA22, SBG17, Sur23, VCD12, ZHP07]. inspirations [SYL^{+}23]. Institute [Hus04, Rei99]. institutional [CWB16]. institutions [Fro99]. Instituto [WBW05]. Instruction [Csi99]. instrument [ATdLCBR02]. Integrated [IWG17, GCS19, GHMG22]. Integrating [BBJ12, BDS05, PZC17]. integration [NMS13]. integrity [MSLN02]. intensive [KGF10, MKK10]. intentional [GCS19]. inter [BRI00, LSP11]. inter-specific [BRI00]. Interacting [dPAGGB16]. Interaction [SMS12, Smi98, ZZ04, BD20]. interactions [ABS15, Ano92f, CCR21a, CCR21b, IJ01, Kul95, OWW04, OCB12, PKC^{+}24, TKB18, dMV20]. intercontinental [VGMAGB02]. interdisciplinary [VBC^{+}23]. interests [CDS16]. interference [OCB12]. intermediate [SMM94]. International [ANL12, Ano93f, Ano94a, Ano94h, Ano94i, Bar95, KGW10, KGW11, WNB12, ENP18, Jua02, Jua10, Nis00]. interplay [TRJ24]. interpretation [HKN18]. interpreting [BLC19]. Interrelationships [CE96]. intertidal [AS95, AF04]. intervals [Cad91]. intestinal [ZRZ20]. intestine [KTW12]. intracoelomic [BEC11, CWN11, CWD11, LR11, Mul11]. Intracohort [NPS17, DPS22]. intraluminal [GELL16]. intraspecific [GLC16, RUL95]. intrinsic [AGJ14]. introduced [ACB08, AKGB11, CTM92, DD15, LMCB^{+}23, RK10, RBK13]. Introduction [Cas93, Hil99, Kat97, ZN04, BEC11, Nel99]. introductions [CM08, HH06]. introgression [PBB00, Utt00]. introgressive [EP00]. invader [PEP12]. invaders [GP15]. Invasion [PBMF12, APLM14, RSB16]. invasions [CAB19, DAN02]. invasive [BLF10, GBOK23, HMQ16, MGB15, MCFC21, MGRRJ21, OCB12, RTT12, SEA22]. Invasiveness [VCZ19, LMCB^{+}23]. inventory [GPS18, LCM22]. invertebrate [GPS18, HTH00, PWB99]. invertebrates [GJA17, HPP15]. investigations [Ara14]. investment [dSSHK21]. Invited [Sny05]. Involved [Shi05, MGL^{+}23, PKF92]. Involvement [HL98b]. involving [Chr96a, Kul02]. Ionian [CCR21a, CCR21b]. Ireland [Har99b]. Irminger [SRGS04]. ISBN [Coc05, MAJ05, Ste05]. Isla [WBW05]. Island [Bag11, Dri05, Ste05, ASS^{+}23, MDM20, BP08]. island-wide [MDM20]. islandica [RR11]. Islands [CBH15, MSR14]. isolated [PDN^{+}24, SBP07]. Isolating [MDM20]. Isolation [OBS08, LKK10, MBP14, SRBS21]. isotope [CGL22, MTC19, SBG17]. Israel [RBK10]. Issue [Han96, Pit98b, Ano99b, Ken95, MMM22b, MMM22a]. Issues [BBS10, HKN18, Jon02, LSP11, Lar96, NMS13]. Italian [PTP14]. Italy [KGW10, KGW11]. iterative [ACST17]. iteroparity [BGBE^{+}23, PM14]. ITQ [Ann96, Arn96, Dav96, Pau96]. ITQs [Arn05, Han96].

J [Bla98, Coc05, Els04, Har99a, Mul09, MAJ05, Rei99, WBW09, Woo98a, dC98]. J. [Web09]. J.-G [Woo98a]. Jack [Dri05]. Jacques [MR18]. James [Zim05]. Japan [WH07, Woo99b]. Japanese [Ara14, SAT11]. japonica [Ara14]. jaw [RL05]. jawed [For95]. Jay [Ste05]. Jeffery [FQSJ23]. Jellyfish [BSM17]. Jens [Fos08]. Jezierska [Nor02]. job [DROCM23]. John [Coc05, SHH95, Soi99, Bur07, LFJ08]. Jonsson [Qui12]. Joseph [Win06]. journey [FÁVG18]. Jr [MBJ12]. Jubilee [Pit98b]. July [Ano94a]. June [Har99b]. just [DROCM23, Pis00]. Juvenile [OWW04, PR05, WGL14, BCK16, DML^{+}24, GCSR09, KTW12, Man94, MC04, NTC12, OCE11, VLD20, ZJDZ10]. juveniles [KKW12, MKK10].

Kapour [Web09]. Karyotype [AVB09, GdMD13, NLFM12, SGF13]. Karyotypic [BVM13, PBMF12, dBdSA12]. Kathleen [KCR17]. Kawanabe [Woo99b]. keeping [BLD19]. keepnets [GAD10]. Kendall [MBJ12, dC98]. Kerala [Sur23]. Kevin [Fos08]. Key [FCH16b, AS96, BOV09a, BOV09b, BLD19, Cha95, CBO19, CJV13, HKN18, NAP22a, NAP22b, FCH16a]. Keys [DC05, Abl06]. killifish [RVRCB11]. Kime [Bro00b]. Kin [BB96]. King [ALJ08, DD18]. Kingdom [Coc05, KCE12]. kisutch [SBN04]. Kit [VCZ19]. knifefish [MCN12]. know [MH20, MC04]. Knowledge [MN98, CHFTV22, CCR^{+}24, CJV13, DT04, DCS20, Duc19, FCP19, FD22, LE24, LF03, MHP12, MBK12, MH20, MSV^{+}23, OAJ14, PZC17, PVJ07, PHH10, REA^{+}23, RSSS23, SK23, TP14, YCT11]. Known [SKD03, RR11]. knowns [PLW17]. Kogia [Bea07]. Kong [MTL21, CS04]. Korean [JPC14]. Koury [WBW09]. Krause [Fos08]. Kullenberg [Bla98]. kutum [ADS11]. Kvarnemo [SW98].

L [Els04, KGW10, KGW11, Rei99, WBW09, dC98]. L. [EDP10, GAD10, GKR10, HKL10, IJ01, KGF10, KKM10, MPK10, MBA15, Mul10, PFL10, RBK10, SZC11, SW10, STS07, WOR09, ZJDZ10]. Labeotropheus [Pau10]. laboratories [PHB14]. Labrador [MDR14, MMG18]. labrax [TRB13]. Lacepède [PV19, BZM12, BVM13]. Lack [LKK10]. lagoon [MTPR15]. lagoons [PBM14]. Lake [MBK12, MLK13, PB98, Tur99a, PVJ07, Ano92q, Ano94d, CCMS12, EDP10, LM93, OMAGL02, Pau10, SN00, WMG00]. Lakes [CBL17, PB98, Soi99, JJC22, Mar93, MLK13, ZMG17, Cra92, HMQ16, HJC09, MSM20]. Laland [Fos08]. Lamna [KWI07]. lamprey [CWO19, EJD18, HMQ16, JMJ19, MSM20, MJM15, PWM23]. lampreys [MBG21, MBD07, MT18]. Lancang [SGD^{+}23, ZDJ19]. land [Mag13, WWS17]. Landings [CR98, CBH15, Mar06, NWG07]. landlocked [CPM14, Jua10]. languages [Ano92b]. Large [Bil02, YKS14, ZDJ19, dPAGGB16, AGM19, BAA18, GK18a, GPS18, HFM19, LS07a, PBG04, REA^{+}23, RFH15, TCS20]. Large-scale [ZDJ19, HFM19]. largemouth [PV19]. largest [RBC16a, RBC16b]. Larimichthys [LLC11]. Larkin [Nor96]. Larry [JBJ06]. larva [HHCM10]. Larvae [MR18, AVA19, CFGG13, DIR20, KKM10, MPK10, MV96, PAT15, SW10, WOR09, YH18]. Larval [Hus04, COGFPV23, HSS21, Mag13, McD09b, MAC02, MLC17, NTC12, SAT11, Sto00]. larviculture [AYN11]. last [CFS^{+}23, JPC14, MBA15, MLK13]. later [MR18]. lateral [MCH95, WT08]. Latimeria [Tho92]. Latin [CD01, DC05, Ano92b]. latitude [SMR11]. Latitudinal [JPC14, TV15]. Laurentian [CBL17, HMQ16, MSM20]. Lawrence [Els04]. Laws [Csi99]. lead [CSS20]. leadership [CDS16]. leading [WMG00]. learn [KLC18]. learned [CCMS12]. Learning [HL98a, CDH^{+}23a, CFS^{+}23, JLC21, KC92, MGRRJ21]. learnt [HB19]. leatherjacket [MS13, RMN11]. leave [SD91]. led [MMJ22]. Lee [GCO03]. length [Bag11, BSL18, MLSP^{+}23, MV96, Sch18]. length-based [MLSP^{+}23, Sch18]. leniusculus [GCSR09]. lentic [PCG19, WJP23]. lepidota [PNC11]. leptodactylus [HH04]. less [CCA^{+}23]. Lessepsian [RMN11]. Lessons [WM96, ADS11, CCMS12, CLX15, HB19, HH06, STW^{+}24]. let [DBR15]. Letter [Ben06, BDD02, TKR06]. level [PFL10, SJB^{+}23]. levels [CMC11, DW09, GVB94, SK02]. LH [PSK12]. Liboy [WBW09]. Lichtenstein [PBG04]. Life [AWC17, GCO03, GP15, JJMD13, MR18, Nor12, Smi91, Soi99, SLK16, VVU22, WGL17, AA20, Ano92g, Ano92q, Ano92n, BCK16, BHP11, CGP^{+}24, CTB16, DROCM23, EJD18, FQSJ23, FC07, GH16, HDI15, JBS02, LE24, LET^{+}23, LFMP21, MBJ12, MT18, NWO16, PQS14, Qui12, RSGS^{+}23a, RSGS^{+}23b, SRD93, SL07, TP14, VEK10, VGMAGB02, VMB07, WK99, WMTL22, WIT14, Woo93]. Life-History [Smi91, AWC17, GP15, Nor12, VVU22, CTB16, FC07, VGMAGB02]. Lifecycle [vGM05]. lifecycles [BGTA19]. Light [CUT07, ODS10]. likely [SLK16]. lima [RVRCB11, RZV12]. limit [LSDRS23, MBP14]. limitations [BM09, WMDS18]. limited [CS04, CDH^{+}23b, TBB^{+}24, WNB12]. Limits [Sän93, CMA15, HSC^{+}24, VW91]. Limnothrissa [Mar93]. Limulus [SBZR17]. line [MCH95, WNB12, WT08]. lineages [EP00]. lines [GCK21]. link [BSKB22]. linkages [PQS14]. linked [ENF^{+}23]. Linking [BJM20, CDC09, HJC09, IJ03, WK99, LSP11]. Links [BCP22, BSWA14]. Linnaeus [MJSOF16, WPD12, vGM05]. lipid [RBPPS16]. lipids [Wie96]. List [MMB22, Car92]. Listening [Jua02]. literacy [KEP22]. Literature [vGM05, BHP11, SL15]. litigation [Ter01]. littoral [AE22]. live [CWP14]. lived [CBB19, RR11]. livelihood [DROCM23]. livelihoods [AFBB23, Lig16]. Living [KM20, Rei93]. lobster [Bag11, HFP14, MP07a, PM09, PMV18, SA12]. lobsters [AVB^{+}23, GGL14, SVRS19]. Local [GSC17, MN98, Woo99a, HWA14, SAC^{+}23]. loci [KKF10]. loggerhead [AA10, MTM22]. loliginid [LS07a, LS07b]. Loligo [MS07, OSB07, STS07]. London [Ste05]. Long [AGT08, Ano92r, Bev98, CCMS12, DZZZ22, MTPR15, OK12, SVM^{+}24b, CBB19, CBSG24, CBL17, GJ07, OK14, OK10, OV24, PWM23, SZC11, SG24, Bag11, BP08, SVM^{+}24a]. long-cultured [SZC11]. long-lived [CBB19]. Long-Term [SVM^{+}24b, AGT08, Ano92r, CCMS12, OK12, CBL17, OK14, OK10, OV24, PWM23, SG24, BP08, SVM^{+}24a]. longest [RR11]. longimanus [YC20]. longline [ENP18, FTH15, GH17, GCM18, GCS20, GCK21, GYH10, Koc01, LSF14, MCL08, MG18, War08]. longliners [WH07]. longlines [CML19]. Longtail [GLM20]. look [Fro99]. Loricariidae [BVMF13, BZM12, BVM13, PASF13, RZV12]. lorigera [HLD07]. lost [SEH24]. lotic [PCG19]. Low [IPT10]. Low-temperature [IPT10]. Lowe [BMF09, CGCG18]. Lower [CBEGR02, VRRCAG02]. Ltd [Coc05]. Lucia [WTC06]. lucioperca [SW12]. Lunas [Mag13]. Lutjanidae [BDS05]. Lutjanus [EBS20]. Lycoteuthidae [HLD07]. Lycoteuthis [HLD07]. lysine [KTW12].

M [Bla98, Har99a, Har98, Nor02, WBW09, Woo99b, SCC09, HHCM10]. Maccullochella [SQR09]. mackerel [AGG03, NRCD^{+}23]. mackerels [JJMD13]. MaClean [Ste05]. MacLennan [Bur07]. macro [KKW12, SJB^{+}23]. macro-level [SJB^{+}23]. macroalgae [PFW16, PFW17]. Macrogard [SZG10]. macrophthalmia [MJM15]. macularius [VRRCAG02]. magnitude [BHD23]. Mahisefid [ADS11]. Mahseer [PBR19, BP16]. main [dMV20]. Maine [SAK14]. mainland [IOHM23]. Maintaining [Lig16, JBJ06]. maintenance [RFH15]. Major [PR05, ANP^{+}23, GBR22, KKW12, SRBS21]. make [MWS22]. makes [BTS97, MKK10]. Making [MKL^{+}23, Ano95f, BJM20]. Malawi [Tur99a, Pau10]. Malaysia [SKW22]. Malaysian [Ara15]. Maldonado [WBW09]. male [Kva98]. males [HKL10, NVA12]. mammal [HB19]. mammalian [HF20]. man [Pie92]. manage [PM19]. managed [AFB15]. Management [Ano93f, Asw05, ALJ08, Bil02, Bla98, BDS05, CBR98, DJ05, De 98, EN00, FB05, Har99b, Smi98, SVM^{+}24a, SVM^{+}24b, ANL12, ATdLCBR02, AH96, Ano92c, Ano92t, Ano93g, Ano94f, Ano94b, Ano94d, Ano95a, Arn96, BGR21, BS03, BBS10, BM09, BDF09, BBW09, BdST16, BCL21, BOD21, CHFTV22, Cad99, CTL17, CCR^{+}24, CDC09, CD01, CMF^{+}24, CM92, CSS20, CVD21, CDH^{+}23b, CGR11, CDS16, CHS18, CJV13, DSU08, Dav96, DC05, DBR15, DDD16, DLP95, ETWE12, Fer94b, FCH16a, FCH16b, GKR10, GCS19, GHMG22, GCB^{+}24, GOP^{+}23, GCUR22, GLM20, HCK21, IH04, IWG17, JZ00, Jia09, Jon07b, Kat97, Knu01, Lar96, LMH21, LC02, LM19, LBS^{+}23, LPC^{+}24, LSH15, MV13, MP16, MBH14, MDV^{+}23, MLN05a, MLN05b, Moo12, MLK13, MRMJJ17, MH20, NMF17, Nor95, OGR17, OK14]. management [OMAGL02, Pat92, Pau97, PZC17, PWB99, PAK^{+}23, REA^{+}23, RSGS^{+}23a, RSGS^{+}23b, RP91, SBSŠ21, SK15, SVC21, SJH22, SVRS19, SAC^{+}23, SLK16, Tan03, TKB18, TBA20, VGA11, VSM12, WP96, WGL17, YMR12, ZMG17, ZKvZ19, vOR15, MAJ05, Sul04, Har98, Ram08]. Management-oriented [De 98]. managers [BLC19]. Managing [Chr96a, Cha95, KFS21]. Máñez [KCR17]. Mangel [Mil99]. mangrove [CGKSP13, IH04, MMB13, Whi17]. mangrove-based [IH04]. manipulated [GKR10]. Manipulation [DJ05, BS03]. map [GPS18, WJP23]. Mapping [PASF13, MPO11, MBdBC13]. Marc [Mil99]. marginatus [CGCG18]. Margins [Bla98]. Mariana [CBH15]. mariculture [FHK04, HP93, LLJC24, RK10]. Marine [CFB14, CWO19, Cus94, GLC98, Hus04, Jon02, Pis00, Ros00, Sto00, TCS20, TKB18, UE02, WBW09, YKS14, AYN11, AFO22, Ano92f, Ano93a, Ano94o, Ano95f, Ber93, BAA18, BBW09, BSKB22, BC03, BOJ16, CA04, CCR21a, CCR21b, CAB19, DIR20, FHvP14, GMP12, GBR22, GJA17, GKC19, GK18b, HB19, HP93, HK14, HCMT24a, HCMT24b, HTH00, HP14, HCvP16, Hun21, IOHM23, Jen00, JLD^{+}24, Jia09, Jon07b, KPN21, KCR17, KBV11, Lar96, LC02, LET^{+}23, LM19, LB23, LBS19, MCFC21, MFV19, MBJ12, MTC19, McD02, MBH14, Mil12, Mis97, MHvH16, MVGW18, NCMT23, OAJ14, OWW04, PAK^{+}23, QM04, RFH15, RP91, RHM11, RWH04, SEA22, SAB^{+}23, SBtIWG12, SVRS19, SLH11, Sot02, SDS15, SJB^{+}23, Sur23, TRJ24, TFP22, TAS10, TS14, VaR92, WZB^{+}24, WMTL22, WGC18]. marine [WSGP22, WW06, YH18, YOC15, ZL23, dMV20, Dri05, JBJ06, LFJ08, Ano93b]. marinus [HMQ16, MSM20]. Maritime [Bla98]. markedly [dSSHK21]. markers [AHW04, BSV11, CH99, Fer94b, SKU20, SZC11, TdSL15, WB94]. marketing [RBK10]. markets [GSC17]. marking [Ano92l, DOP18, WMDS18]. marmorata [PLA07]. marshes [Nor03, Nor06, Whi17]. Martell [MAJ05]. mass [Nor14, SM07, WCP97, WMDS18]. mass-balance [WCP97]. massyae [OJB07]. masu [Jua10]. Mate [RHP99, ANB19, EP00]. maternal [GCSR09]. Mathematical [Ano92s]. mating [Bas93, HII16, HLD07]. Matos [WBW09]. matter [CMF^{+}24, WHS04]. matters [HCB15]. Matthews [Woo98b]. Maturation [COGFPV23, MAP21, ARL12, JJ93, SFO14b, SFO14a, STS07]. mature [BEBFM14, PM14]. maturity [TS14, WMA20]. Mauritius [ASS^{+}23]. maximum [FN02, HPY18, Sch18]. maximum-length [Sch18]. maximus [IJ01]. May [Pit98a]. mbuna [Pau10]. McDonald [Soi99]. meadows [Whi17]. meal [TRG19]. Measurement [EB93]. measurements [SW12]. measures [AGT08, AA10, PAK^{+}23, SQR09]. Mechanics [Ano95g]. mechanisms [AF04, BBS17, KC14, MP07a, MVM99, MBP14, Rob92, SYL^{+}23, SZG10]. mechanistically [HCB15]. mechanosensory [MCH95]. media [SEA22]. mediated [CLB^{+}22]. Mediterranean [CCR21a, CCR21b, CCR^{+}24, CUT07, RMN11, AVB^{+}23, AA10, AS95, AE22, BKB17, CCP13, FCOCNC24, GMP12, MA08, NCMT23, PVL21, PBC12, RT05, RK10, SCC09, SK02, TAS10, TS14, VVU22, VGA11]. meeting [BP08, CSSO02, FD00]. megafauna [LB23, TKB18]. Mekong [SGD^{+}23, KHW09]. melanodermatum [MdAVA11]. melanostomus [dSPPM12]. melas [CTB16, RBK13]. Melatonin [Shi05, BLD19, Kul95, Kul02]. meme [Pau96]. Mendota [Ano94d]. mentella [SRGS04]. Menz [Tur99a]. Merluccius [SSBCL11]. MesoAmerican [BGOW23]. mesopelagic [CGK11]. Meta [GSD18, CNH22, EBS20, LMCB^{+}23, ROP^{+}23, dSSHK21, VCZ19, VLD20, WPF16]. Meta-analysis [GSD18, CNH22, LMCB^{+}23, dSSHK21, VCZ19, VLD20, WPF16]. meta-analytical [EBS20, ROP^{+}23]. metabolic [BCP22, MVM99]. metabolism [DSK11, GP12, RBPPS16, SCC08, WLB13]. metal [EDP10, Nor02]. metamorphosis [EJD18, MP07a, PM09]. Metapopulations [McQ97]. methanesulfonate [CWB11]. method [LS07b, LH18, Sch18]. methodological [JN18]. Methods [BP93, Ano95a, DML^{+}24, DPS22, MKL^{+}23, MSM20, NPS17, SA12, SPB00, Tel09, WFH17]. metrics [GMS17]. Mexican [HPdL02, PJ14]. Mexico [AMVC20, CRRCdL08, CRZHZ02, CGR11, Dia04, EBS20, EBRLGB02, GPS18, LLSDTJ09, Mag13, May02, MSLN02, Nel92, OGR17, RVRCB11, RCCRD03, TJLLC10, VRRCAG02, WM02, CBAVGR02, EGMM02, RCCASG02]. Mg [KKW12]. Michael [ALJ08]. Michigan [Zim05]. micro [SMS12]. micro-organisms [SMS12]. microbiota [TRG19]. microchemistry [PGC19]. microcohort [MP07b]. Microflora [GP12]. Microplastic [WRSG21]. Micropterus [PV19, TP14]. MicroRNA [RNJ16]. Microsatellite [OW97, KKF10, OK14, SKU20]. Microsatellites [WB94, PBB00]. microstructure [Mok93]. Mid [Smi98]. Mid-century [Smi98]. Middle [MKS12, Ano94i, EGMM02, Jon02, MIA^{+}23b, MIA^{+}23a, Mag13]. middle-sized [Ano94i]. Mignucci [WBW09]. Mignucci-Giannoni [WBW09]. migrating [MJM15]. Migration [BOJ16, BB14, CLH21, HK14, JJ93, MMG18, SPS07, SAT11, TØH08]. migrations [Cor02, QM04, WMM92]. migratory [DLP95, EG03, FSSM17, HWA14, SOB20]. Miller [MBJ12, HNSS02]. Milner [Ram08]. Mind [Bro05]. Minho [MBA15]. minimum [CML19]. Minor [PR05, FIØ04]. minuta [LAI05]. miodon [Mar93]. Misidentification [PAT15]. mismatch [HETS23]. Misund [Gla00]. misuse [JNW14]. mitigate [GCK21, MKK10]. Mitigating [BTZ^{+}23, HGC17, Kos09]. mitigation [AA10, DPS22, HB19, LB23, NPS17, TFP22]. mitochondria [JB95]. mitochondria-rich [JB95]. Mitochondrial [CRRCdL08, AVL07, DHG18, SKU20]. mixed [BHP15, RCP19]. mobility [ROP^{+}23]. Mobula [AUdS18]. mobular [AUdS18]. model [CSS20, CCA17, ETWE12, JEL10, Mar08, UIA11, WK99, YCT11, WM02]. Modeling [CSA11, ETWE12, LAWD06, NRCD^{+}23, OGR17, PMVA19, CSA12]. Modelling [CSY12, CAC15, GHF98, HDI15, HSS21, BdST16, MTC19, SP98]. Models [Mil99, PB98, Ano92j, BB14, CTL17, CRPI^{+}22, GBCCO12, KW08, Nee01, SW14, TR94, WCP97, dlCFP19]. modern [MGB15]. modes [MS02]. modification [BS03]. Modifications [Bro00a, MBS17]. modulated [TRG19]. Modulation [DW09]. mojarra [APLL07]. Mola [PHH10]. Molecular [BEBFM14, CH94, CH99, Els04, Fer94a, HRC09, Lin94, Tel09, TRB13, Web02, CBF^{+}24, Fer94b, MAP21, PM94, RBPPS16, SSBCL11, Utt94, WG94]. Mollusca [SBP07, YVR07]. monitor [TMKC24, VaR92]. Monitoring [GPS18, MMB22, NCMT23, SVM^{+}24a, SVM^{+}24b, ADC15, Ano94p, DCS20, DDD16, ENP18, GLG12, GLC16, MT18, SX16, SLH11, WNB12]. Monographs [Mil99, Rei99, RHP99]. Monopterus [Shi05]. MOP [De 98]. Mordaciidae [MBG21]. Moreau [HHCN11]. morphologic [LFdSRM16]. Morphological [PCG19, BOs12, BOV09a, BOV09b, LH18, NLFM12, RBK13, SZC11, SSBCL11]. morphology [AHL12, MR18, OJB07]. Morphometric [RCCRD03, Cad00, DD13a, DD13b, RVRCB11]. morphometrics [SN00]. Morphometry [GGZ10, PTOS^{+}24]. Morphophysiological [MAC02]. Mortality [BB05, Soi99, AVA19, Cad91, CML19, GH17, KM20, MG18, Sch18, SZG10, SM07, VdSD21, VLD20, WH24]. mossambicus [RTT12]. Most [AH23, MKL^{+}23, OV24]. motility [AHL12, PFL10]. movement [BdST16, MC16, MMG18, PHK20, SHS14, SPS07, VEK10, VBC^{+}23, WJP23]. movements [AGM19]. moving [GCS19, HYW13, LM19]. Mozambique [CBO19, KL23, RTT12]. mtDNA [PBB00]. mucosal [ZRZ20]. mucus [She94]. Mugil [WPD12]. Mugilidae [HRC09, WPD12]. mullet [WPD12]. mullets [DHB17]. multi [HSS21, YOC15]. multi-voyage [HSS21]. multi-workshop [YOC15]. multifaceted [AAH98]. multifunctional [Kul02]. multilocus [TdSL15]. multiple [AGT08, HFG07, PRN95]. multispecies [HPY18, Mag95, MLSP^{+}23, SP98, WK99]. murk [dMlTB19]. Murray [SQR09]. Muscle [Joy23, KTW12, Sän93]. mussels [CDC09]. must [Gau01]. myctophid [VRR12]. Myctophidae [CGK11]. mykiss [TRG19]. Mytilidae [CDC09].

N [Bur07, Qui12, Sei08, WBW09, CNH22]. Nacional [WBW05]. Nakamura [Woo99b]. Nakashima [dC98]. namaycush [MBK12]. names [Ano92b, Gil93a]. nano [ZHP07]. nano-scale [ZHP07]. NanoSIMS [ZHP07].

National [ATdLCBR02, HBC13, RCP19]. Nations [Bla98]. Native [Ano94j, ACB08, BGM^{+}24, BO12, CTB16, CTM92, GP15, GH16, HH06, HPdL02, LMCB^{+}23, LMM11, MGB15, MA08, PV19, SL15, XSC15]. natives [GP15]. Natural [Ano94f, AH96, BP16, Cad91, HHCN11, Man94, MAC02]. naturalised [PEP12]. Nautiloidea [PDGBR07]. Nautilus [PDGBR07, SBP07]. Navigation [WA03]. near [MSR14]. near-pristine [MSR14]. nearshore [IH04, MFL21]. nebulosus [RBK13]. necessarily [CCMS12]. Need [AH23, BEC11, ST94]. needed [CCP13]. needs [CCH18, MC12, OGR17, PLW17, Vil03, YCT11]. negative [LSH15]. negatively [BCA^{+}23]. neither [McD02]. Nelson [Win06]. Nelusetta [MS13]. Neotropical [LSB22, SGF13, FSSM17, NKF21, QSBV18, dBdSA12]. Nephrops [AS08, AVB^{+}23, CR08]. nerka [MHP12]. nesisi [Bol07]. net [BCA^{+}23, ZL23]. nets [LH18, MF99]. netukulimk} [MP16]. networks [MVGW18]. Neural [Rob92]. Neuroendocrine [HL98b, PY97]. neurosecretory [Mun95]. Newfoundland [MDR14, MMG18]. newly [RK10]. newly-introduced [RK10]. next [CFS^{+}23, NRJ^{+}23]. Niassa [Tur99a]. niche [BSL18, JJ93, SHHK21]. Nicholas [VAJ10]. Nigel [Ram08]. Nigeria [AY10, TFE07]. Nile [KL23]. Nine [MC16]. Niñno [Csi99]. Nitrogen [HP93, DSK11]. nitrogenous [HP93]. NJ [MAJ05]. No [Ano93e, BB05, Rei99, FN02, Jon07b]. No-take [BB05, Jon07b]. Nodularin [SDG15]. noise [HPP15, dJFS20]. nomenclatural [OJB07]. Non [BO12, BC03, HCNH24, LMM11, MA08, RSB16, RHM11, SL15, XSC15, BCA^{+}23, BBS10, BGM^{+}24, CTB16, DZZZ22, Duc19, GP15, GKC19, GH16, HH06, Kim93, Koc01, KZ20, MWS22, RR11]. non-classical [Kim93]. non-coding [DZZZ22]. Non-indigenous [RSB16, KZ20]. Non-infectious [RHM11]. non-marine [GKC19]. Non-native [BO12, LMM11, MA08, SL15, XSC15, BGM^{+}24, CTB16, GP15, GH16, HH06]. non-oophagous [MWS22]. Non-physical [BC03]. Non-probabilistic [HCNH24]. non-random [Duc19]. non-stationary [BBS10]. non-target [BCA^{+}23, Koc01]. Norris [GdMD13]. Norse [JBJ06]. North [CBL17, KWI07, Nel92, Ste05, Tay99, Ano94j, ABK14, BGR21, BOs12, BM09, CM92, DAN02, Gil93b, Gre93b, JNW14, JZ00, KCE12, MS02, Nel92, Nor03, Nor06, PVJ07, RN04, SAT11, VaR92, WGL14]. Northeast [Har99c, HMN17, VVU22]. northeastern [SSBCL11, TdSL15]. Northern [CCR21a, CCR21b, CBH15, Gou16, HFG07, Lig16, RSB16, STP22, WM96, CBO19]. Northg [BP08]. Northwest [MLK13, dC98]. Northwestern [AMVC20, CRRCdL08, KZ07, OV24, RCCRD03]. norvegicus [AS08, AVB^{+}23]. Norway [Hus04, AVB^{+}23, SA12]. Norwegian [Hus04, CKN21]. Norwich [Ano94a]. Note [DD13b, AF04, GBCCO12]. notomelas [MPO11]. Notonykia [Bol07]. Notorynchus [SRBS21]. nototheniid [KdSD18]. notothenioid [LE24]. nov [Bol07]. novel [GZT13]. November [BP08]. nuclear [CZF22, MKS12]. Number [McD03, BZM12]. Numeric [RZV12]. nursery [Whi17]. nurture [FMM22]. nutrient [NMS13]. Nutrients [AH23]. nutrition [AH23, UIA11, dSKV16]. nutrition-sensitive [AH23]. Nutritional [HF20, LPG17, UIA11]. NW [PTOS^{+}24].

objectives [CBR98, Mus03]. objects [CSSO02]. obligate [SHB21]. obscura [KMF13]. Observations [Est05, BOV09a, BOV09b, JN18, MBK12]. observatories [ADC15]. observed [MSR14]. obstacles [RK10]. obtained [JLD^{+}24]. occur [PM09]. Occurrence [JMJ19, LI03, Jon07a]. Ocean [BNC22, JBX07, NAP22a, NAP22b, PHB14, PAN22, PBS14, Ste05, Ano92f, Ano92j, BB14, Cus94, FMM22, HMH22, HSC^{+}24, KEP22, LPC^{+}24, MS13, MMM22a, MMM22b, NvPV22, PHH10, TFP22, ANL12, BOV09a, BOV09b, CGKSP13, CBO19, DHG18, FHW^{+}24, GOW13, KWM^{+}22, Koc01, LFMP21, Par92, RN04, SKU20, Ste05, TKB18, VRR12, WNB12]. Oceania [Asw05]. oceanic [HYW13, SRBS21, YC20]. oceanography [Bar93]. Oceans [GOP^{+}23, NvPV22, AFO22, GB23, KCR17, KEP22, MMF18]. October [KGW10, KGW11]. octopus [APP^{+}23, DMD21, OBS08, OBS08]. Octopuses [LBS^{+}23]. Odontocete [BHB22]. Oegopsida [Bol07]. off [EBRLGB02, GYH10, MS07, ZFT13]. official [KLC18]. offs [Hal98, RCP19]. offshore [NBF22]. offspring [GCSR09]. oil [Lin94, SOB20]. Olafsen [FIØ04]. old [CH99]. oldest [RBC16a, RBC16b]. Ole [Gla00]. olfaction [Har94]. omega [GSD18, YWS^{+}24]. omega-3 [GSD18, YWS^{+}24]. ommastrephid [FHW^{+}24]. Ommer [Pis00]. omnivores [NKF21]. omnivorous [ZMG17]. Oncorhynchus [CM08, Est05, KFS21, MHP12, QGM15, RN04, SAK14, SBN04, TRG19]. One [DC05, OSC20, ZLF21, CPM14, MBA15, MLK13]. only [GCUR22, Jon07b]. onset [HMN17]. Ontogenetic [CLH21, CFGG13, PBF15, FHW^{+}24, SHPH^{+}22, WOR09]. Ontogeny [Kam02, MDM20, OJB07]. onychoteuthid [Bol07]. Oocyte [TS96]. Oogenesis [PLA07]. oophagous [MWS22]. open [BBS10, DPC12, HK14, JCL07, Mar06]. open/closed [JCL07]. operational [SW95b]. Operations [Ano92t, Ano92x]. opilio [MDR14, MMG18]. opportunities [ANL12, BHD23, GS23, PBR19, PMV18, dMlTB19]. opposite [LPC^{+}24]. optimal [APLL07, DML^{+}24, Mar06, Mar08, WR16]. Optimising [SX16]. Options [DJ05, BDF09, BTW15, PCW17]. oral [MT18]. orange [Jon07a]. order [RG99]. Oreochromis [RTT12]. organ [EM97]. organisations [OSB^{+}23]. organism [UIA11]. organisms [Ano95d, SMS12]. organization [SGF13]. Organogenesis [HHCM10]. organs [PM14, PDGBR07]. oriented [De 98, HJC09]. origin [BVMF13, BSM17, CGL22, EJD18, For95, GGZ10, GCSR09, McD02]. Original [Pau98]. origins [CBA10]. ornamental [COGFPV23, LNT^{+}23, dSPPM12]. ornata [May02]. ornate [May02]. Oscillation [BGR21]. osmoregulation [Kul02]. OSR [Kva98, SW98]. Osteichthyes [BZM12]. other [AGT08, Avi00, EAW22, Nor03, PHK20, RN04, SA12, Utt04]. otherwise [SEH24]. Otolith [Ano94h, Mok93, SG24, WMDS18, EG03, MDM20, PGC19, RSGS^{+}23a, RSGS^{+}23b]. otoliths [SBN04, WDP19]. Our [Pau98, AFO22, Bro05, FMM22, GB23, SVRS19, VBC^{+}23]. outcomes [AFO22, GLC16]. outlook [PMV18]. Over-exploitation [HTH00]. over-quota [BHP15]. Overciew [Fui99]. overfishing [LSB22]. overlapping [ZL23]. Overlooked [BGTA19]. Overview [Hup05, PMV18, AA10, BBV12a, BBV12b, CA04, CBSG24, FCOCNC24, FTH15, Mag95, MSR03, PVL21, SYL^{+}23, Sny05, WH07]. ovulation [PY97, PSK12]. Oxford [Coc05, Els04]. oxidative [VdSD21]. oxygen [BSR19, Nor14].

P [Har99a, Har98, Web09, WBW09, KKW12]. Pacifastacus [GCSR09]. Pacific [Ano94e, RN04, SKU20, ANL12, Ano92n, ABK14, BBJ12, BM09, CLH21, CGKSP13, CGRCM19, CWO19, EBRLGB02, FHW^{+}24, GOW13, Hun21, JMJ19, KFS21, KWI07, LBS^{+}23, LR11, MT14, NWO16, OAJ14, OV24, PDA12, PJ14, PQS14, PFW16, PFW17, QM04, dSSHK21, SAT11, SSBCL11, SFO14b, SFO14a, SW14, TMD^{+}24, WH07, WGL14, WNB12, ZZ04, ZLF21]. paddlefish [BL00, JZ00]. Paedomorphic [LAI05]. pages [Ste05]. pairs [Tay99]. Palauan [MG18]. Palauan-based [MG18]. Pan [FCCS15]. Pan-Atlantic [FCCS15]. Panama [CGRCM19]. pandemic [BPA^{+}23, OSB^{+}23]. Pantojas [WBW09]. paper [Kva98]. Paperback [MAJ05]. Papers [Har99b, Ano91b, Ano91c]. Parachanna [KMF13]. Paradigm [De 98]. paradigms [Cad99]. Paralithodes [DD18]. parameter [AS96]. parameters [Ano92j, GAD10, HHCN11, LMH21]. Paraná [AGO04, BO12]. parasite [BBJ12]. Parasites [PGG07a, PGG07b, WMM92, Ano94m, BHK00, DPC12, Mar95]. Parasitic [Ano95i]. Parasitofauna [DPC12]. Parent [Kam05]. parentage [BHP11]. Parental [AS95, EP00, SW95a, SW95b]. Park [CFB14]. Parodontidae [BSV11]. Partial [JJ93, AVL07]. Partially [HCK21]. participation [CMF^{+}24, DCS20, PRP16]. Participatory [DCS20]. particular [Koc01, RTT12, WH07]. partitioning [CFB14, PBF15, SN00]. partner [DW93]. partnership [PRP16]. partnerships [CDS16, DDD16]. parvipinnis [RVRCB11]. passage [AVA19, CCO20, PDB16, TAD14, WMW18]. Passive [SVM^{+}24a, SVM^{+}24b, WH24, Jua02, KCF14]. Past [LAWD06, MLK13, AMV13, IRPG21, KCR17, SK15, STW^{+}24, TSB22]. Patagonia [ACB08]. Patagonian [PWM23]. path [KMK16, Mar08]. pathogenic [SMS12]. pathogens [JMJ19]. Pathological [RK10]. pathways [BCK16, NLFM12, PAN22, TBB^{+}24]. pattern [HSC20, SPL12, SDJ13]. Patterns [AGM19, Jen00, LKH16, PFW16, SKU20, TV15, TJLLC10, Utt00, Woo98b, ACB08, AF04, APLM14, CRZHZ02, CZF22, EG03, EAW22, FCCS15, FC07, GKC19, LET^{+}23, LLSDTJ09, NWG21, OSB07, PDN^{+}24, SHPH^{+}22, SCC09, SEA22, SPS07, SFO14b, SFO14a, VEK10, PFW17]. Pátzcuaro [OMAGL02]. Paul [GCS19]. Pauly [Ste05]. pay [GCS19]. pbk [Ste05]. Pearl [ZCW19]. Peces [WBW05]. pedigree [OK14]. peelii [SQR09, SQR09]. Pelagic [Ano92u, Gla00, LY07, Tur99a, Ano92r, Ano94i, BAA18, CBR98, CML19, GH17, GCM18, GCS20, GCK21, GYH10, GOW13, MCL08, MIA^{+}23b, MIA^{+}23a, Pat92, WH07, War08]. penaeid [MBS17]. Peninsula [RVRCB11, SAC^{+}23]. Peninsular [SKW22]. PepT1 [KTW12]. perception [ASS^{+}23]. perceptions [BGM^{+}24]. perch [KTW12]. Perciformes [NLFM12, Pau10, PNC11]. Perdido [AMVC20]. Perfect [Ste05]. Performance [KGF10, MLSP^{+}23, DIR20, GKR10, LE24, MFL21, Neh96, VD22, WA03, ZJDZ10]. performances [GGZ10]. perils [Nor12]. period [TAS10, vOR15]. periodicity [OK12]. periphyton [vDBV02]. permanent [OV24]. Persian [BHD23, Moo12]. persistence [CBA10]. personality [KDF13, ZLH^{+}23]. Perspective [FB05, Kam05, BHK00, BZM12, CM92, DZZZ22, Fro99, HETS23, JN18, Kam02, KCE12, PGA^{+}24, PE08, Pou93, TCS20, WK99, WG19, YMR12, dMV20]. Perspectives [DLP95, Utt94, dC98, BSM17, BEC11, BVM13, DW11, Fer94a, JBX07, JZ00, MSR03, NVH21, NRJ^{+}23, PTP14, PHH10, SCC08, SPL12, Utt04, KCR17]. perturbations [HFP14]. Peruvian [Csi99, FSB14]. Peter [GCS19, Nor96]. Peters [RTT12]. petrels [Koc01]. Petromyzon [HMQ16, MSM20]. Petromyzontida [EJD18]. Pfeffer [OJB07]. phages [SMS12]. pharaoh [AVL07]. pharaonis [AVL07]. Pharmacokinetics [RT05]. pharyngeal [RL05]. phenology [MHP12]. phenomena [Avi00]. phenomenon [PE08]. phenotype [SZC11, WPF16]. Philippines [LAWD06]. Phoenix [MSR14]. Pholidoteuthidae [OJB07]. Pholidoteuthis [OJB07]. phosphate [OBS08]. Phylogenetic [May02, BSV11, BVM13, CZF22, JJC22, McD97, McL94, SMNK^{+}23, ZDJ19]. phylogenetics [WMM92, dP99]. phylogeny [CPN11, HRC09]. Phylogeography [AVL07, MOF11]. phylogroups [LKK10]. Physical [Ano94o, Ano92f, BC03, KZ20, VdSD21]. Physicochemical [Nor06]. Physiological [BCL21, LR11, AS08, BBS17, MBP14]. Physiology [Ano94a, HCB15, Ran92, Ano94l, Ano95g, BCP22, CB00, CBP02, DZZZ22, LPG17, ODS10, ZLF21]. phytoplankton [BSWA14]. Piabucus [dSPPM12]. Pickering [Paw99]. piecemeal [GCS19]. pieces [PZC17]. Pierre [Gla00]. Pigmentation [LNT^{+}23, Col10]. pike [PGA^{+}24]. pikeperch [SW12]. Pikitch [Har98]. pilchard [BOs12]. pilchardus [BOs12, CHFTV22]. Pimelodella [GdMD13]. Pimelodidae [MdAVA11]. pineal [EM97]. Pink [RN04, WHS04]. pipefishes [KBV11]. pipelines [BMN21]. piraíba [PBG04]. Pisces [RCCRD03, RTT12]. piscicole [Bil02]. piscine [DW11, Har24]. PIT [VLD20]. Pitcher [Pis00]. pitfalls [UE02]. pituitary [APLL07]. place [McD97, McD10]. places [Pau97]. plains [HK14]. plan [OMAGL02]. Plankton [Bar95]. Planning [HCvP16, CA04, GMP12, SW14]. plans [Bil02, CA04, Mil12]. plasma [ARL12, KdSD18, Nor09]. plasma/serum [Nor09]. plastic [dMV20]. Plasticity [Ree02, AWC17, Joy23, NKF21]. plateau [PBM14]. platforms [SOB20]. plays [GP12]. plea [ETWE12]. Plectropomus [PDN^{+}24, PCW17]. Pleistocene [DHG18]. plenty [GOP^{+}23]. Pleomerism [McD03].

Pleuragramma [BOV09a, BOV09b]. Pleuronectiformes [Bur10]. plight [GLHD^{+}24a, GLHD^{+}24b]. ploidy [PFL10]. plot [Sch18]. plumieri [APLL07]. Point [Jon07b, CSSO02, CGL22, FN02, FD00]. point-of-origin [CGL22]. Point-of-View [Jon07b, FN02]. Points [Kat97, Kva98, MN98, PB98, Ros97, SW98, SP98, Ano93g, RZV12]. Poirino [GGZ10]. Poleward [MTAP22]. policies [ZKvZ19]. Policy [Bro05, CVD21, DROCM23, GC23]. politics [Ter01]. Pollicipes [PTOS^{+}24]. pollutants [DW09]. Pollution [Els04, HP93, JR97, Kim95, TFE07, VaR92, WSGP22]. Polunin [VAJ10]. polyactis [LLC11]. polycentric [CHN18]. polychaete [CCH18]. polymer [HCB11]. polymer-based [HCB11]. polymorphic [AHW04]. polymorphism [MdAVA11, RZV12]. polyphemus [SBZR17]. polyploidy [LI03]. Pomacentridae [RL05]. pomacentrids [COGFPV23]. pompilius [PDGBR07, SBP07]. pond [GKR10, KGF10]. Ponto [GBOK23]. pool [Mar08]. poor [PDA12]. Poorly [SKD03]. Population [BDS05, CM98, Els04, HMQ16, JEL10, Mil99, MMF18, RVRCB11, ROW^{+}23, SHHK21, Utt04, BOs12, BdST16, BB14, CHFTV22, CSA11, CSA12, CBB19, Chr96b, DGV11, DHG18, DD15, GLG12, GLM20, HJC09, HP07, IJ03, Jen00, MCFC21, MDF14, McD10, MOF11, MRMJJ17, NWG21, PDM20, PBB00, Ros00, SRGS04, SW14, TR94, VGA11, VSM12, WMM92]. population-specific [BB14]. Populations [Ano93f, AKGB11, Ano92r, Ano94k, CRRCdL08, CPM14, CFS^{+}23, CTB16, CLB^{+}22, DGV11, DD13a, DD13b, JEL10, KKF10, MPO11, MBH14, MP07b, MBA15, PM19, RCCRD03, RTT12, SA12, SBP07, TMKC24, VSM12, WGL14, YKS14]. porosus [MOF11]. Portugal [SVM^{+}24a, SVM^{+}24b]. Portuguese [BMF09]. position [GH17, SM07]. possibilities [ROL14]. Possible [McD03, Mun95, RT05, AF04, BLD19, NVH21, PCW17, Tel09]. Post [MG18, CBF^{+}24, GMS17, HII16, LSDRS23, PM14]. post-copulatory [HII16]. post-COVID-19 [LSDRS23]. post-genomics [CBF^{+}24]. Post-release [MG18]. post-spawning [PM14]. post-tagging [GMS17]. Potamotrygonidae [LSB22]. Potential [CLX15, FMH07, PB98, SOB20, Tur99a, BD20, BC03, DGV11, DAN02, FGL10, FIØ04, HAC18, IFW05, Kat97, LF13, LLJC24, MJM15, Mus03, PJ08, PJ14, RCP19, RM01, SQR09, Sot02, TMKC24, vDBV02]. potentialities [UE02]. potentials [AY10]. Pots [PVL21]. Poulsen [KCR17]. Powell [Sch18]. power [War08]. powerful [PV19]. pp [Coc05, Zim05]. pp. [Dri05, Hus04, MAJ05, WBW05]. practical [BEC11, CWB16, CRPI^{+}22]. practice [CNT21, DSU08, Gra96, RBC16a, RBC16b, Bur07]. practices [AFO22, Ano93e, BGM^{+}24, CD01, CKN21, MV13, RSSS23, SJH22, WH07, Har99a]. practitioners [SJH22]. prawn [Bro00a]. prebiotics [GDC13]. precision [Vil18]. predation [LS07a]. predator [Chr96a, FM94, PV19]. predators [BAA18, HYW13, YOC15]. predatory [BCM20, Bro94]. predict [MDM20]. predicted [SN00]. Predicting [LSB22, dJFS20, MLSP^{+}23]. prediction [Jen00, OK12]. Preface [KGW10, MJS07, MT03, MMM22a, KGW11, MMM22b]. preference [WR16]. preferences [BGM^{+}24, KL23]. preparation [WDP19]. Prepared [FCP19]. Preparing [HJC20]. prepupae [TRG19]. presence [HCMT24a, HCMT24b, vPKW18]. Present [GLC98, Har11, DD18, MLK13, TSB22]. Press [Dri05, MAJ05, Ste05, Zim05]. pressure [DIR20]. pressures [MTPR15, TAD14]. prevent [CCMS12]. prey [Chr96a, SN00, WW06]. prices [HCMT24a, HCMT24b]. primary [ZMG17, ZL23]. primer [FQSJ23]. Prince [PDA12, WHS04]. Princeton [MAJ05]. Principles [Har99a, Rey93, CDH^{+}23b]. Prionace [KWI07, MG18]. priorities [APP^{+}23, JZ00, Moo12]. priority [RSSS23]. Pristidae [WSC09]. pristine [MSR14, MLK13]. probabilistic [CSS20, HCNH24]. problem [Hal98]. Problems [PB98, Mil12, Rey93]. procedures [CBR98, TTC11]. Proceedings [Hus04, KGW11, KGW10, Ano94a]. process [BTW15, ETWE12, HJC09]. process-oriented [HJC09]. processes [Ano94o, LKH16, LET^{+}23, MAP21, MIA^{+}23b, MIA^{+}23a, VEK10]. processing [Ano92b]. produce [HKL10]. producing [PDGBR07]. product [GGZ10]. Production [Man94, PR05, ADS11, CRPI^{+}22, Gib93, HWA14, HETS23, LNT^{+}23, NRCD^{+}23, OSB^{+}23, PBC12, RL05, RBK10, SOB20, Tan03, WHS04, ZMG17, ZL23, vDBV02]. Productive [LFJ08]. Productivity [Tur99a, BSKB22, HSC^{+}24, SHHK21, WHS04]. Professor [SHH95]. profile [AMVC20]. profiles [KdSD18, ZJDZ10]. progeny [Kam05]. prognosis [Nee01]. Program [Csi99, WBW09, KGF10, MSM20, PRP16, YWS^{+}24]. programming [HF20]. programs [GPS18, Kos09, OK14]. Progress [Mil12, ZRZ20]. projected [MLP17, PGJ15]. promote [CCP13, SQR09]. properties [IJ03]. Property [Arn05, Han05]. proportions [EP00]. proposal [LSDH12]. prospect [KMF13]. prospects [Col10, LSP11, VAJ10]. protected [BBW09, GGZ10, GK18b, HCK21, Jon02, Jon07b, Kel23, LET^{+}23, Sot02, TRJ24]. Protection [RUL95]. protein [ITV98]. proteins [CZF22, DW11, Joy23, KTW12]. Protocol [AVA19]. protocols [AF04, CWB16]. Protozoan [Ano94m]. Proudlove [Wil07]. Proved [CS05]. proven [KM20]. provenance [DMD21]. provide [GCUR22, PHB14]. Providence [BP08]. provides [ZHP07]. providing [PWB99]. provinces [MKS12]. provision [GOP^{+}23]. proxies [IDG16, WLB^{+}23]. proximate [KKW12]. prudent [GC23]. puberty [BLD19]. public [BCK18, KEP22]. publications [BBV12a, BBV12b]. Publisher [Bra99]. Publishing [Els04, Zim05]. Puerto [WBW09]. puerulus [MP07a, PM09]. pufferfishes [NVA12]. pulse [Mag13]. pulsed [WW06, YCT11]. pulsed-flow [YCT11]. pupfish [VRRCAG02]. pups [MWS22]. Purging [Bro05]. Puruvesi [MSV^{+}23]. putitora [BP16]. Putting [Pau97, PZC17]. pygmy [Bea07].

quality [BCO21, BTS97, Neh96, PTOS^{+}24, SQR09]. Quantifying [BMN21, Bro94]. Quantitative [Ano92v, DP12, LCM22]. question [CRD00]. questions [HJC20, JN18, LF03]. quota [BHP15, Neh96, WP96]. quotas [Dia04, Gra96].

R [Coc05, dC98]. race [CBHOH19]. race-to-fish [CBHOH19]. radiation [AA20, LG97]. Rafinesque [IJ01]. Rainbow [TRG19, CCMS12, CTM92]. ranching [BC03, GJA17, LSL03, LZC22, MA03, Mus03, MSR03]. random [AHW04, Duc19]. range [GH16, KCF14, MBP14, PV19, SEA22, WWS17, YKS14]. rangers [DDD16]. ranges [LMCB^{+}23]. ranging [TBA20]. RAPD [AHW04]. rate [GH17, WH24]. rates [BSR19, Bro94, Cad91, KC14]. ratio [CCA17, SW95b]. Rationale [Avi00]. Ray [Mil99, AUdS18, BdST16, PLA07, SHS14, WT08]. Raymond [SHH95]. rays [BSR19, BCK18, CGP^{+}24, PKC^{+}24, VD22]. rDNA [MPO11, RZV12]. Re [CBH15, PQS14, Gou16]. Re-colonization [PQS14]. Re-estimation [CBH15]. reaction [AHL12]. reactions [RFH15]. Reading [RSGS^{+}23a, RSGS^{+}23b]. real [WMA20]. real-world [WMA20]. realities [CWB16, SAC^{+}23]. reality [Chr96b]. really [FN02]. reappraisal [MdAVA11]. rearing [Gib93, Jue95, SAK14]. rearrangements [BVMF13]. Reason [McD03]. reassessment [Tel09]. rebuilding [Mil12]. received [Ano93c, Ano95c, Ano96d, Ano96e, Ano98c, Ano00b]. recognized [Gau01]. recommendations [BTW15, OCE11, SVC21]. Reconstructing [EG03]. reconstruction [SG24]. records [PJ14]. Recovery [WMG00, AGJ14, ABK14, CA04, Jen00, LSH15, MTL21, PDA12, Ros00, SW14, TAD14, YC20]. Recreational [Har99b, SEA22, AAJ21, BPA^{+}23, CCH18, CVD21, GCUR22, GZT13, HJC20, JLC21, MML18, SAB^{+}23, Har99b]. recreational-only [GCUR22]. Recruitment [Nee01, AGO04, Ano92j, Ano94o, HWA14, MIA^{+}23b, MIA^{+}23a, Mye98, VSM12, WK99, vPKW18]. Red [CGR11, DD18, MMB22, EBS20, Fit93, OSC20, MA08]. redistribution [MTAP22]. reduce [Bro00a, HB19, MBS17]. reduces [SZG10]. reducing [WSGP22]. reduction [HCB11, KB21]. redundancy [BCM20]. reed [Whi17]. Reef [CFB14, SKD03, AGJ14, Ara15, BCM20, CWP14, CBH15, CBHOH19, ELC^{+}24, LET^{+}23, MGW93, PBF15, PFW16, PFW17, RP91, STP22, STP22, SBP07]. reefs [Ano92h, HCK21, MSR14, PDN^{+}24, TSB22, TBB^{+}24]. reestablishment [CCMS12]. Reexamination [SW12]. reference [Ano93g, HHCM10, Mar93, Pau10, WH07]. references [DP12]. reflections [CNT21, LMCB^{+}23, SK15]. reform [PAT15]. reforms [MH20]. regarding [BGM^{+}24]. regenerating [BEBFM14]. Regime [Jia09, AGO04]. regimes [ACST17]. region [BGOW23, BLF10, KWI07, LBS^{+}23, NWO16]. regional [CCP13, CNT21, DBT15, GCB^{+}24, GSC17, SBZR17]. regionally [SBtIWG12]. regions [HCvP16, JCL07, SRBS21]. regression [MGB15]. regulate [MGL^{+}23]. regulating [AAJ21]. Regulation [HL98b, IJ01, AGO04, DW11, EB93, GOW13, MVM99, Nor09, PY97]. regulations [ST94]. reintroduction [VCD12]. related [CG00, CCO20, DW09, FC07, HCB11, KL23, Koc01, MTPR15, MH20, SEA22, YCT11]. relation [Ano92x, Jue95, dSPPM12]. relations [ENF^{+}23]. Relationship [WHS04, DBT15]. Relationships [Kam05, KVH98, LG97, Bag11, MBdBC13, May02]. relative [RVRCB11, Vil18]. relativity [CDH^{+}23b]. Release [BB05, MG18]. released [WH24]. relevance [WA03]. reliable [Sch18]. remaining [MLK13]. Remobilizing [MP16]. Remote [Ano94p]. removal [PQS14, QGM15]. reorganization [CZF22]. repeating [AMV13]. repetitive [MBdBC13, PASF13, SGF13]. replacing [GCS19]. Reply [BL02, DCS12, TBK11]. repopulation [MAC02]. report [Ano96f, Ano97e, Ano98d, BJ00, CM97, GCSR09]. reporting [ENP18, MMJ22, TTC11]. representation [dlCFP19]. Reproduction [HHCN11, MLC17, vGM05, AGG03, AGO04, ARL12, AHL12, COGFPV23, GELL16, Kim95, MZ00, WGR20, dJFS20]. Reproductive [Ano94a, CB00, Fle96, MS13, RMN11, ZFT13, CE96, FCCS15, JR97, Kim93, LKK10, PBS14, dSSHK21, TFF09]. reprogramming [MGL^{+}23]. required [CJV13]. requirements [WP96]. Requires [De 98]. Research [Ano94h, BP08, Hus04, SGD^{+}23, STW^{+}24, ZRZ20, AJM22, AFO22, AHW04, Ano92t, BTW15, CCH18, CWN11, CCO20, CDS16, DROCM23, DT04, DDD16, Duc19, FHvP14, GMS17, GS23, HJC20, IRPG21, JZ00, KM20, LSL03, LBS19, MC16, MHP12, MR18, Mis97, MDV^{+}23, MC12, Moo12, NRJ^{+}23, OSB^{+}23, PHH10, RSGS^{+}23a, RSGS^{+}23b, SAN^{+}23, TKB18, VBC^{+}23, YCT11, YMR12]. researchers [BLC19, NRJ^{+}23]. reserve [GLG12]. Reserves [BB05, Dri05, GLC98, AFB15, Jen00, MVGW18, RP91, Ros00, Sto00]. Reservoir [AKGB11, PR05, PMVA19]. reservoirs [dPAGGB16, Kel23]. resident [AF04, PBS14]. residents [OV24]. resilience [OSB^{+}23, SRB^{+}23]. resilient [BPA^{+}23, TFP22]. resistance [YWS^{+}24]. resolve [CH99]. resolving [SPS07]. Resource [Kam08, SN00, AH96, Ano92u, BNC22, BLC19, CCH18, CCO20, CBH15, MP16, SJH22]. resource-sector [SJH22]. resources [Ano94n, ANP^{+}23, AAH98, CA04, Cus94, GCUR22, Lig16, SX16, SE16, VRR12, LFJ08]. respiration [TRB13]. respiratory [Nor14]. Response [DCS12, AGT08, EDP10, LS07a, PDB16, RPA13, SW98, WWS17]. responses [ADC15, BPA^{+}23, LCC19, LR11, OSB^{+}23, TRB13, Els04]. Responsible [GJA17, CNT21]. restocking [WMDS18]. Restoration [Har99a, AGT08, GJA17, MLM19, OGR17, Rey93]. restorative [Lin94]. restore [GLC16]. restored [Mag13]. Restoring [RBC16a, RBC16b]. Restricted [PDM20, MLC17]. retained [GAD10]. retention [CML19, GMS17]. Reticulated [RMN11]. retina [BEBFM14]. Retracted [DD13a, SDJ13]. Retraction [DD13b]. Retrospect [CS05]. Retrospective [CS04, ELC^{+}24]. return [WHS04]. reveal [Duc19, MPO11, PJ14]. revealed [AVB^{+}23]. Revealing [IOHM23]. reveals [MWP^{+}23]. Review [Ano92b, Ano93b, Ano94a, ALJ08, Bar94, BB05, Bil02, Bla98, Bro00b, Bur07, Coc05, Csi99, Dri05, FQSJ23, Fos08, FD00, Gil93a, GH17, Gla00, GCO03, Har99a, Har98, Har99b, Har99c, Hil99, How94, JBJ06, KCR17, KFS21, LAI05, Lor93, LFMP21, LFJ08, MBJ12, Mil99, Mul09, Mul10, MAJ05, NWO16, NMF17, Nor02, Pav05, Paw99, PBG04, Pis00, PGJ15, Pyk05, Qui12, Ram08, Rei99, RHP99, SRGS04, Sei08, Soi99, Tur99a, VAJ10, Web09, Wil07, WBW05, WBW09, Win06, Woo98a, Woo99b, dC98, vGM05, AYN11, AHW04, All11, Ano97a, BHK00, BJM20, BP16, BGBE^{+}23, BTW15, Bro00a, BSKB22, BP08, BOD21, BHP11, CLH21, CWB11, CGK11, CAB19, CCH18, CG00, CGCG18, CGP^{+}24, CWN11, CTB16, CRPI^{+}22, CJV13, DSU08, DW09, DBT15, DBR15, DIR20, DD18, DD13a]. review [DD13b, Els04, EBS20, FGL10, FIØ04, FSB14, GDC13, GBR22, GCSR09, GKC19, GLM20, GH16, HCB11, Har24, HWA14, HPdL02, HPY18, Hun21, Hus04, IPT10, IFW05, IBJ04, JCL07, Jia09, JBS02, Jua10, KC14, KB21, KCF14, Koc01, KK12, KMF13, KZ20, Kul02, LE24, LMCB^{+}23, LKH16, LC02, Lis10, LBS^{+}23, LR11, LB23, MMB13, MFV19, MDF14, MKL^{+}23, MP07a, MBM06, MBK12, MDR14, Nis00, Nor09, OK10, OCE11, PBF15, PAK^{+}23, PM09, PHH10, PHK20, PFW16, PFW17, PKC^{+}24, RK10, ROP^{+}23, RM01, RTT12, SMS12, SCC09, SA12, SAB^{+}23, SHS14, dSSHK21, SL15, SPB00, SQR09, Sny05, SDG15, SDS15, Ste05, Sul04, TKB18, TRB13, VCZ19, VLD20, WHU02, WR16, WTC06, WPD12, WMW18, WW06, WDP19, WPF16, YCT11, YOC15, ZZ04, Zim05]. reviewed [Ano92p, BSV11]. Reviewing [NRCD^{+}23, PDDDE21]. Reviews [Ano95j, Ano97b, Ano98b, Ano96b, Ano96c, Ano97c, Ano97d, Ano98a, BFD91, Han96, HC04, HES95, JCR95, SHM96, SMC95]. revised [Ano93b, HRC09]. revision [VE05]. revisited [McD97, QM04]. Revisiting [PAW17, vPKW18, MR18, WK99, YH18]. revolution [CBF^{+}24]. reynaudii [MS07, OSB07]. Rhamphichthyoidea [CPN11]. rheophilic [AMVV13]. rhino [PKC^{+}24]. Rhizoprionodon [MOF11]. Rhode [BP08]. rhombus [DMK^{+}24]. rhythms [AS08, AVB^{+}23, Ree02]. Rica [WBW05]. Ricefield [Shi05]. rich [JB95, SBtIWG12]. Rico [WBW09]. Rigau [WBW09]. Right [AH23]. Rights [Arn05, Asw05, FB05, Han05, HPL05, Hup05]. Rights-based [Asw05, FB05]. Rineloricaria [RZV12]. Rio [CBEGR02, Mag13, EGMM02, EGMM02, LLSDTJ09, TJLLC10]. riparian [CG00]. risk [Ano93g, CBAVGR02, EEDP18, Gau01, GKC19, MMB22, TRJ24]. risks [BO12, Kos09]. River [LSP11, LQW19, MJSOF16, PBM14, SGD^{+}23, TFE07, VRRCAG02, ZCW19, ZDJ19, BLF10, BSKB22, CBB19, FSSM17, KFS21, Mag13, MKS12, PBMF12, REA^{+}23, SMR11, TØH08, AGO04, BO12, CLX15, MBA15, PBM14, SPL12, WGL17, WH24]. river-floodplain [Mag13]. river-floodplains [CBB19]. Riverine [SAC^{+}23, BTZ^{+}23, PDB16]. rivers [CCA^{+}23, Cha95, GB23, LKH16, LWS17, MV13, MLM19, MSLN02, PQS14, PHK20]. RNAs [DZZZ22]. roach [TV15]. Robbing [GCS19]. Robert [GCO03, HNSS02]. Robertsonian [BVMF13]. Robin [Sei08]. rock [HFP14]. rockfish [BM09]. rocky [AS95, AF04]. Roffe [WBW09]. Role [EN00, BVMF13, BAA18, Bro94, CBO19, Cor02, CHN18, Fer94b, GP12, HFM19, HL11, KCE12, KC92, Mar95, NVH21, RT91, SGF13, TBB^{+}24, Whi17, dlCFP19]. Roles [WZB^{+}24, FCOCNC24, TSB22]. Rosemary [Pis00]. Ross [Mul10]. roughy [Jon07a]. Rounsefell [QM04]. Ruffe [GH16]. Rush [HNSS02]. Russia [ZZ04]. Russian [DD18]. ruthenus [WOR09]. Rutilus [ADS11, TV15].

S [Bla98, MBJ12, Wil07, Win06, Zim05, dC98]. S. [IFW05, SRGS04]. safe [McD09b, MV96, YH18]. Safeguarding [WMTL22, BH17]. sagax [BBJ12, IWG17]. salar [BGTA19, GBR22, Man94, MAP21, SAK14]. Saline [APLL07]. salinities [WTC06]. salinity [DW09, Nor14]. Salmo [BGTA19, Est05, GBR22, Man94, MAP21, SAK14]. salmoides [PV19]. Salmon [BB14, Ano92n, Ano94f, ABK14, BD20, BCK16, CM08, Fle96, Gib93, GBR22, HK14, JNW14, Jua10, Jue95, Kos09, KWI07, Man94, MHP12, MT14, MAP21, OK14, OWW04, QM04, Qui12, RN04, SAK14, SBN04, SW14, TØH08, TAD14, WGL14, WHS04, ZZ04, Jua10]. Salmonid [WHU02, CTL17, Col10, EAW22, McD02, Nis00, QGM15, RHM11, Utt00, VSM12]. Salmonidae [CRRCdL08, LSH15, MAP21, Nor95, RCCRD03]. Salmonids [RN04, ACST17, APLM14, ANB19, BGBE^{+}23, BB96, Col10, JN18, KC14, KFS21, MLM19, McD06, MC04, OCE11, Utt04, VCD12, VLD20, ZLF21]. Salmoninae [Est05, SDS15]. salmonines [SFO14b, SFO14a]. salt [AGT08, Nor03, Nor06, Whi17]. Salvelinus [Est05, MBK12]. same [KL23]. Sampled [MMB22]. sampling [AF04, Ano92c, GZT13, HWA14, HCNH24, MWP^{+}23]. Sander [SW12]. Santo [WBW05]. sarcomeric [Joy23]. Sardina [BOs12, CHFTV22]. sardine [BBS10, CHFTV22, IWG17]. sardines [BBJ12, Hun21]. Sardinops [BBJ12, Gau01, IWG17]. Sarkar [RPA13]. Satellite [CM98]. satisfaction [AAJ21]. sawfish [WSC09]. sawsharks [NWG21]. scabbardfish [BMF09]. scabra [Gou16]. scabripinnis [CMA15, CCA17]. scale [BGOW23, CCA^{+}23, CMF^{+}24, DPV^{+}23, DCS20, ENF^{+}23, EBRLGB02, HFM19, HKTS^{+}23, MTM22, MCD09a, MLSP^{+}23, MMJ22, PRP16, ROW^{+}23, SK23, TKB18, VSM12, YOC15, ZDJ19, ZHP07]. scales [Ano94o]. scallop [TMKC24]. scalloped [GK18a]. scan [HJC20]. Scenarios [BDS05, FCOCNC24]. schistosomiasis [SMM94]. schools [DP12]. Schrank [LF03]. Schwerdtner [KCR17]. Sciaenid [SVM^{+}24a, SVM^{+}24b]. Sciaenidae [VdSD21]. Science [Coc05, Dri05, GCO03, How93, NAP22a, NAP22b, Smi98, Ter01, AFO22, Ano95h, CCO20, CDH^{+}23b, CGR11, HCB15, JBJ06, MBD07, RR11, Ros97, Ter11, TRB13]. science- [CGR11]. Sciences [Paw99, Zim05, Ano93e]. Scientific [MN98, CCP13, Gil93a, LBS19, OAJ14, PWB99, SX16]. scientists [Ano95h]. Scientometric [AMV13, SAN^{+}23, dMV20]. Sclerorhynchidae [WSC09]. scope [HKN18, PBS14, SBtIWG12]. Scophthalmus [IJ01]. scoping [PAK^{+}23]. Scottish [FD22]. Screening [VCZ19]. sculpins [WIT14]. Sea [Asw05, CCR21a, CCR21b, CCR^{+}24, CUT07, RMN11, STS07, WBW09, AA10, GH17, Gou16, GJA17, HMQ16, IPT10, JBJ06, LFM13, LSL03, LZC22, MTM22, MSM20, MA03, Mus03, MSR03, SBG17, TRB13, ADS11, AA10, Ara15, AE22, BCA^{+}23, Bar93, BKB17, DD15, GMP12, KZ07, KCE12, Mad07, MA08, MMG18, NCMT23, OK10, PVL21, PRP16, Ram08, RSSS23, SRGS04, SBP07, TCS20, VaR92, YMR12, Soi99]. seabass [YWS^{+}24]. seabird [GCK21]. seabream [SZC11]. Seafood [HCMT24b, CGL22, FAN22, HETS23, HCvP16, LSDRS23, SRB^{+}23, HCMT24a]. seafood-dependent [HCvP16]. seagrass [Whi17]. seahorse [SAN^{+}23]. seahorses [KBV11]. seals [Pie92]. search [Jon02, WPF16]. Seas [MMM22a, MMM22b, PAN22, ANP^{+}23, DLP95, KLC18, WSGP22, CCP13]. Seasonal [EDP10, NKF21, GBCCO12, HCMT24a, HCMT24b, LLSDTJ09]. seasons [BHS19]. seawater [How93]. Sebastes [BM09]. Second [ALJ08, Sul04, Bar94, How94]. Secor [BOJ16]. Secretion [HL98b]. Sectioned [WDP19]. sector [CVD21, SJH22]. secure [FAN22]. security [NMS13]. sedation [BTW15]. Sedative [BTW15, Mul10]. seed [Tan03]. Seeing [Fit93]. seining [MSV^{+}23]. seismic [HR00]. Selected [DC05, HW24]. selection [Ano91b, Ano91c, MF99, dSKV16]. Selectivity [Mad07, GCM18, GCS20, LH18]. self [dSKV16]. self-selection [dSKV16]. Semen [LLC11]. semi [KGF10]. semi-intensive [KGF10]. séminaires [Woo99a]. seminal [ARL12]. senegalensis [IFW05]. sense [Kot91]. sensitive [AH23]. sensitivity [LSB22]. sensory [Lis10, LB23]. sentiments [SEA22]. sentinels [HP14]. Sepia [AVL07, HFG07]. Sepioteuthis [HP07, MP07b, SM07]. September [KGW10, KGW11]. septicemia [EEDP18]. sequence [AVL07, MPO11, RZV12]. sequences [MBdBC13, SGF13]. sequencing [LW17, RNJ16]. sequential [GC23]. sequentially [Fri04]. sequentially-hermaphroditic [Fri04]. Serbia [LMM11]. Serbian [DPC12]. serial [AAH98]. Series [Woo99b, Csi99]. Serrapinnus [MPO11]. serum [Nor09]. service [HCB15]. SES [CSS20]. Setting [YOC15]. sevengill [SRBS21]. several [HJC09]. Sex [DJ05, Fri04, HL98b, Shi05, Ano95h, BVMF13, CPN11, CCA17, CLB^{+}22, MGB15, MdAVA11, MGL^{+}23, PDDDE21, SW95b, RHP99]. sex-biased [PDDDE21]. Sex-change [Fri04]. sexes [EDP10]. Sexual [SN10, EAW22, JJ93]. sexually [PM14]. SFF [NMS13]. SFFs [NMS13]. shad [CPM14, DML^{+}24, MBA15]. shallow [JJC22, ZMG17]. shame [McD06]. shape [SMNK^{+}23, SBtIWG12]. Shared [BDS05]. Sharing [BBS17, GB23]. Shark [MCL08, MML18, MDV^{+}23, PDM20, BKB17, BdST16, Com93, DHG18, GK18a, HMV17, KWI07, MOF11, MWS22, MC12, MMF18, MSH14, MG18, SHS14, SKW22, YC20]. shark-diving [HMV17]. shark-eat-shark [MWS22]. Sharks [FTH15, BCM20, BSR19, Bre93, BCK18, CFB14, CGP^{+}24, Duc19, LPG17, MWS22, OV24, PJ14, SRBS21, VD22]. sharpnose [MOF11]. Sheffield [Paw99]. shelf [BSWA14, HMN17]. Shelley [WBW09]. shellfish [Ano92b, AFB15, DD15, Vil03]. Shellfisheries [DC05, CD01]. shelves [MMG18]. shift [Jia09, JJ93, SEA22]. shifting [PE08]. shifts [BP08, HSC^{+}24, JPC14, SHPH^{+}22, SJB^{+}23, WWS17, ZL23]. shiner [May02]. shiners [MS02]. Shoal [TP14, RFH15, STP22]. Shoals [DP12]. shock [DW11, ITV98]. shocks [OSB^{+}23]. Short [MSR14, FHvP14, KdSD18, MKK10, PWC06, SW10, SHB21]. Short-term [MSR14, KdSD18, MKK10, PWC06, SW10, SHB21]. Shortcomings [LSDH12]. shortfin [CUT07]. Shoulders [HL98a]. shovelnose [WT08]. show [PDN^{+}24]. shrimp [AAH98, Dia04]. Sicydiinae [WIT14]. sides [LPC^{+}24]. sign [MT18]. signal [GCSR09, MCN12]. signals [Smi92]. signatures [Avi00, CGL22, RBPPS16]. Significance [PGC19, Joy23]. significant [Duc19, GP12]. silky [DHG18, MG18]. Siluriformes [PASF13, BVMF13, BVM13, GdMD13, MdAVA11, RZV12, dBdSA12]. Silurus [AKGB11]. silver [QSBV18]. silverfish [BOV09a, BOV09b]. similar [KL23]. similarities [MKS12]. Simmonds [Bur07]. Simulating [EP00]. simulation [HFP14, JEL10]. Simultaneous [Smi98]. since [CBH15]. sink [McD10]. Sissenwine [Har98]. six [Utt04]. Size [HSC^{+}24, PR05, TS14, BSL18, BGOW23, CBHOH19, FN02, GCM18, MGB15, Mar93, MF99, SM07, WHS04, WMA20, YKS14, ZLF21]. size-at-age [WMA20]. Size-dependent [HSC^{+}24]. size-selection [MF99]. size-selectivity [GCM18]. sized [Ano94i]. sizes [CML19]. skates [BSL18]. Skiftesvik [Hus04]. skin [Col10]. slaughtering [GGZ10]. slope [BMF09]. slow [MRMJJ17]. Small [BGOW23, NMS13, ASS^{+}23, CCA^{+}23, CMF^{+}24, DPV^{+}23, DCS20, ENF^{+}23, EBRLGB02, HKTS^{+}23, JLC21, LFMP21, Mar93, MTM22, MCD09a, MLSP^{+}23, MMJ22, MIA^{+}23b, MIA^{+}23a, Pat92, PRP16, ROW^{+}23, SK23, TKB18]. Small-scale [BGOW23, CCA^{+}23, CMF^{+}24, DPV^{+}23, DCS20, ENF^{+}23, EBRLGB02, HKTS^{+}23, MTM22, MCD09a, MLSP^{+}23, MMJ22, PRP16, ROW^{+}23, SK23, TKB18]. Smartphone [Cal22]. Smith [Rei99, Kva98]. smooth [GK18a]. Smyk [SW12]. snail [SMM94]. snake [Lin94]. snapper [CGR11, EBS20]. Snappers [BDS05, BGOW23]. snapshot [BGHC23]. snout [BSL18]. snow [MDR14, MMG18]. Sobel [Dri05]. Social [Har99b, CSS20, SEA22]. socially [AFB15]. socially-compatible [AFB15]. Society [Soi99, NvPV22]. Socio [BDS05, LBS19, DBR15, MCD09a, SRB^{+}23, TCS20, TBA20]. socio-ecological [MCD09a, SRB^{+}23]. Socio-economic [LBS19, DBR15, TCS20]. Socio-Economics [BDS05]. socio-ecosystems [TBA20]. sockeye [MHP12]. solandri [ZFT13]. solar [OK12]. sole [HHCM10, HHCN11]. Solea [IFW05]. Solitary [Kot91]. solution [BBJ12]. solutions [HAC18]. solved [CHFTV22]. Some [AY10, GIT09, Kos09, Utt04]. soniferous [LCM22]. Sound [PDA12, WHS04, PHK20, RL05, Bag11]. source [McD10]. sources [BGHC23, WZB^{+}24, Whi17]. South [Ara15, HFG07, MTL21, NWG07, OSB07, SE16, ADS11, CJV13, FHvP14, LBS^{+}23, AH23, AVB09, BAA18, CDC09, CBR98, DCS20, JCL07, MGTD18, MS07, MBP14, QSBV18, REA^{+}23, Whi99]. south-east [CJV13, FHvP14]. south-west [LBS^{+}23]. southeast [BSWA14, ENF^{+}23]. southeastern [LLSDTJ09, OWW04]. Southern [Rei99, APLM14, AVB09, BDF09, CCR^{+}24, HP07, HLD07, MC16, McD06, NWG21, PGJ15, BOV09a, BOV09b, CUT07, JBX07, Koc01, MBG21, Par92, PNC11, Tur95, WMW18]. southwestern [TKB18]. sp [Bol07]. space [SHB21]. Spain [BGHC23]. span [LE24]. Spanish [Ano92b]. sparid [WDP19]. Sparidae [MDF14]. sparse [PB98]. Sparus [SZC11]. Spatial [CGKSP13, GCUR22, LLSDTJ09, OSB07, RCP19, VSM12, ZL23, dPAGGB16, CRZHZ02, CJV13, GHF98, GPS18, MWP^{+}23, MFL21, TR94]. Spatiotemporal [LLJC24, NRCD^{+}23]. spatula [MAC02]. Spawning [Est05, MP07b, TAS10, Ara14, APLL07, CLH21, FD22, Gib93, HHCN11, HP07, HTH00, Kat97, MS02, MS07, MLC17, MBK12, PM14, TØH08, WW06, vOR15]. Speak [AUdS18]. spearfishing [SEA22, SAB^{+}23]. Special [Ano99b, Han96, HHCM10, MMM22a, MMM22b, MSV^{+}23, Pau10]. specialisation [MSH14, SHHK21]. specialised [GZT13]. specialists [SHB21]. Speciation [WM02, Pau10]. Species [Ano99b, EN00, MSR03, Pau10, RT05, Tay99, dP99, ABS15, ACST17, BCA^{+}23, BBS17, BLF10, BVMF13, Bol07, BHS19, BOV09a, BOV09b, BCK18, CPN11, CDC09, CH99, CMA15, CSY12, CPM14, Chr96a, CCA17, DD15, EAW22, Fro99, Gil93b, GCM18, GCS20, GK18b, GCUR22, HMQ16, HJC09, HMN17, HCMT24a, HCMT24b, IE06, KdSD18, KDF13, KSK17, Koc01, KZ20, Kul99, LY07, LMM11, LFdSRM16, LSH15, MGB15, MCFC21, MGRRJ21, MdAVA11, May02, MTAP22, MCN12, MBP14, MRMJJ17, NMS13, Nel99, NVA12, PASF13, Pat92, PDN^{+}24, PCG19, PBC12, QSBV18, RK10, RCP19, ROW^{+}23, RBK13, SCC08, SEA22, SOB20, TFE07, TP14, TFF09, Tel09, Tur99b, VdSD21, WGL17, WPD12, WGC18, WMG00, WT08, XSC15, YVR07, dSPPM12]. species- [GCM18]. species-specific [ACST17, BBS17]. specific [ABS15, ACST17, BRI00, BBS17, BB14, CTL17, RBPPS16]. specimens [JLD^{+}24]. speed [VW91]. Spencer [HFG07]. Sperm [ARL12, AHL12, Bea07, CZF22, FÁVG18, HII16]. spermatozoa [Ano92k, PFL10]. Sphyrna [PJ14]. Spiny [MP07a, PM09, DBR15, PMV18]. spiritual [MP16]. Spirulina [RMK24]. sponge [JLD^{+}24]. Spontaneous [SVC21]. sport [Ber93, BO12]. spot [JLC21]. spotted [FIØ04, PLA07]. spp [BM09, BLD19, CG00, Jel22a, Jel22b, KFS21, PBF15, PJ14, PBR19, PCW17, QGM15, SSBCL11]. sprat [OK10]. sprattus [OK10, OK10]. spread [KZ20]. spring [HMN17]. squid [ANP^{+}23, CUT07, DMK^{+}24, FÁVG18, GJ07, Jon07a, LS07a, LS07b, MS07, MP07b, OSB07, PJ08, SKU20, SM07, VMB07]. squids [FHW^{+}24, HII16]. St [WTC06]. Stable [SBG17, CGL22, MTC19, SHPH^{+}22]. Stage [Soi99, YOC15]. stages [AA20, BOV09a, BOV09b, LKH16, SRD93, GCO03]. stalked [PTOS^{+}24]. standard [GKR10, GMS17]. standardize [BEC11]. Standardizing [Bis06]. Standing [HL98a]. starvation [MKK10]. stasis [NLFM12]. State [Ste05, BSM17, CCO20, FCP19, LF03, MHP12, MBD07, SAN^{+}23]. States [Gil93a, ASS^{+}23, Gil93b, HMN17, LM19]. stationary [BBS10]. statistics [KLC18]. Statolith [DMD21]. statoliths [VMB07, ZHP07]. Status [GS23, KWM^{+}22, LSL03, SK15, BDF09, BKB17, CHFTV22, CDH^{+}23a, CGCG18, DD18, EB93, EGMM02, GK18a, Har11, HBC13, HPdL02, KB14, LSP11, LBS19, MC12, OGR17, OJB07, PTP14, PZC17, PBR19, PBC12, SSBCL11, SBZR17, SGD^{+}23, TKB18, VRRCAG02, YC20, YWS^{+}24]. Steatogenini [CPN11]. Steatogenys [CPN11]. steelhead [HK14, Kos09, PM14]. Steenstrup [HLD07]. Steindachneridion [MdAVA11]. Stel [Bla98]. stenolepis [LR11]. Stephanolepis [RMN11]. steps [CNT21, NRJ^{+}23]. sterility [HKL10]. sterlet [WOR09]. Steroids [HL98b, Fri04, Kim93]. Steven [MAJ05]. stewardship [MP16]. stimulation [Har94]. stingrays [LSB22]. Stock [GLG12, Gla00, Har99c, LCC16, MLN05a, MLN05b, Mus03, WP96, ADS11, AMVV13, Ano92j, Ano92s, Ano92v, Ara14, AS96, BBJ12, BM09, CH94, CDH^{+}23b, DGV11, DD13a, DD13b, GJA17, GIT09, IWG17, LSL03, LS07b, LFMP21, MGW93, OK10, OWW04, PZC17, PBG04, PH97, SG24, TCS20, TdSL15, WM96]. stocked [PBB00]. Stocking [PR05, Hil99]. Stocks [BDS05, Pit98a, BDF09, Cad00, CBR98, DSU08, DLP95, GBR22, OK12, dSSHK21]. Stomach [IRPG21]. stop [KZ20]. storage [SMN08]. stories [CDS16]. Story [Pit98a, Jon07b, Rei93]. straddling [ANP^{+}23, DLP95]. strain [PDGBR07]. Strait [SBN04]. stranded [Bea07]. stranding [GLHD^{+}24a, GLHD^{+}24b]. Strategies [Ano93f, Ano92c, AWC17, BB14, DPV^{+}23, Fle96, FC07, GLC16, Jon02, JJMD13, KK12, Kos09, WNB12]. strategy [Ano95f, CCP13, MGRRJ21, McD09b, SAC^{+}23]. straying [KC14]. stream [Ano93e, CTL17, CG00, VSM12, VCD12, YCT11]. stream-dwelling [VSM12, VCD12]. stream-specific [CTL17]. streams [MSLN02, NKF21]. Stress [PWC06, All11, BSR19, CLB^{+}22, DW09, GVB94, KdSD18, ODS10, VdSD21, WLB^{+}23]. Stress-associated [PWC06]. stress-induced [WLB^{+}23]. stress-mediated [CLB^{+}22]. stressor [CBL17]. stressors [GBR22, SGD^{+}23, WGR20]. stricken [LCC19]. stride [VW91]. striped [APLL07, CCR^{+}24]. structural [RZV12]. structure [AMVC20, BBJ12, CHFTV22, DGV11, DHG18, FSSM17, HP07, IWG17, MDF14, MOF11, NWG21, PDN^{+}24, PZC17, PGC19, ROW^{+}23, SRGS04, SGF13, VGA11, WDP19, ZHP07]. structured [ETWE12]. structures [CBHOH19, Vil18]. Structuring [WCP97, SRBS21]. struggling [Ter18]. studies [ANB19, BLF10, BLC19, CWN11, CNT21, DT04, HF20, KCF14, MAC02, MHvH16, MSR03, PNC11, PGC19, Sch18, SPB00, SLH11, TTC11, UIA11]. study [ABS15, AS95, Ano94d, ASS^{+}23, BCK18, CCA17, ENF^{+}23, EBRLGB02, GCSR09, PBB00, Whi99, WM02, ZLH^{+}23]. Studying [JN18, MBM06, dSKV16]. sturgeon [ARL12, AHL12, BL00, CCO20, JEL10, PVJ07]. sturgeons [WH24]. Styles [Smi91]. subadult [WH24]. subantarctic [TBA20]. Subject [Ano92w]. submerged [STP22]. subsea [BMN21]. subsequent [CG00]. subspecific [Utt00]. substantiated [MFL21]. substrate [SM07]. subterranean [Wil07, Wil07]. subtle [MT18]. subtropical [BAA18]. success [AGT08, Ann96, APLM14, HR00, KKM10, LM19, Ter18]. Successes [DC05, COGFPV23, JLC21]. successful [CSS20, Jel22a, Jel22b, PM09]. such [Jel22a, Jel22b]. Suez [MA08]. suggest [BB14]. suggested [CA04]. suggests [CBH15]. suite [GMS17]. suits [ETWE12]. summary [CWD11, OAJ14]. summer [HMN17, Ter01, Ter11, Ter18]. Sun [RPD16]. Sundarbans [MMB13]. sunfish [PHH10]. superciliosus [FCCS15, MMF18]. Superorder [VVU22]. Supplementation [GDC13]. supplements [ROL14]. supply [SRB^{+}23]. Support [SVM^{+}24a, SVM^{+}24b, CDH^{+}23b, FCP19, GHMG22, GC23, OSB^{+}23, TFP22]. supporting [ABS15, KEP22]. surface [PP92, SW12]. surgeonfishes [TMD^{+}24, TSB22]. Surgical [WCD11, BEC11, CMC11, CWN11, CWD11, HL11, OCE11, TTC11]. Surplus [CRPI^{+}22]. survey [Ano95a]. surveying [GZT13, HR00]. Surveys [Ano94n, Ano94b, HCNH24]. survival [GCM18, GCSR09, MHP12, MJM15, TAD14, YH18]. Sustainability [BCK18, CS05, APP^{+}23, BBW09, BH17, CCH18, MLSP^{+}23, PAN22, PAK^{+}23, SRB^{+}23]. Sustainable [HFM19, Lac02, NAP22a, NAP22b, BPA^{+}23, CFS^{+}23, FAN22, KWM^{+}22, Lig16, LCC16, NBF22, PVL21, REA^{+}23, WMW18, vOR15]. swamp [OSC20]. Swim [McD03, dMlTB19]. Swimming [DIR20, VD22, Ano95g, VW91, WA03]. swordfish [CML19]. symbiosis [NVH21, Pou93]. Sympatric [PDN^{+}24, MCN12, PASF13]. Symposium [Ano92x, Ano93f, Ano94a, Ano94h, Bar95, BP08, Har98, Har99b, JBS02, Jua10, Nis00, Soi99, Ano94i, Sol93]. Syndrome [Soi99, KB14, IPT10]. Syngnathidae [MFV19]. Synopsis [RT05, SKD03]. Synthesis [MLM19, CMF^{+}24, CBH15, GCS20, KMK16, MC16, MHP12, MAP21, MLP17, PVJ07, SL15, TP14, VEK10, WRSG21]. system [Ann96, Arn96, BCO21, BVMF13, CRZHZ02, CC22, CSS20, FSSM17, HHCM10, MdAVA11, Mun95, RBC16a, RBC16b, WTC06]. Systematic [VLD20, HPdL02, LB23, WJP23, WPF16]. Systematics [Ano92e, Bol07, Gre93b, Ano92k, Web02]. Systems [Bla98, CM98, Lac02, BC03, CHN18, CE96, FAN22, GIT09, HYW13, KFS21, MCD09a, MBH14, MZ00, PDB16, SP98, SAC^{+}23, TFP22, WP96, WT08, Woo99a].

T [Lac02]. tables [WGL17]. Tacher [WBW09]. Tacher-Roffe [WBW09]. tactics [Bas93]. tag [GMS17, LR11]. tagging [CWN11, DML^{+}24, GMS17, TTC11, VLD20]. tags [BEC11, CMC11, CWN11, CWD11, HL11, LR11, Mul11, OCE11, SMN08, WCD11]. Tagus [SVM^{+}24a, SVM^{+}24b]. take [BB05, GJ07, Jon07b]. Taking [TCS20]. Tana [SN00]. Tanganyika [Ano92q]. tangled [PKC^{+}24]. tannin [KKM10]. Tanzania [IOHM23]. Taranto [CCR21a, CCR21b]. target [BCA^{+}23, Koc01]. targeted [CJV13, RBPPS16]. targeting [BMN21]. targets [Pat92]. Tarpon [CGRCM19]. task [WMA20]. Tasmania [BSWA14]. Tasmanian [HFP14]. taste [KL23, Kot91]. tax [Mar06]. taxa [EP00, GCS19, Smi93]. Taxonomic [VE05, Pau10, PAT15, SSBCL11, ZDJ19]. taxonomy [Tay99]. Technical [HB19, AA10, GBCCO12]. technique [CMC11, MKK10]. techniques [Ano92l, CH99, DOP18, GVB94, HWA14, LNT^{+}23, LR11, MGB15, MBM06, Mul10, NCMT23, PM94, WCD11, WG94, WMDS18, WDP19]. Technological [CVD21]. technology [Bis06, CR08, How93, MT14, TMKC24]. telemetry [AVB^{+}23, Ano94p, BLC19, BOD21, GMS17, KCF14, KM20, LMH21, MBC21, MKL^{+}23, TTC11]. Teleost [HL98b, Kam05, Woo98a, AE22, Bar94, Bas93, CNH22, CBP02, EB93, EM97, Fri04, Joy23, JB95, Mol92, Nor09, SN10, SR91, SW95a, SW95b, Wie96, ZRZ20]. Teleostei [CRRCdL08, CMA15, CCA17, MS02, May02, PASF13, RVRCB11, VE05, WPD12, WM02, dBdSA12]. teleosts [BCM20, FN02, GELL16, MVM99, SW12, SBN04, TS96]. telomere [RZV12]. temperate [JCL07, McD06, Nor03, SQR09, SLH11, Tay99, WIT14, WMW18]. Temperature [MC04, OCB12, BHP11, CNH22, CTL17, DW09, HHCN11, IPT10, KKW12, Nor14, PSK12, WR16]. Temperature-dependent [OCB12]. template [Qui12]. Temporal [AF04, CRZHZ02, Jon07a, SHPH^{+}22, VSM12]. temporarily [JCL07]. Ten [CWB16, CNT21]. Tench [KGW10, KGW11, FGL10, GAD10, GGZ10, GKR10, HKL10, KGF10, KKF10, KKM10, LKK10, MPK10, ODS10, PSK12, PFL10, RBK10, SW10, SZG10, ZJDZ10]. tension [Nor14]. Tenure [Asw05]. Term [SVM^{+}24a, SVM^{+}24b, AGT08, Ano92r, CCMS12, CBSG24, CBL17, CDH^{+}23b, KdSD18, MSR14, MTPR15, MKK10, OK14, OK10, OK12, OV24, PWM23, PWC06, SW10, SG24, SHB21, BP08]. terms [Ano94g]. Territories [MLK13]. Territory [Gou16]. Tesch [Coc05]. testicular [ARL12]. testing [CGL22, KCF14]. Tetraodontiformes [NVA12]. Thailand [STS07]. Their [Ano92c, LFJ08, AGJ14, ADC15, AVA19, Ano92r, Ano95a, ASS^{+}23, BH17, BC03, CG00, Fro99, GKC19, HPdL02, LMM11, MWP^{+}23, MAC02, Moo12, PBS14, SMS12, SYL^{+}23, SOB20, SDG15, dlCFP19]. Them [AH23, GP15]. Theory [Bur07, Woo93, CSSO02, CDH^{+}23b, Gra96, Kat97, Mag95]. there [Cad91, FN02]. thermal [HSC^{+}24, KdSD18, ZLF21]. Thiamine [HAC18]. thinking [HCB15]. Third [Mul10]. Thorpe [Coc05]. Those [AH23]. Threat [HH06, LSH15, dMV20]. Threatened [SKD03, BCK18, GK18b, Jel22a, Jel22b, LSH15, PVJ07, RPD16]. threats [BHD23, BGBE^{+}23, GBR22, KHW09, MV13, MMB22, MJM15, SPL12, SDJ13, TP14, dMlTB19]. three [KKW12, MCN12, Nor14, PBC12, YVR07, LQW19]. thresher [FCCS15, MMF18]. threshold [Kat97]. thresholds [GCB^{+}24]. throughput [RNJ16]. Thunnus [GLM20, NMF17, PZC17, VGA11]. Thymallus [EAW22, Nor95]. thynnus [VGA11]. thyroidal [CE96, EB93]. Thysanoteuthis [DMK^{+}24]. tidepool [AGMB^{+}23]. tides [Gou16]. Tidier [De 98]. tier [BCO21]. tilapia [CB00, RTT12, ZMG17, AY10]. tilapias [KL23]. time [BLD19, Cad91, KKM10, SHPH^{+}22, VdSD21]. time-keeping [BLD19]. Timor [PRP16]. Tinca [EDP10, GAD10, GKR10, HKL10, KGW10, KGW11, KGF10, KKM10, MPK10, MKK10, ODS10, PSK12, PFL10, RBK10, SW10, SZG10, ZJDZ10]. tissue [RBPPS16]. tissue-specific [RBPPS16]. tissues [EDP10, WLB13]. together [PZC17]. tolerance [SYL^{+}23, WR16]. tolerances [Nor06]. tonggol [GLM20]. Tony [Pis00]. took [OSC20]. tool [BS03, DMD21, GHMG22, HCK21, LWS17, MLN05a, MLN05b, SMN08]. tools [JBX07, SBSŠ21, VMB07]. toothed [TBA20]. toothfish [MWP^{+}23]. top [HYW13, MIA^{+}23b, MIA^{+}23a, YOC15]. top-down [MIA^{+}23a]. Tor [BP16, PBR19]. Torgersen [DCS12]. Tori [GCK21]. Torpedo [PLA07]. total [NRCD^{+}23]. tourism [HMV17]. towed [MF99]. Toxicity [MPK10, Nor02, HP93, SCC08, SW10]. Trachurus [AGG03]. tracking [Ano94p, MKL^{+}23]. tract [CFGG13, GP12, WOR09]. trade [COGFPV23, Hal98, IOHM23, MCFC21, RCP19, dMlTB19]. trade-offs [Hal98, RCP19]. traded [SKW22]. tradeoffs [GCS19]. trader [ENF^{+}23]. Traditional [MSV^{+}23, MH20, RBC16a, RBC16b]. Traffic [CUT07]. Training [CWD11]. trait [DIR20]. traits [CTB16, GP15, HII16, LWS17, TFF09, VVU22, ZLH^{+}23]. trajectories [DD15]. trammel [BCA^{+}23]. transactions [Sur23]. transcriptomics [SJH22]. transdisciplinary [NRJ^{+}23]. transfer [LSDH12]. transferable [Gra96]. transform [SMN08]. Transgenic [Pen93, Sin97]. transition [KWI07]. translated [Coc05]. translocated [IE06, LMM11]. Translocation [VCD12, Gau01]. transmission [Mar95, NVH21]. transmitters [TTC11]. Transparent [LAI05]. transport [McD09b]. transportation [DML^{+}24]. transposition [BLF10, PBMF12]. trap [BMN21, KFS21]. trap-and-haul [KFS21]. trapping [MSM20]. traps [MF99]. Trasimeno [EDP10]. Traverse [Zim05]. trawl [CR08, Dia04, Jon07a, Ken95, MTL21]. trawlers [BHB22]. trawling [VdSD21]. trawls [Bro00a, KB21, MBS17]. treatments [PSK12]. Trends [BZM12, Har98, MC12, PGA^{+}24, TTC11, dBdSA12, BAA18, CWN11, HBC13, PFW16, PFW17, SK15, TCS20, VAJ10, dRRGC12]. trepang [Gou16]. tricaine [CWB11]. Trichodinidae [MGTD18]. tridentatus [CWO19, JMJ19]. triploid [FGL10, HKL10, TKR04]. triploidization [HKL10]. tristoechus [CFGG13]. trophic [ABS15, Ano95b, CNH22, JBX07, KL23, Mag13, SHHK21, SK02, TBB^{+}24, WCP97, WK99, YOC15]. Tropical [WIT14, Car92, CGKSP13, CRZHZ02, CBO19, EBRLGB02, How94, Hun21, NWO16, PMVA19, SX16, TMD^{+}24, CGRCM19]. Tropicalization [Pau98]. Troschel [DMK^{+}24]. trouble [Ros97]. Trout [Ram08, Ano94j, CRRCdL08, CCMS12, CM08, CTM92, DAN02, GLC16, KMK16, LPC^{+}24, LSH15, PDN^{+}24, PM14, PBB00, QM04, Qui12, RCCRD03, TRG19]. trouts [HPdL02]. trutta [BGTA19]. Trying [MCFC21]. tshawytscha [SAK14]. tuna [ANL12, CBO19, ENP18, FTH15, GLM20, GSC17, Hal98, HGC17, MRMJJ17, NMF17, PZC17, Pie93, PAT15, VGA11]. Tunas [BH17, BBS17, JJMD13, LFMP21, MLC17]. turbid [SMR11, TBB^{+}24]. turbidity [ROP^{+}23, ZMG17]. turbine [PDB16]. turbot [IJ01]. turf [TBB^{+}24]. turf-based [TBB^{+}24]. Turkey [HH04, HH06, Har11, IE06].

Turkish [AKGB11]. turning [Fit93]. turtle [AA10, GH17, MTM22]. twaite [CPM14]. twenty [BH17, Cad99]. twenty-first [BH17, Cad99]. two [BOV09a, BOV09b, CPN11, FHW^{+}24, KdSD18, NVA12, OK14, OCB12, ROW^{+}23, Utt00, VdSD21, WT08]. type [GH17, GCS20]. types [BEBFM14, BZM12]. typical [SA12].

U [Bla98, MJSOF16]. U.S. [EBS20, GPS18, MCL08, Mil12]. U.S.A. [EGMM02]. ubidiai [VE05]. ubiquitous [WGC18]. ugly [Fro99]. UK [Ano94a, HETS23]. ulcerative [KB14]. ultrastructure [PFL10]. ultraviolet [AA20]. uncertainties [CHFTV22, VGA11]. uncertainty [Ano92v, PGG07a, PGG07b]. uncontrolled [MGRRJ21]. underlying [BBS17, Rob92, SYL^{+}23]. underpinning [MBP14]. understand [BSR19, DMD21]. understanding [DAN02, HPP15, KB14, LCC19, PGG07a, PGG07b, PDB16, SVRS19, VBC^{+}23]. Underwater [Mis97, TMKC24, NCMT23, WFH17]. undulatus [SKD03]. unintended [GCS19]. Uninvited [OV24]. Unique [GCO03, MVGW18]. unitaeniatus [dRRGC12]. United [Coc05, Gil93a, HMN17, Gil93b, KCE12, LM19]. universal [DIR20]. universally [Jel22a, Jel22b]. University [Ano94a, MAJ05]. unlikely [PRP16]. Unravelling [JJC22, SMNK^{+}23]. unreported [CKN21]. unsticking [KKM10]. unsuitable [MKK10]. untapped [MBC21]. until [KM20]. Unveiling [GLHD^{+}24b, GLHD^{+}24a]. update [DZZZ22, SAN^{+}23]. updated [Har24]. upper [BO12, KZ07, PBG04, AGO04, KHW09]. uptake [BSR19]. Upwelling [CM98, Woo99a]. Uruguay [PBM14]. USA [ACST17, Zim05]. Use [Han05, MZ00, OK14, PHK20, SW14, AYN11, BRI00, BNC22, BBS17, BOD21, Cal22, CWB16, DCS20, ENP18, GKC19, GPS18, JNW14, KZ20, LCC16, MT14, Mul11, SHS14, SLH11, SHB21, Dri05]. used [Ano92j, Mad07, TTC11, WDP19]. user [Fro99]. users [SJH22]. uses [WFH17]. Using [BCO21, MBC21, MTC19, CDH^{+}23a, CGL22, HBC13, HSS21, JLD^{+}24, LMH21, MMJ22, MHW07, MT18, NCMT23, TdSL15]. utility [Avi00, BM09, CHN18, Vil18]. utilization [SDJ13, SJB^{+}23, Wie96]. UVR [AA20].

V [VAJ10]. vaccines [ROL14]. Validation [MSLN02]. Valley [MC04]. value [APP^{+}23, HMV17]. vanguards [HP14]. vanishing [Nor12]. Variability [CNH22, CM98, Ano92j, Ano92r, BOs12, CGKSP13, EDP10, KW08, KKF10, MDM20, MFS18, PASF13, SMNK^{+}23, TR94]. Variable [NVA12, CBR98]. Variation [McD03, STP22, AF04, CRRCdL08, JN18, KMK16, LLJC24, Nis00, RVRCB11, RCCRD03, TV15, VEK10, VGMAGB02, WHS04, ZLF21, vPKW18]. variations [War08]. variety [BB14]. vary [dSSHK21]. vasotocin [Kul95]. vasotocin-melatonin [Kul95]. vegetation [CM92]. vendace [MSV^{+}23]. venomous [Har24]. venturing [CAC15]. venus [CCR^{+}24]. Verany [CUT07]. Verde [MTM22]. versus [BO12, FSB14, GH16, HKN18, Han05, JJ93, KDF13, MRMJJ17, NLFM12, Woo99a, ZWW20]. Vertebral [McD03]. Vertebrate [WW06, AJM22]. vertebrates [For95]. vertical [AGM19, LY07]. vessel [GH17]. vestiges [RZV12]. veterinarian [HL11]. viability [JEL10, SW14]. vibrant [CFS^{+}23]. Vibrio [PDGBR07]. Victoria [LM93, WMG00]. Video [MBM06, NCMT23, WFH17]. View [Bev98, Jon07b, MN98, PB98, FN02, Hal98, Kat97, Kva98, PDB16, Ros97, SW98, SP98]. viewpoint [ZRZ20]. viral [EEDP18]. Virtual [Chr96b]. virus [EEDP18]. viscous [MV96, YH18]. vision [GB23]. visual [WLB^{+}23]. vitamins [WZB^{+}24]. vitro [SZC11]. Vivas [WBW09]. viviparous [GELL16, MWS22]. voices [CMF^{+}24]. Vol [Bla98]. Volume [Ano99a, Ano99c, Ano00a, Ano00c, Ano02a, Ano02b, Ano02c, Ano02d, Ano04b, Ano05e, Ano05f]. volumes [Ano92a, Ano92p, Ano92w]. Vortex [JEL10]. vote [WPF16]. vote-counting [WPF16]. voyage [HSS21]. VPA [Mag95]. Vth [KGW10, KGW11]. vulgaris [OBS08]. vulnerability [AGJ14, CFB14, MGB15]. vulnerable [AAH98].

W [MBJ12]. W. [WBW09]. W.J [Woo98b]. wahoo [ZFT13]. waist [GOW13]. Walbaum [BOs12]. Wales [NWG07, SE16]. Walters [MAJ05]. warm [BTW15, JCL07]. warm- [BTW15, JCL07]. Warming [TFP22, HCvP16, HSC^{+}24, PHB14, PBS14]. warning [PHB14]. Warruwi [Gou16]. Washington [Dri05, Ste05]. Washington/Covelo/London [Ste05]. wasp [GOW13]. wasp-waist [GOW13]. Water [Bil02, CCO20, AVB^{+}23, BBS17, CMC11, Gib93, HCB11, NWO16, PP92, PSK12, SD91, SQR09, WJP23, ZMG17]. water-breathing [PP92]. waterfall [PGG07a, PGG07b]. waters [ATdLCBR02, BOD21, DPC12, DD18, FHK04, Gil93b, IE06, IWG17, JPC14, MTL21, Man94, MH20, NWG21, OWW04, PJ14, Rey93, SRGS04, SQR09, STW^{+}24, How94]. Watershed [Har99a]. wave [PTOS^{+}24]. wavelet [SMN08]. way [DROCM23, MA08, RNJ16]. ways [Kel23]. Weak [MKK10]. weather [Cus94]. web [Ano94d, CCR21a, CCR21b, MTC19, PKC^{+}24]. webs [ABS15, Ano94o]. Wedemeyer [Sul04]. wedge [HHCM10, HHCN11]. weight [Bag11, CRD00]. welfare [DCS11, DCS12, GS23, MHW07, TBK11, Mul09]. Werner [GCO03]. west [LBS^{+}23, Hun21]. western [Ano94j, DHG18, DAN02, HSS21, MR18, Nor12, SAT11, TMD^{+}24, KWM^{+}22, KWI07, LKK10, LFdSRM16, SLH11]. Wetherall [Sch18]. wetlands [dSPPM12]. whale [TBA20]. whales [Bea07]. where [PM09]. Which [Duc19]. White [Coc05, BKB17, Com93, HTH00]. whitetip [YC20]. Who [AH23, DBR15, MMM22b, MMM22a]. whole [WDP19]. Wide [RBPPS16, MDM20, TBA20]. wide-ranging [TBA20]. Wide-targeted [RBPPS16]. widely [Utt04]. Widening [HETS23]. wild [Cal22, CRRCdL08, DGV11, DCS11, DCS12, GSD18, KKF10, RCCRD03, SBN04, TBK11, WHS04, ZZ04]. Wildlife [Ano94p, BCA^{+}23, Kel23]. Wilhelm [Coc05]. will [Cad99]. William [PDA12]. Williams [Har99a]. Wilson [dC98]. winter [MBC21, IPT10]. wire [Dav96]. Witeska [Nor02]. within [ENP18, MFS18, PHK20, dSSHK21, SDS15, TØH08]. within-river [TØH08]. without [Bag11]. wolf [McD06]. wolffish [FIØ04]. women [CMF^{+}24]. won’t [Kva98]. Wood [Har99a]. Wootton [Kva98]. Work [Asw05, Mye98, Dia04]. Workshop [Ano93g, JBX07, KGW10, KGW11, LM93, YMR12, ANL12, BP08, Jua02, SBtIWG12, WNB12, YOC15]. World [DC05, Gil93b, GB23, Ano92b, LC02, LCC19, MWS22, MMB22, MLK13, OSC20, PBR19, SRBS21, SEH24, TRJ24, TFP22, Wil07, WMA20, Win06]. worldwide [BHB22, BBV12a, BBV12b]. worms [Ano95i, CCH18]. worth [KGF10]. would [GJ07]. Wrasse [SKD03, MGL^{+}23]. Wrongs [HPL05].

XX [BVMF13]. XX/XY [BVMF13]. XY [BVMF13].

Yangtze [CLX15, LQW19, WGL17, WH24]. year [CRD00, HFG07, HETS23, KCE12, MLM19, VCD12, YWS^{+}24]. Years [Hol98, Ann96, BHS19, CGRCM19, CFS^{+}23, JPC14, LFM13, MR18]. yellow [KTW12, LLC11]. yellowfin [PZC17]. Yellowstone [ACST17]. Yield [PB98, BBW09, HPY18]. yolk [Kam02, Kam08, Wie96]. yolk-feeding [Kam02, Kam08]. young [CRD00, MLM19]. young-of-the-year [CRD00, MLM19]. Yuma [Woo99b].

Zealand [Ann96, Bea07, Bol07, Jon07a]. Zebrafish [UIA11, DOP18]. Zone [Bla98, Tur99a]. Zoogeography [Abl06]. zooplankton [Mar95, OWW04]. Zuiew [Shi05].

References

Alessandro:2010:OLS

[AA10] Lucchetti Alessandro and Sala Antonello. An overview of loggerhead sea turtle (Caretta caretta) bycatch and technical mitigation measures in the Mediterranean Sea. Reviews in Fish Biology and Fisheries, 20(2):141–161, June 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

             link.springer.com/article/10.1007/s11160-009-9126-1.

Alves:2020:EUR

[AA20] Ricardo N. Alves and Susana Agustí. Effect of ultraviolet radiation (UVR) on the life stages of fish. Reviews in Fish Biology and Fisheries, 30(2):335–372, June 2020. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

             link.springer.com/article/10.1007/s11160-020-09603-1.

Armstrong:1998:CRV

[AAH98] Janet Armstrong, David Armstrong, and Ray Hilborn. Crustacean resources are vulnerable to serial depletion — the multifaceted decline of crab and shrimp fisheries in the Greater Gulf of Alaska. Reviews in Fish Biology and Fisheries, 8(2):117– 176, June 1998. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1023/A:1008891412756.

Arostegui:2021:ARR

[AAJ21] Martin C. Arostegui, Christopher M. Anderson, and Abby R. Jahn. Approaches to regulating recreational fisheries: balancing biology with angler satisfaction. Reviews in Fish Biology and Fisheries, 31(3):573–598, September 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-021-09662-y.

Ardren:2014:CBA

[ABK14] William R. Ardren, Craig Busack, and John F. Kocik. Collaboration between Atlantic and Pacific salmon biologists to enhance recovery of endangered salmon in North America. Reviews in Fish Biology and Fisheries, 24(3):685–688, September 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-014-9365-7.

Able:2006:FBH

[Abl06] Kenneth W. Able. Fishes of Bermuda: History, zoogeography, annotated checklist, and identification keys. Reviews in Fish Biology and Fisheries, 16(1):107–108, February 2006. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-006-9000-3.

Abrantes:2015:HSF

[ABS15] Kátya G. Abrantes, Adam Barnett, and Marcus Sheaves. Habitat-specific food webs and trophic interactions supporting coastal-dependent fishery species: an Australian case study. Reviews in Fish Biology and Fisheries, 25(2):337–363, June 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-015-9385-y.

Aigo:2008:DIN

[ACB08] Juana Aigo, Víctor Cussac, and Miguel Battini. Distribution of introduced and native fish in Patagonia (Argentina): patterns and changes in fish assemblages. Reviews in Fish Biology and Fisheries, 18(4):387–408, November 2008. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-007-9080-8.

Al-Chokhachy:2017:ESS

[ACST17] Robert Al-Chokhachy, Adam J. Sepulveda, and Michael T. Tercek. Evaluating species-specific changes in hydrologic regimes: an iterative approach for salmonids in the Greater Yellowstone Area (USA). Reviews in Fish Biology and Fisheries, 27(2):425– 441, June 2017. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-017-9472-3.

Aguzzi:2015:COM

[ADC15] J. Aguzzi, C. Doya, and I. A. Catalán. Coastal observatories for monitoring of fish behaviour and their responses to environmental changes. Reviews in Fish Biology and Fisheries, 25(3): 463–483, September 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-015-9387-9.

Abdolhay:2011:FPS

[ADS11] H. A. Abdolhay, S. K. Daud, and M. K. Abdul Satar. Fingerling production and stock enhancement of mahisefid (Rutilus frisii kutum) lessons for others in the south of Caspian Sea. Reviews in Fish Biology and Fisheries, 21(2):247–257, June 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-010-9163-9.

Arndt:2022:DAL

[AE22] Erik Arndt and Julian Evans. Diel activity of littoral and epipelagic teleost fishes in the Mediterranean Sea. Reviews in Fish Biology and Fisheries, 32(2):497–519, June 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-022-09697-9.

Almada:2004:TVR

[AF04] Vítor C. Almada and Cláudia Faria. Temporal variation of rocky intertidal resident fish assemblages — patterns and possible mechanisms with a note on sampling protocols. Reviews in Fish Biology and Fisheries, 14(2):239–250, June 2004. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-004-6750-7.

Aswani:2015:HHI

[AFB15] Shankar Aswani, Carola F. Flores, and Bernardo R. Broitman. Human harvesting impacts on managed areas: ecological effects of socially-compatible shellfish reserves. Reviews in Fish Biology and Fisheries, 25(1):217–230, March 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-014-9376-4.

Ayilu:2023:BEI

[AFBB23] Raymond K. Ayilu, Michael Fabinyi, Kate Barclay, and Mary Ama Bawa. Blue economy: industrialisation and coastal fishing livelihoods in Ghana. Reviews in Fish Biology and Fisheries, 33(3):801–818, September 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-022-09749-0.

Alexander:2022:EOF

[AFO22] K. A. Alexander, A. Fleming, and E. Ogier. Equity of our future oceans: practices and outcomes in marine science research. Reviews in Fish Biology and Fisheries, 32(1):297–311, March 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-021-09661-z.

Abaunza:2003:GRH

[AGG03] P. Abaunza, L. Gordo, and E. Gallo. Growth and reproduction of horse mackerel, Trachurus trachurus (Carangidae). Reviews in Fish Biology and Fisheries, 13(1):27–61, March 2003. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1026334532390.

Abesamis:2014:IVF

[AGJ14] Rene A. Abesamis, Alison L. Green, and Claro Renato L. Jadloc. The intrinsic vulnerability to fishing of coral reef fishes and their differential recovery in fishery closures. Reviews in Fish Biology and Fisheries, 24(4):1033–1063, December 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-014-9362-x.

Andrzejaczek:2019:PDV

[AGM19] Samantha Andrzejaczek, Adrian C. Gleiss, and Mark G. Meekan. Patterns and drivers of vertical movements of the large fishes of the epipelagic. Reviews in Fish Biology and Fisheries, 29(2):335–354, June 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-019-09555-1.

Andrades:2023:EEC

[AGMB^{+}23] Ryan Andrades, Saúl González-Murcia, Thaddaeus J. Buser, Raphael M. Macieira, Juliana M. Andrade, Hudson T. Pinheiro, Ciro C. Vilar, Caio R. Pimentel, João L. Gasparini, Thaís L. Quintão, Fabíola S. Machado, Gustavo Castellanos-Galindo, Gorgonio Ruiz-Campos, F. Patricio Ojeda, Karen L. Martin, Tommaso Giarrizzo, and Jean-Christophe Joyeux. Ecology, evolution and conservation of tidepool fishes of the Americas. Reviews in Fish Biology and Fisheries, 33(4):1263–1290, December 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09798-z.

Agostinho:2004:FRD

[AGO04] Angelo A. Agostinho, Luiz Carlos Gomes, and Edson K. Okada. Flood regime, dam regulation and fish in the Upper Paraná River: effects on assemblage attributes, reproduction and recruitment. Reviews in Fish Biology and Fisheries, 14(1):11– 19, March 2004. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-004-3551-y.

Able:2008:LTR

[AGT08] K. W. Able, T. M. Grothues, and G. L. Taghon. Longterm response of fishes and other fauna to restoration of former salt hay farms: multiple measures of restoration success. Reviews in Fish Biology and Fisheries, 18(1):65–97, February 2008. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-007-9066-6.

Alverson:1996:BEE

[AH96] D. L. Alverson and Steven E. Hughes. Bycatch: from emotion to effective natural resource management. Reviews in Fish Biology and Fisheries, 6(4):443–462, December 1996. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00164325.

Allegretti:2023:LGR

[AH23] Antonio Allegretti and Christina C. Hicks. ‘Getting the right nutrients to those who need them most’: towards nutritionsensitive governance of fisheries in the Global South. Reviews in Fish Biology and Fisheries, 33(3):561–571, September 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-022-09743-6.

Alavi:2012:SBCb

[AHL12] Sayyed Mohammad Hadi Alavi, Azadeh Hatef, and Otomar Linhart. Sperm biology and control of reproduction in sturgeon: (II) sperm morphology, acrosome reaction, motility and cryopreservation. Reviews in Fish Biology and Fisheries, 22(4): 861–886, December 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-012-9270-x.

Ali:2004:RRA

[AHW04] Bahy A. Ali, Tian-Hua Huang, and Xiao-Mei Wang. A review of random amplified polymorphic DNA (RAPD) markers in fish research. Reviews in Fish Biology and Fisheries, 14(4): 443–453, December 2004. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-005-0815-0.

Ahmad:2022:FGI

[AJM22] Syed Farhan Ahmad, Maryam Jehangir, and Cesar Martins. Fish genomics and its impact on fundamental and applied research of vertebrate biology. Reviews in Fish Biology and Fisheries, 32(2):357–385, June 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-021-09691-7.

Alp:2011:AGE

[AKGB11] Ahmet Alp, Cemil Kara, and Emili García-Berthou. Age and growth of the European catfish (Silurus glanis) in a Turkish reservoir and comparison with introduced populations. Reviews in Fish Biology and Fisheries, 21(2):283–294, June 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-010-9168-4.

Austin:2008:MKR

[ALJ08] E. Austin, S. Lucey, and F. Juanes. Michael King: Review of “Fisheries Biology, Assessment and Management, Second edition”. Reviews in Fish Biology and Fisheries, 18(4):451–452, November 2008. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-008-9090-1.

Allenbach:2011:FAE

[All11] Dawn M. Allenbach. Fluctuating asymmetry and exogenous stress in fishes: a review. Reviews in Fish Biology and Fisheries, 21(3):355–376, September 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-010-9178-2.

Alves:2013:SAF

[AMV13] Diego Corrêa Alves and Carolina Viviana Minte-Vera. Scientometric analysis of freshwater fisheries in Brazil: repeating past errors? Reviews in Fish Biology and Fisheries, 23(1): 113–126, March 2013. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-012-9282-6.

Aguilar-Medrano:2020:IDP

[AMVC20] Rosalía Aguilar-Medrano and María Eugenia Vega-Cendejas. Implications of the depth profile on the functional structure of the fish community of the Perdido Fold Belt, Northwestern Gulf of Mexico. Reviews in Fish Biology and Fisheries, 30(4): 657–680, December 2020. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-020-09615-x.

Alves:2013:HAR

[AMVV13] Diego Corrêa Alves, Carolina Viviana Minte-Vera, and Lilian Paula Vasconcelos. Hydrological attributes and rheophilic freshwater fish: stock assessment. Reviews in Fish Biology and Fisheries, 23(3):375–394, September 2013. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-012-9298-y.

Auld:2019:AMC

[ANB19] Heather L. Auld, David L. G. Noakes, and Michael A. Banks. Advancing mate choice studies in salmonids. Reviews in Fish Biology and Fisheries, 29(2):249–276, June 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-019-09551-5.

Allain:2012:IWO

[ANL12] Valerie Allain, Simon Nicol, and Tim Lawson. International workshop on opportunities for ecosystem approaches to fisheries management in the Pacific Ocean tuna fisheries. Reviews in Fish Biology and Fisheries, 22(1):29–33, March 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-011-9220-z.

Annala:1996:NZI

[Ann96] John H. Annala. New Zealand’s ITQ system: have the first eight years been a success or a failure? Reviews in Fish Biology and Fisheries, 6(1):43–62, March 1996. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00058519.

Anonymous:1991:E

[Ano91a] Anonymous. Editorial. Reviews in Fish Biology and Fisheries, 1(1):1–2, September 1991. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF00042658.

Anonymous:1991:SFPa

[Ano91b] Anonymous. A selection of forthcoming papers. Reviews in Fish Biology and Fisheries, 1(1):92, September 1991. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00042663.

Anonymous:1991:SFPb

[Ano91c] Anonymous. A selection of forthcoming papers. Reviews in Fish Biology and Fisheries, 1(2):182, December 1991. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00157584.

Anonymous:1992:AIV

[Ano92a] Anonymous. Author index to volumes 1 and 2. Reviews in Fish Biology and Fisheries, 2(4):353, December 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00043530.

Anonymous:1992:BRE

[Ano92b] Anonymous. Book review: Elsevier’s dictionary of fishery, processing, fish and shellfish names of the world, in five languages: English, French, Spanish, German and Latin. Reviews in Fish Biology and Fisheries, 2(3):275–277, September 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00045045.

Anonymous:1992:CES

[Ano92c] Anonymous. Catch effort sampling strategies: Their application in freshwater fisheries management. Reviews in Fish Biology and Fisheries, 2(3):273–275, September 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/BF00045044.

Anonymous:1992:C

[Ano92d] Anonymous. Corrigendum. Reviews in Fish Biology and Fisheries, 2(4):358, December 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00043533.

Anonymous:1992:CFS

[Ano92e] Anonymous. Cyprinid fishes: Systematics, biology and exploitation. Reviews in Fish Biology and Fisheries, 2(2):184–186, June 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00042887.

Anonymous:1992:DME

[Ano92f] Anonymous. Dynamics of marine ecosystems: Biologicalphysical interactions in the ocean. Reviews in Fish Biology and Fisheries, 2(1):86–87, March 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00042919.

Anonymous:1992:ELH

[Ano92g] Anonymous. Early life history of fish: an energetics approach. Reviews in Fish Biology and Fisheries, 2(4):348–349, December 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00043528.

Anonymous:1992:EFC

[Ano92h] Anonymous. The ecology of fishes on coral reefs. Reviews in Fish Biology and Fisheries, 2(3):277–278, September 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00045046.

Anonymous:1992:E

[Ano92i] Anonymous. Editorial. Reviews in Fish Biology and Fisheries, 2(4):281–282, December 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00043519.

Anonymous:1992:EOV

[Ano92j] Anonymous. Effects of ocean variability on recruitment and an evaluation of parameters used in stock assessment models. Reviews in Fish Biology and Fisheries, 2(1):88–89, March 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00042920.

Anonymous:1992:FES

[Ano92k] Anonymous. Fish evolution and systematics: Evidence from spermatozoa. Reviews in Fish Biology and Fisheries, 2(1):91– 92, March 1992. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/BF00042922.

Anonymous:1992:FMT

[Ano92l] Anonymous. Fish-marking techniques. Reviews in Fish Biology and Fisheries, 2(4):341–342, December 1992. CODEN RF-BFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00043523.

Anonymous:1992:FA

[Ano92m] Anonymous. Fisheries acoustics. Reviews in Fish Biology and Fisheries, 2(4):339–341, December 1992. CODEN RFBFEA.

ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/BF00043522.

Anonymous:1992:PSL

[Ano92n] Anonymous. Pacific salmon life histories. Reviews in Fish Biology and Fisheries, 2(4):347–348, December 1992. CODEN RF-BFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00043527.

Anonymous:1992:HGE

[Ano92o] Anonymous. Handbook of genotoxic effects and fish chromosomes. Reviews in Fish Biology and Fisheries, 2(3):272–273, September 1992. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/BF00045043.

Anonymous:1992:IBR

[Ano92p] Anonymous. Index of books reviewed in volumes 1 and 2. Reviews in Fish Biology and Fisheries, 2(4):357–358, December 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00043532.

Anonymous:1992:LTL

[Ano92q] Anonymous. Lake Tanganyika and its life. Reviews in Fish Biology and Fisheries, 2(2):180–181, June 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/BF00042885.

Anonymous:1992:LTV

[Ano92r] Anonymous. Long-term variability of pelagic fish populations and their environment. Reviews in Fish Biology and Fisheries, 2(3):269–270, September 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00045041.

Anonymous:1992:MAF

[Ano92s] Anonymous. Mathematical analysis of fish stock dynamics. Reviews in Fish Biology and Fisheries, 2(2):179, June 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00042884.

Anonymous:1992:ORM

[Ano92t] Anonymous. Operations research and management in fishing. Reviews in Fish Biology and Fisheries, 2(1):90–91, March 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00042921.

Anonymous:1992:PFR

[Ano92u] Anonymous. Pelagic fish: The resource and its exploitation. Reviews in Fish Biology and Fisheries, 2(4):342–343, December 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00043524.

Anonymous:1992:QFS

[Ano92v] Anonymous. Quantitative fisheries stock assessment: Choice, dynamics and uncertainty. Reviews in Fish Biology and Fisheries, 2(2):177–178, June 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00042883.

Anonymous:1992:SIV

[Ano92w] Anonymous. Subject index to volumes 1 and 2. Reviews in Fish Biology and Fisheries, 2(4):354–356, December 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00043531.

Anonymous:1992:SFB

[Ano92x] Anonymous. Symposium of fish behaviour in relation to fishing operations. Reviews in Fish Biology and Fisheries, 2(4): 350–352, December 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF00043529.

Anonymous:1993:AHM

[Ano93a] Anonymous. Artificial habitats for marine and freshwater fisheries. Reviews in Fish Biology and Fisheries, 3(1):80–81, March 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00043299.

Anonymous:1993:BRM

[Ano93b] Anonymous. Book review: Marine fisheries of India, first revised edn. Reviews in Fish Biology and Fisheries, 3(1):83– 84, March 1993. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/BF00043301.

Anonymous:1993:BR

[Ano93c] Anonymous. Books received. Reviews in Fish Biology and Fisheries, 3(3):295, September 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00043933.

Anonymous:1993:C

[Ano93d] Anonymous. Corrigendum. Reviews in Fish Biology and Fisheries, 3(3):298, September 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00043935.

Anonymous:1993:IFP

[Ano93e] Anonymous. Impacts of forestry practices on a coastal stream ecosystem, Carnation Creek, British Columbia (Canadian bulletin of fisheries and aquatic sciences no. 223). Reviews in Fish Biology and Fisheries, 3(1):88–89, March 1993. CODEN RF-BFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00043304.

Anonymous:1993:ISM

[Ano93f] Anonymous. International symposium on management strategies for exploited fish populations. Reviews in Fish Biology and Fisheries, 3(3):296–297, September 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/BF00043934.

Anonymous:1993:WRE

[Ano93g] Anonymous. Workshop on risk evaluation and biological reference points for fisheries management. Reviews in Fish Biology and Fisheries, 3(4):366–367, December 1993. CODEN RF-BFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00043385.

Anonymous:1994:BRP

[Ano94a] Anonymous. Book review: Proceedings of the Fourth International Symposium on the Reproductive Physiology of Fish, University of East Anglia, Norwich, UK, 7–12 July 1991. Reviews in Fish Biology and Fisheries, 4(4):489–491, December 1994. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00042893.

Anonymous:1994:CAS

[Ano94b] Anonymous. Creel and angler surveys in fisheries management. Reviews in Fish Biology and Fisheries, 4(1):130–131, March 1994. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00043266.

Anonymous:1994:DGF

[Ano94c] Anonymous. Dr Gerry FitzGerald. Reviews in Fish Biology and Fisheries, 4(2):143–144, June 1994. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00044126.

Anonymous:1994:FWM

[Ano94d] Anonymous. Food web management: a case study of Lake Mendota. Reviews in Fish Biology and Fisheries, 4(1):131– 133, March 1994. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/BF00043267.

Anonymous:1994:FIP

[Ano94e] Anonymous. Fourth Indo-Pacific Fish Conference. Reviews in Fish Biology and Fisheries, 4(2):256–257, June 1994. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00044130.

Anonymous:1994:ASN

[Ano94f] Anonymous. The Atlantic salmon: Natural history, exploitation and future management. Reviews in Fish Biology and Fisheries, 4(2):264–265, June 1994. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF00044135.

Anonymous:1994:GT

[Ano94g] Anonymous. Glossary of terms. Reviews in Fish Biology and Fisheries, 4(3):393–399, September 1994. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/BF00042914.

Anonymous:1994:ISF

[Ano94h] Anonymous. International symposium on fish otolith research and application. Reviews in Fish Biology and Fisheries, 4(1): 124–125, March 1994. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF00043263.

Anonymous:1994:ISM

[Ano94i] Anonymous. International symposium on middle-sized pelagic fish. Reviews in Fish Biology and Fisheries, 4(2):257–258, June 1994. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00044131.

Anonymous:1994:NTW

[Ano94j] Anonymous. Native trout of western North America. Reviews in Fish Biology and Fisheries, 4(4):488–489, December 1994. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00042892.

Anonymous:1994:DEF

[Ano94k] Anonymous. On the dynamics of exploited fish populations. Reviews in Fish Biology and Fisheries, 4(2):259–260, June 1994. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00044132.

Anonymous:1994:PF

[Ano94l] Anonymous. The physiology of fishes. Reviews in Fish Biology and Fisheries, 4(2):261–262, June 1994. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/BF00044133.

Anonymous:1994:PPF

[Ano94m] Anonymous. Protozoan parasites of fishes. Reviews in Fish Biology and Fisheries, 4(1):133–134, March 1994. CODEN RF-BFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00043268.

Anonymous:1994:SFR

[Ano94n] Anonymous. Surveys of fisheries resources. Reviews in Fish Biology and Fisheries, 4(1):128–130, March 1994. CODEN RF-BFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00043265.

Anonymous:1994:WSP

[Ano94o] Anonymous. Webs and scales: Physical and ecological processes in marine fish recruitment. Reviews in Fish Biology and Fisheries, 4(4):486–487, December 1994. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00042891.

Anonymous:1994:WTR

[Ano94p] Anonymous. Wildlife telemetry: Remote monitoring and tracking of animals. Reviews in Fish Biology and Fisheries, 4(2):265– 266, June 1994. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/BF00044136.

Anonymous:1995:ASM

[Ano95a] Anonymous. Angler survey methods and their applications in fisheries management. Reviews in Fish Biology and Fisheries, 5(3):378–380, September 1995. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00043009.

Anonymous:1995:BTF

[Ano95b] Anonymous. Benguela trophic functioning. Reviews in Fish Biology and Fisheries, 5(3):386–387, September 1995. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00043014.

Anonymous:1995:BR

[Ano95c] Anonymous. Books received. Reviews in Fish Biology and Fisheries, 5(1):139–140, March 1995. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF01103370.

Anonymous:1995:GEA

[Ano95d] Anonymous. Genetics and evolution of aquatic organisms. Reviews in Fish Biology and Fisheries, 5(3):385–386, September 1995. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00043013.

Anonymous:1995:GC

[Ano95e] Anonymous. Gillnets and cetaceans. Reviews in Fish Biology and Fisheries, 5(3):391–392, September 1995. CODEN RF-BFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00043017.

Anonymous:1995:GMB

[Ano95f] Anonymous. Global marine biological diversity: a strategy for building conservation into decision making. Reviews in Fish Biology and Fisheries, 5(3):382–383, September 1995. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00043011.

Anonymous:1995:MPA

[Ano95g] Anonymous. Mechanics and physiology of animal swimming. Reviews in Fish Biology and Fisheries, 5(3):377–378, September 1995. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00043008.

Anonymous:1995:SFG

[Ano95h] Anonymous. On the sex of fish and the gender of scientists: Collected essays in fisheries science. Reviews in Fish Biology and Fisheries, 5(3):384–385, September 1995. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/BF00043012.

Anonymous:1995:PWF

[Ano95i] Anonymous. Parasitic worms of fish. Reviews in Fish Biology and Fisheries, 5(3):387–389, September 1995. CODEN RF-BFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00043015.

Anonymous:1995:RFB

[Ano95j] Anonymous. Reviews in fish biology and fisheries — the impact factor. Reviews in Fish Biology and Fisheries, 5(4):455, December 1995. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF01103816.

Anonymous:1996:A

[Ano96a] Anonymous. Announcement. Reviews in Fish Biology and Fisheries, 6(3):361, September 1996. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00122587.

Anonymous:1996:BRa

[Ano96b] Anonymous. Book reviews. Reviews in Fish Biology and Fisheries, 6(2):251–257, June 1996. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00182347.

Anonymous:1996:BRc

[Ano96c] Anonymous. Book reviews. Reviews in Fish Biology and Fisheries, 6(3):363–372, September 1996. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00122588.

Anonymous:1996:BRb

[Ano96d] Anonymous. Books received. Reviews in Fish Biology and Fisheries, 6(2):257–258, June 1996. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00182348.

Anonymous:1996:BRd

[Ano96e] Anonymous. Books received. Reviews in Fish Biology and Fisheries, 6(3):372, September 1996. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00122589.

Anonymous:1996:CR

[Ano96f] Anonymous. Conference report. Reviews in Fish Biology and Fisheries, 6(3):373–374, September 1996. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/BF00122590.

Anonymous:1996:Ea

[Ano96g] Anonymous. Editorial. Reviews in Fish Biology and Fisheries, 6(1):1–3, March 1996. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF00058516.

Anonymous:1996:Eb

[Ano96h] Anonymous. Editorial. Reviews in Fish Biology and Fisheries, 6(2):123, June 1996. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF00182339.

Anonymous:1997:BRb

[Ano97a] Anonymous. Book review. Reviews in Fish Biology and Fisheries, 7(2):291–293, June 1997. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1023/A:1018487727966.

Anonymous:1997:BRa

[Ano97b] Anonymous. Book reviews. Reviews in Fish Biology and Fisheries, 7(1):129–133, March 1997. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1023/A:1018475425241.

Anonymous:1997:BRc

[Ano97c] Anonymous. Book reviews. Reviews in Fish Biology and Fisheries, 7(3):371–385, September 1997. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1023/A:1018448013854.

Anonymous:1997:BRd

[Ano97d] Anonymous. Book reviews. Reviews in Fish Biology and Fisheries, 7(4):493–496, December 1997. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1023/A:1018408432510.

Anonymous:1997:CR

[Ano97e] Anonymous. Conference report. Reviews in Fish Biology and Fisheries, 7(2):295–296, June 1997. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1023/A:1018439812037.

Anonymous:1998:BRa

[Ano98a] Anonymous. Book reviews. Reviews in Fish Biology and Fisheries, 8(1):105–112, March 1998. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1023/A:1017158919306.

Anonymous:1998:BRb

[Ano98b] Anonymous. Book reviews. Reviews in Fish Biology and Fisheries, 8(2):215–222, June 1998. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1023/A:1017139328685.

Anonymous:1998:BRc

[Ano98c] Anonymous. Books received. Reviews in Fish Biology and Fisheries, 8(2):223, June 1998. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1017158427777.

Anonymous:1998:CR

[Ano98d] Anonymous. Conference report. Reviews in Fish Biology and Fisheries, 8(1):113–115, March 1998. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1023/A:1017184504768.

Anonymous:1998:C

[Ano98e] Anonymous. Corrigendum. Reviews in Fish Biology and Fisheries, 8(4):491, December 1998. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1023/A:1017198632046.

Anonymous:1999:AIV

[Ano99a] Anonymous. Author index to volume 9. Reviews in Fish Biology and Fisheries, 9(4):386, December 1999. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1023/A:1017221425751.

Anonymous:1999:SIS

[Ano99b] Anonymous. Special issue: The species concept in fish biology. Reviews in Fish Biology and Fisheries, 9(3):209, September 1999. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1017239415994.

Anonymous:1999:VC

[Ano99c] Anonymous. Volume contents. Reviews in Fish Biology and Fisheries, 9(4):383–385, December 1999. CODEN RFBFEA.

ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1023/A:1017260808913.

Anonymous:2000:AIV

[Ano00a] Anonymous. Author index to volume 10. Reviews in Fish Biology and Fisheries, 10(4):532, December 2000. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1023/A:1016649320253.

Anonymous:2000:BR

[Ano00b] Anonymous. Books received. Reviews in Fish Biology and Fisheries, 10(1):125, March 2000. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1023/A:1017251231132.

Anonymous:2000:VC

[Ano00c] Anonymous. Volume contents. Reviews in Fish Biology and Fisheries, 10(4):529–531, December 2000. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1023/A:1016846527624.

Anonymous:2001:GA

[Ano01] Anonymous. Guide to authors. Reviews in Fish Biology and Fisheries, 11(1):91–94, March 2001. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1023/A:1016683904323.

Anonymous:2002:AIVa

[Ano02a] Anonymous. Author index to volume 11. Reviews in Fish Biology and Fisheries, 11(4):368, December 2002. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1023/A:1021313506299.

Anonymous:2002:AIVb

[Ano02b] Anonymous. Author index to volume 12. Reviews in Fish Biology and Fisheries, 12(4):434, December 2002. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1023/A:1025331825981.

Anonymous:2002:VCa

[Ano02c] Anonymous. Volume contents. Reviews in Fish Biology and Fisheries, 11(4):365–367, December 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1023/A:1021300403445.

Anonymous:2002:VCb

[Ano02d] Anonymous. Volume contents. Reviews in Fish Biology and Fisheries, 12(4):431–433, December 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1023/A:1025356013872.

Anonymous:2004:AI

[Ano04a] Anonymous. Author index. Reviews in Fish Biology and Fisheries, 14(4):504, December 2004. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-005-7700-8.

Anonymous:2004:VC

[Ano04b] Anonymous. Volume contents. Reviews in Fish Biology and Fisheries, 14(4):501–503, December 2004. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-005-7699-x.

Anonymous:2005:AIa

[Ano05a] Anonymous. Author index. Reviews in Fish Biology and Fisheries, 13(4):464, January 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-004-5021-y.

Anonymous:2005:AIb

[Ano05b] Anonymous. Author index. Reviews in Fish Biology and Fisheries, 15(4):425, November 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-006-8368-4.

Anonymous:2005:C

[Ano05c] Anonymous. Corrigendum. Reviews in Fish Biology and Fisheries, 13(4):459, January 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-004-5368-0.

Anonymous:2005:E

[Ano05d] Anonymous. Erratum. Reviews in Fish Biology and Fisheries, 15 (1–2):165, February 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-005-4305-1.

Anonymous:2005:VCa

[Ano05e] Anonymous. Volume contents. Reviews in Fish Biology and Fisheries, 13(4):461–463, January 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-004-5023-9.

Anonymous:2005:VCb

[Ano05f] Anonymous. Volume contents. Reviews in Fish Biology and Fisheries, 15(4):423–424, November 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-006-8367-5.

Arkhipkin:2023:HSF

[ANP^{+}23] A. I. Arkhipkin, Ch. M. Nigmatullin, D. C. Parkyn, A. Winter, and J. Csirke. High seas fisheries: the Achilles’ heel of major straddling squid resources. Reviews in Fish Biology and Fisheries, 33(2):453–474, June 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-022-09733-8.

Avila-Poveda:2007:SAS

[APLL07] Omar Hernando Avila-Poveda and Sandra Liliana Lamouroux-López. Saline acclimation of striped mojarra Eugerres plumieri (Cuvier 1830) and optimal dosage of carp pituitary extract (CPE) to induce spawning. Reviews in Fish Biology and Fisheries, 17(1):11–19, February 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-006-9014-x.

Arismendi:2014:DIS

[APLM14] Ivan Arismendi, Brooke E. Penaluna, and Jorge León-Muñoz. Differential invasion success of salmonids in southern Chile: patterns and hypotheses. Reviews in Fish Biology and Fisheries, 24(3):919–941, September 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-014-9351-0.

Ainsworth:2023:ISP

[APP^{+}23] Gillian B. Ainsworth, Pablo Pita, Cristina Pita, Katina Roumbedakis, Graham J. Pierce, Catherine Longo, Gregory Verutes, Tereza Fonseca, Daniela Castelo, Carlos Montero-Castaño, Julio Valeiras, Francisco Rocha, Laura García dela Fuente, Jose Luis Acuña, M. del Pino Fernández Rueda, Alberto Garazo Fabregat, Alberto Martín-Aristín, and Sebastián Villasante. Identifying sustainability priorities among value chain actors in artisanal common octopus fisheries. Reviews in Fish Biology and Fisheries, 33(3):669–698, September 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09768-5.

Arai:2014:HSG

[Ara14] Takaomi Arai. How have spawning ground investigations of the Japanese eel Anguilla japonica contributed to the stock enhancement? Reviews in Fish Biology and Fisheries, 24 (1):75–88, March 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-013-9318-6.

Arai:2015:DCC

[Ara15] Takaomi Arai. Diversity and conservation of coral reef fishes in the Malaysian South China Sea. Reviews in Fish Biology and Fisheries, 25(1):85–101, March 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-014-9371-9.

Alavi:2012:SBCa

[ARL12] Sayyed Mohammad Hadi Alavi, Marek Rodina, and Otomar Linhart. Sperm biology and control of reproduction in sturgeon: (I) testicular development, sperm maturation and seminal plasma characteristics. Reviews in Fish Biology and Fisheries, 22(3):695–717, September 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-012-9268-4.

Arnason:1996:IFM

[Arn96] Ragnar Arnason. On the ITQ fisheries management system in Iceland. Reviews in Fish Biology and Fisheries, 6(1):63– 90, March 1996. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/BF00058520.

Arnason:2005:PRF

[Arn05] Ragnar Arnason. Property rights in fisheries: Iceland’s experience with ITQs. Reviews in Fish Biology and Fisheries, 15 (3):243–264, August 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-005-5139-6.

Almada:1995:PCR

[AS95] Vitor C. Almada and Ricardo Serrão Santos. Parental care in the rocky intertidal: a case study of adaptation and exaptation in Mediterranean and Atlantic blennies. Reviews in Fish Biology and Fisheries, 5(1):23–37, March 1995. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/BF01103364.

Arreguin-Sanchez:1996:CKP

[AS96] Francisco Arreguín-Sánchez. Catchability: a key parameter for fish stock assessment. Reviews in Fish Biology and Fisheries, 6(2):221–242, June 1996. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF00182344.

Aguzzi:2008:HRA

[AS08] Jacopo Aguzzi and Francesc Sardà. A history of recent advancements on Nephrops norvegicus behavioral and physiological rhythms. Reviews in Fish Biology and Fisheries, 18(2):235– 248, May 2008. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-007-9071-9.

Appadoo:2023:AFS

[ASS^{+}23] Chandani Appadoo, Riad Sultan, Monique Simier, Verena Tandrayen-Ragoobur, and Manuela Capello. Artisanal fishers in small island developing states and their perception of environmental change: the case study of Mauritius. Reviews in Fish Biology and Fisheries, 33(3):611–628, September 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-022-09735-6.

Aswani:2005:CST

[Asw05] Shankar Aswani. Customary sea tenure in Oceania as a case of rights-based fishery management: Does it work? Reviews in Fish Biology and Fisheries, 15(3):285–307, August 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-005-4868-x.

Alvarez-Torres:2002:NFC

[ATdLCBR02] Porfirio Alvarez-Torres, Antonio Díaz de León-Corral, and Enrique Bermúdez-Rodríguez. National fisheries chart 2000: a new instrument for fisheries management in inland waters. Reviews in Fish Biology and Fisheries, 12(2–3):317–326, June 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1025079900993.

Abudaya:2018:SDR

[AUdS18] Mohammed Abudaya, Aylin Ulman, and Giuseppe Notarbartolo di Sciara. Speak of the devil ray (Mobula mobular) fishery in Gaza. Reviews in Fish Biology and Fisheries, 28(1): 229–239, March 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-017-9491-0.

Alves:2019:PAM

[AVA19] Diego Corrêa Alves, Lilian Paula Vasconcelos, and Angelo Antonio Agostinho. Protocol for the assessment of mortality and injuries in fish larvae associated with their downstream passage through hydropower dams. Reviews in Fish Biology and Fisheries, 29(2):501–512, June 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-019-09564-0.

Artoni:2009:KDF

[AVB09] Roberto Ferreira Artoni, Marcelo Ricardo Vicari, and Luiz Antonio Carlos Bertollo. Karyotype diversity and fish conservation of southern field from South Brazil. Reviews in Fish Biology and Fisheries, 19(3):??, September 2009. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-009-9109-2.

Aguzzi:2023:BER

[AVB^{+}23] J. Aguzzi, M. Vigo, N. Bahamon, I. Masmitja, D. Chatzievangelou, N. J. Robinson, J. P. Jónasson, A. Sánchez-Márquez, J. Navarro, and J. B. Company. Burrow emergence rhythms of deep-water Mediterranean Norway lobsters (Nephrops norvegicus) revealed by acoustic telemetry. Reviews in Fish Biology and Fisheries, 33(4):1465–1482, December 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-023-09787-2.

Avise:2000:CGS

[Avi00] John C. Avise. Cytonuclear genetic signatures of hybridization phenomena: Rationale, utility, and empirical examples from fishes and other aquatic animals. Reviews in Fish Biology and Fisheries, 10(3):253–263, September 2000. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1023/A:1016685509431.

Anderson:2007:PPC

[AVL07] Frank E. Anderson, Tooraj Valinassab, and Chung-Cheng Lu. Phylogeography of the pharaoh cuttle Sepia pharaonis based on partial mitochondrial 16S sequence data. Reviews in Fish Biology and Fisheries, 17(2–3):345–352, August 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-007-9042-1.

Augspurger:2017:LHP

[AWC17] Jason M. Augspurger, Manna Warburton, and Gerard P. Closs. Life-history plasticity in amphidromous and catadromous fishes: a continuum of strategies. Reviews in Fish Biology and Fisheries, 27(1):177–192, March 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-016-9463-9.

Ajiboye:2010:SAB

[AY10] O. O. Ajiboye and A. F. Yakubu. Some aspects of biology and aquaculture potentials of Tilapia guineensis (Dumeril) in Nigeria. Reviews in Fish Biology and Fisheries, 20(4):441–455, December 2010. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-009-9152-z.

Ajiboye:2011:RUC

[AYN11] O. O. Ajiboye, A. F. Yakubu, and N. A. Nwogu. A review of the use of copepods in marine fish larviculture. Reviews in Fish Biology and Fisheries, 21(2):225–246, June 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-010-9169-3.

Bornatowski:2018:ERH

[BAA18] Hugo Bornatowski, Ronaldo Angelini, and Alberto F. Amorim. Ecological role and historical trends of large pelagic predators in a subtropical marine ecosystem of the South Atlantic. Reviews in Fish Biology and Fisheries, 28(1):241–259, March 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-017-9492-z.

Baggio:2011:EWL

[Bag11] Michele Baggio. Estimating weight–length relationships without individual weight data: an application to the American lobster (Homarus americanus) fishery of Long Island Sound. Reviews in Fish Biology and Fisheries, 21(4):771–777, December 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-011-9205-y.

Barratt:1993:BSO

[Bar93] L. Barratt. Black Sea oceanography. Reviews in Fish Biology and Fisheries, 3(2):199–200, June 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/BF00045240.

Barlow:1994:BRB

[Bar94] G. W. Barlow. Book review: Behaviour of teleost fishes, second edition. Reviews in Fish Biology and Fisheries, 4(1):126– 128, March 1994. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/BF00043264.

Barange:1995:IIS

[Bar95] Manuel Barange. ICES International Symposium on Fisheries and Plankton Acoustics. Reviews in Fish Biology and Fisheries, 5(4):457–459, December 1995. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF01103817.

Bass:1993:BBH

[Bas93] Andrew H. Bass. From brains to behaviour: hormonal cascades and alternative mating tactics in teleost fishes. Reviews in Fish Biology and Fisheries, 3(2):181–186, June 1993. CODEN RF-BFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00045231.

Brown:1996:KDS

[BB96] Grant E. Brown and Joseph A. Brown. Kin discrimination in salmonids. Reviews in Fish Biology and Fisheries, 6(2):201– 219, June 1996. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/BF00182343.

Bartholomew:2005:RCR

[BB05] Aaron Bartholomew and James A. Bohnsack. A review of catch-and-release angling mortality with implications for notake reserves. Reviews in Fish Biology and Fisheries, 15(1–2): 129–154, February 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-005-2175-1.

Byron:2014:SOM

[BB14] Carrie J. Byron and Brian J. Burke. Salmon ocean migration models suggest a variety of population-specific strategies. Reviews in Fish Biology and Fisheries, 24(3):737–756, September 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-014-9343-0.

Baldwin:2012:IFP

[BBJ12] Rebecca E. Baldwin, Michael A. Banks, and Kym C. Jacobson. Integrating fish and parasite data as a holistic solution for identifying the elusive stock structure of Pacific sardines (Sardinops sagax). Reviews in Fish Biology and Fisheries, 22 (1):137–156, March 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-011-9227-5.

Bakun:2010:IEB

[BBS10] Andrew Bakun, Elizabeth A. Babcock, and Christian J. Salvadeo. Issues of ecosystem-based management of forage fisheries in “open” non-stationary ecosystems: the example of the sardine fishery in the Gulf of California. Reviews in Fish Biology and Fisheries, 20(1):9–29, March 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-009-9118-1.

Bernal:2017:SWC

[BBS17] Diego Bernal, Richard W. Brill, and Holly A. Shiels. Sharing the water column: physiological mechanisms underlying speciesspecific habitat use in tunas. Reviews in Fish Biology and Fisheries, 27(4):843–880, December 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-017-9497-7.

Braga:2012:EFE

[BBV12a] Raul Rennó Braga, Hugo Bornatowski, and Jean Ricardo Simões Vitule. Erratum to: Feeding ecology of fishes: an overview of worldwide publications. Reviews in Fish Biology and Fisheries, 22(4):931–932, December 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https:/

/link.springer.com/article/10.1007/s11160-012-9296-0.

See [BBV12b].

Braga:2012:FEF

[BBV12b] Raul Rennó Braga, Hugo Bornatowski, and Jean Ricardo Simões Vitule. Feeding ecology of fishes: an overview of worldwide publications. Reviews in Fish Biology and Fisheries, 22(4): 915–929, December 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-012-9273-7. See er-

ratum [BBV12a].

Botsford:2009:CSY

[BBW09] Louis W. Botsford, Daniel R. Brumbaugh, and Vidar Wespestad. Connectivity, sustainability, and yield: bridging the gap between conventional fisheries management and marine protected areas. Reviews in Fish Biology and Fisheries, 19(1):69– 95, March 2009. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-008-9092-z.

Bullen:2003:NPF

[BC03] C. R. Bullen and T. J. Carlson. Non-physical fish barrier systems: their development and potential applications to marine ranching. Reviews in Fish Biology and Fisheries, 13(2):201– 212, June 2003. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1023/B:RFBF.0000019481.10670.94.

Baez:2023:WNT

[BCA^{+}23] José C. Báez, Juan A. Camiñas, Raquel Aguilera, Jairo Castro-Gutiérrez, and Raimundo Real. When non-target wildlife species and alien species both affect negatively to an artisanal fishery: the case of trammel net in the Alboran Sea. Reviews in Fish Biology and Fisheries, 33(3):785–799, September 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09759-6.

Bourret:2016:DJC

[BCK16] Samuel L. Bourret, Christopher C. Caudill, and Matthew L. Keefer. Diversity of juvenile Chinook salmon life history pathways. Reviews in Fish Biology and Fisheries, 26(3):375–403, September 2016. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-016-9432-3.

Buckley:2018:STS

[BCK18] Kathryn A. Buckley, David A. Crook, and Peter M. Kyne. Sustainability of threatened species displayed in public aquaria, with a case study of Australian sharks and rays. Reviews in Fish Biology and Fisheries, 28(1):137–151, March 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-017-9501-2.

Brosset:2021:PBF

[BCL21] Pablo Brosset, Steven J. Cooke, and Christophe Lebigre. Physiological biomarkers and fisheries management. Reviews in Fish Biology and Fisheries, 31(4):797–819, December 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-021-09677-5.

Barley:2020:ERB

[BCM20] Shanta C. Barley, Timothy D. Clark, and Jessica J. Meeuwig. Ecological redundancy between coral reef sharks and predatory teleosts. Reviews in Fish Biology and Fisheries, 30(1): 153–172, March 2020. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-019-09588-6.

Benaka:2021:UTC

[BCO21] Lee R. Benaka, Andrea N. Chan, and Noelle A. Olsen. Using a tier classification system to evaluate the quality of bycatch estimates from fisheries. Reviews in Fish Biology and Fisheries, 31(3):737–752, September 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-021-09670-y.

Bailey:2022:LBB

[BCP22] Lauren A. Bailey, Amber R. Childs, and Warren M. Potts. Links between behaviour and metabolic physiology in fishes in the Anthropocene. Reviews in Fish Biology and Fisheries, 32 (2):555–579, June 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-022-09701-2.

Bilous:2020:ASC

[BD20] Miranda Bilous and Karen Dunmall. Atlantic salmon in the Canadian Arctic: potential dispersal, establishment, and interaction with Arctic char. Reviews in Fish Biology and Fisheries, 30(3):463–483, September 2020. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-020-09610-2.

Birstein:2002:LE

[BDD02] Vadim Birstein, Rob DeSalle, and Phaedra Doukakis. Letter to the Editor. Reviews in Fish Biology and Fisheries, 12(1): 107–108, March 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1022665908437. See letter [BL02].

Blaber:2009:ESI

[BDF09] S. J. M. Blaber, C. M. Dichmont, and Fahmi. Elasmobranchs in southern Indonesian fisheries: the fisheries, the status of the stocks and management options. Reviews in Fish Biology and Fisheries, 19(3):367–391, September 2009. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-009-9110-9.

Blaber:2005:SSS

[BDS05] S. J. M. Blaber, C. M. Dichmont, and J. P. Salini. Shared stocks of snappers (Lutjanidae) in Australia and Indonesia: Integrating biology, population dynamics and socio-economics to examine management scenarios. Reviews in Fish Biology and Fisheries, 15(1–2):??, February 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-005-3887-y.

Braccini:2016:IMM

[BdST16] Matias Braccini, Alexandre Aires da Silva, and Ian Taylor. Incorporating movement in the modelling of shark and ray population dynamics: approaches and management implications. Reviews in Fish Biology and Fisheries, 26(1):13–24, March 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-015-9406-x.

Beatson:2007:DPS

[Bea07] Emma Beatson. The diet of pygmy sperm whales, Kogia breviceps, stranded in New Zealand: implications for conservation. Reviews in Fish Biology and Fisheries, 17(2–3):295–303, August 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-007-9039-9.

Bejarano-Escobar:2014:MCC

[BEBFM14] Ruth Bejarano-Escobar, Manuel Blasco, and Javier Francisco-Morcillo. Molecular characterization of cell types in the developing, mature, and regenerating fish retina. Reviews in Fish Biology and Fisheries, 24(1):127–158, March 2014. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-013-9320-z.

Brown:2011:IPE

[BEC11] Richard S. Brown, M. Brad Eppard, and Steven J. Cooke. An introduction to the practical and ethical perspectives on the need to advance and standardize the intracoelomic surgical implantation of electronic tags in fish. Reviews in Fish Biology and Fisheries, 21(1):1–9, March 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-010-9183-5.

Benfey:2006:LE

[Ben06] Tillmann Benfey. Letter to the Editor. Reviews in Fish Biology and Fisheries, 16(1):109–112, February 2006. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-006-9001-2.

Bernatchez:1993:GMS

[Ber93] Louis Bernatchez. Guide to the marine sport fishes of Atlantic Canada and New England. Reviews in Fish Biology and Fisheries, 3(4):385–386, December 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00043395.

Beverton:1998:FFF

[Bev98] Ray Beverton. Fish, fact and fantasy: a long view. Reviews in Fish Biology and Fisheries, 8(3):229–249, September 1998. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1008888411100.

Blaxter:1991:BR

[BFD91] J. H. S. Blaxter, Geoffrey Fryer, and Sophie Des Clers. Book reviews. Reviews in Fish Biology and Fisheries, 1(2):183–197, December 1991. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/BF00157585.

Birnie-Gauvin:2023:RIA

[BGBE^{+}23] Kim Birnie-Gauvin, Xavier Bordeleau, Sindre H. Eldøy, Kristin Bøe, Martin L. Kristensen, Cecilie I. Nilsen, and Robert J. Lennox. A review of iteroparity in anadromous salmonids: biology, threats and implications. Reviews in Fish Biology and Fisheries, 33(4):1005–1025, December 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-023-09773-8.

Blanco-Garrido:2023:FHS

[BGHC23] Francisco Blanco-Garrido, Virgilio Hermoso, and Miguel Clavero. Fishing historical sources: a snapshot of 19th-century freshwater fauna in Spain. Reviews in Fish Biology and Fisheries, 33(4):1353–1369, December 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-022-09753-4.

Banha:2024:APP

[BGM^{+}24] F. Banha, J. Gago, D. Margalejo, J. Feijão, F. Casals, P. M. Anastácio, and F. Ribeiro. Angler’s preferences, perceptions and practices regarding non-native freshwater fish. Reviews in Fish Biology and Fisheries, 34(1):??, ???? 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-023-09819-x.

Baremore:2023:SSF

[BGOW23] Ivy E. Baremore, Rachel T. Graham, Samuel R. D. Owen, and Matthew J. Witt. Small-scale fishing has affected abundance and size distributions of deepwater snappers and groupers in the MesoAmerican region. Reviews in Fish Biology and Fisheries, 33(4):1547–1568, December 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-023-09796-1.

Baez:2021:NAO

[BGR21] José C. Báez, Luis Gimeno, and Raimundo Real. North Atlantic oscillation and fisheries management during global climate change. Reviews in Fish Biology and Fisheries, 31(2):319– 336, June 2021. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-021-09645-z.

Birnie-Gauvin:2019:OAS

[BGTA19] Kim Birnie-Gauvin, Eva B. Thorstad, and Kim Aarestrup. Overlooked aspects of the Salmo salar and Salmo trutta lifecycles. Reviews in Fish Biology and Fisheries, 29(4):749–766, December 2019. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-019-09575-x.

Brill:2017:TTF

[BH17] Richard W. Brill and Alistair J. Hobday. Tunas and their fisheries: safeguarding sustainability in the twenty-first century. Reviews in Fish Biology and Fisheries, 27(4):691–695, December 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-017-9500-3.

Bonizzoni:2022:OCF

[BHB22] Silvia Bonizzoni, Sheryl Hamilton, and Giovanni Bearzi. Odontocete cetaceans foraging behind trawlers, worldwide. Reviews in Fish Biology and Fisheries, 32(3):827–877, September 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-022-09712-z.

Ben-Hasan:2023:APG

[BHD23] Abdulrahman Ben-Hasan and Moslem Daliri. Arabian/Persian Gulf artisanal fisheries: magnitude, threats, and opportunities. Reviews in Fish Biology and Fisheries, 33(3):541–559, September 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-022-09737-4.

Barber:2000:EPF

[BHK00] Iain Barber, Danie Hoare, and Jens Krause. Effects of parasites on fish behaviour: a review and evolutionary perspective. Reviews in Fish Biology and Fisheries, 10(2):131–165, June 2000. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1016658224470.

Burt:2011:IPA

[BHP11] J. M. Burt, S. G. Hinch, and D. A. Patterson. The importance of parentage in assessing temperature effects on fish early life history: a review of the experimental literature. Reviews in Fish Biology and Fisheries, 21(3):377–406, September 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-010-9179-1.

Batsleer:2015:HGQ

[BHP15] J. Batsleer, K. G. Hamon, and J. J. Poos. High-grading and overquota discarding in mixed fisheries. Reviews in Fish Biology and Fisheries, 25(4):715–736, December 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-015-9403-0.

Borcherding:2019:CFA

[BHS19] Jost Borcherding, Katja Heubel, and Svenja Storm. Competition fluctuates across years and seasons in a 6-species-fish community: empirical evidence from the field. Reviews in Fish Biology and Fisheries, 29(3):589–604, September 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-019-09567-x.

Billard:2002:BRD

[Bil02] Roland Billard. Book review: Daniel Gerdeaux (ed.), Gestion piscicole des grands plans d’eau (Fish Management in Large Water Bodies). Reviews in Fish Biology and Fisheries, 11(3): 281–282, September 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1020331818648.

Bishop:2006:SFD

[Bis06] J. Bishop. Standardizing fishery-dependent catch and effort data in complex fisheries with technology change. Reviews in Fish Biology and Fisheries, 16(1):??, February 2006. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-006-0004-9.

Blanchfield:2000:CR

[BJ00] Paul J. Blanchfield and Matthew W. Jones. Conference report. Reviews in Fish Biology and Fisheries, 10(1):119–121, March 2000. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1008945030666.

Bernos:2020:LGF

[BJM20] Thaïs A. Bernos, Ken M. Jeffries, and Nicholas E. Mandrak. Linking genomics and fish conservation decision making: a review. Reviews in Fish Biology and Fisheries, 30(4):587–604, December 2020. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-020-09618-8.

Boldrocchi:2017:DES

[BKB17] G. Boldrocchi, J. Kiszka, and D. Burkholder. Distribution, ecology, and status of the white shark, Carcharodon carcharias, in the Mediterranean Sea. Reviews in Fish Biology and Fisheries, 27(3):515–534, September 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-017-9470-5.

Billard:2000:BCS

[BL00] Roland Billard and Guillaume Lecointre. Biology and conservation of sturgeon and paddlefish. Reviews in Fish Biology and Fisheries, 10(4):355–392, December 2000. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1023/A:1012231526151.

Billard:2002:RBD

[BL02] Roland Billard and Guillaume Lecointre. Reply to Birstein, De-Salle and Doukakis. Reviews in Fish Biology and Fisheries, 12(1):109, March 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1022682604046. See [BDD02].

Blaber:1998:BRC

[Bla98] Stephen J. M. Blaber. Book review: Coastal Zone Management Imperative for Maritime Developing Nations (Coastal Systems and Continental Margins, Vol. 3) B. U. Haq, S. M. Haq, G. Kullenberg and J. H. Stel (eds). Reviews in Fish Biology and Fisheries, 8(4):499–500, December 1998. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1023/A:1008818032408.

Brownscombe:2019:CIF

[BLC19] Jacob W. Brownscombe, Elodie J. I. Lédée, and Steven J. Cooke. Conducting and interpreting fish telemetry studies: considerations for researchers and resource managers. Reviews in Fish Biology and Fisheries, 29(2):369–400, June 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-019-09560-4.

Burgerhout:2019:TKH

[BLD19] Erik Burgerhout, P. Mark Lokman, and Ron P. Dirks. The timekeeping hormone melatonin: a possible key cue for puberty in freshwater eels (Anguilla spp.). Reviews in Fish Biology and Fisheries, 29(1):1–21, March 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-018-9540-3.

Blanco:2010:CIF

[BLF10] Daniel Rodrigues Blanco, Roberto Laridondo Lui, and Orlando Moreira Filho. Characterization of invasive fish species in a river transposition region: evolutionary chromosome studies in the genus Hoplias (Characiformes, Erythrinidae). Reviews in Fish Biology and Fisheries, 20(1):1–8, March 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-009-9116-3.

Berntson:2009:ULG

[BM09] Ewann A. Berntson and Paul Moran. The utility and limitations of genetic data for stock identification and management of North Pacific rockfish (Sebastes spp.). Reviews in Fish Biology and Fisheries, 19(2):??, June 2009. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-008-9101-2.

Bordalo-Machado:2009:FBS

[BMF09] Pedro Bordalo-Machado and Ivone Figueiredo. The fishery for black scabbardfish (Aphanopus carbo Lowe, 1839) in the Portuguese continental slope. Reviews in Fish Biology and Fisheries, 19(1):49–67, March 2009. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-008-9089-7.

Bond:2021:QFA

[BMN21] Todd Bond, Dianne L. McLean, and Stephen J. Newman. Quantifying fishing activity targeting subsea pipelines by commercial trap fishers. Reviews in Fish Biology and Fisheries, 31(4): 1009–1023, December 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-021-09686-4.

Bax:2022:ORU

[BNC22] Narissa Bax, Camilla Novaglio, and Chris G. Carter. Ocean resource use: building the coastal blue economy. Reviews in Fish Biology and Fisheries, 32(1):189–207, March 2022. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-021-09636-0.

Britton:2012:NNF

[BO12] J. Robert Britton and Mário Luís Orsi. Non-native fish in aquaculture and sport fishing in Brazil: economic benefits versus risks to fish diversity in the upper River Paraná Basin. Reviews in Fish Biology and Fisheries, 22(3):555–565, September 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-012-9254-x.

Burnett:2021:FTA

[BOD21] Matthew J. Burnett, Gordon C. O’Brien, and Colleen T. Downs. Fish telemetry in African inland waters and its use in management: a review. Reviews in Fish Biology and Fisheries, 31 (2):337–357, June 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-021-09650-2.

Buscher:2016:DHS

[BOJ16] Elena Buscher, Angeleen M. Olson, and Francis Juanes. David h. Secor: Migration ecology of marine fishes. Reviews in Fish Biology and Fisheries, 26(3):609–610, September 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-016-9423-4.

Bolstad:2007:SDN

[Bol07] K. S. Bolstad. Systematics and distribution of the New Zealand onychoteuthid fauna (Cephalopoda: Oegopsida), including a new species, Notonykia nesisi sp. nov. Reviews in Fish Biology and Fisheries, 17(2–3):305–335, August 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-007-9041-2.

Baibai:2012:FGA

[BOs12] Tarik Baibai, Laila Oukhattar, and Abdelaziz soukri. First global approach: morphological and biological variability in a genetically homogeneous population of the European pilchard, Sardina pilchardus (Walbaum, 1792) in the North Atlantic coast. Reviews in Fish Biology and Fisheries, 22(1):63–80, March 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-011-9223-9.

Bottaro:2009:BAIa

[BOV09a] Massimiliano Bottaro, Diana Oliveri, and Marino Vacchi. Born among the ice: first morphological observations on two developmental stages of the Antarctic silverfish Pleuragramma antarcticum, a key species of the Southern Ocean. Reviews in Fish Biology and Fisheries, 19(2):??, June 2009. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-009-9106-5.

Bottaro:2009:BAIb

[BOV09b] Massimiliano Bottaro, Diana Oliveri, and Marino Vacchi. Born among the ice: first morphological observations on two developmental stages of the Antarctic silverfish Pleuragramma antarcticum, a key species of the Southern Ocean. Reviews in Fish Biology and Fisheries, 19(4):??, December 2009. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-009-9117-2.

Bertmar:1993:MFB

[BP93] Gunnar Bertmar and Lennart Persson. Methods for fish biology. Reviews in Fish Biology and Fisheries, 3(1):82–83, March 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00043300.

Buchsbaum:2008:SRL

[BP08] Robert Buchsbaum and J. Christopher Powell. Symposium review: Long-term shifts in faunal assemblages in eastern Northg American estuaries: a review of a workshop held at the biennial meeting of the Coastal and Estuarine Research Federation (CERF), November 2007, Providence, Rhode Island. Reviews in Fish Biology and Fisheries, 18(4):447–450, November 2008. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-008-9086-x.

Bhatt:2016:EGM

[BP16] Jay P. Bhatt and Maharaj K. Pandit. Endangered golden mahseer Tor putitora Hamilton: a review of natural history. Reviews in Fish Biology and Fisheries, 26(1):25–38, March 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-015-9409-7.

Britton:2023:GRC

[BPA^{+}23] J. Robert Britton, Adrian C. Pinder, Josep Alós, Robert Arlinghaus, Andy J. Danylchuk, Wendy Edwards, Kátia M. F. Freire, Casper Gundelund, Kieran Hyder, Ivan Jarić, Robert Lennox, Wolf-Christian Lewin, Abigail J. Lynch, Stephen R. Midway, Warren M. Potts, Karina L. Ryan, Christian Skov, Harry V. Strehlow, Sean R. Tracey, Jun ichi Tsuboi, Paul A. Venturelli, Jessica L. Weir, Marc Simon Weltersbach, and Steven J. Cooke. Global responses to the COVID-19 pandemic by recreational anglers: considerations for developing more resilient and sustainable fisheries. Reviews in Fish Biology and Fisheries, 33(4): 1095–1111, December 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-023-09784-5.

Bradshaw:1999:PA

[Bra99] Alison J. Bradshaw. Publisher’s announcement. Reviews in Fish Biology and Fisheries, 9(4):275, December 1999. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1023/A:1008971609822.

Bres:1993:BS

[Bre93] M. Bres. The behaviour of sharks. Reviews in Fish Biology and Fisheries, 3(2):133–159, June 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00045229.

Bartley:2000:UIS

[BRI00] D. M. Bartley, K. Rana, and A. J. Immink. The use of inter-specific hybrids in aquaculture and fisheries. Reviews in Fish Biology and Fisheries, 10(3):325–337, September 2000. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1016691725361.

Bromley:1994:RGE

[Bro94] Peter J. Bromley. The role of gastric evacuation experiments in quantifying the feeding rates of predatory fish. Reviews in Fish Biology and Fisheries, 4(1):36–66, March 1994. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00043260.

Broadhurst:2000:MRB

[Bro00a] Matt K. Broadhurst. Modifications to reduce bycatch in prawn trawls: a review and framework for development. Reviews in Fish Biology and Fisheries, 10(1):27–60, March 2000. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1008936820089.

Brooks:2000:BRE

[Bro00b] Suzanne Brooks. Book review: Endocrine Disruption in Fish. David E. Kime. Reviews in Fish Biology and Fisheries, 10 (1):123, March 2000. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1008947014293.

Bromley:2005:PFO

[Bro05] Daniel W. Bromley. Purging the frontier from our mind: Crafting a new fisheries policy. Reviews in Fish Biology and Fisheries, 15(3):217–229, August 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-005-4866-z.

Baine:2003:HMM

[BS03] M. Baine and J. Side. Habitat modification and manipulation as a management tool. Reviews in Fish Biology and Fisheries, 13(2):187–199, June 2003. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/B:RFBF.0000019480.95010.67.

Broadley:2022:GRC

[BSKB22] Andrew Broadley, Ben Stewart-Koster, and Christopher J. Brown. A global review of the critical link between river flows and productivity in marine fisheries. Reviews in Fish Biology and Fisheries, 32(3):805–825, September 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-022-09711-0.

Barbini:2018:CAF

[BSL18] Santiago A. Barbini, David E. Sabadin, and Luis O. Lucifora. Comparative analysis of feeding habits and dietary niche breadth in skates: the importance of body size, snout length, and depth. Reviews in Fish Biology and Fisheries, 28(3):625– 636, September 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-018-9522-5.

Brotz:2017:JFA

[BSM17] Lucas Brotz, Agustín Schiariti, and Hermes Mianzan. Jellyfish fisheries in the Americas: origin, state of the art, and perspectives on new fishing grounds. Reviews in Fish Biology and Fisheries, 27(1):1–29, March 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-016-9445-y.

Bouyoucos:2019:EOU

[BSR19] Ian A. Bouyoucos, Colin A. Simpfendorfer, and Jodie L. Rummer. Estimating oxygen uptake rates to understand stress in sharks and rays. Reviews in Fish Biology and Fisheries, 29 (2):297–311, June 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-019-09553-3.

Bellafronte:2011:CMP

[BSV11] Elisangela Bellafronte, Michelle Orane Schemberger, and Marcelo Ricardo Vicari. Chromosomal markers in Parodontidae: an analysis of new and reviewed data with phylogenetic inferences. Reviews in Fish Biology and Fisheries, 21(3):559– 570, September 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-010-9177-3.

Buchanan:2014:NEL

[BSWA14] P. J. Buchanan, K. M. Swadling, and K. Wild-Allen. New evidence links changing shelf phytoplankton communities to boundary currents in southeast Tasmania. Reviews in Fish Biology and Fisheries, 24(2):427–442, June 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-013-9312-z.

Brooks:1997:EQF

[BTS97] Suzanne Brooks, Charles R. Tyler, and John P. Sumpter. Egg quality in fish: what makes a good egg? Reviews in Fish Biology and Fisheries, 7(4):387–416, December 1997. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1018400130692.

Bowker:2015:SOF

[BTW15] James D. Bowker, Jesse T. Trushenski, and Niccole Wandelear. Sedative options for fish research: a brief review with new data on sedation of warm-, cool-, and coldwater fishes and recommendations for the drug approval process. Reviews in Fish Biology and Fisheries, 25(1):147–163, March 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-014-9374-6.

Bao:2023:MCE

[BTZ^{+}23] Jianghui Bao, William M. Twardek, Chaoshuo Zhang, Weiwei Li, Xiangyuan Mi, Dongxu Zhang, Jinming Wu, Hao Jiang, Steven J. Cooke, and Ming Duan. Mitigating the cumulative effects of hydropower and climate change on riverine fishes. Reviews in Fish Biology and Fisheries, 33(4):915–930, December 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09766-7.

Burwen:2007:BRJ

[Bur07] Debby Burwen. Book review: John E. Simmonds and David N. MacLennan, Fisheries Acoustics: Theory and Practice. Reviews in Fish Biology and Fisheries, 17(4):633–634, November 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-007-9052-z.

Burton:2010:FPC

[Bur10] Derek Burton. Flatfish (Pleuronectiformes) chromatic biology. Reviews in Fish Biology and Fisheries, 20(1):31–46, March 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-009-9119-0.

Bueno:2013:KDH

[BVM13] Vanessa Bueno, Paulo Cesar Venere, and Vladimir Pavan Margarido. Karyotypic diversification in Hypostomus Lacépède, 1803 (Siluriformes, Loricariidae): biogeographical and phylogenetic perspectives. Reviews in Fish Biology and Fisheries, 23 (1):103–112, March 2013. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-012-9280-8.

Blanco:2013:RRR

[BVMF13] Daniel Rodrigues Blanco, Marcelo Ricardo Vicari, and Orlando Moreira-Filho. The role of the Robertsonian rearrangements in the origin of the XX/XY_{1}Y_{2} sex chromosome system and in the chromosomal differentiation in Harttia species (Siluriformes, Loricariidae). Reviews in Fish Biology and Fisheries, 23 (1):127–134, March 2013. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-012-9283-5.

Bueno:2012:TCE

[BZM12] Vanessa Bueno, Claudio Henrique Zawadzki, and Vladimir Pavan Margarido. Trends in chromosome evolution in the genus Hypostomus Lacépède, 1803 (Osteichthyes, Loricariidae): a new perspective about the correlation between diploid number and chromosomes types. Reviews in Fish Biology and Fisheries, 22 (1):241–250, March 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-011-9215-9.

Caddy:2004:ORG

[CA04] J. F. Caddy and D. J. Agnew. An overview of recent global experience with recovery plans for depleted marine resources and suggested guidelines for recovery planning. Reviews in Fish Biology and Fisheries, 14(1):??, March 2004. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-004-3770-2.

Chiesa:2019:GGM

[CAB19] Stefania Chiesa, Ernesto Azzurro, and Giacomo Bernardi. The genetics and genomics of marine fish invasions: a global review.

Reviews in Fish Biology and Fisheries, 29(4):837–859, December 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-019-09586-8.

Coll:2015:MDE

[CAC15] M. Coll, E. Akoglu, and V. Christensen. Modelling dynamic ecosystems: venturing beyond boundaries with the Ecopath approach. Reviews in Fish Biology and Fisheries, 25(2):413– 424, June 2015. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-015-9386-x.

Caddy:1991:DRT

[Cad91] John F. Caddy. Death rates and time intervals: is there an alternative to the constant natural mortality axiom? Reviews in Fish Biology and Fisheries, 1(2):109–138, December 1991. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00157581.

Caddy:1999:FMT

[Cad99] J. F. Caddy. Fisheries management in the twenty-first century: will new paradigms apply? Reviews in Fish Biology and Fisheries, 9(1):1–43, March 1999. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1023/A:1008829909601.

Cadrin:2000:AMI

[Cad00] Steven X. Cadrin. Advances in morphometric identification of fishery stocks. Reviews in Fish Biology and Fisheries, 10 (1):91–112, March 2000. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1008939104413.

Calderwood:2022:SAU

[Cal22] Julia Calderwood. Smartphone application use in commercial wild capture fisheries. Reviews in Fish Biology and Fisheries, 32(4):1063–1083, December 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-022-09727-6.

Carpenter:1992:CLF

[Car92] K. E. Carpenter. Check-list of the fishes of the eastern tropical Atlantic (CLOFETA). Reviews in Fish Biology and Fisheries, 2(2):182–184, June 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF00042886.

Castell:1993:IA

[Cas93] John D. Castell. Introduction to aquaculture. Reviews in Fish Biology and Fisheries, 3(4):380–382, December 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00043392.

Coward:2000:RPF

[CB00] K. Coward and N. R. Bromage. Reproductive physiology of female tilapia broodstock. Reviews in Fish Biology and Fisheries, 10(1):1–25, March 2000. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1008942318272.

Curry:2010:OPA

[CBA10] R. Allen Curry, Louis Bernatchez, and Céline Audet. The origins and persistence of anadromy in brook charr. Reviews in Fish Biology and Fisheries, 20(4):557–570, December 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-010-9160-z.

Contreras-Balderas:2002:FFR

[CBAVGR02] Salvador Contreras-Balderas, Patricia Almada-Villela, and María García-Ramírez. Freshwater fish at risk or extinct in México. Reviews in Fish Biology and Fisheries, 12(2–3):241– 251, June 2002. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1023/A:1025053001155.

Castello:2019:FEA

[CBB19] Leandro Castello, Peter B. Bayley, and Vandick S. Batista. Flooding effects on abundance of an exploited, long-lived fish population in river-floodplains of the Amazon. Reviews in Fish Biology and Fisheries, 29(2):487–500, June 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-019-09559-x.

Contreras-Balderas:2002:FBC

[CBEGR02] Salvador Contreras-Balderas, Robert J. Edwards, and María Elena García-Ramírez. Fish biodiversity changes in the Lower Rio Grande/Rio Bravo, 1953–1996. Reviews in Fish Biology and Fisheries, 12(2–3):219–240, June 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https:

//link.springer.com/article/10.1023/A:1025048106849.

Campbell:2024:BBF

[CBF^{+}24] Matthew A. Campbell, Randy J. Brown, Kevin M. Fraley, Dmitry V. Politov, J. Andrés López, and Martin D. Robards. Biogeography of Beringian fishes after the molecular revolution and into the post-genomics era. Reviews in Fish Biology and Fisheries, 34(1):??, ???? 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-023-09827-x.

Cuetos-Bueno:2015:RES

[CBH15] Javier Cuetos-Bueno and Peter Houk. Re-estimation and synthesis of coral-reef fishery landings in the Commonwealth of the Northern Mariana Islands since the 1950s suggests the decline of a common resource. Reviews in Fish Biology and Fisheries, 25 (1):179–194, March 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-014-9358-6.

Cuetos-Bueno:2019:CER

[CBHOH19] Javier Cuetos-Bueno, Dalia Hernandez-Ortiz, and Peter Houk. Co-evolution of “race-to-fish” dynamics and declining size structures in an expanding commercial coral-reef fishery. Reviews in Fish Biology and Fisheries, 29(1):147–160, March 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-018-9542-1.

Collingsworth:2017:CCL

[CBL17] Paris D. Collingsworth, David B. Bunnell, and Stuart A. Ludsin. Climate change as a long-term stressor for the fisheries of the Laurentian Great Lakes of North America. Reviews in Fish Biology and Fisheries, 27(2):363–391, June 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-017-9480-3.

Chassot:2019:KRN

[CBO19] Emmanuel Chassot, Nathalie Bodin, and David Obura. The key role of the Northern Mozambique Channel for Indian Ocean tropical tuna fisheries. Reviews in Fish Biology and Fisheries, 29(3):613–638, September 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-019-09569-9.

Coward:2002:GPF

[CBP02] K. Coward, N. R. Bromage, and J. Parrington. Gamete physiology, fertilization and egg activation in teleost fish. Reviews in Fish Biology and Fisheries, 12(1):33–58, March 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1022613404123.

Cochrane:1998:MPF

[CBR98] Kevern L. Cochrane, Doug S. Butterworth, and Beatriz A. Roel. Management procedures in a fishery based on highly variable stocks and with conflicting objectives: experiences in the South African pelagic fishery. Reviews in Fish Biology and Fisheries, 8(2):177–214, June 1998. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1008894011847.

Changeux:2024:LTO

[CBSG24] Thomas Changeux, Philippe Boisneau, Nicolas Stolzenberg, and Chloé Goulon. A long term overview of freshwater fisheries in France. Reviews in Fish Biology and Fisheries, 34(1):??, ???? 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09803-5.

Chapuis:2022:ASC

[CC22] Lucille Chapuis and Shaun P. Collin. The auditory system of cartilaginous fishes. Reviews in Fish Biology and Fisheries, 32 (2):521–554, June 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-022-09698-8.

Cornelio:2017:HCC

[CCA17] Diana Cornelio, Jonathan Pena Castro, and Roberto Ferreira Artoni. Hermaphroditism can compensate for the sex ratio in the Astyanax scabripinnis species complex (Teleostei: Characidae): expanding the B chromosome study model. Reviews in Fish Biology and Fisheries, 27(3):681–689, September 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-017-9488-8.

Castello:2023:AAD

[CCA^{+}23] Leandro Castello, Felipe Carvalho, Nelly Ornelle Onana Ateba, Alidor Kankonda Busanga, Amy Ickowitz, and Emmanuel Frimpong. An approach to assess data-less small-scale fisheries: examples from Congo rivers. Reviews in Fish Biology and Fisheries, 33(3):593–610, September 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-023-09770-x.

Cole:2018:RGF

[CCH18] Victoria J. Cole, Rowan C. Chick, and Patricia A. Hutchings. A review of global fisheries for polychaete worms as a resource for recreational fishers: diversity, sustainability and research needs. Reviews in Fish Biology and Fisheries, 28(3):543–565, September 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-018-9523-4.

Carmona-Catot:2012:LTC

[CCMS12] Gerard Carmona-Catot, Peter B. Moyle, and Rachel E. Simmons. Long-term captive breeding does not necessarily prevent reestablishment: lessons learned from Eagle Lake rainbow trout. Reviews in Fish Biology and Fisheries, 22(1):325–342, March 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-011-9230-x.

Cooke:2020:WRD

[CCO20] S. J. Cooke, J. J. Cech, and C. M. O’Connor. Water resource development and sturgeon (Acipenseridae): state of the science and research gaps related to fish passage, entrainment, impingement and behavioural guidance. Reviews in Fish Biology and Fisheries, 30(2):219–244, June 2020. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-020-09596-x.

Coll:2013:SSN

[CCP13] Marta Coll, Philippe Cury, and Workshop Participants. The scientific strategy needed to promote a regional ecosystem-based approach to fisheries in the Mediterranean and Black Seas. Reviews in Fish Biology and Fisheries, 23(4):415–434, December 2013. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-013-9305-y.

Carlucci:2021:ACF

[CCR21a] Roberto Carlucci, Francesca Capezzuto, and Pasquale Ricci. Assessment of cetacean–fishery interactions in the marine food web of the Gulf of Taranto (Northern Ionian Sea, Central Mediterranean Sea). Reviews in Fish Biology and Fisheries, 31(1): 135–156, March 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-020-09623-x. See correction

[CCR21b].

Carlucci:2021:CAC

[CCR21b] Roberto Carlucci, Francesca Capezzuto, and Pasquale Ricci. Correction to: Assessment of cetacean–fishery interactions in the marine food web of the Gulf of Taranto (Northern Ionian Sea, Central Mediterranean Sea). Reviews in Fish Biology and Fisheries, 31(1):157, March 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-020-09630-y. See

[CCR21a].

Carlucci:2024:FAS

[CCR^{+}24] R. Carlucci, D. Cascione, P. Ricci, D. De Padova, V. Dragone, G. Cipriano, and M. Mossa. Fluctuations in abundance of the striped venus clam Chamelea gallina in the southern Adriatic Sea (Central Mediterranean Sea): knowledge, gaps and insights for ecosystem-based fishery management. Reviews in Fish Biology and Fisheries, 34(2):827–848, June 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-024-09840-8.

Castilla:2001:LAB

[CD01] Juan Carlos Castilla and Omar Defeo. Latin American benthic shellfisheries: emphasis on co-management and experimental practices. Reviews in Fish Biology and Fisheries, 11(1):1– 30, March 2001. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1023/A:1014235924952.

Carranza:2009:LFM

[CDC09] Alvar Carranza, Omar Defeo, and Juan Carlos Castilla. Linking fisheries management and conservation in bioengineering species: the case of South American mussels (Mytilidae). Reviews in Fish Biology and Fisheries, 19(3):??, September 2009. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-009-9108-3.

Chen:2023:ACS

[CDH^{+}23a] Jinnan Chen, Chengzhi Ding, Dekui He, Liuyong Ding, Songhao Ji, Tingqi Du, Jingrui Sun, Minrui Huang, and Juan Tao. Assessing the conservation status of Chinese freshwater fish using deep learning. Reviews in Fish Biology and Fisheries, 33(4): 1505–1521, December 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-023-09792-5.

Cope:2023:SAT

[CDH^{+}23b] Jason M. Cope, Natalie A. Dowling, Sybrand A. Hesp, Kristen L. Omori, Pia Bessell-Browne, Leandro Castello, Rowan Chick, Dawn Dougherty, Steven J. Holmes, Richard McGarvey, Daniel Ovando, Josh Nowlis, and Jeremy Prince. The stock assessment theory of relativity: deconstructing the term “data-limited” fisheries into components and guiding principles to support the science of fisheries management. Reviews in Fish Biology and Fisheries, 33(1):241–263, March 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-022-09748-1.

Crook:2016:TDC

[CDS16] David A. Crook, Michael M. Douglas, and Stephan Schnierer. Towards deeper collaboration: stories of indigenous interests, aspirations, partnerships and leadership in aquatic research and management. Reviews in Fish Biology and Fisheries, 26(4): 611–615, December 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-016-9449-7.

Cyr:1996:IBT

[CE96] Daniel G. Cyr and J. G. Eales. Interrelationships between thyroidal and reproductive endocrine systems in fish. Reviews in Fish Biology and Fisheries, 6(2):165–200, June 1996. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00182342.

Ceccarelli:2014:HPV

[CFB14] Daniela M. Ceccarelli, Ashley J. Frisch, and Maria Beger. Habitat partitioning and vulnerability of sharks in the Great Barrier Reef Marine Park. Reviews in Fish Biology and Fisheries, 24 (1):169–197, March 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-013-9324-8.

Comabella:2013:ODD

[CFGG13] Yamilé Comabella, Arlette Hernández Franyutti, and Tsai García-Galano. Ontogenetic development of the digestive tract in Cuban gar (Atractosteus tristoechus) larvae. Reviews in Fish Biology and Fisheries, 23(2):245–260, June 2013. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-012-9289-z.

Cooke:2023:TVF

[CFS^{+}23] Steven J. Cooke, Elizabeth A. Fulton, Warwick H. H. Sauer, Abigail J. Lynch, Jason S. Link, Aaron A. Koning, Joykrushna Jena, Luiz G. M. Silva, Alison J. King, Rachel Kelly, Matthew Osborne, Julia Nakamura, Ann L. Preece, Atsushi Hagiwara, Kerstin Forsberg, Julie B. Kellner, Ilaria Coscia, Sarah Helyar, Manuel Barange, Elizabeth Nyboer, Meryl J. Williams, Ratana Chuenpagdee, Gavin A. Begg, and Bronwyn M. Gillanders. Towards vibrant fish populations and sustainable fisheries that benefit all: learning from the last 30 years to inform the next 30 years. Reviews in Fish Biology and Fisheries, 33 (2):317–347, June 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-023-09765-8.

Collen:2000:GEB

[CG00] P. Collen and R. J. Gibson. The general ecology of beavers (Castor spp.), as related to their influence on stream ecosystems and riparian habitats, and the subsequent effects on fish — a review.

Reviews in Fish Biology and Fisheries, 10(4):439–461, December 2000. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1012262217012.

Condini:2018:RBE

[CGCG18] Mario Vinicius Condini, José Antonio García-Charton, and Alexandre Miranda Garcia. A review of the biology, ecology, behavior and conservation status of the dusky grouper, Epinephelus marginatus (Lowe 1834). Reviews in Fish Biology and Fisheries, 28(2):301–330, June 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-017-9502-1.

Catul:2011:RMF

[CGK11] Venecia Catul, Manguesh Gauns, and P. K. Karuppasamy. A review on mesopelagic fishes belonging to family Myctophidae. Reviews in Fish Biology and Fisheries, 21(3):339–354, September 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-010-9176-4.

Castellanos-Galindo:2013:SVM

[CGKSP13] Gustavo A. Castellanos-Galindo, Uwe Krumme, and Ulrich Saint-Paul. Spatial variability of mangrove fish assemblage composition in the tropical eastern Pacific Ocean. Reviews in Fish Biology and Fisheries, 23(1):69–86, March 2013. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-012-9276-4.

Cusa:2022:FSP

[CGL22] Marine Cusa, Katie St John Glew, and Catherine Longo. A future for seafood point-of-origin testing using DNA and stable isotope signatures. Reviews in Fish Biology and Fisheries, 32 (2):597–621, June 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-021-09680-w.

Constance:2024:RLH

[CGP^{+}24] Julia M. Constance, Erica A. Garcia, Richard D. Pillans, Vinay Udyawer, and Peter M. Kyne. A review of the life history and ecology of euryhaline and estuarine sharks and rays. Reviews in Fish Biology and Fisheries, 34(1):??, ???? 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09807-1.

Cowan:2011:RSM

[CGR11] J. H. Cowan, Jr., C. B. Grimes, and K. A. Rose. Red snapper management in the Gulf of Mexico: science- or faith-based? Reviews in Fish Biology and Fisheries, 21(2):187–204, June 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-010-9165-7.

Castellanos-Galindo:2019:ATT

[CGRCM19] Gustavo A. Castellanos-Galindo, D. Ross Robertson, and Carolina Chong-Montenegro. Atlantic tarpon in the Tropical Eastern Pacific 80 years after it first crossed the Panama Canal. Reviews in Fish Biology and Fisheries, 29(2):401–416, June 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-019-09565-z.

Carvalho:1994:MGS

[CH94] G. R. Carvalho and L. Hauser. Molecular genetics and the stock concept in fisheries. Reviews in Fish Biology and Fisheries, 4(3): 326–350, September 1994. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF00042908.

Carvalho:1999:MMS

[CH99] Gary R. Carvalho and Lorenz Hauser. Molecular markers and the species concept: New techniques to resolve old disputes? Reviews in Fish Biology and Fisheries, 9(4):379–382, December 1999. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1008967532146.

Chadwick:1995:IRK

[Cha95] Michael Chadwick. Index rivers: a key to managing anadromous fish. Reviews in Fish Biology and Fisheries, 5(1):38–51, March 1995. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF01103365.

Caballero-Huertas:2022:CKS

[CHFTV22] Marta Caballero-Huertas, Xènia Frigola-Tepe, and Jordi Viñas. The current knowledge status of the genetic population structure of the European sardine (Sardina pilchardus): uncertainties to be solved for an appropriate fishery management. Reviews in Fish Biology and Fisheries, 32(3):745–763, September 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-022-09704-z.

Cvitanovic:2018:GFT

[CHN18] C. Cvitanovic, A. J. Hobday, and K. L. Nash. Governing fisheries through the critical decade: the role and utility of polycentric systems. Reviews in Fish Biology and Fisheries, 28 (1):1–18, March 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-017-9495-9.

Christensen:1996:MFI

[Chr96a] Villy Christensen. Managing fisheries involving predator and prey species. Reviews in Fish Biology and Fisheries, 6(4):417– 442, December 1996. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF00164324.

Christensen:1996:VPR

[Chr96b] Villy Christensen. Virtual population reality. Reviews in Fish Biology and Fisheries, 6(2):243–247, June 1996. CODEN RF-BFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00182345.

Cucherousset:2018:EBM

[CHS18] Julien Cucherousset, Pavel Horky, and Frédéric Santoul. Ecology, behaviour and management of the European catfish. Reviews in Fish Biology and Fisheries, 28(1):177–190, March 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-017-9507-9.

Curley:2013:RBE

[CJV13] Belinda G. Curley, Alan R. Jordan, and Vanessa C. Valenzuela. A review of the biology and ecology of key fishes targeted by coastal fisheries in south-east Australia: identifying critical knowledge gaps required to improve spatial management. Reviews in Fish Biology and Fisheries, 23(4):435–458, December 2013. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-013-9309-7.

Clegg:2021:AGB

[CKN21] Thomas L. Clegg, Steven J. Kennelly, and Kjell Nedreaas. Applying global best practices for estimating unreported catches in Norwegian fisheries under a discard ban. Reviews in Fish Biology and Fisheries, 31(1):1–23, March 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-020-09624-w.

Crowley:2022:ISM

[CLB^{+}22] Philip H. Crowley, Jacques Labonne, Valérie Bolliet, Françoise Daverat, and Agnès Bardonnet. Implications of stress-mediated environmental sex determination for declining eel populations. Reviews in Fish Biology and Fisheries, 32(4):1157–1186, December 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-022-09730-x.

Carpi:2021:OSM

[CLH21] Piera Carpi, Timothy Loher, and Allan C. Hicks. Ontogenetic and spawning migration of Pacific halibut: a review. Reviews in Fish Biology and Fisheries, 31(4):879–908, December 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-021-09672-w.

Cheng:2015:PED

[CLX15] Fei Cheng, Wei Li, and Songguang Xie. Potential effects of dam cascade on fish: lessons from the Yangtze River. Reviews in Fish Biology and Fisheries, 25(3):569–585, September 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-015-9395-9.

Chilton:1992:BMG

[CM92] Earl W. Chilton II and Maurice I. Muoneke. Biology and management of grass carp (Ctenopharyngodon idella, cyprinidae) for vegetation control: a North American perspective. Reviews in Fish Biology and Fisheries, 2(4):283–320, December 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00043520.

Cochrane:1997:CR

[CM97] Kevern Cochrane and Coleen Moloney. Conference report. Reviews in Fish Biology and Fisheries, 7(1):135–136, March 1997. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1018427509311.

Cole:1998:CPV

[CM98] James Cole and Jacqueline McGlade. Clupeoid population variability, the environment and satellite imagery in coastal upwelling systems. Reviews in Fish Biology and Fisheries, 8(4): 445–471, December 1998. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1008861224731.

Crawford:2008:GIS

[CM08] Stephen S. Crawford and Andrew M. Muir. Global introductions of salmon and trout in the genus Oncorhynchus: 1870–2007. Reviews in Fish Biology and Fisheries, 18(3):313–344, August 2008. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-007-9079-1.

Castro:2015:DAS

[CMA15] Jonathan Pena Castro, Mauricio Osvaldo Moura, and Roberto Ferreira Artoni. Diversity of the Astyanax scabripinnis species complex (Teleostei: Characidae) in the Atlantic Forest, Brazil: species limits and evolutionary inferences. Reviews in Fish Biology and Fisheries, 25(1):231–244, March 2015. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-014-9377-3.

Chomyshyn:2011:EWE

[CMC11] L. Chomyshyn, S. H. McConnachie, and S. J. Cooke. Evaluation of water entry into the coelom and different levels of aseptic technique during surgical implantation of electronic tags in freshwater fish. Reviews in Fish Biology and Fisheries, 21 (1):61–70, March 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-010-9189-z.

Chambon:2024:SWP

[CMF^{+}24] Mouna Chambon, Sara Miñarro, Santiago Alvarez Fernandez, Vincent Porcher, Victoria Reyes-Garcia, Huran Tonalli Drouet, and Patrizia Ziveri. A synthesis of women’s participation in small-scale fisheries management: why women’s voices matter. Reviews in Fish Biology and Fisheries, 34(1):??, ???? 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09806-2.

Coelho:2019:HMS

[CML19] Rui Coelho and Rubén Muñoz-Lechuga. Hooking mortality of swordfish in pelagic longlines: comments on the efficiency of minimum retention sizes. Reviews in Fish Biology and Fisheries, 29(2):453–463, June 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-018-9543-0.

Canseco:2022:VED

[CNH22] J. A. Canseco, E. J. Niklitschek, and C. Harrod. Variability in δ^{13}C and δ^{15}N trophic discrimination factors for teleost fishes: a meta-analysis of temperature and dietary effects. Reviews in Fish Biology and Fisheries, 32(2):313–329, June 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-021-09689-1.

Cooke:2021:TSR

[CNT21] Steven J. Cooke, Elizabeth Nyboer, and William W. Taylor. The ten steps to responsible Inland fisheries in practice: reflections from diverse regional case studies around the globe. Reviews in Fish Biology and Fisheries, 31(4):843–877, December 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-021-09664-w.

Cochran:2005:BRE

[Coc05] Philip A. Cochran. Book review: The Eel, 5th edition. By Friedrich-Wilhelm Tesch, translated from the German by R. J.

White and edited by John E. Thorpe. Blackwell Science, Ltd., Oxford, United Kingdom, 2003. 408 pp. ISBN: 0-632-06389-0. Reviews in Fish Biology and Fisheries, 15(3):175–176, August 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-005-2173-3.

Canedo-Orihuela:2023:MRL

[COGFPV23] Hugo Cañedo-Orihuela, Mayra L. González-Félix, and Martin Perez-Velazquez. Maturation, reproduction, and larval culture of pomacentrids for the ornamental fish trade: successes and challenges. Reviews in Fish Biology and Fisheries, 33(4):1155– 1197, December 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-023-09789-0.

Colihueque:2010:GSS

[Col10] Nelson Colihueque. Genetics of salmonid skin pigmentation: clues and prospects for improving the external appearance of farmed salmonids. Reviews in Fish Biology and Fisheries, 20 (1):71–86, March 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-009-9121-6.

Compagno:1993:GWS

[Com93] L. J. V. Compagno. Great white shark. Reviews in Fish Biology and Fisheries, 3(2):188–191, June 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/BF00045233.

Corten:2002:RCH

[Cor02] Ad Corten. The role of “conservatism” in herring migrations. Reviews in Fish Biology and Fisheries, 11(4):339–361, December 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1021347630813.

Carvalho:1994:E

[CP94] Gary R. Carvalho and Tony J. Pitcher. Editorial. Reviews in Fish Biology and Fisheries, 4(3):269–271, September 1994. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00042905.

Chiesa:2014:GDE

[CPM14] Stefania Chiesa, Armando Piccinini, and Francesco Nonnis Marzano. Genetic data on endangered twaite shad (Clupeidae) assessed in landlocked and anadromous populations: one or more species? Reviews in Fish Biology and Fisheries, 24 (2):659–670, June 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-014-9352-z.

Cardoso:2011:CCT

[CPN11] Adauto Lima Cardoso, Julio Cesar Pieczarka, and Cleusa Yoshiko Nagamachi. Chromosomal characterization of two species of genus Steatogenys (Gymnotiformes: Rhamphichthyoidea: Steatogenini) from the Amazon basin: sex chromosomes and correlations with Gymnotiformes phylogeny. Reviews in Fish Biology and Fisheries, 21(3):613–621, September 2011. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-010-9196-0.

Caddy:1998:CGL

[CR98] J. F. Caddy and P. G. Rodhouse. Cephalopod and groundfish landings: Evidence for ecological change in global fisheries? Reviews in Fish Biology and Fisheries, 8(4):431–444, December 1998. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1008807129366.

Catchpole:2008:GTN

[CR08] T. L. Catchpole and A. S. Revill. Gear technology in Nephrops trawl fisheries. Reviews in Fish Biology and Fisheries, 18(1): 17–31, February 2008. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-007-9061-y.

Craig:1992:HIC

[Cra92] John F. Craig. Human-induced changes in the composition of fish communities in the African Great Lakes. Reviews in Fish Biology and Fisheries, 2(2):93–124, June 1992. CODEN RF-BFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00042880.

Cowan:2000:DDG

[CRD00] J. H. Cowan, Jr., K. A. Rose, and D. R. DeVries. Is densitydependent growth in young-of-the-year fishes a question of critical weight? Reviews in Fish Biology and Fisheries, 10(1):61– 89, March 2000. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1023/A:1008932401381.

Cousido-Rocha:2022:SPM

[CRPI^{+}22] Marta Cousido-Rocha, Maria Grazia Pennino, Francisco Izquierdo, Anxo Paz, Davinia Lojo, Amina Tifoura, Mohamed Yosri Zanni, and Santiago Cerviño. Surplus production models: a practical review of recent approaches. Reviews in Fish Biology and Fisheries, 32(4):1085–1102, December 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-022-09731-w.

Camarena-Rosales:2008:MHV

[CRRCdL08] Faustino Camarena-Rosales, Gorgonio Ruiz-Campos, and Francisco J. García de León. Mitochondrial haplotype variation in wild trout populations (Teleostei: Salmonidae) from northwestern Mexico. Reviews in Fish Biology and Fisheries, 18(1): 33–45, February 2008. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-007-9060-z.

Castillo-Rivera:2002:EST

[CRZHZ02] Manuel Castillo-Rivera, José Alejandro Zavala-Hurtado, and Rocío Zárate. Exploration of spatial and temporal patterns of fish diversity and composition in a tropical estuarine system of Mexico. Reviews in Fish Biology and Fisheries, 12(2– 3):167–177, June 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1025051027676.

Cheung:2004:RED

[CS04] William W. L. Cheung and Yvonne Sadovy. Retrospective evaluation of data-limited fisheries: a case from Hong Kong. Reviews in Fish Biology and Fisheries, 14(2):181–206, June 2004. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-004-5422-y.

Caddy:2005:RAS

[CS05] J. F. Caddy and Tobie Surette. In retrospect the assumption of sustainability for Atlantic fisheries has proved an illusion. Reviews in Fish Biology and Fisheries, 15(4):313–337, November 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-005-5853-0.

Castello:2011:MPD

[CSA11] L. Castello, D. J. Stewart, and C. C. Arantes. Modeling population dynamics and conservation of arapaima in the Amazon. Reviews in Fish Biology and Fisheries, 21(3):623–640, September 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-010-9197-z. See erratum [CSA12].

Castello:2012:EMP

[CSA12] L. Castello, D. J. Stewart, and C. C. Arantes. Erratum to: Modeling population dynamics and conservation of arapaima in the Amazon. Reviews in Fish Biology and Fisheries, 22 (1):375, March 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-011-9241-7. See [CSA11].

Csirke:1999:BRN

[Csi99] Jorge Csirke. Book review: El Niñno and the Peruvian Anchovy Fishery (series: Global Change Instruction Program). Edited by Edward A. Laws. Reviews in Fish Biology and Fisheries, 9 (1):118–121, March 1999. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1008801515441.

Christou:2020:PMD

[CSS20] Maria Christou, Vasiliki Sgardeli, and Konstantinos I. Stergiou. A probabilistic model that determines the social ecological system (SES) attributes that lead to successful discard management. Reviews in Fish Biology and Fisheries, 30(1): 109–119, March 2020. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-020-09593-0.

Castro:2002:GTF

[CSSO02] José J. Castro, José A. Santiago, and Ana T. Santana-Ortega. A general theory on fish aggregation to floating objects: an alternative to the meeting point hypothesis. Reviews in Fish Biology and Fisheries, 11(3):255–277, September 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1020302414472.

Chang:2012:MGC

[CSY12] Yi-Jay Chang, Chi-Lu Sun, and Su-Zan Yeh. Modelling the growth of crustacean species. Reviews in Fish Biology and Fisheries, 22(1):157–187, March 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-011-9228-4.

Copp:2016:RGL

[CTB16] Gordon H. Copp, Ali Serhan Tarkan, and Brian G. Blackwell. A review of growth and life-history traits of native and non-native European populations of black bullhead Ameiurus melas. Reviews in Fish Biology and Fisheries, 26(3):441–469, September 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-016-9436-z.

Carlson:2017:CSS

[CTL17] Andrew K. Carlson, William W. Taylor, and Abigail J. Lynch. Comparing stream-specific to generalized temperature models to guide salmonid management in a changing climate. Reviews in Fish Biology and Fisheries, 27(2):443–462, June 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-017-9467-0.

Crowl:1992:IIB

[CTM92] Todd A. Crowl, Colin R. Townsend, and Angus R. Mcintosh. The impact of introduced brown and rainbow trout on native fish: the case of Australasia. Reviews in Fish Biology and Fisheries, 2(3):217–241, September 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00045038.

Cushing:1994:MCW

[Cus94] D. H. Cushing. Marine climate, weather and fisheries: The effects of weather and climatic changes on fisheries and ocean resources. Reviews in Fish Biology and Fisheries, 4(1):138– 139, March 1994. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/BF00043271.

Ceriola:2007:TLA

[CUT07] Luca Ceriola, Nicola Ungaro, and Francesco Toteda. A “Traffic” light approach for the assessment of the Broadtail shortfin squid Illex coindetii Verany, 1839 in the Southern Adriatic Sea (Central Mediterranean). Reviews in Fish Biology and Fisheries, 17(2–3):??, August 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-006-9019-5.

Cooke:2021:TIR

[CVD21] Steven J. Cooke, Paul Venturelli, and Andy J. Danylchuk. Technological innovations in the recreational fishing sector: implications for fisheries management and policy. Reviews in Fish Biology and Fisheries, 31(2):253–288, June 2021. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-021-09643-1.

Carter:2011:RTM

[CWB11] Kathleen M. Carter, Christa M. Woodley, and Richard S. Brown. A review of tricaine methanesulfonate for anesthesia of fish. Reviews in Fish Biology and Fisheries, 21(1):51–59, March 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-010-9188-0.

Cooke:2016:TPR

[CWB16] Steven J. Cooke, Alexander D. M. Wilson, and Richard S. Brown. Ten practical realities for institutional animal care and use committees when evaluating protocols dealing with fish in the field. Reviews in Fish Biology and Fisheries, 26 (1):123–133, March 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-015-9413-y.

Cooke:2011:TCI

[CWD11] Steven J. Cooke, Glenn N. Wagner, and Katherine A. Deters. Training considerations for the intracoelomic implantation of electronic tags in fish with a summary of common surgical errors. Reviews in Fish Biology and Fisheries, 21(1):11–24, March 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-010-9184-4.

Cooke:2011:ASI

[CWN11] Steven J. Cooke, Christa M. Woodley, and Jennifer L. Nielsen. Advancing the surgical implantation of electronic tags in fish: a gap analysis and research agenda based on a review of trends in intracoelomic tagging effects studies. Reviews in Fish Biology and Fisheries, 21(1):127–151, March 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-010-9193-3.

Clemens:2019:MBP

[CWO19] Benjamin J. Clemens, Laurie Weitkamp, and Alexei M. Orlov. Marine biology of the Pacific lamprey Entosphenus tridentatus. Reviews in Fish Biology and Fisheries, 29(4):767–788, December 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-019-09578-8.

Coker:2014:ILC

[CWP14] Darren J. Coker, Shaun K. Wilson, and Morgan S. Pratchett. Importance of live coral habitat for reef fishes. Reviews in Fish Biology and Fisheries, 24(1):89–126, March 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-013-9319-5.

Contreras:2022:DCC

[CZF22] Pablo Contreras, Mauricio Zamorano, and Jorge G. Farías. Diversity of chromatin condensation patterns, nuclear reorganization, evolution and phylogenetic distribution of sperm nuclear basic proteins in fish. Reviews in Fish Biology and Fisheries, 32(2):331–355, June 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-021-09690-8.

Dunham:2002:AIA

[DAN02] Jason B. Dunham, Susan B. Adams, and Douglas C. Novinger. Alien invasions in aquatic ecosystems: Toward an understanding of brook trout invasions and potential impacts on inland cutthroat trout in western North America. Reviews in Fish Biology and Fisheries, 12(4):373–391, December 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1025338203702.

Davenport:1994:HWD

[Dav94] John Davenport. How and why do flying fish fly? Reviews in Fish Biology and Fisheries, 4(2):184–214, June 1994. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00044128.

Davis:1996:BWB

[Dav96] Anthony Davis. Barbed wire and bandwagons: a comment on ITQ fisheries management. Reviews in Fish Biology and Fisheries, 6(1):97–107, March 1996. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00058522.

deBorba:2012:TKE

[dBdSA12] Rafael Splendore de Borba, Edson Lourenço da Silva, and Anderson Luis Alves. Trends in the karyotypic evolution of the Neotropical catfish family Heptapteridae Bockmann 1998 (Teleostei: Siluriformes). Reviews in Fish Biology and Fisheries, 22(2):509–518, June 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-011-9245-3.

DellApa:2015:WLD

[DBR15] Andrea Dell’Apa, Charles W. Bangley, and Roger A. Rulifson. Who let the dogfish out? A review of management and socio-economic aspects of spiny dogfish fisheries. Reviews in Fish Biology and Fisheries, 25(2):273–295, June 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-014-9379-1.

Deines:2015:RGR

[DBT15] Andrew M. Deines, David B. Bunnell, and William W. Taylor. A review of the global relationship among freshwater fish, autotrophic activity, and regional climate. Reviews in Fish Biology and Fisheries, 25(2):323–336, June 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-015-9384-z.

desClers:1998:BRN

[dC98] Sophie des Clers. Book review: Northwest Atlantic Groundfish: Perspectives on a Fishery Collapse, J. Boreman, B. S. Nakashima, J. A. Wilson and R. L. Kendall (eds). Reviews in Fish Biology and Fisheries, 8(4):497–499, December 1998. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1008808321930.

Defeo:2005:MTO

[DC05] Omar Defeo and Juan Carlos Castilla. More than one bag for the world fishery crisis and keys for co-management successes in selected artisanal Latin American shellfisheries. Reviews in Fish Biology and Fisheries, 15(3):265–283, August 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-005-4865-0.

Diggles:2011:EWA

[DCS11] B. K. Diggles, S. J. Cooke, and W. Sawynok. Ecology and welfare of aquatic animals in wild capture fisheries. Reviews in Fish Biology and Fisheries, 21(4):739–765, December 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-011-9206-x.

Diggles:2012:RTA

[DCS12] B. K. Diggles, S. J. Cooke, and W. Sawynok. Response to Torgersen et al. (2011): Reply to Diggles et al. (2011): Ecology and welfare of aquatic animals in wild capture fisheries. Reviews in Fish Biology and Fisheries, 22(1):367–369, March 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-011-9248-0.

Dias:2020:PMS

[DCS20] Ana Carolina Esteves Dias, Ana Cinti, and Cristiana Simão Seixas. Participatory monitoring of small-scale coastal fisheries in South America: use of fishers’ knowledge and factors affecting participation. Reviews in Fish Biology and Fisheries, 30 (2):313–333, June 2020. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-020-09602-2.

Dwivedi:2013:RAA

[DD13a] A. K. Dwivedi and V. K. Dubey. Retracted article: Advancements in morphometric differentiation: a review on stock identification among fish populations. Reviews in Fish Biology and Fisheries, 23(1):23–39, March 2013. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-012-9279-1.

Dwivedi:2013:RNA

[DD13b] A. K. Dwivedi and V. K. Dubey. Retraction note: Advancements in morphometric differentiation: a review on stock identification among fish populations. Reviews in Fish Biology and Fisheries, 23(4):557, December 2013. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-013-9322-x.

Dvoretsky:2015:CFS

[DD15] Alexander G. Dvoretsky and Vladimir G. Dvoretsky. Commercial fish and shellfish in the Barents Sea: Have introduced crab species affected the population trajectories of commercial fish? Reviews in Fish Biology and Fisheries, 25(2):297–322, June 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-015-9382-1.

Dvoretsky:2018:RKC

[DD18] Alexander G. Dvoretsky and Vladimir G. Dvoretsky. Red king crab (Paralithodes camtschaticus) fisheries in Russian waters: historical review and present status. Reviews in Fish Biology and Fisheries, 28(2):331–353, June 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-017-9510-1.

Dobbs:2016:CRP

[DDD16] Rebecca J. Dobbs, Christy L. Davies, and Peter M. Davies. Collaborative research partnerships inform monitoring and management of aquatic ecosystems by indigenous rangers. Reviews in Fish Biology and Fisheries, 26(4):711–725, December 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-015-9401-2.

DeLaMare:1998:TFM

[De 98] William K. De La Mare. Tidier fisheries management requires a new MOP (Management-oriented Paradigm). Reviews in Fish Biology and Fisheries, 8(3):349–356, September 1998. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1008819416162.

Danancher:2011:GPS

[DGV11] Delphine Danancher and Eva Garcia-Vazquez. Genetic population structure in flatfishes and potential impact of aquaculture and stock enhancement on wild populations in Europe. Reviews in Fish Biology and Fisheries, 21(3):441–462, September 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-011-9198-6.

Durand:2017:DBG

[DHB17] J.-D. Durand, N. Hubert, and P. Borsa. DNA barcoding grey mullets. Reviews in Fish Biology and Fisheries, 27(1):233– 243, March 2017. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-016-9457-7.

Domingues:2018:EPM

[DHG18] Rodrigo R. Domingues, Alexandre W. S. Hilsdorf, and Otto B. F. Gadig. Effects of the Pleistocene on the mitochondrial population genetic structure and demographic history of the silky shark (Carcharhinus falciformis) in the western Atlantic Ocean. Reviews in Fish Biology and Fisheries, 28(1):213–227, March 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-017-9504-z.

Diamond:2004:BQG

[Dia04] Sandra L. Diamond. Bycatch quotas in the Gulf of Mexico shrimp trawl fishery: can they work? Reviews in Fish Biology and Fisheries, 14(2):207–237, June 2004. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-004-7121-0.

Downie:2020:SPM

[DIR20] Adam T. Downie, Björn Illing, and Jodie L. Rummer. Swimming performance of marine fish larvae: review of a universal trait under ecological and environmental pressure. Reviews in Fish Biology and Fisheries, 30(1):93–108, March 2020. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-019-09592-w.

Davey:2005:SDF

[DJ05] Andrew J. H. Davey and Donald J. Jellyman. Sex determination in freshwater eels and management options for manipulation of sex. Reviews in Fish Biology and Fisheries, 15(1– 2):37–52, February 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-005-7431-x.

deJong:2020:PEA

[dJFS20] Karen de Jong, Tonje Nesse Forland, and Lise Doksæter Sivle. Predicting the effects of anthropogenic noise on fish reproduction. Reviews in Fish Biology and Fisheries, 30(2):245–268, June 2020. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-020-09598-9.

delaChesnais:2019:ERC

[dlCFP19] Thibaut de la Chesnais, Elizabeth A. Fulton, and Gretta T. Pecl. The ecological role of cephalopods and their representation in ecosystem models. Reviews in Fish Biology and Fisheries, 29 (2):313–334, June 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-019-09554-2.

Doulman:1995:PMH

[DLP95] David J. Doulman, Indrani Lutchman, and Tony J. Pitcher. Perspectives on the management of high seas fisheries: The UN conference on straddling fish stocks and highly migratory fish stocks. Reviews in Fish Biology and Fisheries, 5(1):103–119, March 1995. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF01103368.

Daryanani:2021:SCN

[DMD21] Divya S. Daryanani, Jasmin C. Martino, and Zoë A. Doubleday. Statolith chemistry: a new tool to understand the ecology and provenance of octopus. Reviews in Fish Biology and Fisheries, 31(4):923–934, December 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-021-09671-x.

Deville:2024:CBC

[DMK^{+}24] Diego Deville, Souta Mori, Kentaro Kawai, Alejandro Escánez, Armando Macali, Fedor Lishchenko, Heather Braid, Jean Githaiga-Mwicigi, Kolliyil S. Mohamed, Kathrin S. R. Bolstad, Kazutaka Miyahara, Chikatoshi Sugimoto, Fernando Á. Fernández-Álvarez, and Gustavo Sanchez. Cryptic biodiversity in the commercial diamondback squid Thysanoteuthis rhombus Troschel 1857. Reviews in Fish Biology and Fisheries, 34(1): ??, ???? 2024. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-023-09813-3.

Deters:2024:DOM

[DML^{+}24] Katherine A. Deters, Robert P. Mueller, Stephanie A. Liss, Jill M. Janak, Huidong Li, Jayson J. Martinez, Ryan A. Harnish, Jun Lu, and Zhiqun Daniel Deng. Development of optimal methods for collection, transportation, holding, handling, and tagging of juvenile American shad. Reviews in Fish Biology and Fisheries, 34(2):731–751, June 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-024-09835-5.

deMitcheson:2019:EMT

[dMlTB19] Yvonne Sadovy de Mitcheson, Allen Wai lun To, and Wing Sum Bud. Emerging from the murk: threats, challenges and opportunities for the global swim bladder trade. Reviews in Fish Biology and Fisheries, 29(4):809–835, December 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-019-09585-9.

deMoura:2020:NTA

[dMV20] Marcelo Santos de Moura and Marcelo Vianna. A new threat: assessing the main interactions between marine fish and plastic debris from a scientometric perspective. Reviews in Fish Biology and Fisheries, 30(4):623–636, December 2020. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-020-09621-z.

Delcourt:2018:IIM

[DOP18] Johann Delcourt, Michaël Ovidio, and Pascal Poncin. Individual identification and marking techniques for zebrafish. Reviews in Fish Biology and Fisheries, 28(4):839–864, December 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-018-9537-y.

Douglas:1993:FC

[Dou93] R. H. Douglas. Fish chemoreception. Reviews in Fish Biology and Fisheries, 3(2):195, June 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00045236.

dePinna:1999:SCP

[dP99] Mário C. C. de Pinna. Species concepts and phylogenetics. Reviews in Fish Biology and Fisheries, 9(4):353–373, December 1999. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1008911414399.

Delcourt:2012:SSB

[DP12] Johann Delcourt and Pascal Poncin. Shoals and schools: back to the heuristic definitions and quantitative references. Reviews in Fish Biology and Fisheries, 22(3):595–619, September 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-012-9260-z.

Affonso:2016:IES

[dPAGGB16] Igor de Paiva Affonso, Luiz Carlos Gomes, and Emili García-Berthou. Interacting effects of spatial gradients and fishing gears on characterization of fish assemblages in large reservoirs. Reviews in Fish Biology and Fisheries, 26(1):71–81, March 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-015-9402-1.

Djikanovic:2012:PFF

[DPC12] Vesna Djikanovic, Momir Paunovic, and Predrag Cakic. Parasitofauna of freshwater fishes in the Serbian open waters: a checklist of parasites of freshwater fishes in Serbian open waters. Reviews in Fish Biology and Fisheries, 22(1):297–324, March 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-011-9226-6.

Duk:2022:CIC

[DPS22] Karolina Duk, Joanna Pajdak, and Józef Szarek. Correction to: Intracohort cannibalism and methods for its mitigation in cultured freshwater fish. Reviews in Fish Biology and Fisheries, 32 (3):1017, September 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-022-09717-8. See [NPS17].

Damasio:2023:AFS

[DPV^{+}23] Ludmila M. A. Damasio, Maria Grazia Pennino, Sebastián Villasante, Adriana Rosa Carvalho, and Priscila F. M. Lopes. Adaptive factors and strategies in small-scale fisheries economies. Reviews in Fish Biology and Fisheries, 33(3):739– 750, September 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-022-09750-7.

Drill:2005:BRJ

[Dri05] Sabrina L. Drill. Book review: Jack Sobel and Craig Dahlgren (eds.), Marine Reserves: a Guide to Science, Design, and Use. Island Press, Washington, D.C., 2004. 383 pp. Reviews in Fish Biology and Fisheries, 15(1–2):159–161, February 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-005-5205-0.

Delgado-Ramirez:2023:FLW

[DROCM23] Claudia E. Delgado-Ramírez, Yoshitaka Ota, and Andrés M. Cisneros-Montemayor. Fishing as a livelihood, a way of life, or just a job: considering the complexity of “fishing communities” in research and policy. Reviews in Fish Biology and Fisheries, 33 (1):265–280, March 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-022-09721-y.

daRosa:2012:ETH

[dRRGC12] Renata da Rosa, Marceléia Rubert, and Lucia Giuliano-Caetano. Evolutionary trends in Hoplerythrinus unitaeniatus (Agassiz 1829) (Characiformes, Erythrinidae). Reviews in Fish Biology and Fisheries, 22(2):467–475, June 2012. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-011-9237-3.

Dolomatov:2011:FNM

[DSK11] S. I. Dolomatov, P. V. Shekk, and M. I. Kryukova. Features of nitrogen metabolism in fishes. Reviews in Fish Biology and Fisheries, 21(4):733–737, December 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-011-9212-z.

daSilva:2016:DSS

[dSKV16] Rodrigo Fortes da Silva, Alexandre Kitagawa, and Francisco Javier Sánchez Vázquez. Dietary self-selection in fish: a new approach to studying fish nutrition and feeding behavior. Reviews in Fish Biology and Fisheries, 26(1):39–51, March 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-015-9410-1.

daSilva:2012:CCO

[dSPPM12] E. Lourenço da Silva, D. Piscor, and P. P. Parise-Maltempi. Cytogenetic characterisation of the ornamental freshwater fish, Piabucus melanostomus (Iguanodectinae) from Brazilian wetlands and its relation with species of Characidae basal group. Reviews in Fish Biology and Fisheries, 22(2):477–484, June 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-011-9238-2.

Schmidt:2021:ABG

[dSSHK21] Thassya C. dos Santos Schmidt, Doug E. Hay, and Olav S. Kjesbu. Adult body growth and reproductive investment vary markedly within and across Atlantic and Pacific herring: a metaanalysis and review of 26 stocks. Reviews in Fish Biology and Fisheries, 31(3):685–708, September 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-021-09665-9.

Dankel:2008:FMP

[DSU08] Dorothy J. Dankel, Dankert W. Skagen, and Øyvind Ulltang. Fisheries management in practice: review of 13 commercially important fish stocks. Reviews in Fish Biology and Fisheries, 18(2):201–233, May 2008. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-007-9068-4.

Dempster:2004:FAD

[DT04] Tim Dempster and Marc Taquet. Fish aggregation device (FAD) research: gaps in current knowledge and future directions for ecological studies. Reviews in Fish Biology and Fisheries, 14 (1):21–42, March 2004. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-004-3151-x.

Ducatez:2019:WSA

[Duc19] Simon Ducatez. Which sharks attract research? Analyses of the distribution of research effort in sharks reveal significant non-random knowledge biases. Reviews in Fish Biology and Fisheries, 29(2):355–367, June 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-019-09556-0.

Dugatkin:1993:FBP

[DW93] Lee Alan Dugatkin and David Sloan Wilson. Fish behaviour, partner choice experiments and cognitive ethology. Reviews in Fish Biology and Fisheries, 3(4):368–372, December 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00043386.

Deane:2009:MFG

[DW09] Eddie E. Deane and Norman Y. S. Woo. Modulation of fish growth hormone levels by salinity, temperature, pollutants and aquaculture related stress: a review. Reviews in Fish Biology and Fisheries, 19(1):97–120, March 2009. CODEN RFBFEA.

ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-008-9091-0.

Deane:2011:APR

[DW11] Eddie E. Deane and Norman Y. S. Woo. Advances and perspectives on the regulation and expression of piscine heat shock proteins. Reviews in Fish Biology and Fisheries, 21(2):153– 185, June 2011. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-010-9164-8.

Deng:2022:LNC

[DZZZ22] Qiuxia Deng, Na Zhao, Chunhua Zhu, and Bo Zhang. Long noncoding RNAs in the physiology of aquaculture animals: a perspective update. Reviews in Fish Biology and Fisheries, 32(4): 1103–1122, December 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-022-09734-7.

Englmaier:2022:GPS

[EAW22] Gernot K. Englmaier, Alexander Antonov, and Steven J. Weiss. General patterns of sexual dimorphism in graylings (Thymallus), with a comparison to other salmonid species. Reviews in Fish Biology and Fisheries, 32(2):645–667, June 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-021-09694-4.

Eales:1993:MRT

[EB93] J. G. Eales and S. B. Brown. Measurement and regulation of thyroidal status in teleost fish. Reviews in Fish Biology and Fisheries, 3(4):299–347, December 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/BF00043383.

Espino-Barr:2002:CTF

[EBRLGB02] Elaine Espino-Barr, Arturo Ruiz-Luna, and Arturo Garcia-Boa. Changes in tropical fish assemblages associated with small-scale fisheries: a case study in the Pacific off central Mexico. Reviews in Fish Biology and Fisheries, 12(4):393–401, December 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1025355102004.

Erisman:2020:MAR

[EBS20] Brad E. Erisman, Derek G. Bolser, and Alexander E. Sacco. A meta-analytical review of the effects of environmental and ecological drivers on the abundance of red snapper (Lutjanus campechanus) in the U.S. Gulf of Mexico. Reviews in Fish Biology and Fisheries, 30(3):437–462, September 2020. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-020-09608-w.

Elia:2010:SVD

[EDP10] Antonia Concetta Elia, Ambrosius Josef Martin Dörr, and Marino Prearo. Seasonal variability of detoxificant response and heavy metal accumulation in tissues of both sexes in Tinca tinca (L.) from Lake Trasimeno. Reviews in Fish Biology and Fisheries, 20(3):425–434, September 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-009-9143-0.

Escobar:2018:IDF

[EEDP18] Luis E. Escobar, Joaquin Escobar-Dodero, and Nicholas B. D. Phelps. Infectious disease in fish: global risk of viral hemorrhagic septicemia virus. Reviews in Fish Biology and Fisheries, 28(3): 637–655, September 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-018-9524-3.

Elsdon:2003:RMP

[EG03] Travis S. Elsdon and Bronwyn M. Gillanders. Reconstructing migratory patterns of fish based on environmental influences on otolith chemistry. Reviews in Fish Biology and Fisheries, 13(3): 217–235, September 2003. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/B:RFBF.0000033071.73952.40.

Edwards:2002:CSF

[EGMM02] Robert J. Edwards, Gary P. Garrett, and Edie Marsh-Matthews. Conservation and status of the fish communities inhabiting the Río Conchos basin and middle Rio Grande, México and U.S.A. Reviews in Fish Biology and Fisheries, 12(2–3):119–132, June 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1025098229262.

Evans:2018:ELP

[EJD18] Thomas M. Evans, Philippe Janvier, and Margaret F. Docker. The evolution of lamprey (Petromyzontida) life history and the origin of metamorphosis. Reviews in Fish Biology and Fisheries, 28(4):825–838, December 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-018-9536-z.

Eggertsen:2024:CRF

[ELC^{+}24] Linda Eggertsen, André L. Luza, César A. M. M. Cordeiro, Cristian Dambros, Carlos E. L. Ferreira, Sergio R. Floeter, Ronaldo B. Francini-Filho, Kátia M. F. Freire, Maria A. Gasalla, Tommaso Giarrizzo, Vinicius J. Giglio, Natalia Hanazaki, Priscila F. M. Lopes, Guilherme O. Longo, Osmar J. Luiz, Rafael A. Magris, Thiago C. Mendes, Hudson T. Pinheiro, Juan P. Quimbayo, José Amorim Reis-Filho, Daniele A. Vila-Nova, and Mariana G. Bender. Complexities of reef fisheries in Brazil: a retrospective and functional approach. Reviews in Fish Biology and Fisheries, 34(1):??, ???? 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-023-09826-y.

Elskus:2004:BRJ

[Els04] Adria A. Elskus. Book review: A. J. Lawrence and K. L. Hemingway (eds). Effects of Pollution on Fish: Molecular Effects and Population Responses. Blackwell Publishing, Oxford 2003. 342pp. Reviews in Fish Biology and Fisheries, 14(2):295– 296, June 2004. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-004-7353-z.

Ekstrom:1997:POT

[EM97] Peter Ekström and Hilmar Meissl. The pineal organ of teleost fishes. Reviews in Fish Biology and Fisheries, 7(2):199–284, June 1997. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1023/A:1018483627058.

Epifanio:2000:RHD

[EN00] John Epifanio and Jennifer Nielsen. The role of hybridization in the distribution, conservation and management of aquatic species. Reviews in Fish Biology and Fisheries, 10(3):245– 251, September 2000. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1016729132297.

Elsler:2023:CSS

[ENF^{+}23] Laura G. Elsler, Muhammad Neil, Sebastian Ferse, Gabriela Navarrete Forero, Marion Glaser, and Maja Schlüter. Compliance in small-scale fisheries is linked to fisher–trader relations: not fishers alone (Southeast Asian case study). Reviews in Fish Biology and Fisheries, 33(3):751–766, September 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09783-6.

Emery:2018:UEM

[ENP18] Timothy J. Emery, Rocio Noriega, and Thomas Peatman. The use of electronic monitoring within tuna longline fisheries: implications for international data collection, analysis and reporting. Reviews in Fish Biology and Fisheries, 28(4):887–907, December 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-018-9533-2.

Epifanio:2000:SEP

[EP00] John Epifanio and David Philipp. Simulating the extinction of parental lineages from introgressive hybridization: the effects of fitness, initial proportions of parental taxa, and mate choice. Reviews in Fish Biology and Fisheries, 10(3):339–354, September 2000. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1016673331459.

Esteve:2005:OSB

[Est05] Manu Esteve. Observations of spawning behaviour in salmoninae: Salmo, Oncorhynchus and Salvelinus. Reviews in Fish Biology and Fisheries, 15(1–2):1–21, February 2005. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-005-7434-7.

Espinoza-Tenorio:2012:WMS

[ETWE12] Alejandro Espinoza-Tenorio, Matthias Wolff, and Ileana Espejel. What model suits ecosystem-based fisheries management? A plea for a structured modeling process. Reviews in Fish Biology and Fisheries, 22(1):81–94, March 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-011-9224-8.

Farmery:2022:FAD

[FAN22] A. K. Farmery, K. Alexander, and B. Nowak. Food for all: designing sustainable and secure future seafood systems. Reviews in Fish Biology and Fisheries, 32(1):101–121, March 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-021-09663-x.

Fernandez-Alvarez:2018:JSS

[FÁVG18] Fernando Á. Fernández-Álvarez, Roger Villanueva, and William F. Gilly. The journey of squid sperm. Reviews in Fish Biology and Fisheries, 28(1):191–199, March 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-017-9498-6.

Fujita:2005:RBF

[FB05] Rod Fujita and Kate Bonzon. Rights-based fisheries management: an environmentalist perspective. Reviews in Fish Biology and Fisheries, 15(3):309–312, August 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-005-4867-y.

Fonseca:2007:FEG

[FC07] V. F. Fonseca and H. N. Cabral. Are fish early growth and condition patterns related to life-history strategies? Reviews in Fish Biology and Fisheries, 17(4):545–564, November 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-007-9054-x.

Fernandez-Carvalho:2015:PAD

[FCCS15] Joana Fernandez-Carvalho, Rui Coelho, and Miguel N. Santos. Pan-Atlantic distribution patterns and reproductive biology of the bigeye thresher, Alopias superciliosus. Reviews in Fish Biology and Fisheries, 25(3):551–568, September 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-015-9389-7.

Frisch:2016:EKA

[FCH16a] Ashley J. Frisch, Darren S. Cameron, and Jean-Paul A. Hobbs. Erratum to: Key aspects of the biology, fisheries and management of Coral grouper. Reviews in Fish Biology and Fisheries, 26(3):327, September 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-016-9428-z. See [FCH16b].

Frisch:2016:KAB

[FCH16b] Ashley J. Frisch, Darren S. Cameron, and Jean-Paul A. Hobbs. Key aspects of the biology, fisheries and management of coral grouper. Reviews in Fish Biology and Fisheries, 26(3):303–325, September 2016. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-016-9427-0. See erratum [FCH16a].

Fernandez-Corredor:2024:OER

[FCOCNC24] Elena Fernández-Corredor, Jazel Ouled-Cheikh, Joan Navarro, and Marta Coll. An overview of the ecological roles of Mediterranean chondrichthyans through extinction scenarios. Reviews in Fish Biology and Fisheries, 34(1):??, ???? 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09822-2.

Fogarty:2019:PCA

[FCP19] Hannah E. Fogarty, Christopher Cvitanovic, and Gretta T. Pecl. Prepared for change? An assessment of the current state of knowledge to support climate adaptation for Australian fisheries. Reviews in Fish Biology and Fisheries, 29(4):877–894, December 2019. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-019-09579-7.

Freon:2000:RFA

[FD00] Pierre Fréon and Laurent Dagorn. Review of fish associative behaviour: Toward a generalisation of the meeting point hypothesis. Reviews in Fish Biology and Fisheries, 10(2):183–207, June 2000. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1016666108540.

Frost:2022:ESG

[FD22] M. Frost and K. Diele. Essential spawning grounds of Scottish herring: current knowledge and future challenges. Reviews in Fish Biology and Fisheries, 32(3):721–744, September 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-022-09703-0.

Ferguson:1994:MGF

[Fer94a] Andrew Ferguson. Molecular genetics in fisheries: current and future perspectives. Reviews in Fish Biology and Fisheries, 4(3): 379–383, September 1994. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF00042911.

Ferguson:1994:RMG

[Fer94b] M. Ferguson. The role of molecular genetic markers in the management of cultured fishes. Reviews in Fish Biology and Fisheries, 4(3):351–373, September 1994. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00042909.

Flajshans:2010:RPT

[FGL10] Martin Flajšhans, David Gela, and Otomar Linhart. A review on the potential of triploid tench for aquaculture. Reviews in Fish Biology and Fisheries, 20(3):317–329, September 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-009-9144-z.

Feng:2004:DMI

[FHK04] Yang Yu Feng, Li Chun Hou, and Chung Ik Kyo. Development of mariculture and its impacts in Chinese coastal waters. Reviews in Fish Biology and Fisheries, 14(1):1–10, March 2004. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-004-3539-7.

Frusher:2014:SHR

[FHvP14] Stewart D. Frusher, Alistair J. Hobday, and E. Ingrid van Putten. The short history of research in a marine climate change hotspot: from anecdote to adaptation in south-east Australia.

Reviews in Fish Biology and Fisheries, 24(2):593–611, June 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-013-9325-7.

Fang:2024:EFI

[FHW^{+}24] Zhou Fang, Peiwu Han, Yan Wang, Jianhua Li, Guanyu Hu, Bilin Liu, and Xinjun Chen. Environmental fluctuation influences the ontogenetic dispersal and distribution of two ommastrephid squids in the Pacific Ocean. Reviews in Fish Biology and Fisheries, 34(1):??, ???? 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-023-09815-1.

Foss:2004:RCP

[FIØ04] Atle Foss, Albert K. Imsland, and Victor Øiestad. A review of the culture potential of spotted wolffish Anarhichas minor Olafsen. Reviews in Fish Biology and Fisheries, 14(2):277–294, June 2004. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-004-8360-9.

Fitzgerald:1993:SRT

[Fit93] Gerard J. Fitzgerald. Seeing red, turning red. Reviews in Fish Biology and Fisheries, 3(3):286–292, September 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00043931.

Fleming:1996:RSA

[Fle96] Ian A. Fleming. Reproductive strategies of Atlantic salmon: ecology and evolution. Reviews in Fish Biology and Fisheries, 6(4):379–416, December 1996. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00164323.

Fuiman:1994:DPD

[FM94] Lee A. Fuiman and Anne E. Magurran. Development of predator defences in fishes. Reviews in Fish Biology and Fisheries, 4 (2):145–183, June 1994. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF00044127.

Ficke:2007:PIG

[FMH07] Ashley D. Ficke, Christopher A. Myrick, and Lara J. Hansen. Potential impacts of global climate change on freshwater fisheries. Reviews in Fish Biology and Fisheries, 17(4):581–613, November 2007. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-007-9059-5.

Fischer:2022:EHG

[FMM22] Mibu Fischer, Kimberley Maxwell, and Kaisu Mustonen. Empowering her guardians to nurture our ocean’s future. Reviews in Fish Biology and Fisheries, 32(1):271–296, March 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-021-09679-3.

Freedman:2002:WTN

[FN02] Jonathan A. Freedman and David L. G. Noakes. Why are there no really big bony fishes? A point-of-view on maximum body size in teleosts and elasmobranchs. Reviews in Fish Biology and Fisheries, 12(4):403–416, December 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1023/A:1025365210414.

Forey:1995:ARF

[For95] P. L. Forey. Agnathans recent and fossil, and the origin of jawed vertebrates. Reviews in Fish Biology and Fisheries, 5(3): 267–303, September 1995. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF00043003.

Foster:2008:BRC

[Fos08] Susan Foster. Book review: Culum Brown, Kevin Laland and Jens Krause, Fish Cognition and Behavior. Reviews in Fish Biology and Fisheries, 18(1):131–132, February 2008. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-007-9069-3.

Field:2023:JHP

[FQSJ23] Kate A. Field, Micah J. Quindazzi, Alejandro Schmill, and Francis Juanes. Book review: Jeffery A. Hutchings: A primer of life histories: ecology, evolution, and application. Reviews in Fish Biology and Fisheries, 33(4):1569–1571, December 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09774-7.

Frisch:2004:SCG

[Fri04] Ashley Frisch. Sex-change and gonadal steroids in sequentiallyhermaphroditic teleost fish. Reviews in Fish Biology and Fisheries, 14(4):481–499, December 2004. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-005-3586-8.

Froese:1999:GBU

[Fro99] Rainer Froese. The good, the bad, and the ugly: a critical look at species and their institutions from a user’s perspective. Reviews in Fish Biology and Fisheries, 9(4):375–378, December 1999. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1008915531238.

Freon:2014:HFV

[FSB14] Pierre Fréon, Juan Carlos Sueiro, and Marilu Bouchon. Harvesting for food versus feed: a review of Peruvian fisheries in a global context. Reviews in Fish Biology and Fisheries, 24 (1):381–398, March 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-013-9336-4.

Ferreira:2017:GSD

[FSSM17] Dhiego Gomes Ferreira, Lenice Souza-Shibatta, and Maristela Cavicchioli Makrakis. Genetic structure and diversity of migratory freshwater fish in a fragmented Neotropical river system. Reviews in Fish Biology and Fisheries, 27(1):209–231, March 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-016-9441-2.

Fredou:2015:SCB

[FTH15] Flávia Lucena Frédou, Mariana Travassos Tolotti, and Fábio Hissa Vieira Hazin. Sharks caught by the Brazilian tuna longline fleet: an overview. Reviews in Fish Biology and Fisheries, 25 (2):365–377, June 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-014-9380-8.

Fuiman:1999:FAB

[Fui99] Lee A. Fuiman. Fishes of Antarctica: a biological overciew. Reviews in Fish Biology and Fisheries, 9(3):272–273, September 1999. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1008920317812.

Gallardo:2010:EDC

[GAD10] José Martín Gallardo, Sayyed Mohammad Hadi Alavi, and Bořek Drozd. External damage and changes in blood parameters in female tench, Tinca tinca (L.) retained in anglers’ keepnets. Reviews in Fish Biology and Fisheries, 20(3):403–408, September 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-009-9149-7.

Gaughan:2001:DTA

[Gau01] Daniel J. Gaughan. Disease-translocation across geographic boundaries must be recognized as a risk even in the absence of disease identification: the case with Australian Sardinops. Reviews in Fish Biology and Fisheries, 11(2):113–123, June 2001. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1015255900836.

Gillanders:2023:WFC

[GB23] Bronwyn M. Gillanders and Gavin A. Begg. The 8th World Fisheries Congress: sharing our oceans and rivers, a vision for the world’s fisheries. Reviews in Fish Biology and Fisheries, 33 (2):311–315, June 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-023-09775-6.

Garcia-Berthou:2012:TNS

[GBCCO12] Emili García-Berthou, Gerard Carmona-Catot, and Derek H. Ogle. A technical note on seasonal growth models. Reviews in Fish Biology and Fisheries, 22(3):635–640, September 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-012-9262-x.

Grabowska:2023:FEF

[GBOK23] Joanna Grabowska, Dagmara Błońska, Markéta Ondračková, and Tomasz Kakareko. The functional ecology of four invasive Ponto–Caspian gobies. Reviews in Fish Biology and Fisheries, 33(4):1329–1352, December 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-023-09801-7.

Gillson:2022:RMS

[GBR22] Jonathan P. Gillson, Tea Bašić, and Ian C. Russell. A review of marine stressors impacting Atlantic salmon Salmo salar, with an assessment of the major threats to English stocks. Reviews in Fish Biology and Fisheries, 32(3):879–919, September 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-022-09714-x.

Gilman:2023:ASE

[GC23] Eric Gilman and Milani Chaloupka. Applying a sequential evidence hierarchy, with caveats, to support prudent fisheries bycatch policy. Reviews in Fish Biology and Fisheries, 33(1): 137–146, March 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-022-09745-4.

Gilman:2024:IFB

[GCB^{+}24] Eric Gilman, Milani Chaloupka, Lyall Bellquist, Heather Bowlby, and Nathan Taylor. Individual and fleetwide bycatch thresholds in regional fisheries management frameworks. Reviews in Fish Biology and Fisheries, 34(1):??, ???? 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09811-5.

Gilman:2021:TLM

[GCK21] Eric Gilman, Milani Chaloupka, and Eric Kingma. Tori lines mitigate seabird bycatch in a pelagic longline fishery. Reviews in Fish Biology and Fisheries, 31(3):653–666, September 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-021-09659-7.

Gilman:2018:EPL

[GCM18] Eric Gilman, Milani Chaloupka, and Michael Musyl. Effects of pelagic longline hook size on species- and size-selectivity and survival. Reviews in Fish Biology and Fisheries, 28(2):417– 433, June 2018. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-017-9509-7.

Gomberg:2003:BRL

[GCO03] Matthew Gomberg, Peter Clarke, and Jeff Ojala. Book review: Lee A. Fuiman and Robert G. Werner (eds.), Fisheries Science: The Unique Contributions of Early Life Stages. Reviews in Fish Biology and Fisheries, 13(3):365–366, September 2003. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/B:RFBF.0000033137.15465.9d.

Gilman:2019:RPP

[GCS19] Eric Gilman, Milani Chaloupka, and Petri Suuronen. Robbing Peter to pay Paul: replacing unintended cross-taxa conflicts with intentional tradeoffs by moving from piecemeal to integrated fisheries bycatch management. Reviews in Fish Biology and Fisheries, 29(1):93–123, March 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-019-09547-1.

Gilman:2020:EPL

[GCS20] Eric Gilman, Milani Chaloupka, and Liming Song. Effect of pelagic longline bait type on species selectivity: a global synthesis of evidence. Reviews in Fish Biology and Fisheries, 30(3): 535–551, September 2020. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-020-09612-0.

Gonzalez:2009:AIC

[GCSR09] Rocío González, Jesus D. Celada, and María Sáez-Royuela. The artificial incubation of crayfish eggs: review and report from an experimental study concerning the effects of offspring origin (maternal or artificial incubation) on the survival and growth of juvenile signal crayfish (Pacifastacus leniusculus, Astacidae). Reviews in Fish Biology and Fisheries, 19(2):167–176, June 2009. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-008-9095-9.

Gray:2022:SME

[GCUR22] Charles A. Gray, M. G. Chapman, A. J. Underwood, and Douglas Rotherham. Spatial management of estuarine fisheries resources: Do recreational-only fished estuaries provide conservation to harvested species? Reviews in Fish Biology and Fisheries, 32(4):1123–1140, December 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-022-09723-w.

Ganguly:2013:SPF

[GDC13] Subha Ganguly, Krushna Chandra Dora, and Supratim Chowdhury. Supplementation of prebiotics in fish feed: a review. Reviews in Fish Biology and Fisheries, 23(2):195–199, June 2013. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-012-9291-5.

Gouveia:2013:CKE

[GdMD13] Juceli Gonzalez Gouveia, Vivian Patrícia Oliveira de Moraes, and Ana Lúcia Dias. Considerations on karyotype evolution in the genera Imparfinis Eigenmann and Norris 1900 and Pimelodella Eigenmann and Eigenmann 1888 (Siluriformes: Heptapteridae). Reviews in Fish Biology and Fisheries, 23 (2):215–227, June 2013. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-012-9286-2.

Guerrero-Estevez:2016:EED

[GELL16] S. M. Guerrero-Estévez and E. López-López. Effects of endocrine disruptors on reproduction in viviparous teleosts with intraluminal gestation. Reviews in Fish Biology and Fisheries, 26(3): 563–587, September 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-016-9443-0.

Green:2014:EEF

[GGL14] Bridget S. Green, Caleb Gardner, and Adrian Linnane. Environmental effects on fished lobsters and crabs. Reviews in Fish Biology and Fisheries, 24(2):613–638, June 2014. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-014-9350-1.

Gasco:2010:MSP

[GGZ10] Laura Gasco, Francesco Gai, and Ivo Zoccarato. Morphometry, slaughtering performances, chemical and fatty acid composition of the protected designation of origin “Golden hump tench of Poirino highland” product. Reviews in Fish Biology and Fisheries, 20(3):357–365, September 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-009-9141-2.

Gutsch:2016:RRG

[GH16] Michelle Gutsch and Joel Hoffman. A review of ruffe (Gymnocephalus cernua) life history in its native versus non-native range. Reviews in Fish Biology and Fisheries, 26(2):213– 233, June 2016. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-016-9422-5.

Gilman:2017:REP

[GH17] Eric Gilman and Hsiang-Wen Huang. Review of effects of pelagic longline hook and bait type on sea turtle catch rate, anatomical hooking position and at-vessel mortality rate. Reviews in Fish Biology and Fisheries, 27(1):43–52, March 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-016-9447-9.

Giske:1998:MSD

[GHF98] Jarl Giske, Geir Huse, and Oyvind Fiksen. Modelling spatial dynamics of fish. Reviews in Fish Biology and Fisheries, 8 (1):57–91, March 1998. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1008864517488.

Gilman:2022:DST

[GHMG22] Eric Gilman, Martin Hall, and E. J. Milner-Gulland. A decision support tool for integrated fisheries bycatch management. Reviews in Fish Biology and Fisheries, 32(2):441–472, June 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-021-09693-5.

Gibson:1993:ASF

[Gib93] R. J. Gibson. The Atlantic salmon in fresh water: spawning, rearing and production. Reviews in Fish Biology and Fisheries, 3(1):39–73, March 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF00043297.

Gilbert:1993:BRC

[Gil93a] Carter R. Gilbert. Book review: Common and scientific names of fishes from the United States and Canada, 5th edition. Reviews in Fish Biology and Fisheries, 3(2):192–194, June 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00045235.

Gilbert:1993:WFI

[Gil93b] Carter R. Gilbert. World fishes important to North Americans, exclusive of species from the continental waters of the United States and Canada. Reviews in Fish Biology and Fisheries, 3 (2):191–192, June 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF00045234.

Guiro:2009:SES

[GIT09] A. Guiro, A. Iggidr, and H. Touré. On the stock estimation for some fishery systems. Reviews in Fish Biology and Fisheries, 19(3):313–327, September 2009. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-009-9104-7.

Grist:2007:HLW

[GJ07] Eric P. M. Grist and George D. Jackson. How long would it take to become a giant squid? Reviews in Fish Biology and Fisheries, 17(2–3):385–399, August 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-007-9046-x.

Grant:2017:RGA

[GJA17] W. Stewart Grant, James Jasper, and Milo Adkison. Responsible genetic approach to stock restoration, sea ranching and stock enhancement of marine fishes and invertebrates. Reviews in Fish Biology and Fisheries, 27(3):615–649, September 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-017-9489-7.

Gallagher:2018:BCS

[GK18a] Austin J. Gallagher and A. Peter Klimley. The biology and conservation status of the large hammerhead shark complex: the great, scalloped, and smooth hammerheads. Reviews in Fish Biology and Fisheries, 28(4):777–794, December 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-018-9530-5.

Gray:2018:BET

[GK18b] Charles A. Gray and Steven J. Kennelly. Bycatches of endangered, threatened and protected species in marine fisheries. Reviews in Fish Biology and Fisheries, 28(3):521–541, September 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-018-9520-7.

Grant:2019:CUP

[GKC19] Michael I. Grant, Peter M. Kyne, and Andrew Chin. Categorising use patterns of non-marine environments by elasmobranchs and a review of their extinction risk. Reviews in Fish Biology and Fisheries, 29(3):689–710, September 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-019-09576-w.

Gela:2010:CPG

[GKR10] David Gela, Martin Kocour, and Marek Rodina. Comparison of performance of genome manipulated and standard tench, Tinca tinca (L.), groups under pond management conditions. Reviews in Fish Biology and Fisheries, 20(3):301–306, September 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-009-9142-1.

Glass:2000:BRD

[Gla00] Chris Glass. Book review: Dynamics of Pelagic Fish Distribution and Behaviour: Effects on Fisheries and Stock Assessment. Pierre Fréeon and Ole Arve Misund. Reviews in Fish Biology and Fisheries, 10(1):124, March 2000. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1023/A:1008928315202.

Guenette:1998:MRB

[GLC98] Sylvie Guénette, Tim Lauck, and Colin Clark. Marine reserves: from Beverton and Holt to the present. Reviews in Fish Biology and Fisheries, 8(3):251–272, September 1998. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1008859130275.

Gil:2016:ESC

[GLC16] J. Gil, J. Labonne, and A. Caudron. Evaluation of strategies to conserve and restore intraspecific biodiversity of brown trout: outcomes from genetic monitoring in the French Alps. Reviews in Fish Biology and Fisheries, 26(1):1–11, March 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-015-9405-y.

Gamez:2012:SEE

[GLG12] M. Gámez, I. López, and J. Garay. Stock estimation, environmental monitoring and equilibrium control of a fish population with reserve area. Reviews in Fish Biology and Fisheries, 22(3): 751–766, September 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-012-9253-y.

Giareta:2024:CUB

[GLHD^{+}24a] Eloisa Pinheiro Giareta, Renata Daldin Leite, Rachel Ann Hauser-Davis, Ana Paula Chaves, Patricia Charvet, and Natascha Wosnick. Correction: Unveiling the batoid plight: insights from global stranding data and future directions. Reviews in Fish Biology and Fisheries, 34(2):669, June 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-024-09841-7. See [GLHD^{+}24b].

Giareta:2024:UBP

[GLHD^{+}24b] Eloisa Pinheiro Giareta, Renata Daldin Leite, Rachel Ann Hauser-Davis, Ana Paula Chaves, Patricia Charvet, and Natascha Wosnick. Unveiling the batoid plight: insights from global stranding data and future directions. Reviews in Fish Biology and Fisheries, 34(2):647–667, June 2024. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.

1007/s11160-024-09837-3. See correction [GLHD^{+}24a].

Griffiths:2020:LTT

[GLM20] Shane P. Griffiths, Duncan Leadbitter, and Mohammed Moazzam. Longtail tuna, Thunnus tonggol (Bleeker, 1851): a global review of population dynamics, ecology, fisheries, and considerations for future conservation and management. Reviews in Fish Biology and Fisheries, 30(1):25–66, March 2020. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-019-09589-5.

Giakoumi:2012:AMC

[GMP12] Sylvaine Giakoumi, Tessa Mazor, and Workshop Participants. Advancing marine conservation planning in the Mediterranean Sea. Reviews in Fish Biology and Fisheries, 22(4):943–949, December 2012. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-012-9272-8.

Gerber:2017:SSP

[GMS17] Kayla M. Gerber, Martha E. Mather, and Joseph M. Smith. A suite of standard post-tagging evaluation metrics can help assess tag retention for field-based fish telemetry research. Reviews in Fish Biology and Fisheries, 27(3):651–664, September 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-017-9484-z.

Goethel:2023:OPC

[GOP^{+}23] Daniel R. Goethel, Kristen L. Omori, André E. Punt, Patrick D. Lynch, Aaron M. Berger, Carryn L. de Moor, Éva E. Plagányi, Jason M. Cope, Natalie A. Dowling, Richard McGarvey, Ann L. Preece, James T. Thorson, Milani Chaloupka, Sarah Gaichas, Eric Gilman, Sybrand A. Hesp, Catherine Longo, Nan Yao, and Richard D. Methot. Oceans of plenty? Challenges, advancements, and future directions for the provision of evidencebased fisheries management advice. Reviews in Fish Biology and Fisheries, 33(2):375–410, June 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-022-09726-7.

Gould:2016:CTR

[Gou16] Jackie Gould. Caught in the tides: the (re)development of a trepang (sea cucumber, Holothuria scabra) industry at Warruwi, Northern Territory. Reviews in Fish Biology and Fisheries, 26(4):617–628, December 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-015-9400-3.

Griffiths:2013:CWW

[GOW13] Shane P. Griffiths, Robert J. Olson, and George M. Watters. Complex wasp-waist regulation of pelagic ecosystems in the Pacific Ocean. Reviews in Fish Biology and Fisheries, 23(4): 459–475, December 2013. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-012-9301-7.

Ganguly:2012:MFD

[GP12] S. Ganguly and A. Prasad. Microflora in fish digestive tract plays significant role in digestion and metabolism. Reviews in Fish Biology and Fisheries, 22(1):11–16, March 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-011-9214-x.

Grabowska:2015:LHT

[GP15] Joanna Grabowska and Mirosław Przybylski. Life-history traits of non-native freshwater fish invaders differentiate them from natives in the Central European bioregion. Reviews in Fish Biology and Fisheries, 25(1):165–178, March 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-014-9375-5.

Gruss:2018:MPU

[GPS18] Arnaud Grüss, Holly A. Perryman, and Theodore S. Switzer. Monitoring programs of the U.S. Gulf of Mexico: inventory, development and use of a large monitoring database to map fish and invertebrate spatial distributions. Reviews in Fish Biology and Fisheries, 28(4):667–691, December 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-018-9525-2.

Grafton:1996:ITQ

[Gra96] R. Quentin Grafton. Individual transferable quotas: theory and practice. Reviews in Fish Biology and Fisheries, 6(1):5– 20, March 1996. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/BF00058517.

Greenwood:1993:CGR

[Gre93a] P. H. Greenwood. Catalog of the genera of recent fishes. Reviews in Fish Biology and Fisheries, 3(2):187–188, June 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00045232.

Greenwood:1993:SHE

[Gre93b] P. H. Greenwood. Systematics, historical ecology and North American freshwater fishes. Reviews in Fish Biology and Fisheries, 3(4):373–375, December 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00043387.

Grunow:2023:SAO

[GS23] Bianka Grunow and Sebastian M. Strauch. Status assessment and opportunities for improving fish welfare in animal experimental research according to the 3R-Guidelines. Reviews in Fish Biology and Fisheries, 33(4):1075–1093, December 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09781-8.

Guillotreau:2017:LRG

[GSC17] Patrice Guillotreau, Dale Squires, and Guillermo A. Compeán. Local, regional and global markets: what drives the tuna fisheries? Reviews in Fish Biology and Fisheries, 27(4):909–929, December 2017. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-016-9456-8.

Gladyshev:2018:MAF

[GSD18] Michail I. Gladyshev, Nadezhda N. Sushchik, and Yury Yu Dgebuadze. Meta-analysis of factors associated with omega-3 fatty acid contents of wild fish. Reviews in Fish Biology and Fisheries, 28(2):277–299, June 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-017-9511-0.

Gamperl:1994:ECS

[GVB94] A. K. Gamperl, M. M. Vijayan, and R. G. Boutilier. Experimental control of stress hormone levels in fishes: techniques and applications. Reviews in Fish Biology and Fisheries, 4 (2):215–255, June 1994. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF00044129.

Griffiths:2010:EEL

[GYH10] Shane P. Griffiths, Jock W. Young, and Michael G. Hinton. Ecological effects of longline fishing and climate change on the pelagic ecosystem off eastern Australia. Reviews in Fish Biology and Fisheries, 20(2):239–272, June 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-009-9157-7.

Griffiths:2013:ENS

[GZT13] Shane P. Griffiths, Mitchell T. Zischke, and Sharon Tickell. Efficacy of novel sampling approaches for surveying specialised recreational fisheries. Reviews in Fish Biology and Fisheries, 23(3):395–413, September 2013. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-012-9299-x.

Harder:2018:TDF

[HAC18] Avril M. Harder, William R. Ardren, and Mark R. Christie. Thiamine deficiency in fishes: causes, consequences, and potential solutions. Reviews in Fish Biology and Fisheries, 28(4): 865–886, December 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-018-9538-x.

Hall:1996:B

[Hal96] Martin A. Hall. On bycatches. Reviews in Fish Biology and Fisheries, 6(3):319–352, September 1996. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/BF00122585.

Hall:1998:EVT

[Hal98] Martin A. Hall. An ecological view of the tuna–dolphin problem: impacts and trade-offs. Reviews in Fish Biology and Fisheries, 8(1):1–34, March 1998. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1008854816580.

Hannesson:1996:IES

[Han96] Rögnvaldur Hannesson. On ITQs: an essay for the special issue of reviews in fish biology and fisheries. Reviews in Fish Biology and Fisheries, 6(1):91–96, March 1996. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/BF00058521.

Hannesson:2005:RBF

[Han05] Rögnvaldur Hannesson. Rights based fishing: Use rights versus property rights to fish. Reviews in Fish Biology and Fisheries, 15 (3):231–241, August 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-005-4870-3.

Hara:1994:DCS

[Har94] Toshiaki J. Hara. The diversity of chemical stimulation in fish olfaction and gustation. Reviews in Fish Biology and Fisheries, 4(1):1–35, March 1994. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF00043259.

Hart:1998:BRG

[Har98] Paul J. B. Hart. Book review: Global Trends: Fisheries Management (AFS Symposium 20) E. K. Pikitch, D. D. Huppert and M. P. Sissenwine (eds). Reviews in Fish Biology and Fisheries, 8(4):494–495, December 1998. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1023/A:1008868522838.

Harper:1999:BRW

[Har99a] David Harper. Book review: Watershed Restoration: Principles and Practices. Edited by J. E. Williams, C. A. Wood and M. P. Dombeck. Reviews in Fish Biology and Fisheries, 9(1): 122–123, March 1999. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1008809732280.

Hart:1999:BRR

[Har99b] Paul J. B. Hart. Book review: Recreational Fisheries: Social, Economic and Management Aspects (Papers from the EIFAC Symposium on Recreational Fisheries, held in Dublin, Ireland, from 11 to 14 June 1997). Reviews in Fish Biology and Fisheries, 9(3):273–274, September 1999. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1023/A:1008972301882.

Hart:1999:RNF

[Har99c] Paul J. B. Hart. Review of Northeast fishery stock assessments. Reviews in Fish Biology and Fisheries, 9(3):270–271, September 1999. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1008916216903.

Harlioglu:2011:PSF

[Har11] Ayşe Gül Harlioğlu. Present status of fisheries in Turkey. Reviews in Fish Biology and Fisheries, 21(4):667–680, December 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-011-9204-z.

Harris:2024:PAU

[Har24] Richard J. Harris. The piscine arsenal: an updated review of venomous fishes. Reviews in Fish Biology and Fisheries, 34 (2):539–574, June 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-023-09828-w.

Hamilton:2019:TMR

[HB19] Sheryl Hamilton and G. Barry Baker. Technical mitigation to reduce marine mammal bycatch and entanglement in commercial fishing gear: lessons learnt and future directions. Reviews in Fish Biology and Fisheries, 29(2):223–247, June 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-019-09550-6.

Hatry:2013:SFC

[HBC13] Charles Hatry, Thomas R. Binder, and Steven J. Cooke. The status of fishways in Canada: trends identified using the national CanFishPass database. Reviews in Fish Biology and Fisheries, 23(3):271–281, September 2013. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-012-9293-3.

Harrington:2004:BR

[HC04] Gretchen A. Harrington and David Cross. Book reviews. Reviews in Fish Biology and Fisheries, 14(1):145–146, March 2004. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-004-3538-8.

Harnish:2011:RPB

[HCB11] Ryan A. Harnish, Alison H. Colotelo, and Richard S. Brown. A review of polymer-based water conditioners for reduction of handling-related injury. Reviews in Fish Biology and Fisheries, 21(1):43–49, March 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-010-9187-1.

Horodysky:2015:PSF

[HCB15] Andrij Z. Horodysky, Steven J. Cooke, and Richard W. Brill. Physiology in the service of fisheries science: Why thinking mechanistically matters. Reviews in Fish Biology and Fisheries, 25(3):425–447, September 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-015-9393-y.

Hall:2021:PPA

[HCK21] April E. Hall, Darren S. Cameron, and Michael J. Kingsford. Partially protected areas as a management tool on inshore reefs. Reviews in Fish Biology and Fisheries, 31(3):631–651, September 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-021-09654-y.

Hirt-Chabbert:2024:CSA

[HCMT24a] Jorge A. Hirt-Chabbert, Alejandro S. Mechaly, and Ciro Tapia. Correction: Seafood in Argentina: marine fish species, seasonal presence and prices. Reviews in Fish Biology and Fisheries, 34 (2):775–779, June 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-024-09851-5. See [HCMT24b].

Hirt-Chabbert:2024:SAM

[HCMT24b] Jorge A. Hirt-Chabbert, Alejandro S. Mechaly, and Ciro Tapia. Seafood in Argentina: marine fish species, seasonal presence and prices. Reviews in Fish Biology and Fisheries, 34(2): 753–774, June 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-024-09836-4. See correction

[HCMT24a].

Howarth:2024:NPS

[HCNH24] Andrew Howarth, Steven J. Cooke, Vivian M. Nguyen, and Len M. Hunt. Non-probabilistic surveys and sampling in the human dimensions of fisheries. Reviews in Fish Biology and Fisheries, 34(2):597–622, June 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-023-09831-1.

Hobday:2016:PAC

[HCvP16] Alistair J. Hobday, Kevern Cochrane, and E. Ingrid van Putten. Planning adaptation to climate change in fast-warming marine regions with seafood-dependent coastal communities. Reviews in Fish Biology and Fisheries, 26(2):249–264, June 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-016-9419-0.

Higgins:2015:MGF

[HDI15] R. M. Higgins, H. Diogo, and E. J. Isidro. Modelling growth in fish with complex life histories. Reviews in Fish Biology and Fisheries, 25(3):449–462, September 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-015-9388-8.

Hilborn:1995:BR

[HES95] Ray Hilborn, Martin Esseen, and J. R. Sargent. Book reviews. Reviews in Fish Biology and Fisheries, 5(4):461–473, December 1995. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF01103818.

Harrison:2023:WMB

[HETS23] Luke O. J. Harrison, Georg H. Engelhard, Ruth H. Thurstan, and Anna M. Sturrock. Widening mismatch between UK seafood production and consumer demand: a 120-year perspective. Reviews in Fish Biology and Fisheries, 33(4):1387– 1408, December 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-023-09776-5.

Hou:2020:NPF

[HF20] Zhenxin Hou and Lee A. Fuiman. Nutritional programming in fishes: insights from mammalian studies. Reviews in Fish Biology and Fisheries, 30(1):67–92, March 2020. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-019-09590-y.

Hall:2007:EMY

[HFG07] Karina C. Hall, Anthony J. Fowler, and Michael C. Geddes. Evidence for multiple year classes of the giant Australian cuttlefish Sepia apama in northern Spencer Gulf, South Australia. Reviews in Fish Biology and Fisheries, 17(2–3):??, August 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-007-9045-y.

Haas:2019:BFR

[HFM19] Bianca Haas, Aysha Fleming, and Jeffrey McGee. Big fishing: the role of the large-scale commercial fishing industry in achieving sustainable development goal 14. Reviews in Fish Biology and Fisheries, 29(1):161–175, March 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-018-09546-8.

Hamon:2014:ABF

[HFP14] Katell G. Hamon, Stewart D. Frusher, and André E. Punt. Adaptive behaviour of fishers to external perturbations: simulation of the Tasmanian rock lobster fishery. Reviews in Fish Biology and Fisheries, 24(2):577–592, June 2014. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-013-9302-1.

Hall:2017:MBT

[HGC17] Martin Hall, Eric Gilman, and Erin Carruthers. Mitigating bycatch in tuna fisheries. Reviews in Fish Biology and Fisheries, 27(4):881–908, December 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-017-9478-x.

Harlioglu:2004:HFC

[HH04] Muzaffer Mustafa Harlioğlu and Ayşe Gül Harlioğlu. The harvest of freshwater crayfish, Astacus leptodactylus (Eschscholtz, 1823) in Turkey. Reviews in Fish Biology and Fisheries, 14(4): 415–419, December 2004. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-005-0812-3.

Harlioglu:2006:TNN

[HH06] Muzaffer Mustafa Harlioğlu and Ayse Gül Harlioğlu. Threat of non-native crayfish introductions into Turkey: global lessons. Reviews in Fish Biology and Fisheries, 16(2):171–181, May 2006. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-006-9010-1.

Herrera:2010:OHD

[HHCM10] Marcelino Herrera, Ismael Hachero-Cruzado, and Juan Miguel Mancera. Organogenesis and histological development of the wedge sole Dicologoglossa cuneata M. larva with special reference to the digestive system. Reviews in Fish Biology and Fisheries, 20(4):489–497, December 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-010-9161-y.

Herrera:2011:RWS

[HHCN11] Marcelino Herrera, Ismael Hachero-Cruzado, and José Ignacio Navas. Reproduction of the wedge sole (Dicologoglossa cuneata Moreau) in captivity: spawning parameters and influence of the natural temperature. Reviews in Fish Biology and Fisheries, 21(3):517–526, September 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-010-9172-8.

Hirohashi:2016:CAT

[HII16] Noritaka Hirohashi, Tomohiro Iida, and Yoko Iwata. Complex adaptive traits between mating behaviour and post-copulatory sperm behaviour in squids. Reviews in Fish Biology and Fisheries, 26(3):601–607, September 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-016-9434-1.

Hilborn:1999:BRS

[Hil99] Ray Hilborn. Book review: Stocking and Introduction of Fish. Edited by I. G. Cowx. Reviews in Fish Biology and Fisheries, 9 (1):121–122, March 1999. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1008839713624.

Hayes:2009:LFP

[HJC09] Daniel Hayes, Michael Jones, and Nicholas Collins. Linking fish population dynamics to habitat conditions: insights from the application of a process-oriented approach to several Great Lakes species. Reviews in Fish Biology and Fisheries, 19(3): 295–312, September 2009. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-009-9103-8.

Holder:2020:PCF

[HJC20] Peter E. Holder, Amanda L. Jeanson, and Steven J. Cooke. Preparing for a changing future in recreational fisheries: 100 research questions for global consideration emerging from a horizon scan. Reviews in Fish Biology and Fisheries, 30(1): 137–151, March 2020. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-020-09595-y.

Hayes:2014:CEM

[HK14] Sean A. Hayes and John F. Kocik. Comparative estuarine and marine migration ecology of Atlantic salmon and steelhead: blue highways and open plains. Reviews in Fish Biology and Fisheries, 24(3):757–780, September 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-014-9348-8.

Hulak:2010:DTP

[HKL10] Martin Hulak, Vojtech Kaspar, and Otomar Linhart. Does triploidization produce functional sterility of triploid males of tench Tinca tinca (L.)? Reviews in Fish Biology and Fisheries, 20(3):307–315, September 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-009-9139-9.

Halsey:2018:EKI

[HKN18] Lewis G. Halsey, Shaun S. Killen, and Tommy Norin. Exploring key issues of aerobic scope interpretation in ectotherms: absolute versus factorial. Reviews in Fish Biology and Fisheries, 28 (2):405–415, June 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-018-9516-3.

Hallwass:2023:LDA

[HKTS^{+}23] Gustavo Hallwass, Friedrich W. Keppeler, Luís H. Tomazoni-Silva, Ivan A. Alves, Victoria J. Isaac, Morgana C. Almeida, and Renato A. M. Silvano. ‘Disentangling’ the advantages from gillnets in freshwater small-scale fisheries in the Brazilian Amazon. Reviews in Fish Biology and Fisheries, 33(3):853–874, September 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09771-w.

Hilborn:1998:SSG

[HL98a] Ray Hilborn and Martin Liermann. Standing on the shoulders of giants: Learning from experience in fisheries. Reviews in Fish Biology and Fisheries, 8(3):273–283, September 1998. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1008877912528.

Holloway:1998:NRG

[HL98b] A. C. Holloway and J. F. Leatherland. Neuroendocrine regulation of growth hormone secretion in teleost fishes with emphasis on the involvement of gonadal sex steroids. Reviews in Fish Biology and Fisheries, 8(4):409–429, December 1998. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1008824723747.

Harms:2011:VRS

[HL11] Craig A. Harms and Gregory A. Lewbart. The veterinarian’s role in surgical implantation of electronic tags in fish. Reviews in Fish Biology and Fisheries, 21(1):25–33, March 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-010-9185-3.

Hoving:2007:GMS

[HLD07] H. J. T. Hoving, M. R. Lipinski, and M. D. Durholtz. Growth and mating of southern African Lycoteuthis lorigera (Steenstrup, 1875) (Cephalopoda; Lycoteuthidae). Reviews in Fish Biology and Fisheries, 17(2–3):259–270, August 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-006-9031-9.

Haas:2022:FOG

[HMH22] Bianca Haas, Mary Mackay, and Marcus Haward. The future of ocean governance. Reviews in Fish Biology and Fisheries, 32 (1):253–270, March 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-020-09631-x.

Henderson:2017:ESO

[HMN17] M. Elisabeth Henderson, Katherine E. Mills, and Janet A. Nye. Effects of spring onset and summer duration on fish species distribution and biomass along the Northeast United States continental shelf. Reviews in Fish Biology and Fisheries, 27 (2):411–424, June 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-017-9487-9.

Hansen:2016:PES

[HMQ16] Michael J. Hansen, Charles P. Madenjian, and Bernardo R. Quintella. Population ecology of the sea lamprey (Petromyzon marinus) as an invasive species in the Laurentian Great Lakes and an imperiled species in Europe. Reviews in Fish Biology and Fisheries, 26(3):509–535, September 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-016-9440-3.

Huveneers:2017:EVS

[HMV17] Charlie Huveneers, Mark G. Meekan, and Gabriel M. S. Vianna. The economic value of shark-diving tourism in Australia. Reviews in Fish Biology and Fisheries, 27(3):665–680, September 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-017-9486-x.

Hendrickson:2002:DDR

[HNSS02] Dean A. Hendrickson, Steven M. Norris, and Juan Jacobo Schmitter-Soto. Dedication to Dr. Robert “Bob” Rush Miller. Reviews in Fish Biology and Fisheries, 12(2–3):113–118, June 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1025035119416.

Holt:1998:FY

[Hol98] Sidney J. Holt. Fifty years on. Reviews in Fish Biology and Fisheries, 8(3):357–366, September 1998. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1023/A:1008808804372.

Howe:1993:CSF

[How93] Jeffrey C. Howe. Captive seawater fishes: Science and technology. Reviews in Fish Biology and Fisheries, 3(2):198–199, June 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00045239.

Howe:1994:BRF

[How94] Jeffrey C. Howe. Book review: Fishes of The Bahamas and adjacent tropical waters, second edition. Reviews in Fish Biology and Fisheries, 4(1):136–138, March 1994. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/BF00043270.

Handy:1993:NPM

[HP93] R. D. Handy and M. G. Poxton. Nitrogen pollution in mariculture: toxicity and excretion of nitrogenous compounds by marine fish. Reviews in Fish Biology and Fisheries, 3(3):205– 241, September 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF00043929.

Hibberd:2007:ECF

[HP07] Ty Hibberd and Gretta T. Pecl. Effects of commercial fishing on the population structure of spawning southern calamary (Sepioteuthis australis). Reviews in Fish Biology and Fisheries, 17(2– 3):207–221, August 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-006-9028-4.

Hobday:2014:IGM

[HP14] Alistair J. Hobday and Gretta T. Pecl. Identification of global marine hotspots: sentinels for change and vanguards for adaptation action. Reviews in Fish Biology and Fisheries, 24(2):415– 425, June 2014. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-013-9326-6.

Hendrickson:2002:MNT

[HPdL02] Dean A. Hendrickson, Héctor Espinosa Pérez, and Francisco J. García de León. Mexican native trouts: a review of their history and current systematic and conservation status. Reviews in Fish Biology and Fisheries, 12(2–3):273–316, June 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1025062415188.

Hilborn:2005:FRF

[HPL05] Ray Hilborn, Julia K. Parrish, and Kate Litle. Fishing rights or fishing wrongs? Reviews in Fish Biology and Fisheries, 15 (3):191–199, August 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-005-5138-7.

Hawkins:2015:IGU

[HPP15] Anthony D. Hawkins, Ann E. Pembroke, and Arthur N. Popper. Information gaps in understanding the effects of noise on fishes and invertebrates. Reviews in Fish Biology and Fisheries, 25 (1):39–64, March 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-014-9369-3.

Hoshino:2018:EME

[HPY18] Eriko Hoshino, Sean Pascoe, and Satoshi Yamazaki. Estimating maximum economic yield in multispecies fisheries: a review. Reviews in Fish Biology and Fisheries, 28(2):261–276, June 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-017-9508-8.

Hirst:2000:IGS

[HR00] Andrew G. Hirst and Paul G. Rodhouse. Impacts of geophysical seismic surveying on fishing success. Reviews in Fish Biology and Fisheries, 10(1):113–118, March 2000. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1023/A:1008987014736.

Heras:2009:MPM

[HRC09] Sandra Heras, María Inés Roldán, and Mariano González Castro. Molecular phylogeny of Mugilidae fishes revised. Reviews in Fish Biology and Fisheries, 19(2):217–231, June 2009. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-008-9100-3.

He:2020:DPE

[HSC20] Dekui He, Xiaoyun Sui, and Yiyu Chen. Diversity, pattern and ecological drivers of freshwater fish in China and adjacent areas. Reviews in Fish Biology and Fisheries, 30(2):387–404, June 2020. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-020-09600-4.

Holland:2024:SDT

[HSC^{+}24] Owen J. Holland, Callum Smythe, Timothy D. Clark, Norman L. C. Ragg, Julie Mondon, Patricia Corbett, and Adam D. Miller. Size-dependent thermal limits in Australian hybrid abalone: implications for productivity shifts with ocean warming. Reviews in Fish Biology and Fisheries, 34(1):??, ???? 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09812-4.

Hinchliffe:2021:MDL

[HSS21] Charles Hinchliffe, James A. Smith, and Iain M. Suthers. Modelling the distribution of larval fish in a western boundary current using a multi-voyage database. Reviews in Fish Biology and Fisheries, 31(2):399–415, June 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-021-09647-x.

Hobday:2000:EBS

[HTH00] Alistair J. Hobday, Mia J. Tegner, and Peter L. Haaker. Overexploitation of a broadcast spawning marine invertebrate: Decline of the white abalone. Reviews in Fish Biology and Fisheries, 10(4):493–514, December 2000. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1023/A:1012274101311.

Hunnam:2021:BET

[Hun21] Kimberley Hunnam. The biology and ecology of tropical marine sardines and herrings in Indo-West Pacific fisheries: a review. Reviews in Fish Biology and Fisheries, 31(3):449–484, September 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-021-09649-9.

Huppert:2005:OFR

[Hup05] Daniel D. Huppert. An overview of fishing rights. Reviews in Fish Biology and Fisheries, 15(3):201–215, August 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-005-4869-9.

Huston:2004:BRH

[Hus04] Mark Huston. Book review: Howard I. Browman, and Anne Berit Skiftesvik (eds.), The Big Fish Bang — Proceedings of the 26th Annual Larval Fish Conference. Norwegian Institute of Marine Research, Bergen, Norway. 476 pp. Reviews in Fish Biology and Fisheries, 14(1):147–148, March 2004. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-004-2732-z.

Harrison:2024:GFA

[HW24] Trevor D. Harrison and Alan K. Whitfield. A global functional analysis of fish-estuary associations and selected environmental factors. Reviews in Fish Biology and Fisheries, 34 (2):805–825, June 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-024-09839-1.

Harrison:2014:RGE

[HWA14] Andrew J. Harrison, Alan M. Walker, and Miran W. Aprahamian. A review of glass eel migratory behaviour, sampling techniques and abundance estimates in estuaries: implications for assessing recruitment, local production and exploitation. Reviews in Fish Biology and Fisheries, 24(4):967–983, December 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-014-9356-8.

Hobday:2013:CIO

[HYW13] Alistair J. Hobday, Jock W. Young, and Kevin C. Weng. Climate impacts and oceanic top predators: moving from impacts to adaptation in oceanic systems. Reviews in Fish Biology and Fisheries, 23(4):537–546, December 2013. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-013-9311-0.

Iwanowicz:2004:BR

[IBJ04] Luke Iwanowicz, Leah Brown, and Francis Juanes. Book review. Reviews in Fish Biology and Fisheries, 14(1):149–151, March 2004. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-004-4456-5.

Izzo:2016:FPE

[IDG16] Christopher Izzo, Zoë A. Doubleday, and Bronwyn M. Gillanders. Fish as proxies of ecological and environmental change. Reviews in Fish Biology and Fisheries, 26(3):265–286, September 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-016-9424-3.

Innal:2006:EET

[IE06] Deniz Innal and Füsun Erk’akan. Effects of exotic and translocated fish species in the inland waters of Turkey. Reviews in Fish Biology and Fisheries, 16(1):39–50, February 2006. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-006-9005-y.

Imsland:2005:RCP

[IFW05] A. K. Imsland, A. Foss, and P. White. A review of the culture potential of Solea solea and S. senegalensis. Reviews in Fish Biology and Fisheries, 13(4):379–408, January 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-004-1632-6.

Islam:2004:MBC

[IH04] Md. Shahidul Islam and Mahfuzul Haque. The mangrove-based coastal and nearshore fisheries of Bangladesh: ecology, exploitation and management. Reviews in Fish Biology and Fisheries, 14(2):153–180, June 2004. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-004-3769-8.

Imsland:2001:RGT

[IJ01] Albert K. Imsland and Thor M. Jonassen. Regulation of growth in turbot (Scophthalmus maximus Rafinesque) and Atlantic halibut (Hippoglossus hippoglossus L.): aspects of environment × genotype interactions. Reviews in Fish Biology and Fisheries, 11(1):71–90, March 2001. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1014240430779.

Imsland:2003:LPG

[IJ03] Albert K. Imsland and Ólöf Dóra Bartels Jónsdóttir. Linking population genetics and growth properties of Atlantic cod. Reviews in Fish Biology and Fisheries, 13(1):1–26, March 2003. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1026373509576.

Ibanez:2010:CC

[IK10] Christian M. Ibáñez and Friedemann Keyl. Cannibalism in cephalopods. Reviews in Fish Biology and Fisheries, 20(1): 123–136, March 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-009-9129-y.

Ibengwe:2023:RHM

[IOHM23] Lilian J. Ibengwe, Paul O. Onyango, Aloyce S. Hepelwa, and Prosper L. Mfilinge. Revealing the hidden marine dagaa crossborder trade in mainland Tanzania. Reviews in Fish Biology and Fisheries, 33(3):717–738, September 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-023-09769-4.

Ibarz:2010:LTC

[IPT10] Antoni Ibarz, Francesc Padrós, and Lluís Tort. Low-temperature challenges to gilthead sea bream culture: review of cold-induced alterations and ‘Winter Syndrome’. Reviews in Fish Biology and Fisheries, 20(4):539–556, December 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-010-9159-5.

Ibanez:2021:SCA

[IRPG21] Christian M. Ibáñez, Rodrigo Riera, and M. Cecilia Pardo-Gandarillas. Stomach content analysis in cephalopods: past research, current challenges, and future directions. Reviews in Fish Biology and Fisheries, 31(3):505–522, September 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-021-09653-z.

Iwama:1998:HSP

[ITV98] George K. Iwama, Philip T. Thomas, and Mathilakath M. Vijayan. Heat shock protein expression in fish. Reviews in Fish Biology and Fisheries, 8(1):35–56, March 1998. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1023/A:1008812500650.

Izzo:2017:IAD

[IWG17] Christopher Izzo, Tim M. Ward, and Bronwyn M. Gillanders. Integrated approach to determining stock structure: implications for fisheries management of sardine, Sardinops sagax, in Australian waters. Reviews in Fish Biology and Fisheries, 27 (1):267–284, March 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-017-9468-z.

Jurss:1995:FMR

[JB95] Karl Jürss and Ralf Bastrop. The function of mitochondria-rich cells (chloride cells) in teleost gills. Reviews in Fish Biology and Fisheries, 5(2):235–255, June 1995. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00179758.

Jessen:2006:BRE

[JBJ06] B. Jessen, J. Black, and F. Juanes. Book review: Elliott A. Norse and Larry B. Crowder (eds): Review of “Marine Conservation Biology: the science of maintaining the sea’s biodiversity”. Reviews in Fish Biology and Fisheries, 16(2):229–231, May 2006. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-006-9006-x.

Juanes:2002:SRB

[JBS02] Francis Juanes, Jeff Buckel, and Fred Scharf. Symposium review: Biology, ecology and life history of bluefish. Reviews in Fish Biology and Fisheries, 12(4):429–430, December 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1025344418159.

Jackson:2007:ANT

[JBX07] G. D. Jackson, P. Bustamante, and J. C. Xavier. Applying new tools to cephalopod trophic dynamics and ecology: perspectives from the Southern Ocean Cephalopod Workshop, February 2–3, 2006. Reviews in Fish Biology and Fisheries, 17(2– 3):79–99, August 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-007-9055-9.

James:2007:FCT

[JCL07] Nicola C. James, Paul D. Cowley, and Steve J. Lamberth. Fish communities in temporarily open/closed estuaries from the warm- and cool-temperate regions of South Africa: a review. Reviews in Fish Biology and Fisheries, 17(4):565–580, November 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-007-9057-7.

Jonsson:1995:BR

[JCR95] Bror Jonsson, Patrick Colgan, and Gordon McG. Reid. Book reviews. Reviews in Fish Biology and Fisheries, 5(2):256– 265, June 1995. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/BF00179759.

Jaric:2010:PVA

[JEL10] I. Jarić, T. Ebenhard, and M. Lenhardt. Population viability analysis of the Danube sturgeon populations in a vortex simulation model. Reviews in Fish Biology and Fisheries, 20 (2):219–237, June 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-009-9151-0.

Jellyman:2022:CEH

[Jel22a] Donald J. Jellyman. Correction to: An enigma: how can freshwater eels (Anguilla spp.) be such a successful genus yet be universally threatened? Reviews in Fish Biology and Fisheries, 32(2):719, June 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-021-09675-7. See [Jel22b].

Jellyman:2022:EHC

[Jel22b] Donald J. Jellyman. An enigma: how can freshwater eels (Anguilla spp.) be such a successful genus yet be universally threatened? Reviews in Fish Biology and Fisheries, 32(2): 701–718, June 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-021-09658-8. See correction

[Jel22a].

Jennings:2000:PPP

[Jen00] Simon Jennings. Patterns and prediction of population recovery in marine reserves. Reviews in Fish Biology and Fisheries, 10 (2):209–231, June 2000. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1016619102955.

Jiao:2009:RSM

[Jia09] Yan Jiao. Regime shift in marine ecosystems and implications for fisheries management, a review. Reviews in Fish Biology and Fisheries, 19(2):177–191, June 2009. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-008-9096-8.

Jonsson:1993:PMN

[JJ93] Bror Jonsson and Nina Jonsson. Partial migration: niche shift versus sexual maturation in fishes. Reviews in Fish Biology and Fisheries, 3(4):348–365, December 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/BF00043384.

Jia:2022:UFC

[JJC22] Yintao Jia, Yihang Jiang, and Yifeng Chen. Unravelling fish community assembly in shallow lakes: Insights from functional and phylogenetic diversity. Reviews in Fish Biology and Fisheries, 32(2):623–644, June 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-021-09688-2.

Juan-Jorda:2013:LDL

[JJMD13] Maria José Juan-Jordá, Iago Mosqueira, and Nicholas K. Dulvy. Life in 3-D: life history strategies in tunas, mackerels and bonitos. Reviews in Fish Biology and Fisheries, 23(2):135–155, June 2013. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-012-9284-4.

Jeanson:2021:BSA

[JLC21] Amanda L. Jeanson, A. J. Lynch, and S. J. Cooke. A bright spot analysis of inland recreational fisheries in the face of climate change: learning about adaptation from small successes. Reviews in Fish Biology and Fisheries, 31(2):181–200, June 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-021-09638-y.

Jeunen:2024:CAF

[JLD^{+}24] Gert-Jan Jeunen, Miles Lamare, Jennifer Devine, Stefano Mariani, Sadie Mills, Jackson Treece, Sara Ferreira, and Neil J. Gemmell. Characterizing Antarctic fish assemblages using eDNA obtained from marine sponge bycatch specimens. Reviews in Fish Biology and Fisheries, 34(1):??, ???? 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-023-09805-3.

Jackson:2019:OPP

[JMJ19] Aaron D. Jackson, Mary L. Moser, and Jeffrey C. Jolley. Occurrence of pathogens in Pacific lamprey (Entosphenus tridentatus). Reviews in Fish Biology and Fisheries, 29(3):653–668, September 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-019-09572-0.

Johnsson:2018:SBV

[JN18] Jörgen I. Johnsson and Joacim Näslund. Studying behavioural variation in salmonids from an ecological perspective: observations questions methodological considerations. Reviews in Fish Biology and Fisheries, 28(4):795–823, December 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-018-9532-3.

Jane:2014:UMA

[JNW14] Stephen F. Jane, Keith H. Nislow, and Andrew R. Whiteley. The use (and misuse) of archaeological salmon data to infer historical abundance in North America with a focus on New England. Reviews in Fish Biology and Fisheries, 24(3):943–954, September 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-013-9337-3.

Jones:2002:MPA

[Jon02] Peter J. S. Jones. Marine protected area strategies: issues, divergences and the search for middle ground. Reviews in Fish Biology and Fisheries, 11(3):197–216, September 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1020327007975.

Jones:2007:HTD

[Jon07a] M. R. L. Jones. Historic trawl data and recent information infers temporal change in the occurrence of squid in the diet of orange roughy (Hoplostethus atlanticus Collett) in New Zealand. Reviews in Fish Biology and Fisheries, 17(2–3):493–499, August 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-006-9038-2.

Jones:2007:PVA

[Jon07b] P. J. S. Jones. Point-of-view: Arguments for conventional fisheries management and against no-take marine protected areas: only half of the story? Reviews in Fish Biology and Fisheries, 17 (1):31–43, February 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-006-9016-8.

Joyce:2023:MGP

[Joy23] William Joyce. Muscle growth and plasticity in teleost fish: the significance of evolutionarily diverse sarcomeric proteins. Reviews in Fish Biology and Fisheries, 33(4):1311–1327, December 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09800-8.

Jung:2014:LSD

[JPC14] Sukgeun Jung, Ig-Chan Pang, and Hyung Kee Cha. Latitudinal shifts in the distribution of exploited fishes in Korean waters during the last 30 years: a consequence of climate change. Reviews in Fish Biology and Fisheries, 24(2):443–462, June 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-013-9310-1.

Jones:1997:EPR

[JR97] Jackie C. Jones and John D. Reynolds. Effects of pollution on reproductive behaviour of fishes. Reviews in Fish Biology and Fisheries, 7(4):463–491, December 1997. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1023/A:1018456315671.

Juanes:2002:LFI

[Jua02] Francis Juanes. Listening to fish: an international workshop on the application of passive acoustics in fisheries. Reviews in Fish Biology and Fisheries, 12(1):105–106, March 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1022698020541.

Juanes:2010:SRI

[Jua10] Francis Juanes. Symposium review: International symposium on Formosa: landlocked salmon and masu salmon. Reviews in Fish Biology and Fisheries, 20(2):273–275, June 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-009-9156-8.

Juell:1995:BAS

[Jue95] Jon-Erik Juell. The behaviour of Atlantic salmon in relation to efficient cage-rearing. Reviews in Fish Biology and Fisheries, 5(3):320–335, September 1995. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00043005.

Jennings:2000:EBP

[JZ00] Cecil A. Jennings and Steven J. Zigler. Ecology and biology of paddlefish in North America: historical perspectives, management approaches, and research priorities. Reviews in Fish Biology and Fisheries, 10(2):167–181, June 2000. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1016633604301.

Kamler:2002:OYF

[Kam02] Ewa Kamler. Ontogeny of yolk-feeding fish: an ecological perspective. Reviews in Fish Biology and Fisheries, 12(1):79–103, March 2002. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1023/A:1022603204337.

Kamler:2005:PEP

[Kam05] Ewa Kamler. Parent–egg–progeny relationships in teleost fishes: an energetics perspective. Reviews in Fish Biology and Fisheries, 15(4):??, November 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-006-0002-y.

Kamler:2008:RAY

[Kam08] Ewa Kamler. Resource allocation in yolk-feeding fish. Reviews in Fish Biology and Fisheries, 18(2):??, May 2008. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-007-9070-x.

Katsukawa:1997:PVI

[Kat97] Toshio Katsukawa. Points of view: Introduction of spawning potential: improvement in the threshold management theory. Reviews in Fish Biology and Fisheries, 7(2):285–289, June 1997. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1018435711128.

Kamilya:2014:EUS

[KB14] Dibyendu Kamilya and Arunjyoti Baruah. Epizootic ulcerative syndrome (EUS) in fish: history and current status of understanding. Reviews in Fish Biology and Fisheries, 24(1):

369–380, March 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-013-9335-5.

Kennelly:2021:RBR

[KB21] Steven J. Kennelly and Matt K. Broadhurst. A review of bycatch reduction in demersal fish trawls. Reviews in Fish Biology and Fisheries, 31(2):289–318, June 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-021-09644-0.

Kleiber:2011:ISP

[KBV11] D. Kleiber, L. K. Blight, and A. C. J. Vincent. The importance of seahorses and pipefishes in the diet of marine animals. Reviews in Fish Biology and Fisheries, 21(2):205–223, June 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-010-9167-5.

Kieffer:1992:RLF

[KC92] James D. Kieffer and Patrick W. Colgan. The role of learning in fish behaviour. Reviews in Fish Biology and Fisheries, 2(2):125–143, June 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF00042881.

Keefer:2014:HSA

[KC14] Matthew L. Keefer and Christopher C. Caudill. Homing and straying by anadromous salmonids: a review of mechanisms and rates. Reviews in Fish Biology and Fisheries, 24(1):333– 368, March 2014. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-013-9334-6.

Kerby:2012:UKR

[KCE12] Tina K. Kerby, William W. L. Cheung, and Georg H. Engelhard. The United Kingdom’s role in North Sea demersal fisheries: a hundred year perspective. Reviews in Fish Biology and Fisheries, 22(3):621–634, September 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-012-9261-y.

Kessel:2014:RDR

[KCF14] S. T. Kessel, S. J. Cooke, and A. T. Fisk. A review of detection range testing in aquatic passive acoustic telemetry studies. Reviews in Fish Biology and Fisheries, 24(1):199–218, March 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-013-9328-4.

Kelly:2017:BRK

[KCR17] Rachel Kelly, Richard S. Cottrell, and Yannick Rousseau. Book review: Kathleen Schwerdtner Máñez and Bo Poulsen (eds): Perspectives on oceans past: a handbook of marine environmental history. Reviews in Fish Biology and Fisheries, 27(1): 285–286, March 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-016-9462-x.

Kang:2013:EFF

[KDF13] Bin Kang, Junming Deng, and Yan Feng. Explaining freshwater fish biogeography: history versus environment versus species personality. Reviews in Fish Biology and Fisheries, 23(4):523– 536, December 2013. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-013-9314-x.

Kandalski:2018:EST

[KdSD18] Priscila Krebsbach Kandalski, Maria Rosa Dmengeon Pedreiro de Souza, and Lucélia Donatti. Effects of short-term thermal stress on the plasma biochemical profiles of two Antarctic nototheniid species. Reviews in Fish Biology and Fisheries, 28(4): 925–940, December 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-018-9535-0.

Kelkar:2023:ICF

[Kel23] Nachiket Kelkar. Inland capture fisheries, dam reservoirs, and protected areas for wildlife conservation in India: conflicts and ways forward. Reviews in Fish Biology and Fisheries, 33(1): 201–220, March 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-023-09758-7.

Kennelly:1995:IBA

[Ken95] Steven J. Kennelly. The issue of bycatch in Australia’s demersal trawl fisheries. Reviews in Fish Biology and Fisheries, 5 (2):213–234, June 1995. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF00179757.

Kelly:2022:COS

[KEP22] Rachel Kelly, Karen Evans, and Gretta T. Pecl. Connecting to the oceans: supporting ocean literacy and public engagement. Reviews in Fish Biology and Fisheries, 32(1):123–143, March 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-020-09625-9.

Kock:2021:RTH

[KFS21] Tobias J. Kock, John W. Ferguson, and Carl B. Schreck. Review of trap-and-haul for managing Pacific salmonids (Oncorhynchus spp.) in impounded river systems. Reviews in Fish Biology and Fisheries, 31(1):53–94, March 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-020-09627-7.

Kocour:2010:PDT

[KGF10] Martin Kocour, David Gela, and Martin Flajšhans. Performance of different tench, Tinca tinca (L.), groups under semi-intensive pond conditions: it is worth establishing a coordinated breeding program. Reviews in Fish Biology and Fisheries, 20(3): 345–355, September 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-009-9140-3.

Kamler:2010:PPV

[KGW10] Ewa Kamler, Laura Gasco, and Jacek Wolnicki. Preface: Proceedings of the Vth International Workshop on Biology and Culture of the Tench (Tinca tinca) (L. 1875), Ceresole d’Alba (Italy), September 29–October 3, 2008. Reviews in Fish Biology and Fisheries, 20(3):277–278, September 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-009-9147-9. See erratum [KGW11].

Kamler:2011:EPP

[KGW11] Ewa Kamler, Laura Gasco, and Jacek Wolnicki. Erratum to: Preface: Proceedings of the Vth International Workshop on Biology and Culture of the Tench (Tinca tinca) (L. 1875), Ceresole d’Alba (Italy), September 29–October 3, 2008. Reviews in Fish Biology and Fisheries, 21(2):295, June 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-009-9154-x. See [KGW10].

Kang:2009:FFU

[KHW09] Bin Kang, Daming He, and Yunfei Wu. Fish and fisheries in the Upper Mekong: current assessment of the fish community, threats and conservation. Reviews in Fish Biology and Fisheries, 19(4):??, December 2009. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-009-9114-5.

Kime:1993:CNC

[Kim93] David E. Kime. ‘Classical’ and ‘non-classical’ reproductive steroids in fish. Reviews in Fish Biology and Fisheries, 3 (2):160–180, June 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF00045230.

Kime:1995:EPR

[Kim95] David E. Kime. The effects of pollution on reproduction in fish. Reviews in Fish Biology and Fisheries, 5(1):52–95, March 1995. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF01103366.

Korwin-Kossakowski:2012:FHS

[KK12] Michał Korwin-Kossakowski. Fish hatching strategies: a review. Reviews in Fish Biology and Fisheries, 22(1):225–240, March 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-011-9233-7.

Kohlmann:2010:GVD

[KKF10] Klaus Kohlmann, Petra Kersten, and Martin Flajšhans. Genetic variability and differentiation of wild and cultured tench populations inferred from microsatellite loci. Reviews in Fish Biology and Fisheries, 20(3):279–288, September 2010. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-009-9138-x.

Kujawa:2010:ETC

[KKM10] Roman Kujawa, Dariusz Kucharczyk, and Andrzej Mamcarz. The effect of tannin concentration and egg unsticking time on the hatching success of tench Tinca tinca (L.) larvae. Reviews in Fish Biology and Fisheries, 20(3):339–343, September 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-009-9136-z.

Kamler:2012:EDT

[KKW12] Ewa Kamler, Rafał Kamiński, and Jakub Wałowski. Effects of diet and temperature on condition, proximate composition and three major macro elements, Ca, P and Mg, in barbel Barbus barbus juveniles. Reviews in Fish Biology and Fisheries, 22(3): 767–777, September 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-012-9256-8.

Kasumyan:2023:TPS

[KL23] Alexander Kasumyan and Alexandra Levina. Are the taste preferences similar in closely related fish of the same trophic category? A case of Nile and Mozambique tilapias. Reviews in Fish Biology and Fisheries, 33(4):1371–1386, December 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09763-w.

Kang:2018:FCS

[KLC18] Bin Kang, Min Liu, and Shao-Bo Chen. Fisheries in Chinese seas: What can we learn from controversial official fisheries statistics? Reviews in Fish Biology and Fisheries, 28(3):503– 519, September 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-018-9518-1.

Klinard:2020:LUP

[KM20] Natalie V. Klinard and Jordan K. Matley. Living until proven dead: addressing mortality in acoustic telemetry research. Reviews in Fish Biology and Fisheries, 30(3):485–499, September 2020. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-020-09613-z.

Kpogue:2013:RBE

[KMF13] Diane N. S. Kpogue, Guy A. Mensah, and Emile D. Fiogbe. A review of biology, ecology and prospect for aquaculture of Parachanna obscura. Reviews in Fish Biology and Fisheries, 23(1):41–50, March 2013. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-012-9281-7.

Kovach:2016:ICV

[KMK16] Ryan P. Kovach, Clint C. Muhlfeld, and Jeffrey L. Kershner. Impacts of climatic variation on trout: a global synthesis and path forward. Reviews in Fish Biology and Fisheries, 26 (2):135–151, June 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-015-9414-x.

Knudsen:2001:FGD

[Knu01] E. Eric Knudsen. Fishing grounds: Defining a new era for American fisheries management. Reviews in Fish Biology and Fisheries, 11(2):169, June 2001. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1015258416765.

Kock:2001:DIF

[Koc01] Karl-Hermann Kock. The direct influence of fishing and fisheryrelated activities on non-target species in the Southern Ocean with particular emphasis on longline fishing and its impact on albatrosses and petrels — a review. Reviews in Fish Biology and Fisheries, 11(1):31–56, March 2001. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1023/A:1014207719529.

Kostow:2009:FCE

[Kos09] Kathryn Kostow. Factors that contribute to the ecological risks of salmon and steelhead hatchery programs and some mitigating strategies. Reviews in Fish Biology and Fisheries, 19 (1):9–31, March 2009. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-008-9087-9.

Kotrschal:1991:SCC

[Kot91] Kurt Kotrschal. Solitary chemosensory cells — taste, common chemical sense or what? Reviews in Fish Biology and Fisheries, 1(1):3–22, September 1991. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF00042659.

Kang:2021:CCI

[KPN21] Bin Kang, Gretta T. Pecl, and Wentao Niu. Climate change impacts on China’s marine ecosystems. Reviews in Fish Biology and Fisheries, 31(3):599–629, September 2021. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-021-09668-6.

Klein:2017:ECC

[KSK17] Emily S. Klein, Sarah L. Smith, and Jacob P. Kritzer. Effects of climate change on four New England groundfish species. Reviews in Fish Biology and Fisheries, 27(2):317–338, June 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-016-9444-z.

Kwasek:2012:EDD

[KTW12] Karolina Kwasek, Genciana Terova, and Macdonald Wick. The effect of dietary dipeptide lysine–glycine on growth, muscle proteins, and intestine PepT1 gene expression in juvenile yellow perch. Reviews in Fish Biology and Fisheries, 22(3):797–812, September 2012. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-012-9266-6.

Kulczykowska:1995:AVM

[Kul95] Ewa Kulczykowska. Arginine vasotocin-melatonin interactions in fish: a hypothesis. Reviews in Fish Biology and Fisheries, 5 (1):96–102, March 1995. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF01103367.

Kullander:1999:FSH

[Kul99] Sven O. Kullander. Fish species — how and why. Reviews in Fish Biology and Fisheries, 9(4):325–352, December 1999. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1008959313491.

Kulczykowska:2002:RMH

[Kul02] Ewa Kulczykowska. A review of the multifunctional hormone melatonin and a new hypothesis involving osmoregulation. Reviews in Fish Biology and Fisheries, 11(4):321–330, December 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1021348822635.

Kvarnemo:1998:PVW

[Kva98] Charlotta Kvarnemo. Points of view: Why male cannibalism won’t cause a female-biased OSR — a comment on Smith and Wootton’s paper. Reviews in Fish Biology and Fisheries, 8 (1):93–98, March 1998. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1008816601559.

Kotrschal:1998:FBE

[KVH98] K. Kotrschal, M. J. Van Staaden, and R. Huber. Fish brains: Evolution and anvironmental relationships. Reviews in Fish Biology and Fisheries, 8(4):373–408, December 1998. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1008839605380.

Keyl:2008:EVF

[KW08] Friedemann Keyl and Matthias Wolff. Environmental variability and fisheries: what can models do? Reviews in Fish Biology and Fisheries, 18(3):273–299, August 2008. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-007-9075-5.

Kubodera:2007:FHB

[KWI07] Tsunemi Kubodera, Hikaru Watanabe, and Taro Ichii. Feeding habits of the blue shark, Prionace glauca, and salmon shark, Lamna ditropis, in the transition region of the Western North Pacific. Reviews in Fish Biology and Fisheries, 17 (2–3):??, August 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-006-9020-z.

Kadagi:2022:SCS

[KWM^{+}22] N. I. Kadagi, N. Wambiji, B. Mann, D. Parker, R. Daly, P. Thoya, D. A. M. Rato, J. Halafo, L. Gaspare, E. A. Sweke, S. Ahmed, S. B. Raseta, M. Osore, J. Maina, S. Glaser, R. Ahrens, and U. R. Sumaila. Status and challenges for sustainable billfish fisheries in the Western Indian Ocean. Reviews in Fish Biology and Fisheries, 32(4):1035–1061, December 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-022-09725-8.

Katugin:2007:DCU

[KZ07] Oleg N. Katugin and Nikolai N. Zuev. Distribution of cephalopods in the upper epipelagic northwestern Bering Sea in autumn. Reviews in Fish Biology and Fisheries, 17(2–3): 283–294, August 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-007-9040-3.

Krieg:2020:RUP

[KZ20] Raphael Krieg and Armin Zenker. A review of the use of physical barriers to stop the spread of non-indigenous crayfish species. Reviews in Fish Biology and Fisheries, 30(3):423–435, September 2020. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-020-09606-y.

Lackey:2002:ACS

[Lac02] Robert T. Lackey. Anthony T. Charles, sustainable fishery systems. Reviews in Fish Biology and Fisheries, 11(4):363–364, December 2002. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1023/A:1021326610318.

LaMesa:2005:TGA

[LAI05] Mario La Mesa, Enrico Arneri, and M. Iglesias. The transparent goby, Aphia minuta review of biology and fisheries of a paedomorphic European fish. Reviews in Fish Biology and Fisheries, 15(1–2):89–109, February 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-005-1613-4.

Larkin:1996:CIM

[Lar96] P. A. Larkin. Concepts and issues in marine ecosystem management. Reviews in Fish Biology and Fisheries, 6(2):139–164, June 1996. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00182341.

Lynch:2017:E

[LAS17] Abigail J. Lynch, Rebecca G. Asch, and Warwick H. H. Sauer. Editorial. Reviews in Fish Biology and Fisheries, 27(2):293– 296, June 2017. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-017-9483-0.

Lachica-Alino:2006:PFF

[LAWD06] Lualhati Lachica-Aliño, Matthias Wolff, and Laura T. David. Past and future fisheries modeling approaches in the Philippines. Reviews in Fish Biology and Fisheries, 16(2):??, May 2006. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-006-9008-8.

Lucas:2023:SRS

[LB23] Sol Lucas and Per Berggren. A systematic review of sensory deterrents for bycatch mitigation of marine megafauna. Reviews in Fish Biology and Fisheries, 33(1):1–33, March 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-022-09736-5.

Lucifora:2019:SED

[LBS19] Luis O. Lucifora, Santiago A. Barbini, and David E. Sabadin. Socio-economic development, scientific research, and exploitation explain differences in conservation status of marine and freshwater chondrichthyans among countries. Reviews in Fish Biology and Fisheries, 29(4):951–964, December 2019. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-019-09584-w.

Loganimoce:2023:OSW

[LBS^{+}23] Epeli M. Loganimoce, Kelly T. Brown, Rusila Savou, Jokim V. Kitolelei, Max Tukana, Paul C. Southgate, and Monal M.

Lal. Octopuses in the south-west Pacific region: a review of fisheries, ecology, cultural importance and management. Reviews in Fish Biology and Fisheries, 33(4):977–1003, December 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09772-9.

Leiva:2002:RWM

[LC02] Germán E. Leiva and Juan C. Castilla. A review of the world marine gastropod fishery: evolution of catches, management and the Chilean experience. Reviews in Fish Biology and Fisheries, 11(4):283–300, December 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1023/A:1021368216294.

Lorenzen:2016:SAI

[LCC16] K. Lorenzen, I. G. Cowx, and S. J. Cooke. Stock assessment in inland fisheries: a foundation for sustainable use and conservation. Reviews in Fish Biology and Fisheries, 26(3):405– 440, September 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-016-9435-0.

Lennox:2019:TBU

[LCC19] Robert J. Lennox, David A. Crook, and Steven J. Cooke. Toward a better understanding of freshwater fish responses to an increasingly drought-stricken world. Reviews in Fish Biology and Fisheries, 29(1):71–92, March 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-018-09545-9.

Looby:2022:QIG

[LCM22] Audrey Looby, Kieran Cox, and Charles W. Martin. A quantitative inventory of global soniferous fish diversity. Reviews in Fish Biology and Fisheries, 32(2):581–595, June 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-022-09702-1.

LaMesa:2024:ACK

[LE24] Mario La Mesa and Joseph T. Eastman. Assessing current knowledge and future challenges of age determination, life span and growth performance in notothenioid fishes: a review.

Reviews in Fish Biology and Fisheries, 34(2):575–596, June 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09829-9.

Lima:2023:IMP

[LET^{+}23] André L. R. Lima, Linda M. Eggertsen, Jessyca L. S. Teixeira, Alexandre Schiavetti, Fabiana C. Félix-Hackradt, and Carlos W. Hackradt. The influence of marine protected areas on the patterns and processes in the life cycle of reef fishes. Reviews in Fish Biology and Fisheries, 33(4):893–913, December 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09761-y.

Lecomte-Finiger:2003:GAS

[LF03] R. Lecomte-Finiger. The genus Anguilla Schrank, 1798: current state of knowledge and questions. Reviews in Fish Biology and Fisheries, 13(3):265–279, September 2003. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1023/B:RFBF.0000033072.
03829.6d.

Ladich:2013:AEP

[LF13] Friedrich Ladich and Richard R. Fay. Auditory evoked potential audiometry in fish. Reviews in Fish Biology and Fisheries, 23(3): 317–364, September 2013. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-012-9297-z.

Lima-Filho:2016:EDW

[LFdSRM16] Paulo Augusto Lima-Filho, Ricardo de Souza Rosa, and Wagner Franco Molina. Evolutionary diversification of Western Atlantic Bathygobius species based on cytogenetic, morphologic and DNA barcode data. Reviews in Fish Biology and Fisheries, 26(1):109–121, March 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-015-9411-0.

Lucey:2008:JCR

[LFJ08] S. Lucey, A. Franklin, and F. Juanes. John Caddy, review of “Marine Habitat and Cover: Their Importance for Productive Coastal Fishery Resources”. Reviews in Fish Biology and Fisheries, 18(4):445–446, November 2008. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-008-9083-0.

Lescrauwaet:2013:HEY

[LFM13] Ann-Katrien Lescrauwaet, Nancy Fockedey, and Jan Mees. Hundred and eighty years of fleet dynamics in the Belgian sea fisheries. Reviews in Fish Biology and Fisheries, 23(2):229– 243, June 2013. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-012-9287-1.

Lucena-Fredou:2021:RLH

[LFMP21] Flávia Lucena-Frédou, Bruno Mourato, and Maite Pons. Review of the life history, fisheries, and stock assessment for small tunas in the Atlantic Ocean. Reviews in Fish Biology and Fisheries, 31(3):709–736, September 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-021-09666-8.

Lund:1997:RCB

[LG97] Richard Lund and Eileen D. Grogan. Relationships of the Chimaeriformes and the basal radiation of the Chondrichthyes. Reviews in Fish Biology and Fisheries, 7(1):65–123, March 1997. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1018471324332.

Lobyrev:2018:MGM

[LH18] Feodor Lobyrev and Matthew J. Hoffman. A morphological and geometric method for estimating the selectivity of gill nets. Reviews in Fish Biology and Fisheries, 28(4):909–924, December 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-018-9534-1.

Leggatt:2003:OPF

[LI03] Rosalind A. Leggatt and George K. Iwama. Occurrence of polyploidy in the fishes. Reviews in Fish Biology and Fisheries, 13(3):237–246, September 2003. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1023/B:RFBF.0000033049.
00668.fe.

Ligtermoet:2016:MCH

[Lig16] Emma Ligtermoet. Maintaining customary harvesting of freshwater resources: sustainable indigenous livelihoods in the floodplains of northern Australia. Reviews in Fish Biology and Fisheries, 26(4):649–678, December 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-016-9429-y.

Lincoln:1994:MGA

[Lin94] Rich Lincoln. Molecular genetics applications in fisheries: snake oil or restorative? Reviews in Fish Biology and Fisheries, 4(3): 389–392, September 1994. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF00042913.

Lisney:2010:RSB

[Lis10] Thomas J. Lisney. A review of the sensory biology of chimaeroid fishes (Chondrichthyes; Holocephali). Reviews in Fish Biology and Fisheries, 20(4):571–590, December 2010. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-010-9162-x.

Lechner:2016:PPD

[LKH16] Aaron Lechner, Hubert Keckeis, and Paul Humphries. Patterns and processes in the drift of early developmental stages of fish in rivers: a review. Reviews in Fish Biology and Fisheries, 26(3): 471–489, September 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-016-9437-y.

Lajbner:2010:LRI

[LKK10] Zdeněk Lajbner, Klaus Kohlmann, and Petr Kotlík. Lack of reproductive isolation between the Western and Eastern phylogroups of the tench. Reviews in Fish Biology and Fisheries, 20(3):289–300, September 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-009-9137-y.

Le:2011:SCY

[LLC11] M. H. Le, H. K. Lim, and Y. J. Chang. Semen cryopreservation of yellow croaker Larimichthys polyactis. Reviews in Fish Biology and Fisheries, 21(4):789–797, December 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-011-9209-7.

Liu:2024:SVC

[LLJC24] Xiangwei Liu, Yue Liu, Ziyu Jiang, and Ling Cao. Spatiotemporal variation of China’s mariculture potential under climate change. Reviews in Fish Biology and Fisheries, 34(1):??, ???? 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09814-2.

Lopez-Lopez:2009:SSD

[LLSDTJ09] Eugenia López-López, J. Elías Sedeño-Díaz, and Patricia Trujillo-Jiménez. Spatial and seasonal distribution patterns of fish assemblages in the Río Champotón, southeastern Mexico. Reviews in Fish Biology and Fisheries, 19(2):127–142, June 2009. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-008-9093-y.

Lowe-Mcconnell:1993:WBF

[LM93] Rosemary Lowe-Mcconnell. Workshop on biodiversity, fisheries and the future of Lake Victoria. Reviews in Fish Biology and Fisheries, 3(2):201–203, June 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00045241.

Link:2019:CCM

[LM19] Jason S. Link and Anthony R. Marshak. Characterizing and comparing marine fisheries ecosystems in the United States: determinants of success in moving toward ecosystem-based fisheries management. Reviews in Fish Biology and Fisheries, 29 (1):23–70, March 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-018-9544-z.

Latorre:2023:RMA

[LMCB^{+}23] Dani Latorre, Guillem Masó, Carlos Cano-Barbacil, José M. Zamora-Marin, David Almeida, Lorenzo Vilizzi, J. Robert Britton, Alejandra Cruz, Carlos Fernández-Delgado, Anni G. González-Rojas, Rafael Miranda, Francesc Rubio-Gracia, Ali Serhan Tarkan, Mar Torralva, Anna Vila-Gispert, Gordon H.

Copp, and Filipe Ribeiro. A review and meta-analysis of the environmental biology of bleak Alburnus alburnus in its native and introduced ranges, with reflections on its invasiveness. Reviews in Fish Biology and Fisheries, 33(4):931–975, December 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09767-6.

Lees:2021:EDP

[LMH21] K. J. Lees, M. A. MacNeil, and N. E. Hussey. Estimating demographic parameters for fisheries management using acoustic telemetry. Reviews in Fish Biology and Fisheries, 31(1):25– 51, March 2021. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-020-09626-8.

Lenhardt:2011:NNT

[LMM11] Mirjana Lenhardt, Goran Markovic, and Zoran Markovic. Nonnative and translocated fish species in Serbia and their impact on the native ichthyofauna. Reviews in Fish Biology and Fisheries, 21(3):407–421, September 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-010-9180-8.

Lau:2023:PET

[LNT^{+}23] Cher Chien Lau, Siti Azizah Mohd Nor, Min Pau Tan, Yik Sung Yeong, Li Lian Wong, Yves Van de Peer, Patrick Sorgeloos, and Muhd Danish-Daniel. Pigmentation enhancement techniques during ornamental fish production. Reviews in Fish Biology and Fisheries, 33(4):1027–1048, December 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-023-09777-4.

Lorenzen:1993:BRF

[Lor93] Kai Lorenzen. Book review: Fish diseases, 5th edition. Reviews in Fish Biology and Fisheries, 3(4):379–380, December 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00043391.

Losee:2024:ATO

[LPC^{+}24] James P. Losee, Daniel Palm, Andrew Claiborne, Gabe Madel, Lo Persson, Thomas P. Quinn, Tomas Brodin, and Gustav Hellström. Anadromous trout from opposite sides of the globe: biology, ocean ecology, and management of anadromous brown and cutthroat trout. Reviews in Fish Biology and Fisheries, 34(1):??, ???? 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-023-09824-0.

Leigh:2017:NPS

[LPG17] Samantha C. Leigh, Yannis Papastamatiou, and Donovan P. German. The nutritional physiology of sharks. Reviews in Fish Biology and Fisheries, 27(3):561–585, September 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-017-9481-2.

Liu:2019:BDF

[LQW19] Xiongjun Liu, Jiajun Qin, and Xiaoping Wu. Biodiversity decline of fish assemblages after the impoundment of the Three Gorges Dam in the Yangtze River Basin, China. Reviews in Fish Biology and Fisheries, 29(1):177–195, March 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-019-09548-0.

Loher:2011:PRP

[LR11] Timothy Loher and Renee Rensmeyer. Physiological responses of Pacific halibut, Hippoglossus stenolepis, to intracoelomic implantation of electronic archival tags, with a review of tag implantation techniques employed in flatfishes. Reviews in Fish Biology and Fisheries, 21(1):97–115, March 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-010-9192-4.

Lipinski:2007:DLC

[LS07a] M. R. Lipinski and M. Soule. Disintegration of a large concentration of loliginid squid as a response to predation. Reviews in Fish Biology and Fisheries, 17(2–3):477–485, August 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-006-9026-6.

Lipinski:2007:NDM

[LS07b] Marek R. Lipinski and Michael A. Soule. A new direct method of stock assessment of the loliginid squid. Reviews in Fish Biology and Fisheries, 17(2–3):437–453, August 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-007-9050-1.

Lucifora:2022:PCS

[LSB22] Luis O. Lucifora, Pablo A. Scarabotti, and Santiago A. Barbini. Predicting and contextualizing sensitivity to overfishing in neotropical freshwater stingrays (Chondrichthyes: Potamotrygonidae). Reviews in Fish Biology and Fisheries, 32(2):669– 686, June 2022. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-021-09696-2.

Linh-Son:2012:SCF

[LSDH12] Nguyen Linh-Son and To Dieu-Hang. Shortcomings of current fish transfer functions and a proposal for a new transfer function. Reviews in Fish Biology and Fisheries, 22(4):933–942, December 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-012-9278-2.

Lindley:2023:ILS

[LSDRS23] Jade Lindley, Emily De Sousa, Zoe Doubleday, and Patrick Reis-Santos. Innovation to limit seafood fraud post-COVID-19. Reviews in Fish Biology and Fisheries, 33(2):501–512, June 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-022-09747-2.

Lokkeborg:2014:TME

[LSF14] Svein Løkkeborg, Sten Ivar Siikavuopio, and Keno Ferter. Towards more efficient longline fisheries: fish feeding behaviour, bait characteristics and development of alternative baits. Reviews in Fish Biology and Fisheries, 24(4):985–1003, December 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-014-9360-z.

Lusardi:2015:TEN

[LSH15] Robert A. Lusardi, Molly R. Stephens, and Josh M. Hull. Threat evolution: negative feedbacks between management action and species recovery in threatened trout (Salmonidae). Reviews in Fish Biology and Fisheries, 25(3):521–535, September 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-015-9394-x.

Liao:2003:SRS

[LSL03] I. C. Liao, M. S. Su, and E. M. Leaño. Status of research in stock enhancement and sea ranching. Reviews in Fish Biology and Fisheries, 13(2):Article 151, June 2003. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1023/B:RFBF.0000019477.
97551.73.

Lakra:2011:RIL

[LSP11] W. S. Lakra, U. K. Sarkar, and A. Pandey. River inter linking in India: status, issues, prospects and implications on aquatic ecosystems and freshwater fish diversity. Reviews in Fish Biology and Fisheries, 21(3):463–479, September 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-011-9199-5.

Li:2017:AGS

[LW17] Yan-He Li and Han-Ping Wang. Advances of genotypingby-sequencing in fisheries and aquaculture. Reviews in Fish Biology and Fisheries, 27(3):535–559, September 2017. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-017-9473-2.

Lima:2017:FTA

[LWS17] Ana Carolina Lima, Frederick J. Wrona, and Amadeu M. V. M. Soares. Fish traits as an alternative tool for the assessment of impacted rivers. Reviews in Fish Biology and Fisheries, 27 (1):31–42, March 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-016-9446-x.

Lansdell:2007:PCE

[LY07] Matt Lansdell and Jock Young. Pelagic cephalopods from eastern Australia: species composition, horizontal and vertical distribution determined from the diets of pelagic fishes. Reviews in Fish Biology and Fisheries, 17(2–3):??, August 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-006-9024-8.

Liu:2022:CDS

[LZC22] Shurong Liu, Xijie Zhou, and Ling Cao. Characterizing the development of sea ranching in China. Reviews in Fish Biology and Fisheries, 32(3):783–803, September 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-022-09709-8.

Mokhtar:2003:FSR

[MA03] M. B. Mokhtar and A. Awaluddin. Framework for sea ranching. Reviews in Fish Biology and Fisheries, 13(2):213–217, June 2003. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/B:RFBF.0000019482.06550.80.

Mavruk:2008:NNF

[MA08] Sinan Mavruk and Dursun Avsar. Non-native fishes in the Mediterranean from the Red Sea, by way of the Suez Canal. Reviews in Fish Biology and Fisheries, 18(3):251–262, August 2008. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-007-9073-7.

Mendoza:2002:MSA

[MAC02] R. Mendoza, C. Aguilera, and R. Castro. Morphophysiological studies on alligator gar (Atractosteus spatula) larval development as a basis for their culture and repopulation of their natural habitats. Reviews in Fish Biology and Fisheries, 12(2– 3):133–142, June 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1025047914814.

Madsen:2007:SFG

[Mad07] Niels Madsen. Selectivity of fishing gears used in the Baltic Sea cod fishery. Reviews in Fish Biology and Fisheries, 17(4):

517–544, November 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-007-9053-y.

Magnusson:1995:OMV

[Mag95] Kjartan G. Magnússon. An overview of the multispecies VPA — theory and applications. Reviews in Fish Biology and Fisheries, 5(2):195–212, June 1995. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF00179756.

Magana:2013:FPT

[Mag13] Hugo A. Magaña. Flood pulse trophic dynamics of larval fishes in a restored arid-land, river-floodplain, Middle Rio Grande, Los Lunas, New Mexico. Reviews in Fish Biology and Fisheries, 23(4):507–521, December 2013. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-013-9313-y.

Murt:2005:BRC

[MAJ05] John Murt, Katie Anderson, and Francis Juanes. Book review: Carl J. Walters and Steven J. D. Martell, Fisheries Ecology and Management, Princeton University Press, Princeton, NJ 2004, ISBN 0-691-11545-1 Paperback, 423 pp. Reviews in Fish Biology and Fisheries, 15(1–2):155–157, February 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-005-0057-1.

Martinez-Alvarez:2005:ADF

[MÁMS05] Rosa M. Martínez-Álvarez, Amalia E. Morales, and Ana Sanz. Antioxidant defenses in fish: Biotic and abiotic factors. Reviews in Fish Biology and Fisheries, 15(1–2):75–88, February 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-005-7846-4.

Mann:1994:PJA

[Man94] R. H. K. Mann. Production of juvenile Atlantic salmon, Salmo salar, in natural waters. Reviews in Fish Biology and Fisheries, 4(2):262–263, June 1994. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF00044134.

Mobley:2021:MAS

[MAP21] Kenyon B. Mobley, Tutku Aykanat, and Craig R. Primmer. Maturation in Atlantic salmon (Salmo salar, Salmonidae): a synthesis of ecological, genetic, and molecular processes. Reviews in Fish Biology and Fisheries, 31(3):523–571, September 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-021-09656-w.

Marshall:1993:BAC

[Mar93] B. E. Marshall. Biology of the African clupeid Limnothrissa miodon with reference to its small size in artificial lakes. Reviews in Fish Biology and Fisheries, 3(1):17–38, March 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00043296.

Marcogliese:1995:RZT

[Mar95] David J. Marcogliese. The role of zooplankton in the transmission of helminth parasites to fish. Reviews in Fish Biology and Fisheries, 5(3):336–371, September 1995. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/BF00043006.

Marutani:2006:OAO

[Mar06] Teruhiko Marutani. On open access and optimal landings tax. Reviews in Fish Biology and Fisheries, 16(2):115–124, May 2006. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-006-9007-9.

Marutani:2008:OPD

[Mar08] Teruhiko Marutani. On the optimal path in the dynamic pool model for a fishery. Reviews in Fish Biology and Fisheries, 18 (2):133–141, May 2008. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-007-9065-7.

Mayden:2002:PRE

[May02] Richard L. Mayden. Phylogenetic relationships of the enigmatic ornate shiner, Cyprinella ornata, a species endemic to Mexico (Teleostei: Cyprinidae). Reviews in Fish Biology and Fisheries, 12(2–3):339–347, June 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1023/A:1025006625743.

Mota:2015:NIB

[MBA15] Micaela Mota, Ana Bio, and Carlos Antunes. New insights into biology and ecology of the Minho River Allis shad (Alosa alosa L.): contribution to the conservation of one of the last European shad populations. Reviews in Fish Biology and Fisheries, 25 (2):395–412, June 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-015-9383-0.

Marsden:2021:UUT

[MBC21] J. E. Marsden, P. J. Blanchfield, and S. J. Cooke. Using untapped telemetry data to explore the winter biology of freshwater fish. Reviews in Fish Biology and Fisheries, 31(1):115– 134, March 2021. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-021-09634-2.

Moser:2007:CCL

[MBD07] Mary L. Moser, JoAnne M. Butzerin, and Douglas B. Dey. Capture and collection of lampreys: the state of the science. Reviews in Fish Biology and Fisheries, 17(1):45–56, February 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-006-9037-3.

Martins:2013:DER

[MBdBC13] Nícolas Fernandes Martins, Luiz Antonio Carlos Bertollo, and Marcelo de Bello Cioffi. Differentiation and evolutionary relationships in Erythrinus erythrinus (Characiformes, Erythrinidae): comparative chromosome mapping of repetitive sequences. Reviews in Fish Biology and Fisheries, 23(2):261– 269, June 2013. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-012-9292-4.

Miller:2021:SHL

[MBG21] Allison K. Miller, Cindy Baker, and Neil J. Gemmell. The Southern Hemisphere lampreys (Geotriidae and Mordaciidae). Reviews in Fish Biology and Fisheries, 31(2):201–232, June 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-021-09639-x.

Melnychuk:2014:ACF

[MBH14] Michael C. Melnychuk, Jeannette A. Banobi, and Ray Hilborn. The adaptive capacity of fishery management systems for confronting climate change impacts on marine populations. Reviews in Fish Biology and Fisheries, 24(2):561–575, June 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-013-9307-9.

Martinez:2012:BMA

[MBJ12] Sarah Martinez, Melissa Belcher, and Francis Juanes. Bruce S. Miller and Arthur W. Kendall, Jr.: Review of “Early life history of marine fishes”. Reviews in Fish Biology and Fisheries, 22 (1):371–373, March 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-011-9217-7.

Muir:2012:LCS

[MBK12] A. M. Muir, C. T. Blackie, and C. C. Krueger. Lake charr Salvelinus namaycush spawning behaviour: new field observations and a review of current knowledge. Reviews in Fish Biology and Fisheries, 22(3):575–593, September 2012. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-012-9258-6.

Mueller:2006:VAC

[MBM06] Robert P. Mueller, Richard S. Brown, and Larry Moulton. Video and acoustic camera techniques for studying fish under ice: a review and comparison. Reviews in Fish Biology and Fisheries, 16(2):213–226, May 2006. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-006-9011-0.

Morley:2014:GIP

[MBP14] S. A. Morley, M. Belchier, and L. S. Peck. Geographic isolation and physiological mechanisms underpinning species distributions at the range limit hotspot of South Georgia. Reviews in Fish Biology and Fisheries, 24(2):485–492, June 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-013-9308-8.

McHugh:2017:CAG

[MBS17] Matthew J. McHugh, Matt K. Broadhurst, and David J. Sterling. Choosing anterior-gear modifications to reduce the global environmental impacts of penaeid trawls. Reviews in Fish Biology and Fisheries, 27(1):111–134, March 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-016-9459-5.

Myrick:2004:TEJ

[MC04] Christopher A. Myrick and Joseph J. Cech, Jr. Temperature effects on juvenile anadromous salmonids in California’s Central Valley: what don’t we know? Reviews in Fish Biology and Fisheries, 14(1):113–123, March 2004. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-004-2739-5.

Molina:2012:TSB

[MC12] Juan M. Molina and Steven J. Cooke. Trends in shark bycatch research: current status and research needs. Reviews in Fish Biology and Fisheries, 22(3):719–737, September 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-012-9269-3.

Maggs:2016:NDF

[MC16] J. Q. Maggs and P. D. Cowley. Nine decades of fish movement research in southern Africa: a synthesis of research and findings from 1928 to 2014. Reviews in Fish Biology and Fisheries, 26(3): 287–302, September 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-016-9425-2.

McDowall:1997:EDF

[McD97] R. M. McDowall. The evolution of diadromy in fishes (revisited) and its place in phylogenetic analysis. Reviews in Fish Biology and Fisheries, 7(4):443–462, December 1997. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1023/A:1018404331601.

McDowall:2002:OSF

[McD02] R. M. McDowall. The origin of the salmonid fishes: marine, freshwater, … or neither? Reviews in Fish Biology and Fisheries, 11(3):171–179, September 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1023/A:1020370328194.

McDowall:2003:VVN

[McD03] R. M. McDowall. Variation in vertebral number in galaxiid fishes, how fishes swim and a possible reason for pleomerism. Reviews in Fish Biology and Fisheries, 13(3):247–263, September 2003. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/B:RFBF.0000033121.97066.c1.

McDowall:2006:CWC

[McD06] R. M. McDowall. Crying wolf, crying foul, or crying shame: alien salmonids and a biodiversity crisis in the southern cooltemperate galaxioid fishes? Reviews in Fish Biology and Fisheries, 16(3–4):233–422, November 2006. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-006-9017-7.

McClanahan:2009:HSS

[MCD09a] Timothy R. McClanahan, Juan Carlos Castilla, and Omar Defeo. Healing small-scale fisheries by facilitating complex socioecological systems. Reviews in Fish Biology and Fisheries, 19 (1):33–47, March 2009. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-008-9088-8.

McDowall:2009:EHS

[McD09b] R. M. McDowall. Early hatch: a strategy for safe downstream larval transport in amphidromous gobies. Reviews in Fish Biology and Fisheries, 19(1):??, March 2009. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-008-9085-y.

McDowall:2010:WAE

[McD10] R. M. McDowall. Why be amphidromous: expatrial dispersal and the place of source and sink population dynamics? Reviews in Fish Biology and Fisheries, 20(1):87–100, March 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-009-9125-2.

Malpica-Cruz:2021:TCP

[MCFC21] Luis Malpica-Cruz, Stuart Fulton, and Isabelle M. Côté. Trying to collapse a population for conservation: commercial trade of a marine invasive species by artisanal fishers. Reviews in Fish Biology and Fisheries, 31(3):667–683, September 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-021-09660-0.

Montgomery:1995:BML

[MCH95] John Montgomery, Sheryl Coombs, and Matthew Halstead. Biology of the mechanosensory lateral line in fishes. Reviews in Fish Biology and Fisheries, 5(4):399–416, December 1995. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF01103813.

McLennan:1994:PAE

[McL94] Deborah A. McLennan. A phylogenetic approach to the evolution of fish behaviour. Reviews in Fish Biology and Fisheries, 4(4):430–460, December 1994. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00042889.

Mandelman:2008:SBD

[MCL08] John W. Mandelman, Peter W. Cooper, and Kerry M. Lagueux. Shark bycatch and depredation in the U.S. Atlantic pelagic longline fishery. Reviews in Fish Biology and Fisheries, 18 (4):??, November 2008. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-008-9084-z.

Milhomem:2012:CES

[MCN12] Susana Suely Rodrigues Milhomem, William Gareth Richard Crampton, and Cleusa Yoshiko Nagamachi. Chromosomal and electric signal diversity in three sympatric electric knifefish species (Gymnotus, gymnotidae) from the Central Amazon floodplain. Reviews in Fish Biology and Fisheries, 22(2):485– 497, June 2012. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-011-9239-1.

McQuinn:1997:MAH

[McQ97] Ian H. McQuinn. Metapopulations and the Atlantic herring. Reviews in Fish Biology and Fisheries, 7(3):297–329, September 1997. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1018491828875.

Matoso:2011:CPS

[MdAVA11] Daniele Aparecida Matoso, Vera Maria Fonseca de Almeida Val, and Roberto Ferreira Artoni. Chromosomal polymorphism in Steindachneridion melanodermatum Garavello, 2005 (Siluriformes, Pimelodidae): a reappraisal the existence of sex chromosome system in the species. Reviews in Fish Biology and Fisheries, 21(3):??, September 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-011-9201-2.

Marengo:2014:RBF

[MDF14] Michel Marengo, Eric D. H. Durieux, and Patrice Francour. A review of biology, fisheries and population structure of Dentex dentex (Sparidae). Reviews in Fish Biology and Fisheries, 24(4): 1065–1088, December 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-014-9363-9.

Macdonald:2020:IIO

[MDM20] Jed I. Macdonald, Russell N. Drysdale, and Gudrún Marteinsdóttir. Isolating the influence of ontogeny helps predict island-wide variability in fish otolith chemistry. Reviews in Fish Biology and Fisheries, 30(1):173–202, March 2020. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-019-09591-x.

Mullowney:2014:RFC

[MDR14] Darrell R. J. Mullowney, Earl G. Dawe, and George A. Rose. A review of factors contributing to the decline of Newfoundland and Labrador snow crab (Chionoecetes opilio). Reviews in Fish Biology and Fisheries, 24(2):639–657, June 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-014-9349-7.

Mitchell:2023:SDF

[MDV^{+}23] J. D. Mitchell, J. M. Drymon, J. Vardon, P. G. Coulson, C. A. Simpfendorfer, S. B. Scyphers, S. M. Kajiura, K. Hoel, S. Williams, K. L. Ryan, A. Barnett, M. R. Heupel, A. Chin, M. Navarro, T. Langlois, M. J. Ajemian, E. Gilman, E. Prasky, and G. Jackson. Shark depredation: future directions in research and management. Reviews in Fish Biology and Fisheries, 33(2):475–499, June 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-022-09732-9.

Millar:1999:ESS

[MF99] Russell B. Millar and Robert J. Fryer. Estimating the size-selection curves of towed gears, traps, nets and hooks. Reviews in Fish Biology and Fisheries, 9(1):89–116, March 1999. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1008838220001.

Miranda:2021:FGP

[MFL21] L. E. Miranda, Nicky M. Faucheux, and Kurt M. Lakin. Fishing gear performance nearshore is substantiated by spatial analyses. Reviews in Fish Biology and Fisheries, 31(4):977–987, December 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-021-09683-7.

Meager:2018:BDA

[MFS18] Justin J. Meager, Anders Fernö, and Jon Egil Skjæraasen. The behavioural diversity of Atlantic cod: insights into variability within and between individuals. Reviews in Fish Biology and Fisheries, 28(1):153–176, March 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-017-9505-y.

Manning:2019:RDF

[MFV19] C. G. Manning, S. J. Foster, and A. C. J. Vincent. A review of the diets and feeding behaviours of a family of biologically diverse marine fishes (family Syngnathidae). Reviews in Fish Biology and Fisheries, 29(2):197–221, June 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-019-09549-z.

Musyl:2018:PRF

[MG18] Michael K. Musyl and Eric L. Gilman. Post-release fishing mortality of blue (Prionace glauca) and silky shark (Carcharhinus falciformes) from a palauan-based commercial longline fishery. Reviews in Fish Biology and Fisheries, 28(3):567–586, September 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-018-9517-2.

Magellan:2015:ISS

[MGB15] Kit Magellan and Emili García-Berthou. Influences of size and sex on invasive species aggression and native species vulnerability: a case for modern regression techniques. Reviews in Fish Biology and Fisheries, 25(3):537–549, September 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-015-9391-0.

Muncaster:2023:GIS

[MGL^{+}23] S. Muncaster, A. Goikoetxea, P. M. Lokman, C. E. De Farias e Moraes, E. L. Damsteegt, J. Edgecombe, N. J. Gemmell, and E. V. Todd. Genes involved in sex differentiation, epigenetic reprogramming, and cell fate regulate sex change in a wrasse. Reviews in Fish Biology and Fisheries, 33(1):281–294, March 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-022-09755-2.

Martinez-Gonzalez:2021:DLA

[MGRRJ21] Ángel Trinidad Martínez-González, Víctor Manuel Ramírez-Rivera, and David Antonio Gómez Jáuregui. Deep learning algorithm as a strategy for detection an invasive species in uncontrolled environment. Reviews in Fish Biology and Fisheries, 31(4):909–922, December 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-021-09667-7.

Maciel:2018:TCF

[MGTD18] Patricia Oliveira Maciel, Fabiana Garcia, and Marcos Tavares-Dias. Trichodinidae in commercial fish in South America. Reviews in Fish Biology and Fisheries, 28(1):33–56, March 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-017-9490-1.

Medley:1993:CRF

[MGW93] P. A. Medley, G. Gaudian, and S. Wells. Coral reef fisheries stock assessment. Reviews in Fish Biology and Fisheries, 3(3): 242–285, September 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF00043930.

Mustonen:2020:HKA

[MH20] Tero Mustonen and Noora Huusari. How to know about waters? Finnish traditional knowledge related to waters and implications for management reforms. Reviews in Fish Biology and Fisheries, 30(4):699–718, December 2020. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-020-09619-7.

Martins:2012:CEG

[MHP12] Eduardo G. Martins, Scott G. Hinch, and David A. Patterson. Climate effects on growth, phenology, and survival of sockeye salmon (Oncorhynchus nerka): a synthesis of the current state of knowledge and future research directions. Reviews in Fish Biology and Fisheries, 22(4):887–914, December 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-012-9271-9.

Montanari:2016:IEB

[MHvH16] Stefano R. Montanari, Jean-Paul A. Hobbs, and Lynne van Herwerden. The importance of ecological and behavioural data in studies of hybridisation among marine fishes. Reviews in Fish Biology and Fisheries, 26(2):181–198, June 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-016-9420-7.

Moltschaniwskyj:2007:EWC

[MHW07] N. A. Moltschaniwskyj, K. Hall, and K. Warnke. Ethical and welfare considerations when using cephalopods as experimental animals. Reviews in Fish Biology and Fisheries, 17(2–3): 455–476, August 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-007-9056-8.

Moyano:2023:CMB

[MIA^{+}23a] Marta Moyano, Björn Illing, Anna Akimova, Katharina Alter, Valerio Bartolino, Gregor Börner, Catriona Clemmesen, Annegret Finke, Tomas Gröhsler, Paul Kotterba, Lina Livdane, Felix Mittermayer, Dorothee Moll, Lena von Nordheim, Myron A. Peck, Matthias Schaber, and Patrick Polte. Caught in the middle: bottom-up and top-down processes impacting recruitment in a small pelagic fish. Reviews in Fish Biology and Fisheries, 33(1):55–84, March 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-022-09739-2. See correction

[MIA^{+}23b].

Moyano:2023:CCM

[MIA^{+}23b] Marta Moyano, Björn Illing, Anna Akimova, Katharina Alter, Valerio Bartolino, Gregor Börner, Catriona Clemmesen, Annegret Finke, Tomas Gröhsler, Paul Kotterba, Lina Livdane, Felix Mittermayer, Dorothee Moll, Lena von Nordheim, Myron A. Peck, Matthias Schaber, and Patrick Polte. Correction: Caught in the middle: bottom up and top down processes impacting recruitment in a small pelagic fish. Reviews in Fish Biology and Fisheries, 33(1):85–87, March 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-022-09754-3. See

[MIA^{+}23a].

Millar:1999:BRE

[Mil99] Russell Millar. Book review: The Ecological Detective: Confronting Models with Data (Monographs in Population Biology, 28). By Ray Hilborn and Marc Mangel. Reviews in Fish Biology and Fisheries, 9(1):117–118, March 1999. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1023/A:1008861700420.

Milazzo:2012:PPU

[Mil12] Matteo J. Milazzo. Progress and problems in U.S. marine fisheries rebuilding plans. Reviews in Fish Biology and Fisheries, 22 (1):273–296, March 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-011-9219-5.

Misund:1997:UAM

[Mis97] Ole Arve Misund. Underwater acoustics in marine fisheries and fisheries research. Reviews in Fish Biology and Fisheries, 7(1):1–34, March 1997. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1018476523423.

Moser:2015:BPT

[MJM15] Mary L. Moser, Aaron D. Jackson, and Robert P. Mueller. Behavior and potential threats to survival of migrating lamprey ammocoetes and macrophthalmia. Reviews in Fish Biology and Fisheries, 25(1):103–116, March 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-014-9372-8.

Moltschaniwskyj:2007:P

[MJS07] Natalie Moltschaniwskyj, George Jackson, and Jayson Semmens. Preface. Reviews in Fish Biology and Fisheries, 17(2– 3):77–78, August 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-007-9049-7.

Mendes-Junior:2016:BEE

[MJSOF16] Raimundo Nonato G. Mendes-Junior, Júlio César Sá-Oliveira, and Stephen F. Ferrari. Biology of the electric eel, Electrophorus electricus, Linnaeus, 1766 (Gymnotiformes: Gymnotidae) on the floodplain of the ú River, eastern Amazonia. Reviews in Fish Biology and Fisheries, 26(1):83–91, March 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-015-9407-9.

Myszkowski:2010:WCG

[MKK10] Leszek Myszkowski, Ewa Kamler, and Sławomir Kwiatkowski. Weak compensatory growth makes short-term starvation an unsuitable technique to mitigate body deformities of Tinca tinca juveniles in intensive culture. Reviews in Fish Biology and Fisheries, 20(3):381–388, September 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-009-9134-1.

Matley:2023:MMA

[MKL^{+}23] J. K. Matley, N. V. Klinard, S. M. Larocque, M. F. McLean, J. W. Brownscombe, G. D. Raby, V. M. Nguyen, and A. P. Barbosa Martins. Making the most of aquatic animal tracking: a review of complementary methods to bolster acoustic telemetry. Reviews in Fish Biology and Fisheries, 33(1):35–54, March 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-022-09738-3.

Matamoros:2012:DNM

[MKS12] Wilfredo A. Matamoros, Brian R. Kreiser, and Jacob F. Schaefer. A delineation of Nuclear Middle America biogeographical provinces based on river basin faunistic similarities. Reviews in Fish Biology and Fisheries, 22(1):351–365, March 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-011-9232-8.

Muhling:2017:RLB

[MLC17] Barbara A. Muhling, John T. Lamkin, and Raul Laiz Carrion. Reproduction and larval biology in tunas, and the importance of restricted area spawning grounds. Reviews in Fish Biology and Fisheries, 27(4):697–732, December 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-017-9471-4.

Muir:2013:PPF

[MLK13] A. M. Muir, D. M. Leonard, and C. C. Krueger. Past, present and future of fishery management on one of the world’s last remaining pristine great lakes: Great Bear Lake, Northwest Territories, Canada. Reviews in Fish Biology and Fisheries, 23(3): 293–315, September 2013. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-012-9295-1.

Marttila:2019:SHR

[MLM19] Maare Marttila, Pauliina Louhi, and Timo Muotka. Synthesis of habitat restoration impacts on young-of-the-year salmonids in boreal rivers. Reviews in Fish Biology and Fisheries, 29(3): 513–527, September 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-019-09557-z.

Molony:2005:SEFa

[MLN05a] Brett W. Molony, R. Lenanton, and J. Norriss. Stock enhancement as a fisheries management tool. Reviews in Fish Biology and Fisheries, 13(4):409–432, January 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-005-1886-7.

Molony:2005:SEFb

[MLN05b] Brett W. Molony, R. Lenanton, and J. Norriss. Stock enhancement as a fisheries management tool. Reviews in Fish Biology and Fisheries, 13(4):409–432, January 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-004-1886-z.

Myers:2017:GSD

[MLP17] Bonnie J. E. Myers, Abigail J. Lynch, and Craig P. Paukert. Global synthesis of the documented and projected effects of climate change on inland fishes. Reviews in Fish Biology and Fisheries, 27(2):339–361, June 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-017-9476-z.

Medeiros-Leal:2023:PLB

[MLSP^{+}23] Wendell Medeiros-Leal, Régis Santos, Ualerson I. Peixoto, Morgan Casal-Ribeiro, Ana Novoa-Pabon, Michael F. Sigler, and Mário Pinho. Performance of length-based assessment in predicting small-scale multispecies fishery sustainability. Reviews in Fish Biology and Fisheries, 33(3):819–852, September 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09764-9.

Mandal:2013:RID

[MMB13] Banani Mandal, Arunava Mukherjee, and Samir Banerjee. A review on the ichthyofaunal diversity in mangrove based estuary of sundarbans. Reviews in Fish Biology and Fisheries, 23(3): 365–374, September 2013. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-012-9300-8.

Miranda:2022:MER

[MMB22] Rafael Miranda, Imanol Miqueleiz, and Monika Böhm. Monitoring extinction risk and threats of the world’s fishes based on the Sampled Red List Index. Reviews in Fish Biology and Fisheries, 32(3):975–991, September 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-022-09710-1.

Morales:2018:PGB

[MMF18] Millke Jasmine Arminini Morales, Fernando Fernandes Mendonça, and Fausto Foresti. Population genetics of the bigeye thresher shark Alopias superciliosus in the Atlantic and Indian Oceans: implications for conservation. Reviews in Fish Biology and Fisheries, 28(4):941–951, December 2018. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-018-9531-4.

Mullowney:2018:DSC

[MMG18] Darrell Mullowney, Corey Morris, and Svetlana Goryanina. Dynamics of snow crab (Chionoecetes opilio) movement and migration along the Newfoundland and Labrador and Eastern Barents Sea continental shelves. Reviews in Fish Biology and Fisheries, 28(2):435–459, June 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-017-9513-y.

Mendo:2022:ADI

[MMJ22] T. Mendo, J. Mendo, and M. A. James. Assessing discards in an illegal small-scale fishery using fisher-led reporting. Reviews in Fish Biology and Fisheries, 32(3):963–974, September 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-022-09708-9.

Mitchell:2018:SDC

[MML18] J. D. Mitchell, D. L. McLean, and T. J. Langlois. Shark depredation in commercial and recreational fisheries. Reviews in Fish Biology and Fisheries, 28(4):715–748, December 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-018-9528-z.

Mustonen:2022:CWO

[MMM22a] Tero Mustonen, Kimberley H. Maxwell, and Eero Murtomäki. Correction to: Who is the ocean? Preface to the Future Seas 2030 special issue. Reviews in Fish Biology and Fisheries, 32(1):17, March 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-021-09657-9. See [MMM22b].

Mustonen:2022:WOP

[MMM22b] Tero Mustonen, Kimberley H. Maxwell, and Eero Murtomäki. Who is the ocean? Preface to the Future Seas 2030 special issue. Reviews in Fish Biology and Fisheries, 32(1):9–16, March 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-021-09655-x. See correction [MMM22a].

Mackinson:1998:PVC

[MN98] Steven Mackinson and Leif Nottestad. Points of view: Combining local and scientific knowledge. Reviews in Fish Biology and Fisheries, 8(4):481–490, December 1998. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1023/A:1008847106984.

Mendonca:2011:PGP

[MOF11] Fernando F. Mendonça, Claudio Oliveira, and Fausto Foresti. Phylogeography and genetic population structure of Caribbean sharpnose shark Rhizoprionodon porosus. Reviews in Fish Biology and Fisheries, 21(4):799–814, December 2011. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-011-9210-1.

Moksness:1993:OME

[Mok93] Erlend Moksness. Otolith microstructure examination and analysis. Reviews in Fish Biology and Fisheries, 3(3):293–294, September 1993. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/BF00043932.

Moller:1992:ETF

[Mol92] Peter Moller. Electrocommunication in teleost fishes: Behavior and experiments. Reviews in Fish Biology and Fisheries, 2 (1):84–85, March 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF00042918.

Moore:2012:EPA

[Moo12] Alec B. M. Moore. Elasmobranchs of the Persian (Arabian) Gulf: ecology, human aspects and research priorities for their improved management. Reviews in Fish Biology and Fisheries, 22(1):35–61, March 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-011-9222-x.

McWilliam:2007:SLD

[MP07a] Paulette S. McWilliam and Bruce F. Phillips. Spiny lobster development: mechanisms inducing metamorphosis to the puerulus: a review. Reviews in Fish Biology and Fisheries, 17(4): 615–632, November 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-007-9067-5.

Moltschaniwskyj:2007:SAS

[MP07b] Natalie A. Moltschaniwskyj and Gretta T. Pecl. Spawning aggregations of squid (Sepioteuthis australis) populations: a continuum of ‘microcohorts’. Reviews in Fish Biology and Fisheries, 17(2–3):??, August 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-006-9025-7.

McMillan:2016:RNI

[MP16] L. Jane McMillan and Kerry Prosper. Remobilizing netukulimk: indigenous cultural and spiritual connections with resource stewardship and fisheries management in Atlantic Canada. Reviews in Fish Biology and Fisheries, 26(4):629–647, December 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-016-9433-2.

Machova:2010:TDE

[MPK10] Jana Máchová, Miroslav Prokeš, and Hana Kroupová. Toxicity of Diazinon 60 EC for embryos and larvae of tench, Tinca tinca (L.). Reviews in Fish Biology and Fisheries, 20(3):409– 415, September 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-009-9148-8.

Mariguela:2011:RCM

[MPO11] T. C. Mariguela, L. R. S. Paiva, and C. Oliveira. 5S rDNA chromosomal mapping and COI sequence analysis reveal differentiation among distinct populations of a characid fish Serrapinnus notomelas. Reviews in Fish Biology and Fisheries, 21(4): 779–788, December 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-011-9208-8.

Miller:2018:LHM

[MR18] Michael J. Miller and Tony Robinet. Life history and morphology of eel larvae in the Gulf of Guinea of western Africa: revisiting Jacques Blache’s research (1960–1977) 40 years later. Reviews in Fish Biology and Fisheries, 28(2):355–379, June 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-017-9512-z.

Murua:2017:FVS

[MRMJJ17] Hilario Murua, Enrique Rodriguez-Marin, and María Jose Juan-Jordá. Fast versus slow growing tuna species: age, growth, and implications for population dynamics and fisheries management. Reviews in Fish Biology and Fisheries, 27(4):733–773, December 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-017-9474-1.

Mayden:2002:CSB

[MS02] Richard L. Mayden and Andrew M. Simons. Crevice spawning behavior in Dionda dichroma, with comments on the evolution of spawning modes in North American shiners (Teleostei: Cyprinidae). Reviews in Fish Biology and Fisheries, 12(2– 3):327–337, June 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1025056809814.

Melo:2007:DDS

[MS07] Y. Melo and W. H. H. Sauer. Determining the daily spawning cycle of the chokka squid, Loligo reynaudii off the South African Coast. Reviews in Fish Biology and Fisheries, 17(2– 3):247–257, August 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-006-9034-6.

Miller:2013:RCO

[MS13] M. E. Miller and J. Stewart. Reproductive characteristics of the ocean leatherjacket, Nelusetta ayraudi. Reviews in Fish Biology and Fisheries, 23(1):87–101, March 2013. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-012-9277-3.

Munroe:2014:DSE

[MSH14] S. E. M. Munroe, C. A. Simpfendorfer, and M. R. Heupel. Defining shark ecological specialisation: concepts, context, and examples. Reviews in Fish Biology and Fisheries, 24(1):317– 331, March 2014. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-013-9333-7.

Mercado-Silva:2002:VFB

[MSLN02] Norman Mercado-Silva, John D. Lyons, and Martina Medina Nava. Validation of a fish-based index of biotic integrity for streams and rivers of central Mexico. Reviews in Fish Biology and Fisheries, 12(2–3):179–191, June 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1023/A:1025099711746.

Miehls:2020:FBT

[MSM20] Scott Miehls, Paul Sullivan, and Rodney McDonald. The future of barriers and trapping methods in the sea lamprey (Petromyzon marinus) control program in the Laurentian Great Lakes. Reviews in Fish Biology and Fisheries, 30(1):1–24, March 2020. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-019-09587-7.

Mustafa:2003:SSS

[MSR03] S. Mustafa, S. Saad, and R. A. Rahman. Species studies in sea ranching: an overview and economic perspectives. Reviews in Fish Biology and Fisheries, 13(2):Article 165, June 2003. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/B:RFBF.0000019478.17950.ab.

Mangubhai:2014:STC

[MSR14] Sangeeta Mangubhai, Ayron M. Strauch, and Randi D. Rotjan. Short-term changes of fish assemblages observed in the nearpristine reefs of the Phoenix Islands. Reviews in Fish Biology and Fisheries, 24(2):505–518, June 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-013-9327-5.

Mustonen:2023:TKS

[MSV^{+}23] Tero Mustonen, Antoine Scherer, Brie Van Dam, Stefan Milkowski, and Noora Huusari. Traditional knowledge in special fisheries: the case of Puruvesi vendace and seining. Reviews in Fish Biology and Fisheries, 33(3):649–667, September 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-022-09728-5.

Mustafa:2003:P

[MT03] Saleem Mustafa and Nobuhiko Taniguchi. Preface. Reviews in Fish Biology and Fisheries, 13(2):3–4, June 2003. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1023/B:RFBF.0000019578.32449.8f.

Maynard:2014:UHT

[MT14] Desmond J. Maynard and Joan G. Trial. The use of hatchery technology for the conservation of Pacific and Atlantic salmon. Reviews in Fish Biology and Fisheries, 24(3):803–817, September 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-013-9341-7.

Mustonen:2018:BLL

[MT18] Tero Mustonen and Tarmo Tossavainen. Brook lampreys of life: towards holistic monitoring of boreal aquatic habitats using ‘subtle signs’ and oral histories. Reviews in Fish Biology and Fisheries, 28(3):657–665, September 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-018-9527-0.

Melbourne-Thomas:2022:PBA

[MTAP22] Jess Melbourne-Thomas, Asta Audzijonyte, and Gretta T. Pecl. Poleward bound: adapting to climate-driven species redistribution. Reviews in Fish Biology and Fisheries, 32(1):231–251, March 2022. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-021-09641-3.

McCormack:2019:USI

[MTC19] Stacey A. McCormack, Rowan Trebilco, and Andrew Constable. Using stable isotope data to advance marine food web modelling. Reviews in Fish Biology and Fisheries, 29(2):277–296, June 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-019-09552-4.

Mak:2021:IRD

[MTL21] Yanny K. Y. Mak, Lily S. R. Tao, and Kenneth M. Y. Leung. Initial recovery of demersal fish communities in coastal waters of Hong Kong, South China, following a trawl ban. Reviews in Fish Biology and Fisheries, 31(4):989–1007, December 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-021-09685-5.

Martins:2022:ELS

[MTM22] Samir Martins, Manjula Tiwari, and Adolfo Marco. Evaluating loggerhead sea turtle (Caretta caretta) bycatch in the small-scale fisheries of Cabo Verde. Reviews in Fish Biology and Fisheries, 32(3):1001–1015, September 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-022-09718-7.

Marcos:2015:LTE

[MTPR15] Concepción Marcos, Inmaculada Torres, and Angel Pérez-Ruzafa. Long term evolution of fisheries in a coastal lagoon related to changes in lagoon ecology and human pressures. Reviews in Fish Biology and Fisheries, 25(4):689–713, December 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-015-9397-7.

Mulcahy:2009:BRJ

[Mul09] Daniel M. Mulcahy. Book review: E. J. Branson (ed): Fish Welfare. Reviews in Fish Biology and Fisheries, 19(1):125– 126, March 2009. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-008-9098-6.

Mulcahy:2010:BRG

[Mul10] Daniel M. Mulcahy. Book review: L. G. Ross and B. Ross: Anaesthetic and sedative techniques for aquatic animals (Third Edition). Reviews in Fish Biology and Fisheries, 20(1):139– 140, March 2010. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-009-9124-3.

Mulcahy:2011:AUD

[Mul11] Daniel M. Mulcahy. Antibiotic use during the intracoelomic implantation of electronic tags into fish. Reviews in Fish Biology and Fisheries, 21(1):83–96, March 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-010-9190-6.

Munn:1993:IDF

[Mun93] C. B. Munn. Infectious diseases of fish. Reviews in Fish Biology and Fisheries, 3(4):377–378, December 1993. CODEN RF-BFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00043390.

Munro:1995:PFC

[Mun95] A. D. Munro. Possible functions of the caudal neurosecretory system. Reviews in Fish Biology and Fisheries, 5(4):447–454, December 1995. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/BF01103815.

Mustafa:2003:SES

[Mus03] S. Mustafa. Stock enhancement and sea ranching: objectives and potential. Reviews in Fish Biology and Fisheries, 13(2): 141–149, June 2003. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/B:RFBF.0000019476.87730.3b.

Muller:1996:ISH

[MV96] Ulrike K. Müller and John J. Videler. Inertia as a ‘safe harbour’: do fish larvae increase length growth to escape viscous drag? Reviews in Fish Biology and Fisheries, 6(3):353–360, September 1996. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00122586.

Maceda-Veiga:2013:TCF

[MV13] Alberto Maceda-Veiga. Towards the conservation of freshwater fish: Iberian rivers as an example of threats and management practices. Reviews in Fish Biology and Fisheries, 23 (1):1–22, March 2013. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-012-9275-5.

Munguia-Vega:2018:EGD

[MVGW18] Adrian Munguia-Vega, Alison L. Green, and Amy Hudson Weaver. Ecological guidelines for designing networks of marine reserves in the unique biophysical environment of the Gulf of California. Reviews in Fish Biology and Fisheries, 28(4): 749–776, December 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-018-9529-y.

Mommsen:1999:CTD

[MVM99] Thomas P. Mommsen, Mathilakath M. Vijayan, and Thomas W. Moon. Cortisol in teleosts: dynamics, mechanisms of action, and metabolic regulation. Reviews in Fish Biology and Fisheries, 9(3):211–268, September 1999. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1023/A:1008924418720.

Maschette:2023:CSR

[MWP^{+}23] D. Maschette, S. Wotherspoon, A. Polanowski, B. Deagle, D. Welsford, and P. Ziegler. Circumpolar sampling reveals high genetic connectivity of Antarctic toothfish across their spatial distribution. Reviews in Fish Biology and Fisheries, 33(1): 295–310, March 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-023-09756-9.

Miller:2022:SES

[MWS22] Erin Miller, Christy N. Wails, and James Sulikowski. It’s a shark-eat-shark world, but does that make for bigger pups? A comparison between oophagous and non-oophagous viviparous sharks. Reviews in Fish Biology and Fisheries, 32(4):1019– 1033, December 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-022-09707-w.

Myers:1998:WDE

[Mye98] Ransom A. Myers. When do environment–recruitment correlations work? Reviews in Fish Biology and Fisheries, 8(3): 285–305, September 1998. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1008828730759.

Mylonas:2000:UGD

[MZ00] Constantinos C. Mylonas and Yonathan Zohar. Use of GnRHadelivery systems for the control of reproduction in fish. Reviews in Fish Biology and Fisheries, 10(4):463–491, December 2000. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1012279814708.

Nash:2022:CDA

[NAP22a] Kirsty L. Nash, Karen Alexander, and Gretta T. Pecl. Correction to: Developing achievable alternate futures for key challenges during the UN Decade of Ocean Science for Sustainable Development. Reviews in Fish Biology and Fisheries, 32(1): 37, March 2022. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-021-09635-1. See [NAP22b].

Nash:2022:DAA

[NAP22b] Kirsty L. Nash, Karen Alexander, and Gretta T. Pecl. Developing achievable alternate futures for key challenges during the UN Decade of Ocean Science for Sustainable Development. Reviews in Fish Biology and Fisheries, 32(1):19–36, March 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-020-09629-5. See correction [NAP22a].

Novaglio:2022:DAT

[NBF22] Camilla Novaglio, Narissa Bax, and Elizabeth A. Fulton. Deep aspirations: towards a sustainable offshore Blue Economy. Reviews in Fish Biology and Fisheries, 32(1):209–230, March 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-020-09628-6.

Nalmpanti:2023:MMF

[NCMT23] Melina Nalmpanti, Anna Chrysafi, Jessica J. Meeuwig, and Athanassios C. Tsikliras. Monitoring marine fishes using underwater video techniques in the Mediterranean Sea. Reviews in Fish Biology and Fisheries, 33(4):1291–1310, December 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09799-y.

Needle:2001:RMD

[Nee01] Coby L. Needle. Recruitment models: diagnosis and prognosis. Reviews in Fish Biology and Fisheries, 11(2):95–111, June 2001. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1015208017674.

Neher:1996:FQQ

[Neh96] Philip A. Neher. Fishing quota quality and fishery performance. Reviews in Fish Biology and Fisheries, 6(1):113–116, March 1996. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00058524.

Nelson:1992:FGF

[Nel92] Joseph S. Nelson. A field guide to freshwater fishes: North America North of Mexico. Reviews in Fish Biology and Fisheries, 2(3):279–280, September 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00045047.

Nelson:1999:EIS

[Nel99] Joseph S. Nelson. Editorial and introduction: The species concept in fish biology. Reviews in Fish Biology and Fisheries, 9(4): 277–280, December 1999. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1008976411673.

Nielsen:2002:F

[Nie02] Jennifer L. Nielsen. Foreword. Reviews in Fish Biology and Fisheries, 12(2–3):111, June 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1023/A:1025093902577.

Nislow:2000:SRI

[Nis00] Keith H. Nislow. Symposium review: International symposium on the implications of salmonid growth variation. Reviews in Fish Biology and Fisheries, 10(4):521–527, December 2000. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1012234511093.

Neves:2021:SFP

[NKF21] Mayara P. Neves, Pavel Kratina, and Clarice B. Fialho. Seasonal feeding plasticity can facilitate coexistence of dominant omnivores in Neotropical streams. Reviews in Fish Biology and Fisheries, 31(2):417–432, June 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-021-09648-w.

Neto:2012:DEP

[NLFM12] C. C. Motta Neto, P. A. Lima-Filho, and W. F. Molina. Differentiated evolutionary pathways in Haemulidae (Perciformes): karyotype stasis versus morphological differentiation. Reviews in Fish Biology and Fisheries, 22(2):457–465, June 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-011-9236-4.

Nikolic:2017:RAT

[NMF17] Natacha Nikolic, Gilles Morandeau, and Alain Fonteneau. Review of albacore tuna, Thunnus alalunga, biology, fisheries and management. Reviews in Fish Biology and Fisheries, 27(4): 775–810, December 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-016-9453-y.

Nandi:2013:SFF

[NMS13] Sudarshana Nandi, Sandip Majumder, and Surjya Kumar Saikia. Small freshwater fish species (SFFs) culture: issues from nutrient security, carp-SFF integration and feeding ecology. Reviews in Fish Biology and Fisheries, 23(3):283–291, September 2013. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-012-9294-2.

Northcote:1995:CBM

[Nor95] Thomas G. Northcote. Comparative biology and management of Arctic and European grayling (Salmonidae, Thymallus). Reviews in Fish Biology and Fisheries, 5(2):141–194, June 1995. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00179755.

Northcote:1996:PAL

[Nor96] T. G. Northcote. Peter anthony larkin. Reviews in Fish Biology and Fisheries, 6(4):375–378, December 1996. CODEN RF-BFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00164322.

Norris:2002:BRJ

[Nor02] David O. Norris. Book review: Jezierska, B. and Witeska, M. (2001) Metal Toxicity to Fish. Reviews in Fish Biology and Fisheries, 11(3):279, September 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1023/A:1020380730689.

Nordlie:2003:FCE

[Nor03] Frank G. Nordlie. Fish communities of estuarine salt marshes of eastern North America, and comparisons with temperate estuaries of other continents. Reviews in Fish Biology and Fisheries, 13(3):281–325, September 2003. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1023/B:RFBF.0000033050.
51699.84.

Nordlie:2006:PET

[Nor06] Frank G. Nordlie. Physicochemical environments and tolerances of cyprinodontoid fishes found in estuaries and salt marshes of eastern North America. Reviews in Fish Biology and Fisheries, 16(1):51–106, February 2006. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-006-9003-0.

Nordlie:2009:EIR

[Nor09] Frank G. Nordlie. Environmental influences on regulation of blood plasma/serum components in teleost fishes: a review. Reviews in Fish Biology and Fisheries, 19(4):481–564, December 2009. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-009-9131-4.

Nordlie:2012:LHC

[Nor12] Frank G. Nordlie. Life-history characteristics of eleotrid fishes of the western hemisphere, and perils of life in a vanishing environment. Reviews in Fish Biology and Fisheries, 22(1): 189–224, March 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-011-9229-3.

Nordlie:2014:IBM

[Nor14] Frank G. Nordlie. Influences of body mass, temperature, oxygen tension, and salinity on respiratory oxygen consumption of cyprinodontoid fishes of three families. Reviews in Fish Biology and Fisheries, 24(1):269–315, March 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-013-9331-9.

Naumowicz:2017:ICM

[NPS17] Karolina Naumowicz, Joanna Pajdak, and Józef Szarek. Intracohort cannibalism and methods for its mitigation in cultured freshwater fish. Reviews in Fish Biology and Fisheries, 27 (1):193–208, March 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-017-9465-2. See correction

[DPS22].

Nunez-Riboni:2023:RIS

[NRCD^{+}23] Ismael Núñez-Riboni, Gersom Costas, Rabea Diekmann, Jens Ulleweit, and Matthias Kloppmann. Reviewing and improving spatiotemporal modeling approaches for mackerel’s total annual egg production. Reviews in Fish Biology and Fisheries, 33(4): 1523–1546, December 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-023-09795-2.

Nyboer:2023:GCN

[NRJ^{+}23] Elizabeth A. Nyboer, Andrea J. Reid, Amanda L. Jeanson, Rachel Kelly, Mary Mackay, Jenny House, Sarah M. Arnold, Paul W. Simonin, Mary Grace C. Sedanza, Emma D. Rice, T. E. Angela L. Quiros, Andrea Pierucci, Kelly Ortega-Cisneros, Julia N. Nakamura, Valentina Melli, Stella Mbabazi, Mariana S. L. Martins, Anne Brigette B. Ledesma, Clara Obregón, Chepkemboi K. Labatt, Andrew N. Kadykalo, Michael Heldsinger, Madeline E. Green, Jessica L. Fuller, Milagros Franco-Meléndez, Matthew J. Burnett, Jessica A. Bolin, Solange Andrade-Vera, and Steven J. Cooke. Goals, challenges, and next steps in transdisciplinary fisheries research: perspectives and experiences from early-career researchers. Reviews in Fish Biology and Fisheries, 33(2):349–374, June 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-022-09719-6.

Nunn:2012:FEL

[NTC12] A. D. Nunn, L. H. Tewson, and I. G. Cowx. The foraging ecology of larval and juvenile fishes. Reviews in Fish Biology and Fisheries, 22(2):377–408, June 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-011-9240-8.

Noleto:2012:VBC

[NVA12] Rafael Bueno Noleto, Marcelo Ricardo Vicari, and Roberto Ferreira Artoni. Variable B chromosomes frequencies between males and females of two species of pufferfishes (Tetraodontiformes). Reviews in Fish Biology and Fisheries, 22(1):343–349, March 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-011-9231-9.

Narvaez:2021:NPR

[NVH21] Pauline Narvaez, David Brendan Vaughan, and Kate Suzanne Hutson. New perspectives on the role of cleaning symbiosis in the possible transmission of fish diseases. Reviews in Fish Biology and Fisheries, 31(2):233–251, June 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-021-09642-2.

Nash:2022:OSF

[NvPV22] Kirsty L. Nash, Ingrid van Putten, and Joanna Vince. Oceans and society: feedbacks between ocean and human health. Reviews in Fish Biology and Fisheries, 32(1):161–187, March 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-021-09669-5.

Nottage:2007:CDC

[NWG07] Jonathan D. Nottage, Ronald J. West, and Ken Graham. Cephalopod diversity in commercial fisheries landings of New South Wales, Australia. Reviews in Fish Biology and Fisheries, 17(2–3):271–281, August 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-006-9032-8.

Nevatte:2021:CPP

[NWG21] Ryan J. Nevatte, Jane E. Williamson, and Michael R. Gillings. Contrasting patterns of population structure in commercially fished sawsharks from southern Australian waters. Reviews in Fish Biology and Fisheries, 31(2):359–379, June 2021. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-021-09640-4.

Newman:2016:RLH

[NWO16] Stephen J. Newman, Ashley J. Williams, and Joseph M. O’Malley. Review of the life history characteristics, ecology and fisheries for deep-water tropical demersal fish in the Indo-Pacific region. Reviews in Fish Biology and Fisheries, 26(3): 537–562, September 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-016-9442-1.

Okey:2014:ECC

[OAJ14] Thomas A. Okey, Hussein M. Alidina, and Sabine Jessen. Effects of climate change on Canada’s Pacific marine ecosystems: a summary of scientific knowledge. Reviews in Fish Biology and Fisheries, 24(2):519–559, June 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-014-9342-1.

Oukhattar:2008:ICG

[OBS08] Laila Oukhattar, Tarik Baibai, and Abdelaziz Soukri. Isolation and characterization of glyceraldehyde-3-phosphate dehydrogenase from the common octopus (Octopus vulgaris Cuvier, 1797). Reviews in Fish Biology and Fisheries, 18(3):263–271, August 2008. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-007-9074-6.

Oyugi:2012:TDF

[OCB12] Dalmas O. Oyugi, Julien Cucherousset, and J. Robert Britton. Temperature-dependent feeding interactions between two invasive fishes competing through interference and exploitation. Reviews in Fish Biology and Fisheries, 22(2):499–508, June 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-011-9243-5.

Oldenburg:2011:HJS

[OCE11] Eric W. Oldenburg, Alison H. Colotelo, and M. Brad Eppard. Holding of juvenile salmonids for surgical implantation of electronic tags: a review and recommendations. Reviews in Fish Biology and Fisheries, 21(1):35–42, March 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-010-9186-2.

Owen:2010:LCI

[ODS10] Matthew A. G. Owen, Simon J. Davies, and Katherine A. Sloman. Light colour influences the behaviour and stress physiology of captive tench (Tinca tinca). Reviews in Fish Biology and Fisheries, 20(3):375–380, September 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-009-9150-1.

OFarrell:2017:EMG

[OGR17] Halie O’Farrell, Arnaud Grüss, and Kenneth A. Rose. Ecosystem modeling in the Gulf of Mexico: current status and future needs to address ecosystem-based fisheries management and restoration activities. Reviews in Fish Biology and Fisheries, 27(3):587–614, September 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-017-9482-1.

OShea:2007:NSO

[OJB07] S. O’Shea, G. Jackson, and K. S. Bolstad. The nomenclatural status, ontogeny and morphology of Pholidoteuthis massyae (Pfeffer, 1912) new comb (Cephalopoda: Pholidoteuthidae). Reviews in Fish Biology and Fisheries, 17(2–3):425–435, August 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-007-9047-9.

Ojaveer:2010:ELT

[OK10] Evald Ojaveer and Margers Kalejs. Ecology and long-term forecasting of sprat (Sprattus sprattus balticus) stock in the Baltic Sea: a review. Reviews in Fish Biology and Fisheries, 20 (2):203–217, June 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-009-9130-5.

Ojaveer:2012:LTP

[OK12] Evald Ojaveer and Margers Kalejs. Long-term prediction on Baltic fish stocks based on periodicity of solar activity. Reviews in Fish Biology and Fisheries, 22(3):683–693, September 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-012-9264-8.

OReilly:2014:UMD

[OK14] Patrick T. O’Reilly and Christine C. Kozfkay. Use of microsatellite data and pedigree information in the genetic management of two long-term salmon conservation programs. Reviews in Fish Biology and Fisheries, 24(3):819–848, September 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-014-9347-9.

Orbe-Mendoza:2002:LPF

[OMAGL02] Alma Araceli Orbe-Mendoza, Javier Acevedo-García, and John Lyons. Lake Pátzcuaro fishery management plan. Reviews in Fish Biology and Fisheries, 12(2–3):207–217, June 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1025087705940.

Olyott:2007:SPB

[OSB07] L. J. H. Olyott, W. H. H. Sauer, and A. J. Booth. Spatial patterns in the biology of the chokka squid, Loligo reynaudii on the Agulhas Bank, South Africa. Reviews in Fish Biology and Fisheries, 17(2–3):159–172, August 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-006-9027-5.

Ogier:2023:IIC

[OSB^{+}23] Emily M. Ogier, David C. Smith, Sian Breen, Caleb Gardner, Daniel J. Gaughan, Harry K. Gorfine, Alistair J. Hobday, Natalie Moltschaniwskyj, Ryan Murphy, Thor Saunders, Mike Steer, and James Woodhams. Initial impacts of the COVID-19 pandemic on Australian fisheries production, research organisations and assessment: shocks, responses and implications for decision support and resilience. Reviews in Fish Biology and Fisheries, 33(2):513–534, June 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-023-09760-z.

Oficialdegui:2020:OCH

[OSC20] Francisco J. Oficialdegui, Marta I. Sánchez, and Miguel Clavero. One century away from home: how the red swamp crayfish took over the world. Reviews in Fish Biology and Fisheries, 30 (1):121–135, March 2020. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-020-09594-z.

Orlov:2024:UGP

[OV24] Alexei M. Orlov and Igor V. Volvenko. Uninvited guests and permanent residents: long-term changes in the distribution and abundance of the five most common sharks in the northwestern Pacific. Reviews in Fish Biology and Fisheries, 34(2):703– 729, June 2024. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-024-09834-6.

OConnell:1997:MDF

[OW97] Michael O’Connell and Jonathan M. Wright. Microsatellite DNA in fishes. Reviews in Fish Biology and Fisheries, 7(3): 331–363, September 1997. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1018443912945.

Orsi:2004:JCS

[OWW04] Joseph A. Orsi, Alex C. Wertheimer, and Bruce L. Wing. Juvenile chum salmon consumption of zooplankton in marine waters of southeastern Alaska: a bioenergetics approach to implications of hatchery stock interactions. Reviews in Fish Biology and Fisheries, 14(3):335–359, September 2004. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-004-3813-8.

Petza:2023:CAB

[PAK^{+}23] Dimitra Petza, Panagiotis Anastopoulos, Stefanos Kalogirou, Marta Coll, Serge Garcia, Michel Kaiser, Nikoletta Koukourouvli, Irene Lourdi, Jake Rice, Marija Sciberras, and Stelios Katsanevakis. Contribution of area-based fisheries management measures to fisheries sustainability and marine conservation: a global scoping review. Reviews in Fish Biology and Fisheries, 33(4):1049–1073, December 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-023-09780-9.

Pecl:2022:FSP

[PAN22] Gretta T. Pecl, Karen A. Alexander, and Kirsty L. Nash. Future Seas 2030: pathways to sustainability for the UN Ocean Decade and beyond. Reviews in Fish Biology and Fisheries, 32(1):1–7, March 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-022-09705-y.

Parkes:1992:FSO

[Par92] Graeme Parkes. Fishes of the Southern Ocean. Reviews in Fish Biology and Fisheries, 2(4):344–345, December 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00043525.

Pansonato-Alves:2013:MFR

[PASF13] José Carlos Pansonato-Alves, Érica Alves Serrano, and Fausto Foresti. Mapping five repetitive DNA classes in sympatric species of Hypostomus (Teleostei: Siluriformes: Loricariidae): analysis of chromosomal variability. Reviews in Fish Biology and Fisheries, 23(4):477–489, December 2013. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-013-9303-0.

Patterson:1992:FSP

[Pat92] Kenneth Patterson. Fisheries for small pelagic species: an empirical approach to management targets. Reviews in Fish Biology and Fisheries, 2(4):321–338, December 1992. CODEN RF-BFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00043521.

Puncher:2015:MBT

[PAT15] Gregory Neils Puncher, Francisco Alemany, and Fausto Tinti. Misidentification of bluefin tuna larvae: a call for caution and taxonomic reform. Reviews in Fish Biology and Fisheries, 25(3): 485–502, September 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-015-9390-1.

Pauly:1996:IAB

[Pau96] Daniel Pauly. ITQ: the assumptions behind a meme. Reviews in Fish Biology and Fisheries, 6(1):109–112, March 1996. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00058523.

Pauly:1997:PFM

[Pau97] Daniel Pauly. Putting fisheries management back in places. Reviews in Fish Biology and Fisheries, 7(1):125–127, March 1997. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1018423408402.

Pauly:1998:BOO

[Pau98] Daniel Pauly. Beyond our original horizons: the tropicalization of Beverton and Holt. Reviews in Fish Biology and Fisheries, 8(3):307–334, September 1998. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1023/A:1008863215253.

Pauers:2010:SCS

[Pau10] Michael J. Pauers. Species concepts, speciation, and taxonomic change in the Lake Malawi mbuna, with special reference to the genus Labeotropheus Ahl 1927 (Perciformes: Cichlidae). Reviews in Fish Biology and Fisheries, 20(2):187–202, June 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-009-9128-z.

Pavey:2005:BR

[Pav05] Scott A. Pavey. Book review. Reviews in Fish Biology and Fisheries, 15(1–2):163, February 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-005-7203-7.

Pawson:1999:BRB

[Paw99] Mike Pawson. Book review: Biology of Farmed Fish (Sheffield Biological Sciences, 1). Edited by K. D. Black and A. D. Pickering. Reviews in Fish Biology and Fisheries, 9(2):206, June 1999. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1008912112838.

Pereira:2017:RCF

[PAW17] Larissa Strictar Pereira, Angelo Antonio Agostinho, and Kirk O. Winemiller. Revisiting cannibalism in fishes. Reviews in Fish Biology and Fisheries, 27(3):499–513, September 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-017-9469-y.

Pitcher:1998:PVP

[PB98] Tony J. Pitcher and Alida Bundy. Points of view: The problems of estimating potential yield in data-sparse lake fisheries: a new family of approximate models with examples from African lakes. Reviews in Fish Biology and Fisheries, 8(4):473–480, December 1998. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1008892007501.

Poteaux:2000:AMM

[PBB00] C. Poteaux, P. Berrebi, and F. Bonhomme. Allozymes, mtDNA and microsatellites study introgression in a stocked trout population in France. Reviews in Fish Biology and Fisheries, 10(3): 281–292, September 2000. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1016687714621.

Pombo:2012:SFF

[PBC12] Ana Pombo, Manuel Blasco, and Vicente Climent. The status of farmed fish hearts: an alert to improve health and production in three Mediterranean species. Reviews in Fish Biology and Fisheries, 22(3):779–789, September 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-012-9259-5.

Pereira:2015:ODC

[PBF15] Pedro Henrique Cipresso Pereira, Breno Barros, and Beatrice Padovani Ferreira. Ontogenetic diet changes and food partitioning of Haemulon spp. coral reef fishes, with a review of the genus diet. Reviews in Fish Biology and Fisheries, 25(1): 245–260, March 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-014-9378-2.

Petrere:2004:RLC

[PBG04] Miguel Petrere, Jr., Ronaldo Borges Barthem, and Bernardo Corrales Gómez. Review of the large catfish fisheries in the upper Amazon and the stock depletion of piraíba (Brachyplatystoma filamentosum Lichtenstein). Reviews in Fish Biology and Fisheries, 14(4):403–414, December 2004. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-004-8362-7.

Paiz:2014:EBA

[PBM14] Leonardo Marcel Paiz, Lucas Baumgartner, and Vladimir Pavan Margarido. Evolutionary and biogeographical approach on Australoheros angiru (Cichlidae) from lagoons in a dividing plateau between the basins of the Iguassu River and the Uruguay River, Brazil. Reviews in Fish Biology and Fisheries, 24(1):399–407, March 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-013-9321-y.

Peres:2012:IDH

[PBMF12] Wellington Adriano Moreira Peres, Luiz Antonio Carlos Bertollo, and Orlando Moreira-Filho. Invasion, dispersion and hybridization of fish associated to river transposition: karyotypic evidence in Astyanax “bimaculatus group” (Characiformes: Characidae). Reviews in Fish Biology and Fisheries, 22(2):519–526, June 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-011-9246-2.

Pinder:2019:MTS

[PBR19] Adrian C. Pinder, J. Robert Britton, and Rajeev Raghavan. Mahseer (Tor spp.) fishes of the world: status, challenges and opportunities for conservation. Reviews in Fish Biology and Fisheries, 29(2):417–452, June 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-019-09566-y.

Potts:2014:OWA

[PBS14] Warren M. Potts, Anthony J. Booth, and Warwick H. H. Sauer. Ocean warming affects the distribution and abundance of resident fishes by changing their reproductive scope. Reviews in Fish Biology and Fisheries, 24(2):493–504, June 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-013-9329-3.

Perazzo:2019:MDB

[PCG19] Giselle Xavier Perazzo, Fabiano Corrêa, and Adriana Gava. Morphological differences between an artificial lentic and adjacent lotic environments in a characid species. Reviews in Fish Biology and Fisheries, 29(4):935–949, December 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-019-09582-y.

Pratchett:2017:ECC

[PCW17] Morgan S. Pratchett, Darren S. Cameron, and David H. Williamson. Effects of climate change on coral grouper (Plectropomus spp.) and possible adaptation options. Reviews in Fish Biology and Fisheries, 27(2):297–316, June 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-016-9455-9.

Pearson:2012:HCD

[PDA12] Walter H. Pearson, Richard B. Deriso, and Jack W. Anderson. Hypotheses concerning the decline and poor recovery of Pacific herring in Prince William Sound, Alaska. Reviews in Fish Biology and Fisheries, 22(1):95–135, March 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-011-9225-7.

Pracheil:2016:FEV

[PDB16] Brenda M. Pracheil, C. R. DeRolph, and M. S. Bevelhimer. A fish-eye view of riverine hydropower systems: the current understanding of the biological response to turbine passage. Reviews in Fish Biology and Fisheries, 26(2):153–167, June 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-015-9416-8.

Phillips:2021:RGE

[PDDDE21] Nicole M. Phillips, Floriaan Devloo-Delva, and Toby S. Daly-Engel. Reviewing the genetic evidence for sex-biased dispersal in elasmobranchs. Reviews in Fish Biology and Fisheries, 31(4): 821–841, December 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-021-09673-9.

Pernice:2007:IVS

[PDGBR07] M. Pernice, D. Destoumieux-Garzón, and R. Boucher-Rodoni. Identification of a Vibrio strain producing antimicrobial agents in the excretory organs of Nautilus pompilius (Cephalopoda: Nautiloidea). Reviews in Fish Biology and Fisheries, 17(2–3): 197–205, August 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-006-9029-3.

Pinhal:2020:RCP

[PDM20] Danillo Pinhal, Rodrigo R. Domingues, and Cesar Martins. Restricted connectivity and population genetic fragility in a globally endangered hammerhead shark. Reviews in Fish Biology and Fisheries, 30(3):501–517, September 2020. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-020-09607-x.

Payet:2024:SSC

[PDN^{+}24] Samuel D. Payet, Joseph D. DiBattista, Stephen J. Newman, Kelvin J. Rushworth, Corey B. Wakefield, Richard D. Evans, and Michael J. Travers. Sympatric species of coral trout (Plectropomus) show contrasting patterns of genomic structure across isolated atoll reefs. Reviews in Fish Biology and Fisheries, 34(1):??, ???? 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-023-09809-z.

Pinnegar:2008:SBP

[PE08] John K. Pinnegar and Georg H. Engelhard. The ‘shifting baseline’ phenomenon: a global perspective. Reviews in Fish Biology and Fisheries, 18(1):1–16, February 2008. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-007-9058-6.

Penman:1993:TF

[Pen93] David Penman. Transgenic fish. Reviews in Fish Biology and Fisheries, 3(4):382–383, December 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/BF00043393.

Perdikaris:2012:CGG

[PEP12] Costas Perdikaris, Anna Ergolavou, and Ioannis Paschos. Carassius gibelio Greece: the dominant naturalised invader of freshwaters. Reviews in Fish Biology and Fisheries, 22(1):17–27, March 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-011-9216-8.

Persson:1993:EBE

[Per93] Lennart Persson. Environmental biology of European cyprinids. Reviews in Fish Biology and Fisheries, 3(2):197–198, June 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00045238.

Psenicka:2010:IPL

[PFL10] Martin Pšenička, Martin Flajšhans, and Otomar Linhart. The influence of ploidy level on ultrastructure and motility of tench Tinca tinca (L.) spermatozoa. Reviews in Fish Biology and Fisheries, 20(3):331–338, September 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-009-9135-0.

Puk:2016:PTC

[PFW16] Laura D. Puk, Sebastian C. A. Ferse, and Christian Wild. Patterns and trends in coral reef macroalgae browsing: a review of browsing herbivorous fishes of the Indo-Pacific. Reviews in Fish Biology and Fisheries, 26(1):53–70, March 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-015-9412-z. See erratum [PFW17].

Puk:2017:EPT

[PFW17] Laura D. Puk, Sebastian C. A. Ferse, and Christian Wild. Erratum to: Patterns and trends in coral reef macroalgae browsing: a review of browsing herbivorous fishes of the Indo-Pacific. Reviews in Fish Biology and Fisheries, 27(1):287–292, March 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-016-9439-9. See [PFW16].

Paiz:2024:TCE

[PGA^{+}24] Leonardo Marcel Paiz, Mariane Gavazzoni, Gabrielle Jovana Antoniazi, Lucas Baumgärtner, Weferson Júnio da Graça, Eliana Feldberg, Roberto Laridondo Lui, and Vladimir Pavan Margarido. Trends in chromosome evolution in Crenicichlina (Cichliformes, Cichlidae, Cichlinae): a new perspective based on the recent classification of the pike cichlids. Reviews in Fish Biology and Fisheries, 34(2):849–866, June 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-024-09842-6.

Pracheil:2019:SOC

[PGC19] Brenda M. Pracheil, Robert George, and Bryan C. Chakoumakos. Significance of otolith calcium carbonate crystal structure diversity to microchemistry studies. Reviews in Fish Biology and Fisheries, 29(3):569–588, September 2019. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-019-09561-3.

Pascual:2007:PCFa

[PGG07a] S. Pascual, A. Gonzáez, and A. Guerra. Parasites and cephalopod fisheries uncertainty: towards a waterfall understanding. Reviews in Fish Biology and Fisheries, 17(2–3):139–144, August 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-006-9021-y.

Pascual:2007:PCFb

[PGG07b] S. Pascual, A. González, and A. Guerra. Parasites and cephalopod fisheries uncertainty: towards a waterfall understanding. Reviews in Fish Biology and Fisheries, 17(4):635, November 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-007-9062-x.

Potts:2015:RPI

[PGJ15] W. M. Potts, A. Götz, and N. James. Review of the projected impacts of climate change on coastal fishes in southern Africa. Reviews in Fish Biology and Fisheries, 25(4):603–630, December 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-015-9399-5.

Punt:1997:FSA

[PH97] André E. Punt and Ray Hilborn. Fisheries stock assessment and decision analysis: the Bayesian approach. Reviews in Fish Biology and Fisheries, 7(1):35–63, March 1997. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1023/A:1018419207494.

Pecl:2014:OWH

[PHB14] Gretta T. Pecl, Alistair J. Hobday, and Amanda E. Bates. Ocean warming hotspots provide early warning laboratories for climate change impacts. Reviews in Fish Biology and Fisheries, 24(2):409–413, June 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-014-9355-9.

Pope:2010:BEO

[PHH10] Edward C. Pope, Graeme C. Hays, and Jonathan D. R. Houghton. The biology and ecology of the ocean sunfish Mola mola: a review of current knowledge and future research perspectives. Reviews in Fish Biology and Fisheries, 20(4):471– 487, December 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-009-9155-9.

Popper:2020:USG

[PHK20] Arthur N. Popper, Anthony D. Hawkins, and Justin Krebs. Use of sound to guide the movement of eels and other fishes within rivers: a critical review. Reviews in Fish Biology and Fisheries, 30(4):605–622, December 2020. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-020-09620-0.

Pierce:1992:HSM

[Pie92] Graham J. Pierce. Harp seals, man and ice. Reviews in Fish Biology and Fisheries, 2(3):270–271, September 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00045042.

Pierce:1993:DTI

[Pie93] Graham J. Pierce. Dolphins and the tuna industry. Reviews in Fish Biology and Fisheries, 3(4):386–387, December 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00043396.

Pister:2000:BRJ

[Pis00] Edwin P. Pister. Book review: Just Fish: Ethics and Canadian Marine Fisheries. Edited by Harold Coward, Rosemary Ommer, and Tony Pitcher. Reviews in Fish Biology and Fisheries, 10 (2):243–244, June 2000. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1016625400509.

Pitcher:1998:CSF

[Pit98a] Tony J. Pitcher. A cover story: Fisheries may drive stocks to extinction. Reviews in Fish Biology and Fisheries, 8(3):367– 370, September 1998. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1008804029850.

Pitcher:1998:EBH

[Pit98b] Tony J. Pitcher. Editorial: The Beverton and Holt Jubilee Issue, 1947–1997. Reviews in Fish Biology and Fisheries, 8(3): 225–227, September 1998. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1008876613921.

Pecl:2008:PIC

[PJ08] Gretta T. Pecl and George D. Jackson. The potential impacts of climate change on inshore squid: biology, ecology and fisheries. Reviews in Fish Biology and Fisheries, 18(4):373–385, November 2008. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-007-9077-3.

Perez-Jimenez:2014:HRR

[PJ14] Juan Carlos Pérez-Jiménez. Historical records reveal potential extirpation of four hammerhead sharks (Sphyrna spp.) in Mexican Pacific waters. Reviews in Fish Biology and Fisheries, 24 (2):671–683, June 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-014-9353-y.

Pytka:2024:TWG

[PKC^{+}24] J. M. Pytka, P. M. Kyne, J. K. Carlson, N. Wosnick, and R. W. Jabado. A tangled web: global review of fishing interactions with rhino rays. Reviews in Fish Biology and Fisheries, 34(1):??, ???? 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09821-3.

Perry:1992:CCI

[PKF92] S. F. Perry, R. Kinkead, and R. Fritsche. Are circulating catecholamines involved in the control of breathing by fishes? Reviews in Fish Biology and Fisheries, 2(1):65–83, March 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00042917.

Prisco:2007:OSR

[PLA07] Marina Prisco, Annamaria Liguoro, and Piero Andreuccetti. Oogenesis in the spotted ray Torpedo marmorata. Reviews in Fish Biology and Fisheries, 17(1):1–10, February 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-006-9013-y.

Paukert:2017:DGA

[PLW17] Craig P. Paukert, Abigail J. Lynch, and Ian J. Winfield. Designing a global assessment of climate change on inland fishes and fisheries: knowns and needs. Reviews in Fish Biology and Fisheries, 27(2):393–409, June 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-017-9477-y.

Park:1994:DMG

[PM94] Linda K. Park and Paul Moran. Developments in molecular genetic techniques in fisheries. Reviews in Fish Biology and Fisheries, 4(3):272–299, September 1994. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00042906.

Phillips:2009:SLD

[PM09] Bruce Frank Phillips and Paulette S. McWilliam. Spiny lobster development: where does successful metamorphosis to the puerulus occur?: a review. Reviews in Fish Biology and Fisheries, 19(2):193–215, June 2009. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-008-9099-5.

Penney:2014:HAO

[PM14] Zachary L. Penney and Christine M. Moffitt. Histological assessment of organs in sexually mature and post-spawning steelhead trout and insights into iteroparity. Reviews in Fish Biology and Fisheries, 24(3):781–801, September 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-013-9338-2.

Putland:2019:ADM

[PM19] R. L. Putland and A. F. Mensinger. Acoustic deterrents to manage fish populations. Reviews in Fish Biology and Fisheries, 29(4):789–807, December 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-019-09583-x.

Pla:2021:HFI

[PMP21] Susanna Pla, Francesc Maynou, and Francesc Piferrer. Hermaphroditism in fish: incidence, distribution and associations with abiotic environmental factors. Reviews in Fish Biology and Fisheries, 31(4):935–955, December 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-021-09681-9.

Plaganyi:2018:OOO

[PMV18] Éva E. Plagányi, Richard McGarvey, and Cecilia Villanueva. Overview, opportunities and outlook for Australian spiny lobster fisheries. Reviews in Fish Biology and Fisheries, 28(1):57– 87, March 2018. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-017-9493-y.

Philippsen:2019:AFI

[PMVA19] Juliana Strieder Philippsen, Carolina V. Minte-Vera, and Ronaldo Angelini. Assessing fishing impacts in a tropical reservoir through an ecosystem modeling approach. Reviews in Fish Biology and Fisheries, 29(1):125–146, March 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-018-9539-9.

Perazzo:2011:CSC

[PNC11] Giselle Perazzo, Rafael Bueno Noleto, and Marta Margarete Cestari. Chromosomal studies in Crenicichla lepidota and Australoheros facetus (Cichlidae, Perciformes) from extreme Southern Brazil. Reviews in Fish Biology and Fisheries, 21(3):509– 515, September 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-010-9170-x.

Poulin:1993:CPC

[Pou93] Robert Poulin. A cleaner perspective on cleaning symbiosis. Reviews in Fish Biology and Fisheries, 3(1):75–79, March 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00043298.

Palzenberger:1992:GSA

[PP92] Margit Palzenberger and Hannes Pohla. Gill surface area of water-breathing freshwater fish. Reviews in Fish Biology and Fisheries, 2(3):187–216, September 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/BF00045037.

Pess:2014:RCA

[PQS14] G. R. Pess, T. P. Quinn, and R. Saunders. Re-colonization of Atlantic and Pacific rivers by anadromous fishes: linkages between life history and the benefits of barrier removal. Reviews in Fish Biology and Fisheries, 24(3):881–900, September 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-013-9339-1.

Piska:2005:IJS

[PR05] Ravi Shankar Piska and A. Madhusudhan Rao. Impact of juvenile stocking size on the major carp production in a minor reservoir, Bibinagar, India. Reviews in Fish Biology and Fisheries, 15(3):167–173, August 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-005-3888-x.

Perez:1995:EMH

[PRN95] Julio Pérez, Kent Rylander, and Mauro Nirchio. The evolution of multiple haemoglobins in fishes. Reviews in Fish Biology and Fisheries, 5(3):304–319, September 1995. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/BF00043004.

Prescott:2016:UPF

[PRP16] James Prescott, James Riwu, and Andhika Prasetyo. An unlikely partnership: fishers’ participation in a small-scale fishery data collection program in the Timor Sea. Reviews in Fish Biology and Fisheries, 26(4):679–692, December 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-015-9417-7.

Podhorec:2012:EWT

[PSK12] Peter Podhorec, Magdalena Socha, and Jan Kouril. The effects of water temperature and hormone treatments on circulating LH and ovulation in tench (Tinca tinca). Reviews in Fish Biology and Fisheries, 22(3):791–796, September 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-012-9265-7.

Penas-Torramilans:2024:ICW

[PTOS^{+}24] Raquel Peñas-Torramilans, Raquel Outeiral, José Santiago, Elsa Vázquez, and Nicolas Weidberg. Influence of a changing wave climate on the quality and morphometry of the stalked barnacle Pollicipes pollicipes (Gmelin, 1789), along the coasts of NW Iberia. Reviews in Fish Biology and Fisheries, 34(2):781– 804, June 2024. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-024-09838-2.

Parisi:2014:CSF

[PTP14] Giuliana Parisi, Genciana Terova, and Antonio Pais. Current status and future perspectives of Italian finfish aquaculture. Reviews in Fish Biology and Fisheries, 24(1):15–73, March 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-013-9317-7.

Pereira:2019:LBM

[PV19] Felipe Walter Pereira and Jean Ricardo Simões Vitule. The largemouth bass Micropterus salmoides (Lacepède, 1802): impacts of a powerful freshwater fish predator outside of its native range. Reviews in Fish Biology and Fisheries, 29(3):639– 652, September 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-019-09570-2.

Peterson:2007:EBL

[PVJ07] Douglas L. Peterson, Paul Vecsei, and Cecil A. Jennings. Ecology and biology of the lake sturgeon: a synthesis of current knowledge of a threatened North American Acipenseridae. Reviews in Fish Biology and Fisheries, 17(1):59–76, February 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-006-9018-6.

Petetta:2021:PAS

[PVL21] Andrea Petetta, Massimo Virgili, and Alessandro Lucchetti. Pots as alternative and sustainable fishing gears in the Mediterranean Sea: an overview. Reviews in Fish Biology and Fisheries, 31(4):773–795, December 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-021-09676-6.

Perry:1999:FPS

[PWB99] R. Ian Perry, Carl J. Walters, and James A. Boutillier. A framework for providing scientific advice for the management of new and developing invertebrate fisheries. Reviews in Fish Biology and Fisheries, 9(2):125–150, June 1999. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https:

//link.springer.com/article/10.1023/A:1008946522213.

Portz:2006:SAI

[PWC06] Donald E. Portz, Christa M. Woodley, and Joseph J. Cech, Jr. Stress-associated impacts of short-term holding on fishes. Reviews in Fish Biology and Fisheries, 16(2):125–170, May 2006. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-006-9012-z.

Phillips:2023:DHL

[PWM23] Richard A. Phillips, Claire M. Waluda, and Allison K. Miller. Distribution, hosts and long-term decline in abundance of the Patagonian lamprey inferred from diet assessment of albatrosses. Reviews in Fish Biology and Fisheries, 33(4):1443–1464, December 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09786-3.

Peter:1997:NRO

[PY97] R. E. Peter and K. L. Yu. Neuroendocrine regulation of ovulation in fishes: basic and applied aspects. Reviews in Fish Biology and Fisheries, 7(2):173–197, June 1997. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1023/A:1018431610220.

Pyke:2005:RBG

[Pyk05] Graham H. Pyke. A review of the biology of Gambusia affinis and G. holbrooki. Reviews in Fish Biology and Fisheries, 15(4): 339–365, November 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-006-6394-x.

Pecoraro:2017:PAP

[PZC17] C. Pecoraro, I. Zudaire, and E. Chassot. Putting all the pieces together: integrating current knowledge of the biology, ecology, fisheries status, stock structure and management of yellowfin tuna (Thunnus albacares). Reviews in Fish Biology and Fisheries, 27(4):811–841, December 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-016-9460-z.

Quinones:2015:DRA

[QGM15] Rebecca M. Quiñones, Theodore E. Grantham, and Peter B. Moyle. Dam removal and anadromous salmonid (Oncorhynchus spp.) conservation in California. Reviews in Fish Biology and Fisheries, 25(1):195–215, March 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-014-9359-5.

Quinn:2004:AMM

[QM04] Thomas P. Quinn and Katherine W. Myers. Anadromy and the marine migrations of Pacific salmon and trout: Rounsefell revisited. Reviews in Fish Biology and Fisheries, 14(4): 421–442, December 2004. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-005-0802-5.

Queiroz-Sousa:2018:BEB

[QSBV18] Jamile Queiroz-Sousa, Eduardo Meneguzzi Brambilla, and Jean Ricardo Simões Vitule. Biology, ecology and biogeography of the South American silver croaker, an important Neotropical fish species in South America. Reviews in Fish Biology and Fisheries, 28(4):693–714, December 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-018-9526-1.

Quinn:2012:BRB

[Qui12] Thomas P. Quinn. Book review: B. Jonsson and N. Jonsson: Ecology of Atlantic salmon and brown trout: habitat as a template for life histories. Reviews in Fish Biology and Fisheries, 22(3):847–848, September 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-012-9255-9.

Rainboth:1994:IFI

[Rai94] Walter Rainboth. Inland fishes of India and adjacent countries. Reviews in Fish Biology and Fisheries, 4(1):135–136, March 1994. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00043269.

Ramey:2008:BRG

[Ram08] Andrew Ramey. Book review: Graeme Harris and Nigel Milner (eds): Review of Sea Trout: Biology, Conservation, & Management. Reviews in Fish Biology and Fisheries, 18(4):443– 444, November 2008. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-007-9081-7.

Rankin:1992:PFF

[Ran92] J. Cliff Rankin. Physiology and form of fish circulation. Reviews in Fish Biology and Fisheries, 2(4):345–347, December 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00043526.

Rose:2016:ERH

[RBC16a] Denis Rose, Damein Bell, and David A. Crook. Erratum to: Restoring habitat and cultural practice in Australia’s oldest and largest traditional aquaculture system. Reviews in Fish Biology and Fisheries, 26(4):737, December 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-016-9448-8. See [RBC16b].

Rose:2016:RHC

[RBC16b] Denis Rose, Damein Bell, and David A. Crook. Restoring habitat and cultural practice in Australia’s oldest and largest traditional aquaculture system. Reviews in Fish Biology and Fisheries, 26(3):589–600, September 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-016-9426-1. See er-

ratum [RBC16a].

Rothbard:2010:BPM

[RBK10] Shmuel Rothbard, Inbar Biton, and Zeev Kulikovski. Breeding, production and marketing of golden tench (Tinca tinca (L.)) in Gan Shmuel Fish Breeding Center, Israel. Reviews in Fish Biology and Fisheries, 20(3):367–373, September 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-009-9132-3.

Rutkayova:2013:AMB

[RBK13] Jitka Rutkayová, Roman Biskup, and Ján Koščo. Ameiurus melas (black bullhead): morphological characteristics of new introduced species and its comparison with Ameiurus nebulosus (brown bullhead). Reviews in Fish Biology and Fisheries, 23 (1):51–68, March 2013. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-012-9274-6.

Rimoldi:2016:WTG

[RBPPS16] Simona Rimoldi, Laura Benedito-Palos, and Jaume Pérez-Sánchez. Wide-targeted gene expression infers tissue-specific molecular signatures of lipid metabolism in fed and fasted fish. Reviews in Fish Biology and Fisheries, 26(1):93–108, March 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-015-9408-8.

Ruiz-Campos:2002:ADC

[RCCASG02] Gorgonio Ruiz-Campos, José Luis Castro-Aguirre, and Sergio Sánchez-Gonzáles. An annotated distributional checklist of the freshwater fish from Baja California Sur, México. Reviews in Fish Biology and Fisheries, 12(2–3):143–155, June 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1025067202218.

Ruiz-Campos:2003:MVW

[RCCRD03] Gorgonio Ruiz-Campos, Faustino Camarena-Rosales, and Jorge De La Rosa-Vélez. Morphometric variation of wild trout populations from northwestern Mexico (Pisces: Salmonidae). Reviews in Fish Biology and Fisheries, 13(1):91–110, March 2003. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1026367113735.

Robert:2019:SDD

[RCP19] Marianne Robert, Julia Calderwood, and Lionel Pawlowski. Spatial distribution of discards in mixed fisheries: species tradeoffs, potential spatial avoidance and national contrasts. Reviews in Fish Biology and Fisheries, 29(4):917–934, December 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-019-09581-z.

Rabuffetti:2023:FKL

[REA^{+}23] Ana P. Rabuffetti, Luis A. Espínola, Pablo Amsler, Patricio Ferreira, Elie Abrial, Martin C. M. Blettler, and Mario L. Amsler. Fishers’ knowledge on a large floodplain river in South America.

Contributions for sustainable management of inland fisheries. Reviews in Fish Biology and Fisheries, 33(3):573–592, September 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-022-09722-x.

Reebs:2002:PDC

[Ree02] Stephan G. Reebs. Plasticity of diel and circadian activity rhythms in fishes. Reviews in Fish Biology and Fisheries, 12(4): 349–371, December 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1025371804611.

Reid:1993:LFS

[Rei93] Gordon McGregor Reid. Living fossil: The story of the coelacanth. Reviews in Fish Biology and Fisheries, 3(1):84–86, March 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00043302.

Reid:1999:BRB

[Rei99] Gordon McGregor Reid. Book review: Biology and Ecology of Fishes in Southern African Estuaries (Ichthyological Monographs of the J. L. B. Smith Institute of Ichthyology, No. 2). Reviews in Fish Biology and Fisheries, 9(3):271–272, September 1999. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1008968200973.

Reynolds:1993:EFW

[Rey93] Colin Reynolds. Eutrophication of fresh waters: Principles, problems and restoration. Reviews in Fish Biology and Fisheries, 3(4):376–377, December 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00043389.

Rieucau:2015:TFE

[RFH15] Guillaume Rieucau, Anders Fernö, and Nils Olav Handegard. Towards of a firmer explanation of large shoal formation, maintenance and collective reactions in marine fish. Reviews in Fish Biology and Fisheries, 25(1):21–37, March 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-014-9367-5.

Rizzotti:1999:FHO

[RG99] Martino Rizzotti and Flavio Gioppato. Fish haemoglobins: the order Clupeiformes. Reviews in Fish Biology and Fisheries, 9 (1):71–87, March 1999. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1008881730480.

Rodger:2011:NIG

[RHM11] Hamish D. Rodger, Louise Henry, and Susan O. Mitchell. Non-infectious gill disorders of marine salmonid fish. Reviews in Fish Biology and Fisheries, 21(3):423–440, September 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-010-9182-6.

Rodd:1999:BRS

[RHP99] F. Helen Rodd, Kimberly A. Hughes, and Trevor E. Pitcher. Book review: Sex, Color, and Mate Choice in Guppies (Monographs in Behavior and Ecology). Anne E. Houde. Reviews in Fish Biology and Fisheries, 9(2):203–204, June 1999. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1008950613746.

Rigos:2010:PON

[RK10] George Rigos and Pantelis Katharios. Pathological obstacles of newly-introduced fish species in Mediterranean mariculture: a review. Reviews in Fish Biology and Fisheries, 20(1):47– 70, March 2010. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-009-9120-7.

Rice:2005:PJA

[RL05] Aaron N. Rice and Phillip S. Lobel. The pharyngeal jaw apparatus of the Cichlidae and Pomacentridae: function in feeding and sound production. Reviews in Fish Biology and Fisheries, 13 (4):433–444, January 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-004-8794-0.

Rueda:2001:RBP

[RM01] F. M. Rueda and F. M. Martínez. A review on the biology and potential aquaculture of Dentex dentex. Reviews in Fish Biology and Fisheries, 11(1):57–70, March 2001. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1023/A:1014276700138.

Rana:2024:SFI

[RMK24] Masud Rana, Shovon Mandal, and Sk. Kabita. Spirulina in fish immunity development: find the black box. Reviews in Fish Biology and Fisheries, 34(2):623–646, June 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09832-0.

Rim:2011:RBL

[RMN11] Zouari-Ktari Rim and Bradai Mohamed-Nejmeddine. Reproductive biology of the Lessepsian reticulated leatherjacket Stephanolepis diaspros (Fraser — Brünner, 1940) in the Gulf of Gabes (Eastern Mediterranean Sea). Reviews in Fish Biology and Fisheries, 21(3):641–648, September 2011. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-011-9202-1.

Ruggerone:2004:ECD

[RN04] Gregory T. Ruggerone and Jennifer L. Nielsen. Evidence for competitive dominance of pink salmon (Oncorhynchus gorbuscha) over other salmonids in the North Pacific Ocean. Reviews in Fish Biology and Fisheries, 14(3):371–390, September 2004. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-004-6927-0.

Rasal:2016:MAF

[RNJ16] Kiran Dashrath Rasal, Priyanka C. Nandanpawar, and Pallipuram Jayasankar. MicroRNA in aquaculture fishes: a way forward with high-throughput sequencing and a computational approach. Reviews in Fish Biology and Fisheries, 26(2):199– 212, June 2016. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-016-9421-6.

Roberts:1992:NMU

[Rob92] B. L. Roberts. Neural mechanisms underlying escape behaviour in fishes. Reviews in Fish Biology and Fisheries, 2(3):243– 266, September 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF00045039.

Ringo:2014:AVD

[ROL14] Einar Ringø, Rolf Erik Olsen, and Hélène L. Lauzon. Application of vaccines and dietary supplements in aquaculture: possibilities and challenges. Reviews in Fish Biology and Fisheries, 24(4):1005–1032, December 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-014-9361-y.

Rodrigues:2023:MAR

[ROP^{+}23] Jean N. Rodrigues, Jean C. G. Ortega, Danielle K. Petsch, Andre A. Padial, Dieison A. Moi, and Bruno R. S. Figueiredo. A meta-analytical review of turbidity effects on fish mobility. Reviews in Fish Biology and Fisheries, 33(4):1113–1127, December 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09785-4.

Rose:1997:PVT

[Ros97] George A. Rose. Points of view: The trouble with fisheries science! Reviews in Fish Biology and Fisheries, 7(3):365– 370, September 1997. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1018495929784.

Rosenberg:2000:MRP

[Ros00] Andrew A. Rosenberg. Marine reserves and population recovery or how do closed areas affect exploited population dynamics. Reviews in Fish Biology and Fisheries, 10(4):519–520, December 2000. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1012256015195.

Roman:2023:PSH

[ROW^{+}23] Salvador Román, Celia Olabarria, Nicolás Weidberg, Marta Román, and Elsa Vázquez. Population structure and habitat assessment for two commercial clam species exploited in smallscale fisheries. Reviews in Fish Biology and Fisheries, 33(4): 1483–1504, December 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-023-09791-6.

Roberts:1991:MRE

[RP91] Callum M. Roberts and Nicholas V. C. Polunin. Are marine reserves effective in management of reef fisheries? Reviews in Fish Biology and Fisheries, 1(1):65–91, September 1991. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00042662.

Raghavan:2013:FBI

[RPA13] Rajeev Raghavan, Siby Philip, and Anvar Ali. Freshwater biodiversity of India: a response to Sarkar et al. (2013). Reviews in Fish Biology and Fisheries, 23(4):547–554, December 2013. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-013-9315-9. See [SPL12].

Raghavan:2016:FBA

[RPD16] Rajeev Raghavan, Siby Philip, and Neelesh Dahanukar. Fishery, biology, aquaculture and conservation of the threatened Asian Sun catfish. Reviews in Fish Biology and Fisheries, 26 (2):169–180, June 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-016-9418-1.

Ridgway:2011:AIL

[RR11] I. D. Ridgway and C. A. Richardson. Arctica islandica: the longest lived non colonial animal known to science. Reviews in Fish Biology and Fisheries, 21(3):297–310, September 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-010-9171-9.

Rakauskas:2016:NIF

[RSB16] Vytautas Rakauskas, Saulius Stakėnas, and Egidijus Bukelskis. Non-indigenous fish in the northern branch of the central European invasion corridor. Reviews in Fish Biology and Fisheries, 26(3):491–508, September 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-016-9438-x.

Reis-Santos:2023:CRB

[RSGS^{+}23a] Patrick Reis-Santos, Bronwyn M. Gillanders, Anna M. Sturrock, Christopher Izzo, Dion S. Oxman, Jessica A. Lueders-Dumont, Karin Hüssy, Susanne E. Tanner, Troy Rogers, Zoë A.

Doubleday, Allen H. Andrews, Clive Trueman, Deirdre Brophy, Jason D. Thiem, Lee J. Baumgartner, Malte Willmes, Ming-Tsung Chung, Patrick Charapata, Rachel C. Johnson, Stephen Trumble, Yvette Heimbrand, Karin E. Limburg, and Benjamin D. Walther. Correction: Reading the biomineralized book of life: expanding otolith biogeochemical research and applications for fisheries and ecosystem-based management. Reviews in Fish Biology and Fisheries, 33(2):451–452, June 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-022-09729-4. See [RSGS^{+}23b].

Reis-Santos:2023:RBB

[RSGS^{+}23b] Patrick Reis-Santos, Bronwyn M. Gillanders, Anna M. Sturrock, Christopher Izzo, Dion S. Oxman, Jessica A. Lueders-Dumont, Karin Hüssy, Susanne E. Tanner, Troy Rogers, Zoë A. Doubleday, Allen H. Andrews, Clive Trueman, Deirdre Brophy, Jason D. Thiem, Lee J. Baumgartner, Malte Willmes, Ming-Tsung Chung, Patrick Charapata, Rachel C. Johnson, Stephen Trumble, Yvette Heimbrand, Karin E. Limburg, and Benjamin D. Walther. Reading the biomineralized book of life: expanding otolith biogeochemical research and applications for fisheries and ecosystem-based management. Reviews in Fish Biology and Fisheries, 33(2):411–449, June 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-022-09720-z. See

correction [RSGS^{+}23a].

Rasekhi:2023:IFK

[RSSS23] Sare Rasekhi, Abolfazl Sharifian, Mohammadreza Shahraki, and Renato A. M. Silvano. Indigenous fishers’ knowledge on fish behavior, fishing practices and climatic conditions in a conservation priority coastal ecosystem in the Caspian Sea. Reviews in Fish Biology and Fisheries, 33(3):629–648, September 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-022-09746-3.

Randall:1991:ERC

[RT91] David J. Randall and E. W. Taylor. Evidence of a role for catecholamines in the control of breathing in fish. Reviews in Fish Biology and Fisheries, 1(2):139–157, December 1991. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00157582.

Rigos:2005:AAM

[RT05] G. Rigos and G. M. Troisi. Antibacterial agents in Mediterranean finfish farming: a synopsis of drug pharmacokinetics in important euryhaline fish species and possible environmental implications. Reviews in Fish Biology and Fisheries, 15(1– 2):53–73, February 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-005-7850-8.

Russell:2012:RBE

[RTT12] D. J. Russell, P. A. Thuesen, and F. E. Thomson. A review of the biology, ecology, distribution and control of Mozambique tilapia, Oreochromis mossambicus (Peters 1852) (Pisces: Cichlidae) with particular emphasis on invasive Australian populations. Reviews in Fish Biology and Fisheries, 22(3):533– 554, September 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-011-9249-z.

Ryman:1995:PIB

[RUL95] Nils Ryman, Fred Utter, and Linda Laikre. Protection of intraspecific biodiversity of exploited fishes. Reviews in Fish Biology and Fisheries, 5(4):417–446, December 1995. CODEN RF-BFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF01103814.

Reyes-Valdez:2011:PMV

[RVRCB11] Claudia Alejandra Reyes-Valdez, Gorgonio Ruiz-Campos, and Giacomo Bernardi. Population morphometric variation of the endemic freshwater killifish, Fundulus lima (Teleostei: Fundulidae), and its coastal relative F. parvipinnis from the Baja California Peninsula, Mexico. Reviews in Fish Biology and Fisheries, 21(3):543–558, September 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-010-9175-5.

Roessig:2004:EGC

[RWH04] Julie M. Roessig, Christa M. Woodley, and Lara J. Hansen. Effects of global climate change on marine and estuarine fishes and fisheries. Reviews in Fish Biology and Fisheries, 14(2):251– 275, June 2004. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-004-6749-0.

Rosa:2012:NSC

[RZV12] Kamila Oliveira Rosa, Kaline Ziemniczak, and Marcelo Ricardo Vicari. Numeric and structural chromosome polymorphism in Rineloricaria lima (Siluriformes: Loricariidae): fusion points carrying 5S rDNA or telomere sequence vestiges. Reviews in Fish Biology and Fisheries, 22(3):739–749, September 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-011-9250-6.

Sarda:2012:RBC

[SA12] Francisco Sardà and Jacopo Aguzzi. A review of burrow counting as an alternative to other typical methods of assessment of Norway lobster populations. Reviews in Fish Biology and Fisheries, 22(2):409–422, June 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-011-9242-6.

Sbragaglia:2023:GRM

[SAB^{+}23] Valerio Sbragaglia, Robert Arlinghaus, Daniel T. Blumstein, Hugo Diogo, Vinicius J. Giglio, Ana Gordoa, Fraser Andrew Januchowski-Hartley, Martín Laporta, Steven J. Lindfield, Josep Lloret, Bruce Mann, Daryl McPhee, José A. C. C. Nunes, Pablo Pita, Mafalda Rangel, O. Kennedy Rhoades, Leonardo A. Venerus, and Sebastián Villasante. A global review of marine recreational spearfishing. Reviews in Fish Biology and Fisheries, 33(4):1199–1222, December 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-023-09790-7.

Stratoudakis:2023:RCM

[SAC^{+}23] Yorgos Stratoudakis, Carlos Antunes, Cláudia Correia, Ana Filipa Belo, and Pedro R. Almeida. Riverine communities and management systems for anadromous fisheries in the Iberian Peninsula: global strategy, local realities. Reviews in Fish Biology and Fisheries, 33(3):875–892, September 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-022-09742-7.

Stark:2014:CRC

[SAK14] Eric J. Stark, Ernest J. Atkinson, and Christine C. Kozfkay. Captive rearing for Chinook salmon (Oncorhynchus tshawytscha) and Atlantic salmon (Salmo salar): the Idaho and Maine experiences. Reviews in Fish Biology and Fisheries, 24(3):849– 880, September 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-014-9346-x.

Sanger:1993:LAF

[Sän93] Alexandra M. Sänger. Limits to the acclimation of fish muscle. Reviews in Fish Biology and Fisheries, 3(1):1–15, March 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00043295.

Segaran:2023:ASS

[SAN^{+}23] Thirukanthan Chandra Segaran, Hani Amir Aouissi, Mohd Iqbal Mohd Noor, Mohd Effendy Abd Wahid, Fathurrahman Lananan, Alexandru-Ionut Petrisor, and Mohamad Nor Azra. Assessing the state of seahorse research through scientometric analysis: an update. Reviews in Fish Biology and Fisheries, 33(4):1237–1262, December 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-023-09794-3.

Shinoda:2011:ELD

[SAT11] Akira Shinoda, Jun Aoyama, and Katsumi Tsukamoto. Evaluation of the larval distribution and migration of the Japanese eel in the western North Pacific. Reviews in Fish Biology and Fisheries, 21(3):591–611, September 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-010-9195-1.

Shipley:2017:SIA

[SBG17] Oliver. N. Shipley, Edward J. Brooks, and R. Dean Grubbs. Stable isotope analysis in deep-sea chondrichthyans: recent challenges, ecological insights, and future directions. Reviews in Fish Biology and Fisheries, 27(3):481–497, September 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-017-9466-1.

Sweeting:2004:COT

[SBN04] R. M. Sweeting, R. J. Beamish, and C. M. Neville. Crystalline otoliths in teleosts: Comparisons between hatchery and wild Coho salmon (Oncorhynchus kisutch) in the Strait of Georgia. Reviews in Fish Biology and Fisheries, 14(3):361–369, September 2004. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-005-3793-3.

Sinclair:2007:GDI

[SBP07] Billy Sinclair, Leica Briskey, and Graham Pegg. Genetic diversity of isolated populations of Nautilus pompilius (Mollusca, Cephalopoda) in the Great Barrier Reef and Coral Sea. Reviews in Fish Biology and Fisheries, 17(2–3):223–235, August 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-006-9030-x.

Sabolic:2021:IEB

[SBSŠ21] Iva Sabolić, Miguel Baltazar-Soares, and Anamaria Štambuk. Incorporating evolutionary based tools in cephalopod fisheries management. Reviews in Fish Biology and Fisheries, 31(3): 485–503, September 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-021-09652-0.

Shin:2012:GSR

[SBtIWG12] Yunne-Jai Shin, Alida Bundy, and the IndiSeas Working Group. Global in scope and regionally rich: an IndiSeas workshop helps shape the future of marine ecosystem indicators. Reviews in Fish Biology and Fisheries, 22(3):835–845, September 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-012-9252-z.

Smith:2017:CSA

[SBZR17] David R. Smith, H. Jane Brockmann, and Jaime Zaldívar-Rae. Conservation status of the American horseshoe crab, (Limulus polyphemus): a regional assessment. Reviews in Fish Biology and Fisheries, 27(1):135–175, March 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-016-9461-y.

Santacroce:2008:AAS

[SCC08] Maria Pia Santacroce, M. C. Conversano, and G. Crescenzo. Aflatoxins in aquatic species: metabolism, toxicity and perspectives. Reviews in Fish Biology and Fisheries, 18(1):99–130, February 2008. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-007-9064-8.

Sarda:2009:BPE

[SCC09] F. Sardà, J. B. Company, and M. Coll. Biological patterns and ecological indicators for Mediterranean fish and crustaceans below 1,000 m: a review. Reviews in Fish Biology and Fisheries, 19(3):329–347, September 2009. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-009-9105-6.

Schwamborn:2018:HRP

[Sch18] Ralf Schwamborn. How reliable are the Powell–Wetherall plot method and the maximum-length approach? Implications for length-based studies of growth and mortality. Reviews in Fish Biology and Fisheries, 28(3):587–605, September 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-018-9519-0.

Sayer:1991:AFW

[SD91] M. D. J. Sayer and J. Davenport. Amphibious fish: why do they leave water? Reviews in Fish Biology and Fisheries, 1(2): 159–181, December 1991. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF00157583.

Sotton:2015:NCR

[SDG15] Benoît Sotton, Isabelle Domaizon, and Jean Guillard. Nodularin and cylindrospermopsin: a review of their effects on fish. Reviews in Fish Biology and Fisheries, 25(1):1–19, March 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-014-9366-6.

Sarkar:2013:RAF

[SDJ13] U. K. Sarkar, V. K. Dubey, and J. K. Jena. Retracted article: Freshwater fish biodiversity of India: pattern, utilization, importance, threats and challenges. Reviews in Fish Biology and Fisheries, 23(4):555, December 2013. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https:/

/link.springer.com/article/10.1007/s11160-013-9306-x.

See [SPL12].

Spares:2015:CRM

[SDS15] Aaron D. Spares, Michael J. Dadswell, and Michael J. W. Stokesbury. A critical review of marine adaptability within the anadromous salmoninae. Reviews in Fish Biology and Fisheries, 25(3): 503–519, September 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-015-9392-z.

Schnierer:2016:CAH

[SE16] S. Schnierer and H. Egan. Composition of the Aboriginal harvest of fisheries resources in coastal New South Wales, Australia. Reviews in Fish Biology and Fisheries, 26(4):693–709, December 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-016-9452-z.

Sbragaglia:2022:RAS

[SEA22] Valerio Sbragaglia, Lucía Espasandín, and Robert Arlinghaus. Recreational angling and spearfishing on social media: insights on harvesting patterns, social engagement and sentiments related to the distributional range shift of a marine invasive species. Reviews in Fish Biology and Fisheries, 32(2):687– 700, June 2022. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-022-09699-7.

Ssempijja:2024:ALO

[SEH24] Drake Ssempijja, Haraldur Arnar Einarsson, and Pingguo He. Abandoned, lost, and otherwise discarded fishing gear in world’s inland fisheries. Reviews in Fish Biology and Fisheries, 34 (2):671–683, June 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-024-09843-5.

Seitz:2008:BRR

[Sei08] Andrew C. Seitz. Book review: Robin N. Gibson (Ed.), Flatfishes: Biology and Exploitation. Reviews in Fish Biology and Fisheries, 18(2):249–250, May 2008. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-007-9072-8.

Sloat:2014:EEP

[SFO14a] Matthew R. Sloat, Dylan J. Fraser, and Haley A. Ohms. Ecological and evolutionary patterns of freshwater maturation in Pacific and Atlantic salmonines. Reviews in Fish Biology and Fisheries, 24(3):689–707, September 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-014-9344-z. See er-

ratum [SFO14b].

Sloat:2014:EEE

[SFO14b] Matthew R. Sloat, Dylan J. Fraser, and Haley A. Ohms. Erratum to: Ecological and evolutionary patterns of freshwater maturation in Pacific and Atlantic salmonines. Reviews in Fish Biology and Fisheries, 24(3):709–711, September 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-014-9354-x. See [SFO14a].

Smolinski:2024:OBL

[SG24] Szymon Smoliński and Julita Gutkowska. Otolith biochronology for the long-term reconstruction of growth and stock dynamics of fish. Reviews in Fish Biology and Fisheries, 34(1):??, ???? 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09820-4.

Sun:2023:RSL

[SGD^{+}23] Jingrui Sun, Shams M. Galib, Liuyong Ding, Juan Tao, Chengzhi Ding, and Daming He. Research status of the Lancang–Mekong River Basin: fish and environmental stressors. Reviews in Fish Biology and Fisheries, 33(1):89–109, March 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-022-09740-9.

Schneider:2013:CEN

[SGF13] Carlos Henrique Schneider, Maria Claudia Gross, and Eliana Feldberg. Chromosomal evolution of neotropical cichlids: the role of repetitive DNA sequences in the organization and structure of karyotype. Reviews in Fish Biology and Fisheries, 23 (2):201–214, June 2013. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-012-9285-3.

Streit:2021:HFH

[SHB21] Robert P. Streit, Christopher R. Hemingson, and David R. Bellwood. How flexible are habitat specialists? Short-term space use in obligate coral-dwelling damselfishes. Reviews in Fish Biology and Fisheries, 31(2):381–398, June 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-021-09646-y.

Shephard:1994:FFM

[She94] Kerry L. Shephard. Functions for fish mucus. Reviews in Fish Biology and Fisheries, 4(4):401–429, December 1994. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00042888.

Shelton:1995:PRJ

[SHH95] R. G. J. Shelton, A. D. Hawkins, and Sidney Holt. Professor Raymond (John Heaphy) Beverton, CBE, FRS, FIBiol. Reviews in Fish Biology and Fisheries, 5(4):393–397, December 1995. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF01103812.

Sanchez-Hernandez:2021:PNB

[SHHK21] Javier Sánchez-Hernández, Brian Hayden, and Kimmo K. Kahilainen. Population niche breadth and individual trophic specialisation of fish along a climate-productivity gradient. Reviews in Fish Biology and Fisheries, 31(4):1025–1043, December 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-021-09687-3.

Shi:2005:MIS

[Shi05] Qiong Shi. Melatonin is involved in sex change of the ricefield eel, Monopterus albus Zuiew. Reviews in Fish Biology and Fisheries, 15(1–2):??, February 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-005-7848-2.

Saunders:1996:BR

[SHM96] Richard L. Saunders, Paul J. B. Hart, and Gary K. Meffe. Book reviews. Reviews in Fish Biology and Fisheries, 6(4): 463–477, December 1996. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF00164326.

Sanchez-Hernandez:2022:ETP

[SHPH^{+}22] Javier Sánchez-Hernández, Sebastian Prati, Eirik Haugstvedt Henriksen, Aslak Smalås, Rune Knudsen, Anders Klemetsen, and Per-Arne Amundsen. Exploring temporal patterns in fish feeding ecology: Are ontogenetic dietary shifts stable over time? Reviews in Fish Biology and Fisheries, 32(4):1141–1155, December 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-022-09724-9.

Schlaff:2014:IEF

[SHS14] Audrey M. Schlaff, Michelle R. Heupel, and Colin A. Simpfendorfer. Influence of environmental factors on shark and ray movement, behaviour and habitat use: a review. Reviews in Fish Biology and Fisheries, 24(4):1089–1103, December 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-014-9364-8.

Sinha:1993:EC

[Sin93] V. R. P. Sinha. Eel culture. Reviews in Fish Biology and Fisheries, 3(4):375–376, December 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00043388.

Sin:1997:TF

[Sin97] Frank Y. T. Sin. Transgenic fish. Reviews in Fish Biology and Fisheries, 7(4):417–441, December 1997. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1023/A:1018452214763.

Surathkal:2023:IUS

[SJB^{+}23] Prasanna Surathkal, Amalendu Jyotishi, Ramchandra Bhatta, Joeri Scholtens, Derek Johnson, Gargi Mondal, and Priya Gupta. Implications of utilization shifts of marine fish in India: a macro-level empirical analysis. Reviews in Fish Biology and Fisheries, 33(3):767–783, September 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-022-09752-5.

Semeniuk:2022:ITP

[SJH22] C. A. D. Semeniuk, K. M. Jeffries, and D. D. Heath. Innovating transcriptomics for practitioners in freshwater fish management and conservation: best practices across diverse resourcesector users. Reviews in Fish Biology and Fisheries, 32(3): 921–939, September 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-022-09715-w.

Stergiou:2002:FHT

[SK02] Konstantinos I. Stergiou and Vasiliki S. Karpouzi. Feeding habits and trophic levels of Mediterranean fish. Reviews in Fish Biology and Fisheries, 11(3):217–254, September 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1020556722822.

Samy-Kamal:2015:SFE

[SK15] Mohamed Samy-Kamal. Status of fisheries in Egypt: reflections on past trends and management challenges. Reviews in Fish Biology and Fisheries, 25(4):631–649, December 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-015-9404-z.

Silvano:2023:AKA

[SK23] Renato Azevedo Matias Silvano and John Kurien. Advancing knowledge about great small-scale fisheries. Reviews in Fish Biology and Fisheries, 33(3):535–539, September 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09797-0.

Sadovy:2003:HWC

[SKD03] Y. Sadovy, M. Kulbicki, and T. J. Donaldson. The humphead wrasse, Cheilinus undulatus: Synopsis of a threatened and poorly known giant coral reef fish. Reviews in Fish Biology and Fisheries, 13(3):327–364, September 2003. CODEN RFBFEA.

ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1023/B:RFBF.0000033122.
90679.97.

Sanchez:2020:PMM

[SKU20] Gustavo Sanchez, Kentaro Kawai, and Tetsuya Umino. Patterns of mitochondrial and microsatellite DNA markers describe historical and contemporary dynamics of the Humboldt squid Dosidicus gigas in the Eastern Pacific Ocean. Reviews in Fish Biology and Fisheries, 30(3):519–533, September 2020. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-020-09609-9.

Seah:2022:DBT

[SKW22] Ying Giat Seah, Caroline Kibat, and Benjamin J. Wainwright. DNA barcoding of traded shark fins in Peninsular Malaysia. Reviews in Fish Biology and Fisheries, 32(3):993–999, September 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-022-09713-y.

Strugnell:2007:BLD

[SL07] Jan M. Strugnell and Annie R. Lindgren. A barcode of life database for the Cephalopoda? Considerations and concerns. Reviews in Fish Biology and Fisheries, 17(2–3):337–344, August 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-007-9043-0.

Schofield:2015:NNF

[SL15] Pamela J. Schofield and William F. Loftus. Non-native fishes in freshwaters: a literature review and synthesis. Reviews in Fish Biology and Fisheries, 25(1):117–145, March 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-014-9373-7.

Smale:2011:FFU

[SLH11] Dan A. Smale, Timothy J. Langlois, and Euan S. Harvey. From fronds to fish: the use of indicators for ecological monitoring in marine benthic ecosystems, with case studies from temperate Western Australia. Reviews in Fish Biology and Fisheries, 21(3):

311–337, September 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-010-9173-7.

Stromme:2016:LCH

[SLK16] Tore Strømme, Marek R. Lipinski, and Paulus Kainge. Life cycle of hake and likely management implications. Reviews in Fish Biology and Fisheries, 26(2):235–248, June 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-015-9415-9.

Steer:2007:EEP

[SM07] M. A. Steer and N. A. Moltschaniwskyj. The effects of egg position, egg mass size, substrate and biofouling on embryo mortality in the squid Sepioteuthis australis. Reviews in Fish Biology and Fisheries, 17(2–3):??, August 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-006-9023-9.

Sevenster:1995:BR

[SMC95] P. Sevenster, C. B. Munn, and L. J. V. Compagno. Book reviews. Reviews in Fish Biology and Fisheries, 5(1):120–138, March 1995. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF01103369.

Smith:1991:ALH

[Smi91] John W. Smith. Alternative life-history styles of fishes. Reviews in Fish Biology and Fisheries, 1(1):93–108, September 1991. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00042664.

Smith:1992:ASF

[Smi92] R. Jan F. Smith. Alarm signals in fishes. Reviews in Fish Biology and Fisheries, 2(1):33–63, March 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/BF00042916.

Smith:1993:CEE

[Smi93] Carl Smith. Cannibalism: Ecology and evolution among diverse taxa. Reviews in Fish Biology and Fisheries, 3(1):86–87, March 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00043303.

Smith:1998:SCA

[Smi98] Tim D. Smith. “Simultaneous and complementary advances”: Mid-century expectations of the interaction of fisheries science and management. Reviews in Fish Biology and Fisheries, 8(3): 335–348, September 1998. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1008823503867.

Slootweg:1994:BCS

[SMM94] R. Slootweg, E. A. Malek, and F. S. McCullough. The biological control of snail intermediate hosts of schistosomiasis by fish. Reviews in Fish Biology and Fisheries, 4(1):67–90, March 1994. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00043261.

Subbey:2008:TAI

[SMN08] Sam Subbey, Kathrine Michalsen, and Geir Kjetil Nilsen. A tool for analyzing information from data storage tags: the continuous wavelet transform (CWT). Reviews in Fish Biology and Fisheries, 18(3):301–312, August 2008. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-007-9078-2.

Sanchez-Marquez:2023:UPE

[SMNK^{+}23] Antoni Sánchez-Márquez, Joan Navarro, Antigoni Kaliontzopoulou, Marc Farré, Morag Taite, Oscar Escolar, Roger Villanueva, A. Louise Allcock, and Fernando Á. Fernández-Álvarez. Unravelling the phylogenetic and ecological drivers of beak shape variability in cephalopods. Reviews in Fish Biology and Fisheries, 33(1):221–239, March 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-022-09744-5.

Seitz:2011:EFH

[SMR11] Andrew C. Seitz, Katie Moerlein, and Amanda E. Rosenberger. Ecology of fishes in a high-latitude, turbid river with implications for the impacts of hydrokinetic devices. Reviews in Fish Biology and Fisheries, 21(3):481–496, September 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-011-9200-3.

Sanmukh:2012:IFP

[SMS12] Swapnil G. Sanmukh, Dilip B. Meshram, and Sandhya Swaminathan. Interaction of fishes with pathogenic micro-organisms and application of phages for their control: a review. Reviews in Fish Biology and Fisheries, 22(3):567–574, September 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-012-9257-7.

Sibbing:2000:RPL

[SN00] Ferdinand A. Sibbing and Leo A. J. Nagelkerke. Resource partitioning by Lake Tana barbs predicted from fish morphometrics and prey characteristics. Reviews in Fish Biology and Fisheries, 10(4):393–437, December 2000. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1023/A:1012270422092.

Sandra:2010:SDD

[SN10] Guerrero-Estévez Sandra and Moreno-Mendoza Norma. Sexual determination and differentiation in teleost fish. Reviews in Fish Biology and Fisheries, 20(1):101–121, March 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-009-9123-4.

Snyder:2005:IOC

[Sny05] Darrel E. Snyder. Invited overview: conclusions from a review of electrofishing and its harmful effects on fish. Reviews in Fish Biology and Fisheries, 13(4):445–453, January 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-004-1095-9.

Snodgrass:2020:PIO

[SOB20] Derke J. G. Snodgrass, Eric S. Orbesen, and Craig A. Brown. Potential impacts of oil production platforms and their function as fish aggregating devices on the biology of highly migratory fish species. Reviews in Fish Biology and Fisheries, 30(3): 405–422, September 2020. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-020-09605-z.

Soivio:1999:BRE

[Soi99] Antti Soivio. Book review: Early Life Stage Mortality Syndrome in Fishes of the Great Lakes and Baltic Sea (American Fisheries Society Symposium 21). Edited by Gordon McDonald, John D. Fitzsimons and Dale C. Honeyfield. Reviews in Fish Biology and Fisheries, 9(2):205, June 1999. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1023/A:1008998529676.

Solomon:1993:FBS

[Sol93] D. J. Solomon. Fisheries bioengineering symposium. Reviews in Fish Biology and Fisheries, 3(2):196–197, June 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00045237.

Soto:2002:PIG

[Sot02] Cristina G. Soto. The potential impacts of global climate change on marine protected areas. Reviews in Fish Biology and Fisheries, 11(3):181–195, September 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1023/A:1020364409616.

Stefansson:1998:PVF

[SP98] Gunnar Stefansson and Olafur K. Palsson. Points of view: a framework for multispecies modelling of arcto-boreal systems. Reviews in Fish Biology and Fisheries, 8(1):101–104, March 1998. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1008820702467.

Scribner:2000:HFF

[SPB00] Kim T. Scribner, Kevin S. Page, and Meredith L. Bartron. Hybridization in freshwater fishes: a review of case studies and cytonuclear methods of biological inference. Reviews in Fish Biology and Fisheries, 10(3):293–323, September 2000. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1016642723238.

Sarkar:2012:FFB

[SPL12] U. K. Sarkar, A. K. Pathak, and W. S. Lakra. Freshwater fish biodiversity in the River Ganga (India): changing pattern, threats and conservation perspectives. Reviews in Fish Biology and Fisheries, 22(1):251–272, March 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.

1007/s11160-011-9218-6. See response [RPA13] and retraction [SDJ13].

Semmens:2007:ARC

[SPS07] Jayson M. Semmens, Gretta T. Pecl, and Paul W. Shaw. Approaches to resolving cephalopod movement and migration patterns. Reviews in Fish Biology and Fisheries, 17(2–3):??, August 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-007-9048-8.

Sierp:2009:BRB

[SQR09] Michael T. Sierp, Jian G. Qin, and Friedrich Recknagel. Biomanipulation: a review of biological control measures in eutrophic waters and the potential for Murray cod Maccullochella peelii peelii to promote water quality in temperate Australia. Reviews in Fish Biology and Fisheries, 19(2):143–165, June 2009. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-008-9094-x.

Smith:1991:CTF

[SR91] Carl Smith and Peter Reay. Cannibalism in teleost fish. Reviews in Fish Biology and Fisheries, 1(1):41–64, September 1991. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00042661.

Subramaniam:2023:SER

[SRB^{+}23] Roshni C. Subramaniam, Mélodie Ruwet, Fabio Boschetti, Simon Fielke, Aysha Fleming, Rosa Mar Dominguez-Martinez, Éva Plagányi, Peggy Schrobback, and Jessica Melbourne-Thomas. The socio-ecological resilience and sustainability implications of seafood supply chain disruption. Reviews in Fish Biology and Fisheries, 33(4):1129–1154, December 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09788-1.

Schmidt-Roach:2021:EHI

[SRBS21] Alicia C. J. Schmidt-Roach, Christine C. Bruels, and Mahmood S. Shivji. Evidence of historical isolation and genetic structuring among broadnose sevengill sharks (Notorynchus cepedianus) from the world’s major oceanic regions. Reviews in Fish Biology and Fisheries, 31(2):433–447, June 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-021-09651-1.

Sayer:1993:FAE

[SRD93] M. D. J. Sayer, J. P. Reader, and T. R. K. Dalziel. Freshwater acidification: effects on the early life stages of fish. Reviews in Fish Biology and Fisheries, 3(2):95–132, June 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00045228.

Saborido-Rey:2004:RPS

[SRGS04] Fran Saborido-Rey, Dolores Garabana, and Vladimir Shibanov. Review of the population structure and ecology of S. mentella in the Irminger Sea and adjacent waters. Reviews in Fish Biology and Fisheries, 14(4):455–479, December 2004. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-005-3585-9.

Silva-Segundo:2011:CTS

[SSBCL11] Claudia A. Silva-Segundo, Mariela Brito-Chavarria, and Francisco J. García-De León. Clarifying the taxonomic status of Merluccius spp. in the northeastern Pacific: a combined morphological and molecular approach. Reviews in Fish Biology and Fisheries, 21(2):259–282, June 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-010-9166-6.

Skurdal:1994:DWN

[ST94] J. Skurdal and T. Taugbøl. Do we need harvest regulations for European crayfish? Reviews in Fish Biology and Fisheries, 4(4): 461–485, December 1994. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF00042890.

Stergiou:2005:BRD

[Ste05] Konstantinos I. Stergiou. Book review: Daniel Pauly and Jay MaClean, In a Perfect Ocean — The State of Fisheries and Ecosystems in the North Atlantic Ocean. Island Press (Washington/Covelo/London), 2003. $25 pbk (175 pages, 28 figures in color) ISBN 1-55963-324-7. Reviews in Fish Biology and Fisheries, 13(4):455–457, January 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-004-4457-4.

Stobutzki:2000:MRC

[Sto00] Ilona C. Stobutzki. Marine reserves and the complexity of larval dispersal. Reviews in Fish Biology and Fisheries, 10(4): 515–518, December 2000. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1012241618520.

Scott:2022:VAD

[STP22] Molly E. Scott, Sterling B. Tebbett, and Morgan S. Pratchett. Variation in abundance, diversity and composition of coral reef fishes with increasing depth at a submerged shoal in the northern Great Barrier Reef. Reviews in Fish Biology and Fisheries, 32(3):941–962, September 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-022-09716-9.

Sukramongkol:2007:AML

[STS07] Natinee Sukramongkol, Kotaro Tsuchiya, and Susumu Segawa. Age and maturation of Loligo duvauceli and L. chinensis from Andaman Sea of Thailand. Reviews in Fish Biology and Fisheries, 17(2–3):237–246, August 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-006-9033-7.

Syahailatua:2024:CIR

[STW^{+}24] Augy Syahailatua, Muhammad Taufik, Karsono Wagiyo, Hagi Y. Sugeha, Charles P. H. Simanjuntak, Sam Wouthuyzen, Michael J. Miller, and Jun Aoyama. A century of ichthyoplankton research in Indonesian waters: lessons from the past, challenges for the future. Reviews in Fish Biology and Fisheries, 34(1):??, ???? 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-023-09802-6.

Sullivan:2004:BRG

[Sul04] Joseph R. Sullivan. Book review: Gary Wedemeyer (ed.), Fish Hatchery Management, Second Edition. 2001. Reviews in Fish Biology and Fisheries, 14(2):297–300, June 2004. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-004-6925-2.

Suresh:2023:CCT

[Sur23] A. Suresh. Contextualising credit transactions in artisanal marine fishing: insights from Kerala, India. Reviews in Fish Biology and Fisheries, 33(3):699–715, September 2023. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-023-09782-7.

Schreier:2021:SAA

[SVC21] Andrea D. Schreier, Joel P. Van Eenennaam, and Jamie Crossman. Spontaneous autopolyploidy in the Acipenseriformes, with recommendations for management. Reviews in Fish Biology and Fisheries, 31(2):159–180, June 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-021-09637-z.

Stratoudakis:2024:CLP

[SVM^{+}24a] Yorgos Stratoudakis, Manuel Vieira, João Pedro Marques, Maria Clara P. Amorim, Paulo J. Fonseca, and Bernardo R. Quintella. Correction: Long-term passive acoustic monitoring to support adaptive management in a Sciaenid fishery (Tagus Estuary, Portugal). Reviews in Fish Biology and Fisheries, 34(2):867, June 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-024-09833-7. See [SVM^{+}24b].

Stratoudakis:2024:LTP

[SVM^{+}24b] Yorgos Stratoudakis, Manuel Vieira, João Pedro Marques, Maria Clara P. Amorim, Paulo J. Fonseca, and Bernardo R. Quintella. Long-term passive acoustic monitoring to support adaptive management in a Sciaenid fishery (Tagus Estuary, Portugal). Reviews in Fish Biology and Fisheries, 34(1):??, ???? 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.

1007/s11160-023-09825-z. See correction [SVM^{+}24a].

Silva:2019:AOU

[SVRS19] Catarina N. S. Silva, Cecilia Villacorta-Rath, and Jan M. Strugnell. Advancing our understanding of the connectivity, evolution and management of marine lobsters through genetics. Reviews in Fish Biology and Fisheries, 29(3):669–687, September 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-019-09573-z.

Smith:1995:CPC

[SW95a] Carl Smith and Robert J. Wootton. The costs of parental care in teleost fishes. Reviews in Fish Biology and Fisheries, 5(1):7–22, March 1995. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF01103363.

Smith:1995:EBC

[SW95b] Carl Smith and Robert J. Wootton. The effect of brood cannibalism on the operational sex ratio in parental teleost fishes. Reviews in Fish Biology and Fisheries, 5(3):372–376, September 1995. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00043007.

Smith:1998:PVC

[SW98] Carl Smith and Robert J. Wootton. Points of view: Cannibalism and OSR — a response to kvarnemo. Reviews in Fish Biology and Fisheries, 8(1):99–100, March 1998. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1023/A:1008868618397.

Sikorska:2010:CCT

[SW10] Justyna Sikorska and Jacek Wolnicki. Cadmium and copper toxicity to tench Tinca tinca (L.) larvae after a short-term exposure. Reviews in Fish Biology and Fisheries, 20(3):417–423, September 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-009-9145-y.

Satora:2012:RBS

[SW12] Leszek Satora and Nicholas C. Wegner. Reexamination of the Byczkowska–Smyk gill surface area data for European teleosts, with new measurements on the pikeperch, Sander lucioperca. Reviews in Fish Biology and Fisheries, 22(1):1–9, March 2012. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-011-9213-y.

Sweka:2014:UPV

[SW14] John A. Sweka and Thomas C. Wainwright. Use of population viability analysis models for Atlantic and Pacific salmon recovery planning. Reviews in Fish Biology and Fisheries, 24(3): 901–917, September 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-013-9332-8.

Saunders:2016:OMT

[SX16] Thor Saunders and Simon Xuereb. Optimising the monitoring of tropical aquatic resources through the development of indigenous scientific capability. Reviews in Fish Biology and Fisheries, 26(4):727–736, December 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-016-9451-0.

Shen:2023:OMU

[SYL^{+}23] Yawei Shen, Weiwei You, Xuan Luo, Ying Lu, Miaoqin Huang, and Caihuan Ke. An overview of the mechanisms underlying hypoxia tolerance differences in aquatic animals and their inspirations for aquaculture. Reviews in Fish Biology and Fisheries, 33(4):1223–1236, December 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-023-09793-4.

Santacroce:2011:EHD

[SZC11] Maria Pia Santacroce, Valentina Zacchino, and Gerardo Centoducati. Expression of a highly differentiated phenotype and hepatic functionality markers in gilthead seabream (Sparus aurata L.) long-cultured hepatocytes: first morphological and functional in vitro characterization. Reviews in Fish Biology and Fisheries, 21(3):571–590, September 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-010-9181-7.

Siwicki:2010:DMR

[SZG10] Andrzej K. Siwicki, Zdzisław Zakȩś, and Edward Gła̧bski. Dietary macrogard reduces Aeromonas hydrophila mortality in tench (Tinca tinca) through the activation of cellular and humoral defence mechanisms. Reviews in Fish Biology and Fisheries, 20(3):435–439, September 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-009-9133-2.

Trumbo:2014:IHP

[TAD14] Bradly A. Trumbo, Martin L. Ahmann, and Z. D. Deng. Improving hydroturbine pressures to enhance salmon passage survival and recovery. Reviews in Fish Biology and Fisheries, 24(3): 955–965, September 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-013-9340-8.

Taniguchi:2003:GFB

[Tan03] N. Taniguchi. Genetic factors in broodstock management for seed production. Reviews in Fish Biology and Fisheries, 13(2): Article 177, June 2003. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/B:RFBF.0000019479.49749.fe.

Tsikliras:2010:SPM

[TAS10] Athanassios C. Tsikliras, Efthimia Antonopoulou, and Konstantinos I. Stergiou. Spawning period of Mediterranean marine fishes. Reviews in Fish Biology and Fisheries, 20(4):499– 538, December 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-010-9158-6.

Taylor:1999:SPN

[Tay99] Eric B. Taylor. Species pairs of north temperate freshwater fishes: Evolution, taxonomy, and conservation. Reviews in Fish Biology and Fisheries, 9(4):299–324, December 1999. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1008955229420.

Tixier:2020:AIT

[TBA20] Paul Tixier, Paul Burch, and John P. Y. Arnould. Assessing the impact of toothed whale depredation on socio-ecosystems and fishery management in wide-ranging subantarctic fisheries. Reviews in Fish Biology and Fisheries, 30(1):203–217, March 2020. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-020-09597-w.

Tebbett:2024:LRH

[TBB^{+}24] Sterling B. Tebbett, David R. Bellwood, Tahlia Bassett, Michael V. W. Cuttler, Molly Moustaka, Shaun K. Wilson, Helen F. Yan, and Richard D. Evans. The limited role of herbivorous fishes and turf-based trophic pathways in the functioning of turbid coral reefs. Reviews in Fish Biology and Fisheries, 34(1):??, ???? 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09823-1.

Torgersen:2011:RDA

[TBK11] Thomas Torgersen, Marc B. M. Bracke, and Tore S. Kristiansen. Reply to Diggles et al. (2011): Ecology and welfare of aquatic animals in wild capture fisheries. Reviews in Fish Biology and Fisheries, 21(4):767–769, December 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-011-9221-y.

Teh:2020:TSL

[TCS20] Louise S. L. Teh, Tim Cashion, and U. Rashid Sumaila. Taking stock: a Large Marine Ecosystem perspective of socioeconomic and ecological trends in East China Sea fisheries. Reviews in Fish Biology and Fisheries, 30(2):269–292, June 2020. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-020-09599-8.

Torres:2015:DCI

[TdSL15] Rodrigo A. Torres, Filipe Alberto dos Santos, and Rosangela P. R. Lessa. Disentangling the controversial identity of the halfbeak stock (Hemiramphus brasiliensis and H. balao) from northeastern Brazil using multilocus DNA markers. Reviews in Fish Biology and Fisheries, 25(2):379–394, June 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-015-9381-2.

Teletchea:2009:MIM

[Tel09] Fabrice Teletchea. Molecular identification methods of fish species: reassessment and possible applications. Reviews in Fish Biology and Fisheries, 19(3):??, September 2009. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-009-9107-4.

Terceiro:2001:SFC

[Ter01] Mark Terceiro. The summer flounder chronicles: Science, politics, and litigation, 1975–2000. Reviews in Fish Biology and Fisheries, 11(2):125–168, June 2001. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1023/A:1015260005887.

Terceiro:2011:SFC

[Ter11] Mark Terceiro. The summer flounder chronicles II: new science, new controversy, 2001–2010. Reviews in Fish Biology and Fisheries, 21(4):681–712, December 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-011-9207-9.

Terceiro:2018:SFC

[Ter18] Mark Terceiro. The summer flounder chronicles III: struggling with success, 2011–2016. Reviews in Fish Biology and Fisheries, 28(2):381–404, June 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-017-9506-x.

Tawari-Fufeyin:2007:FSD

[TFE07] P. Tawari-Fufeyin and S. A. Ekaye. Fish species diversity as indicator of pollution in Ikpoba River, Benin City, Nigeria. Reviews in Fish Biology and Fisheries, 17(1):21–30, February 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-006-9015-9.

Teletchea:2009:CAR

[TFF09] Fabrice Teletchea, Alexis Fostier, and Pascal Fontaine. Comparative analysis of reproductive traits in 65 freshwater fish species: application to the domestication of new fish species. Reviews in Fish Biology and Fisheries, 19(4):403–430, December 2009. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-008-9102-1.

Trebilco:2022:WWC

[TFP22] Rowan Trebilco, Aysha Fleming, and Gretta T. Pecl. Warming world, changing ocean: mitigation and adaptation to support resilient marine systems. Reviews in Fish Biology and Fisheries, 32(1):39–63, March 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-021-09678-4.

Thomson:1992:BLC

[Tho92] Keith S. Thomson. The biology of Latimeria chalumnae and evolution of coelacanths. Reviews in Fish Biology and Fisheries, 2(3):267–268, September 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00045040.

Trujillo-Jimenez:2010:PDF

[TJLLC10] Patricia Trujillo-Jiménez, Eugenia López-López, and Julio A. Camargo. Patterns in the distribution of fish assemblages in Río Amacuzac, Mexico: influence of abiotic factors and biotic factors. Reviews in Fish Biology and Fisheries, 20(4): 457–469, December 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-009-9153-y.

Temple:2018:MMI

[TKB18] Andrew J. Temple, Jeremy J. Kiszka, and Per Berggren. Marine megafauna interactions with small-scale fisheries in the southwestern Indian Ocean: a review of status and challenges for research and management. Reviews in Fish Biology and Fisheries, 28(1):89–115, March 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-017-9494-x.

Tiwary:2004:BTF

[TKR04] Basant K. Tiwary, R. Kirubagaran, and Arun K. Ray. The biology of triploid fish. Reviews in Fish Biology and Fisheries, 14(4): 391–402, December 2004. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-004-8361-8.

Tiwary:2006:LE

[TKR06] Basant K. Tiwary, R. Kirubagaran, and Arun K. Ray. Letter to the Editor. Reviews in Fish Biology and Fisheries, 16(1): 113–114, February 2006. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-006-9002-1.

Taylor:2024:CDS

[TMD^{+}24] Brett M. Taylor, Andrew J. K. McInnis, Mari Deinhart, Ka’ohinani Kawahigashi, and John Gourley. Comparative demography of surgeonfishes from the tropical western Pacific. Reviews in Fish Biology and Fisheries, 34(1):??, ???? 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09816-0.

Tahmasbian:2024:UHI

[TMKC24] Iman Tahmasbian, Matthew N. McMillan, Jonathan Kok, and Anthony J. Courtney. Underwater hyperspectral imaging technology has potential to differentiate and monitor scallop populations. Reviews in Fish Biology and Fisheries, 34(1):??, ???? 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09817-z.

Thorstad:2008:FAW

[TØH08] Eva B. Thorstad, Finn Økland, and Tor G. Heggberget. Factors affecting the within-river spawning migration of Atlantic salmon, with emphasis on human impacts. Reviews in Fish Biology and Fisheries, 18(4):345–371, November 2008. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-007-9076-4.

Taylor:2014:SBL

[TP14] Andrew T. Taylor and Douglas L. Peterson. Shoal bass life history and threats: a synthesis of current knowledge of a Micropterus species. Reviews in Fish Biology and Fisheries, 24 (1):159–167, March 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-013-9323-9.

Tyler:1994:IVS

[TR94] Jeffrey A. Tyler and Kenneth A. Rose. Individual variability and spatial heterogeneity in fish population models. Reviews in Fish Biology and Fisheries, 4(1):91–123, March 1994. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00043262.

Terova:2013:MCG

[TRB13] Genciana Terova, Simona Rimoldi, and Giovanni Bernardini. Molecular cloning and gene expression analysis in aquaculture science: a review focusing on respiration and immune responses in European sea bass (Dicentrarchus labrax). Reviews in Fish Biology and Fisheries, 23(2):175–194, June 2013. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-012-9290-6.

Terova:2019:RTO

[TRG19] Genciana Terova, Simona Rimoldi, and Laura Gasco. Rainbow trout (Oncorhynchus mykiss) gut microbiota is modulated by insect meal from Hermetia illucens prepupae in the diet. Reviews in Fish Biology and Fisheries, 29(2):465–486, June 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-019-09558-y.

Torres-Romero:2024:ERW

[TRJ24] Erik Joaquín Torres-Romero and Juan Carlos Pérez Jiménez. Extinction risk of the world’s chondrichthyan fishes: a global assessment of the interplay between anthropogenic factors and marine protected areas. Reviews in Fish Biology and Fisheries, 34(2):685–701, June 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-023-09830-2.

Tyler:1996:OGD

[TS96] C. R. Tyler and J. P. Sumpter. Oocyte growth and development in teleosts. Reviews in Fish Biology and Fisheries, 6(3): 287–318, September 1996. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF00122584.

Tsikliras:2014:SMM

[TS14] Athanassios C. Tsikliras and Konstantinos I. Stergiou. Size at maturity of Mediterranean marine fishes. Reviews in Fish Biology and Fisheries, 24(1):219–268, March 2014. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-013-9330-x.

Tebbett:2022:FRS

[TSB22] Sterling B. Tebbett, Alexandre C. Siqueira, and David R. Bellwood. The functional roles of surgeonfishes on coral reefs: past, present and future. Reviews in Fish Biology and Fisheries, 32 (2):387–439, June 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-021-09692-6.

Thiem:2011:TRT

[TTC11] Jason D. Thiem, Mark K. Taylor, and Steven J. Cooke. Trends in the reporting of tagging procedures for fish telemetry studies that have used surgical implantation of transmitters: a call for more complete reporting. Reviews in Fish Biology and Fisheries, 21(1):117–126, March 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-010-9194-2.

Turner:1995:CGF

[Tur95] George F. Turner. A complete guide to the freshwater fishes of Southern Africa. Reviews in Fish Biology and Fisheries, 5(3): 380–382, September 1995. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/BF00043010.

Turner:1999:BRF

[Tur99a] George F. Turner. Book review: The Fishery Potential and Productivity of the Pelagic Zone of Lake Malawi / Niassa. Edited by Andrew Menz. Reviews in Fish Biology and Fisheries, 9 (2):207–208, June 1999. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1008908026525.

Turner:1999:WFS

[Tur99b] George F. Turner. What is a fish species? Reviews in Fish Biology and Fisheries, 9(4):281–297, December 1999. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1023/A:1008903228512.

Tarkan:2015:PLC

[TV15] Ali Serhan Tarkan and Lorenzo Vilizzi. Patterns, latitudinal clines and countergradient variation in the growth of roach Rutilus rutilus (Cyprinidae) in its Eurasian area of distribution. Reviews in Fish Biology and Fisheries, 25(4):587–602, December 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-015-9398-6.

Utter:2002:MAG

[UE02] Fred Utter and John Epifanio. Marine aquaculture: Genetic potentialities and pitfalls. Reviews in Fish Biology and Fisheries, 12(1):59–77, March 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1022644021870.

Ulloa:2011:ZMO

[UIA11] Pilar E. Ulloa, Patricia Iturra, and Cristian Araneda. Zebrafish as a model organism for nutrition and growth: towards comparative studies of nutritional genomics applied to aquacultured fishes. Reviews in Fish Biology and Fisheries, 21(4): 649–666, December 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-011-9203-0.

Utter:1994:PMG

[Utt94] Fred M. Utter. Perspectives of molecular genetics and fisheries into the 21st century. Reviews in Fish Biology and Fisheries, 4(3):374–378, September 1994. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00042910.

Utter:2000:PSA

[Utt00] Fred Utter. Patterns of subspecific anthropogenic introgression in two salmonid genera. Reviews in Fish Biology and Fisheries, 10(3):265–279, September 2000. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1023/A:1016686415022.

Utter:2004:PGC

[Utt04] Fred Utter. Population genetics, conservation and evolution in salmonids and other widely cultured fishes: some perspectives over six decades. Reviews in Fish Biology and Fisheries, 14 (1):125–144, March 2004. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-004-3768-9.

VanFleet:2010:BRN

[VAJ10] Tyler E. Van Fleet, Elizabeth Austin, and Francis Juanes. Book review: Nicholas V. C. Polunin (ed): Review of “Aquatic ecosystems: trends and global prospects”. Reviews in Fish Biology and Fisheries, 20(1):137–138, March 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-009-9122-5.

Vethaak:1992:FDM

[VaR92] A. D. Vethaak and T. ap Rheinallt. Fish disease as a monitor for marine pollution: the case of the North Sea. Reviews in Fish Biology and Fisheries, 2(1):1–32, March 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00042915.

Verhelst:2023:EOU

[VBC^{+}23] Pieterjan Verhelst, Rein Brys, Steven J. Cooke, Ine Pauwels, Mehis Rohtla, and Jan Reubens. Enhancing our understanding of fish movement ecology through interdisciplinary and crossboundary research. Reviews in Fish Biology and Fisheries, 33 (1):111–135, March 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-022-09741-8.

Vincenzi:2012:TSD

[VCD12] Simone Vincenzi, Alain J. Crivelli, and Giulio A. De Leo. Translocation of stream-dwelling salmonids in headwaters: insights from a 15-year reintroduction experience. Reviews in Fish Biology and Fisheries, 22(2):437–455, June 2012. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-011-9235-5.

Vilizzi:2019:GRM

[VCZ19] Lorenzo Vilizzi, Gordon H. Copp, and Yiwen Zeng. A global review and meta-analysis of applications of the freshwater fish invasiveness screening kit. Reviews in Fish Biology and Fisheries, 29(3):529–568, September 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-019-09562-2.

Vilmar:2022:SPS

[VD22] Matilda Vilmar and Valentina Di Santo. Swimming performance of sharks and rays under climate change. Reviews in Fish Biology and Fisheries, 32(3):765–781, September 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-022-09706-x.

vanDam:2002:PFP

[vDBV02] Anne A. van Dam, Malcolm C. M. Beveridge, and Marc C. J. Verdegem. The potential of fish production based on periphyton. Reviews in Fish Biology and Fisheries, 12(1):1–31, March 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1022639805031.

Viana:2021:EBT

[VdSD21] Douglas Viana, Maria Rosa Dmengeon Pedreiro de Souza, and Lucélia Donatti. The effect of bottom trawling time on mortality, physical damage and oxidative stress in two Sciaenidae species. Reviews in Fish Biology and Fisheries, 31(4):957–975, December 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-021-09682-8.

Velez-Espino:2005:TRE

[VE05] Luis A. Vélez-Espino. Taxonomic revision, ecology and endangerment categorization of the Andean catfish Astroblepus ubidiai (Teleostei: Astroblepidae). Reviews in Fish Biology and Fisheries, 13(4):367–378, January 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-004-1096-8.

Velez-Espino:2010:SEP

[VEK10] Luis A. Vélez-Espino and Marten A. Koops. A synthesis of the ecological processes influencing variation in life history and movement patterns of American eel: towards a global assessment. Reviews in Fish Biology and Fisheries, 20(2):163– 186, June 2010. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-009-9127-0.

Vinas:2011:FUA

[VGA11] Jordi Viñas, Ana Gordoa, and Rosa M. Araguas. Facts and uncertainties about the genetic population structure of Atlantic bluefin tuna (Thunnus thynnus) in the Mediterranean. implications for fishery management. Reviews in Fish Biology and Fisheries, 21(3):527–541, September 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-010-9174-6.

vanGinneken:2005:EEA

[vGM05] Vincent J. T. van Ginneken and Gregory E. Maes. The European eel (Anguilla anguilla, Linnaeus), its lifecycle, evolution and reproduction: a literature review. Reviews in Fish Biology and Fisheries, 15(4):367–398, November 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-006-0005-8.

Vila-Gispert:2002:GLH

[VGMAGB02] A. Vila-Gispert, R. Moreno-Amich, and E. García-Berthou. Gradients of life-history variation: an intercontinental comparison of fishes. Reviews in Fish Biology and Fisheries, 12(4): 417–427, December 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1025352026974.

Villena:2003:ANF

[Vil03] Alberto J. Villena. Applications and needs of fish and shellfish cell culture for disease control in aquaculture. Reviews in Fish Biology and Fisheries, 13(1):111–140, March 2003. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1026304212673.

Vilizzi:2018:ADC

[Vil18] Lorenzo Vilizzi. Age determination in common carp Cyprinus carpio: history, relative utility of ageing structures, precision and accuracy. Reviews in Fish Biology and Fisheries, 28(3): 461–484, September 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-018-9514-5.

Vollset:2020:SRM

[VLD20] Knut Wiik Vollset, Robert J. Lennox, and Ian Dohoo. Systematic review and meta-analysis of PIT tagging effects on mortality and growth of juvenile salmonids. Reviews in Fish Biology and Fisheries, 30(4):553–568, December 2020. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-020-09611-1.

Villanueva:2007:AIE

[VMB07] Roger Villanueva, Natalie A. Moltschaniwskyj, and Anna Bozzano. Abiotic influences on embryo growth: statoliths as experimental tools in the squid early life history. Reviews in Fish Biology and Fisheries, 17(2–3):??, August 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-006-9022-x.

vanOverzee:2015:EFD

[vOR15] Harriët M. J. van Overzee and Adriaan D. Rijnsdorp. Effects of fishing during the spawning period: implications for sustainable management. Reviews in Fish Biology and Fisheries, 25 (1):65–83, March 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-014-9370-x.

vanPoorten:2018:RBH

[vPKW18] Brett van Poorten, Josh Korman, and Carl Walters. Revisiting Beverton–Holt recruitment in the presence of variation in food availability. Reviews in Fish Biology and Fisheries, 28(3): 607–624, September 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-018-9521-6.

Vipin:2012:DMR

[VRR12] P. M. Vipin, Renju Ravi, and M. P. Remesan. Distribution of myctophid resources in the Indian Ocean. Reviews in Fish Biology and Fisheries, 22(2):423–436, June 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-011-9244-4.

Varela-Romero:2002:DHC

[VRRCAG02] Alejandro Varela-Romero, Gorgonio Ruiz-Campos, and Jorge Alaníz-García. Distribution, habitat and conservation status of desert pupfish (Cyprinodon macularius) in the Lower Colorado River Basin, Mexico. Reviews in Fish Biology and Fisheries, 12 (2–3):157–165, June 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1025006920052.

Vincenzi:2012:STS

[VSM12] Simone Vincenzi, William Hallowell Satterthwaite, and Marc Mangel. Spatial and temporal scale of density-dependent body growth and its implications for recruitment, population dynamics and management of stream-dwelling salmonid populations. Reviews in Fish Biology and Fisheries, 22(3):813–825, September 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-011-9247-1.

Villagra:2022:LHT

[VVU22] Damian Villagra, Noemi Van Bogaert, and Sven Sebastian Uhlmann. Life-history traits of batoids (superorder Batoidea) in the Northeast Atlantic and the Mediterranean. Reviews in Fish Biology and Fisheries, 32(2):473–495, June 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-021-09695-3.

Videler:1991:FSS

[VW91] J. J. Videler and C. S. Wardle. Fish swimming stride by stride: speed limits and endurance. Reviews in Fish Biology and Fisheries, 1(1):23–40, September 1991. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00042660.

Wolter:2003:NIF

[WA03] C. Wolter and R. Arlinghaus. Navigation impacts on freshwater fish assemblages: the ecological relevance of swimming performance. Reviews in Fish Biology and Fisheries, 13(1):63– 89, March 2003. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1023/A:1026350223459.

Ward:2008:EEH

[War08] Peter Ward. Empirical estimates of historical variations in the catchability and fishing power of pelagic longline fishing gear. Reviews in Fish Biology and Fisheries, 18(4):409–426, November 2008. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-007-9082-6.

Waltrick:2012:EDE

[WAS12] Daniela Waltrick, Cynthia Awruch, and Colin Simpfendorfer. Embryonic diapause in the elasmobranchs. Reviews in Fish Biology and Fisheries, 22(4):849–859, December 2012. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-012-9267-5.

Wright:1994:MGM

[WB94] Jonathan M. Wright and Paul Bentzen. Microsatellites: genetic markers for the future. Reviews in Fish Biology and Fisheries, 4(3):384–388, September 1994. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00042912.

Williams:2005:BRG

[WBW05] Ernesth Williams, Jr. and Lucy Bunkley-Williams. Book review: Ginger Garrison, Peces de la Isla del Coco Fishes. Instituto Nacional de Biodiversidad (INBio), Santo Domingo, Costa Rica, 2000. 392 pp. Reviews in Fish Biology and Fisheries, 15 (3):177–187, August 2005. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-005-2171-5.

Williams:2009:BRW

[WBW09] Ernest H. Williams, Jr. and Lucy Bunkley-Williams. Book review: W. Beller, M. A. Casellas, M. J. Cerame-Vivas, L. Duffy, J. El Koury, P. A. Gilabert, J. A. González Liboy, M. Hernández Avila, N. Maldonado, C. A. Matos, A. A. Mignucci-Giannoni, E. Pantojas Garcia, J. J. Rigau, D. Shelley and M. Tacher-Roffe: Puerto Rico and the Sea — 1999: an Action Program for Marine Affairs. Reviews in Fish Biology and Fisheries, 19 (1):121–123, March 2009. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-008-9097-7.

Wagner:2011:SIT

[WCD11] Glenn N. Wagner, Steven J. Cooke, and Katherine A. Deters. Surgical implantation techniques for electronic tags in fish. Reviews in Fish Biology and Fisheries, 21(1):71–81, March 2011. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-010-9191-5.

Walters:1997:SDM

[WCP97] Carl Walters, Villy Christensen, and Daniel Pauly. Structuring dynamic models of exploited ecosystems from trophic massbalance assessments. Reviews in Fish Biology and Fisheries, 7 (2):139–172, June 1997. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1018479526149.

Winkler:2019:SWO

[WDP19] Alexander Claus Winkler, Murray Ian Duncan, and Warren Mason Potts. Sectioned or whole otoliths? A global review of hard structure preparation techniques used in ageing sparid fishes. Reviews in Fish Biology and Fisheries, 29(3):605–611, September 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-019-09571-1.

Webb:2002:MSG

[Web02] Shane A. Webb. Molecular systematics of the genus Allodontichthys (Cyprinodontiformes: Goodeidae). Reviews in Fish Biology and Fisheries, 12(2–3):193–205, June 2002. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1023/A:1025099709739.

Webster:2009:BRJ

[Web09] Carl Webster. Book review: J. E. P. Cyrino, D. P. Bureau, B. G. Kapour (eds): Feeding and digestive functions of fishes. Reviews in Fish Biology and Fisheries, 19(2):261–263, June 2009. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-009-9113-6.

Whitmarsh:2017:WBB

[WFH17] Sasha K. Whitmarsh, Peter G. Fairweather, and Charlie Huveneers. What is Big BRUVver up to? Methods and uses of baited underwater video. Reviews in Fish Biology and Fisheries, 27(1):53–73, March 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-016-9450-1.

Ward:1994:AMG

[WG94] Robert D. Ward and Peter M. Grewe. Appraisal of molecular genetic techniques in fisheries. Reviews in Fish Biology and Fisheries, 4(3):300–325, September 1994. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00042907.

Wang:2019:AEE

[WG19] Yingnan Wang and Baocheng Guo. Adaption to extreme environments: a perspective from fish genomics. Reviews in Fish Biology and Fisheries, 29(4):735–747, December 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-019-09577-9.

Whitfield:2018:CCH

[WGC18] Alan K. Whitfield, Gareth N. Grant, and Paul D. Cowley. Causes and consequences of human induced impacts on a ubiquitous estuary-dependent marine fish species. Reviews in Fish Biology and Fisheries, 28(1):19–31, March 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-017-9499-5.

Weitkamp:2014:JSE

[WGL14] Laurie A. Weitkamp, Graham Goulette, and Christine Lipsky. Juvenile salmon in estuaries: comparisons between North American Atlantic and Pacific salmon populations. Reviews in Fish Biology and Fisheries, 24(3):713–736, September 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-014-9345-y.

Wang:2017:LTE

[WGL17] Teng Wang, Xin Gao, and Huan-Zhang Liu. Life tables and elasticity analyses of Yangtze River fish species with implications for conservation and management. Reviews in Fish Biology and Fisheries, 27(1):255–266, March 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-016-9464-8.

Wheeler:2020:ASI

[WGR20] Carolyn R. Wheeler, Connor R. Gervais, and Jodie L. Rummer. Anthropogenic stressors influence reproduction and development in elasmobranch fishes. Reviews in Fish Biology and Fisheries, 30(2):373–386, June 2020. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-020-09604-0.

Ward:2007:OHC

[WH07] Peter Ward and Sheree Hindmarsh. An overview of historical changes in the fishing gear and practices of pelagic longliners, with particular reference to Japan’s Pacific fleet. Reviews in Fish Biology and Fisheries, 17(4):501–516, November 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-007-9051-0.

Wang:2024:PDH

[WH24] Luhai Wang and Zhenli Huang. Passive drifting and high mortality rate of released subadult Chinese sturgeons in the Yangtze River. Reviews in Fish Biology and Fisheries, 34(1):??, ???? 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09804-4.

Whitfield:1999:IAE

[Whi99] Alan K. Whitfield. Ichthyofaunal assemblages in estuaries: a South African case study. Reviews in Fish Biology and Fisheries, 9(2):151–186, June 1999. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1008994405375.

Whitfield:2017:RSM

[Whi17] Alan K. Whitfield. The role of seagrass meadows, mangrove forests, salt marshes and reed beds as nursery areas and food sources for fishes in estuaries. Reviews in Fish Biology and Fisheries, 27(1):75–110, March 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-016-9454-x.

Wertheimer:2004:RSR

[WHS04] Alex C. Wertheimer, William R. Heard, and William W. Smoker. Relationship of size at return with environmental variation, hatchery production, and productivity of wild pink salmon in Prince William Sound, Alaska: does size matter? Reviews in Fish Biology and Fisheries, 14(3):321–334, September 2004. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-004-2942-4.

Wang:2002:SIR

[WHU02] Shizhen Wang, Jeffrey J. Hard, and Fred Utter. Salmonid inbreeding: a review. Reviews in Fish Biology and Fisheries, 11(4): 301–319, December 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1021330500365.

Wiegand:1996:CAU

[Wie96] Murray D. Wiegand. Composition, accumulation and utilization of yolk lipids in teleost fish. Reviews in Fish Biology and Fisheries, 6(3):259–286, September 1996. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00122583.

Williams:1996:FAF

[Wil96] Meryl J. Williams. Fourth Asian Fisheries Forum. Reviews in Fish Biology and Fisheries, 6(2):248–250, June 1996. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL

https://link.springer.com/article/10.1007/BF00182346.

Wilkens:2007:BRG

[Wil07] Horst Wilkens. Book review: Graham S. Proudlove, Subterranean fishes of the world. An account of the subterranean (hypogean) fishes described to 2003 with a bibliography 1541– 2004. Reviews in Fish Biology and Fisheries, 17(1):57, February 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-006-9035-5.

Winterbottom:2006:RFW

[Win06] Richard Winterbottom. Review of “Fishes of the World” by Joseph S. Nelson. Reviews in Fish Biology and Fisheries, 16 (2):227–228, May 2006. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-006-9004-z.

Wisenden:1999:ACF

[Wis99] Brian D. Wisenden. Alloparental care in fishes. Reviews in Fish Biology and Fisheries, 9(1):45–70, March 1999. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1023/A:1008865801329.

Watanabe:2014:TTF

[WIT14] Shun Watanabe, Midori Iida, and Katsumi Tsukamoto. Tropical and temperate freshwater amphidromy: a comparison between life history characteristics of sicydiinae, ayu, sculpins and galaxiids. Reviews in Fish Biology and Fisheries, 24(1):1–14, March 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-013-9316-8.

Williamson:2023:DAE

[WJP23] Michael J. Williamson, David M. P. Jacoby, and Adam T. Piper. The drivers of anguillid eel movement in lentic water bodies: a systematic map. Reviews in Fish Biology and Fisheries, 33 (1):147–174, March 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-022-09751-6.

Walters:1999:LRT

[WK99] Carl Walters and Josh Korman. Linking recruitment to trophic factors: revisiting the Beverton–Holt recruitment model from a life history and multispecies perspective. Reviews in Fish Biology and Fisheries, 9(2):187–202, June 1999. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1023/A:1008991021305.

Weil:2013:CMD

[WLB13] Claudine Weil, Florence Lefèvre, and Jerôme Bugeon. Characteristics and metabolism of different adipose tissues in fish.

Reviews in Fish Biology and Fisheries, 23(2):157–173, June 2013. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-012-9288-0.

Wosnick:2023:BVS

[WLB^{+}23] Natascha Wosnick, Renata Daldin Leite, Samuel Balanin, Ana Paula Chaves, Eduardo Rufino de Senna Gastal, Rachel Ann Hauser-Davis, and Eloísa Pinheiro Giareta. Behavioral and visual stress-induced proxies in elasmobranchs. Reviews in Fish Biology and Fisheries, 33(1):175–199, March 2023. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-023-09757-8.

Walters:1996:LSA

[WM96] Carl Walters and Jean-Jacques Maguire. Lessons for stock assessment from the northern cod collapse. Reviews in Fish Biology and Fisheries, 6(2):125–137, June 1996. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/BF00182340.

Wood:2002:SAG

[WM02] Robert M. Wood and Richard L. Mayden. Speciation and anagenesis in the genus Cyprinella of Mexico (Teleostei: Cyprinidae): a case study of Model III allopatric speciation. Reviews in Fish Biology and Fisheries, 12(2–3):253–271, June 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1025018118339.

Wootton:2020:EMS

[WMA20] Henry F. Wootton, John R. Morrongiello, and Asta Audzijonyte. Estimating maturity from size-at-age data: Are real-world fisheries datasets up to the task? Reviews in Fish Biology and Fisheries, 30(4):681–697, December 2020. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-020-09617-9.

Warren-Myers:2018:OMM

[WMDS18] Fletcher Warren-Myers, Tim Dempster, and Stephen E. Swearer. Otolith mass marking techniques for aquaculture and restocking: benefits and limitations. Reviews in Fish Biology and Fisheries, 28(3):485–501, September 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/s11160-018-9515-4.

Witte:2000:RCS

[WMG00] F. Witte, B. S. Msuku, and T. Goldschmidt. Recovery of cichlid species in Lake Victoria: an examination of factors leading to differential extinction. Reviews in Fish Biology and Fisheries, 10(2):233–241, June 2000. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1016677515930.

Williams:1992:PBI

[WMM92] H. H. Williams, K. MacKenzie, and A. M. McCarthy. Parasites as biological indicators of the population biology, migrations, diet, and phylogenetics of fish. Reviews in Fish Biology and Fisheries, 2(2):144–176, June 1992. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/BF00042882.

Ward:2022:SML

[WMTL22] Delphi Ward, Jessica Melbourne-Thomas, and Cayne Layton. Safeguarding marine life: conservation of biodiversity and ecosystems. Reviews in Fish Biology and Fisheries, 32 (1):65–100, March 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-022-09700-3.

Wilkes:2018:FPD

[WMW18] Martin A. Wilkes, Morwenna Mckenzie, and J. Angus Webb. Fish passage design for sustainable hydropower in the temperate Southern Hemisphere: an evidence review. Reviews in Fish Biology and Fisheries, 28(1):117–135, March 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-017-9496-8.

Williams:2012:IWD

[WNB12] Ashley J. Williams, Simon J. Nicol, and Mike Batty. International workshop on developing strategies for monitoring datalimited deepwater demersal line fisheries in the Pacific Ocean. Reviews in Fish Biology and Fisheries, 22(2):527–531, June 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-011-9234-6.

Wootton:1993:ELH

[Woo93] R. J. Wootton. The evolution of life histories: Theory and analysis. Reviews in Fish Biology and Fisheries, 3(4):384–385, December 1993. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00043394.

Wootton:1995:FB

[Woo95] R. J. Wootton. Fish bioenergetics. Reviews in Fish Biology and Fisheries, 5(3):389–390, September 1995. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://

link.springer.com/article/10.1007/BF00043016.

Wootton:1998:BRB

[Woo98a] R. J. Wootton. Book review: Behavioural Ecology of Teleost Fishes J.-G. J. Godin (ed.). Reviews in Fish Biology and Fisheries, 8(4):493–494, December 1998. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1023/A:1008872606908.

Wootton:1998:PFF

[Woo98b] R. J. Wootton. Patterns in freshwater fish ecology W.J. Matthews. Reviews in Fish Biology and Fisheries, 8(4):495– 497, December 1998. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1023/A:1008891703274.

Wooster:1999:GVL

[Woo99a] Warren S. Wooster. Global versus Local Changes in Upwelling Systems (Collection Colloques et séminaires). Reviews in Fish Biology and Fisheries, 9(3):269–270, September 1999. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1008902700065.

Wootton:1999:BRF

[Woo99b] R. J. Wootton. Book review: Fish Biology in Japan: an Anthology in Honour of Hiroya Kawanabe (Developments in Environmental Biology of Fishes 18). Series Editor (Eugene K. Balon). Edited by M. Yuma, I. Nakamura and K. D. Fausch.

Reviews in Fish Biology and Fisheries, 9(2):208–209, June 1999. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1023/A:1008998525616.

Wegner:2009:ODD

[WOR09] Arleta Wegner, Teresa Ostaszewska, and Wojciech Rożek. The ontogenetic development of the digestive tract and accessory glands of sterlet (Acipenser ruthenus L.) larvae during endogenous feeding. Reviews in Fish Biology and Fisheries, 19(4):??, December 2009. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-009-9111-8.

Walters:1996:SIR

[WP96] Carl Walters and Peter H. Pearse. Stock information requirements for quota management systems in commercial fisheries. Reviews in Fish Biology and Fisheries, 6(1):21–42, March 1996. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/BF00058518.

Whitfield:2012:GRC

[WPD12] A. K. Whitfield, J. Panfili, and J.-D. Durand. A global review of the cosmopolitan flathead mullet Mugil cephalus Linnaeus 1758 (Teleostei: Mugilidae), with emphasis on the biology, genetics, ecology and fisheries aspects of this apparent species complex. Reviews in Fish Biology and Fisheries, 22(3):641–681, September 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-012-9263-9.

Wringe:2016:SCF

[WPF16] Brendan F. Wringe, Craig F. Purchase, and Ian A. Fleming. In search of a “cultured fish phenotype”: a systematic review, meta-analysis and vote-counting analysis. Reviews in Fish Biology and Fisheries, 26(3):351–373, September 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-016-9431-4.

Westhoff:2016:GRF

[WR16] J. T. Westhoff and A. E. Rosenberger. A global review of freshwater crayfish temperature tolerance, preference, and optimal growth. Reviews in Fish Biology and Fisheries, 26(3): 329–349, September 2016. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-016-9430-5.

Wootton:2021:MFG

[WRSG21] Nina Wootton, Patrick Reis-Santos, and Bronwyn M. Gillanders. Microplastic in fish — a global synthesis. Reviews in Fish Biology and Fisheries, 31(4):753–771, December 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-021-09684-6.

Wueringer:2009:BEE

[WSC09] Barbara E. Wueringer, Lyle Squire, Jr., and Shaun P. Collin. The biology of extinct and extant sawfish (Batoidea: Sclerorhynchidae and Pristidae). Reviews in Fish Biology and Fisheries, 19(4):??, December 2009. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-009-9112-7.

Willis:2022:CSR

[WSGP22] Kathryn A. Willis, Catarina Serra-Gonçalves, and Peter S. Puskic. Cleaner seas: reducing marine pollution. Reviews in Fish Biology and Fisheries, 32(1):145–160, March 2022. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-021-09674-8.

Wueringer:2008:CLL

[WT08] Barbara E. Wueringer and Ian R. Tibbetts. Comparison of the lateral line and ampullary systems of two species of shovelnose ray. Reviews in Fish Biology and Fisheries, 18(1):47–64, February 2008. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-007-9063-9.

Whitfield:2006:FSS

[WTC06] Alan K. Whitfield, Ricky H. Taylor, and Digby P. Cyrus. Fishes and salinities in the St Lucia estuarine system — a review. Reviews in Fish Biology and Fisheries, 16(1):??, February 2006. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-006-0003-x.

Willson:2006:VEP

[WW06] Mary F. Willson and Jamie N. Womble. Vertebrate exploitation of pulsed marine prey: a review and the example of spawning herring. Reviews in Fish Biology and Fisheries, 16(2):183– 200, May 2006. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-006-9009-7.

Whitney:2017:FRS

[WWS17] James E. Whitney, Joanna B. Whittier, and Angela L. Strecker. Forecasted range shifts of arid-land fishes in response to climate change. Reviews in Fish Biology and Fisheries, 27(2):463– 479, June 2017. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-017-9479-9.

Wang:2024:RSB

[WZB^{+}24] Lin Wang, Hancheng Zhao, Ran Bi, Xiaohan Chen, Zhendong Lyu, and Wenhua Liu. Roles and sources of B vitamins in the marine ecosystem. Reviews in Fish Biology and Fisheries, 34(1):??, ???? 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-023-09818-y.

Xiong:2015:NNF

[XSC15] Wen Xiong, Xiaoyun Sui, and Yifeng Chen. Non-native freshwater fish species in China. Reviews in Fish Biology and Fisheries, 25(4):651–687, December 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-015-9396-8.

Young:2020:BCS

[YC20] Chelsey N. Young and John K. Carlson. The biology and conservation status of the oceanic whitetip shark (Carcharhinus longimanus) and future directions for recovery. Reviews in Fish Biology and Fisheries, 30(2):293–312, June 2020. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-020-09601-3.

Young:2011:HRP

[YCT11] Paciencia S. Young, Joseph J. Cech, Jr., and Lisa C. Thompson. Hydropower-related pulsed-flow impacts on stream fishes: a brief review, conceptual model, knowledge gaps, and research needs. Reviews in Fish Biology and Fisheries, 21(4):713–731, December 2011. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-011-9211-0.

Yavno:2018:DVF

[YH18] Stan Yavno and Roi Holzman. Do viscous forces affect survival of marine fish larvae? Revisiting the ‘safe harbour’ hypothesis. Reviews in Fish Biology and Fisheries, 28(1):201–212, March 2018. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-017-9503-0.

Yemane:2014:ACD

[YKS14] Dawit Yemane, Stephen P. Kirkman, and Toufiek Samaai. Assessing changes in the distribution and range size of demersal fish populations in the Benguela Current Large Marine Ecosystem. Reviews in Fish Biology and Fisheries, 24(2):463–483, June 2014. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-014-9357-7.

Young:2012:WEF

[YMR12] Jock W. Young, A. David McKinnon, and Anthony J. Richardson. Workshop on the ecosystem and fisheries of the Coral Sea: an Australian perspective on research and management. Reviews in Fish Biology and Fisheries, 22(3):827–834, September 2012. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-011-9251-5.

Young:2015:SSG

[YOC15] J. W. Young, R. J. Olson, and C. A. Choy. Setting the stage for a global-scale trophic analysis of marine top predators: a multi-workshop review. Reviews in Fish Biology and Fisheries, 25(1):261–272, March 2015. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-014-9368-4.

Young:2007:NGT

[YVR07] R. E. Young, M. Vecchione, and C. F. E. Roper. A new genus and three new species of decapodiform cephalopods (Mollusca:

Cephalopoda). Reviews in Fish Biology and Fisheries, 17(2– 3):353–365, August 2007. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-007-9044-z.

Yue:2024:IGO

[YWS^{+}24] G. H. Yue, L. Wang, F. Sun, Z. T. Yang, J. Wong, Y. F. Wen, H. Y. Pang, M. Lee, S. T. Yeo, B. Liang, K. Chen, H. S. Lim, and J. H. Jiang. Improving growth, omega-3 contents, and disease resistance of Asian seabass: status of a 20-year familybased breeding program. Reviews in Fish Biology and Fisheries, 34(1):??, ???? 2024. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-023-09810-6.

Zhou:2002:DGD

[ZCT02] Rongjia Zhou, Hanhua Cheng, and Terrence R. Tiersch. Differential genome duplication and fish diversity. Reviews in Fish Biology and Fisheries, 11(4):331–337, December 2002. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1023/A:1021395506705.

Zeng:2019:ECC

[ZCW19] Zeyu Zeng, William W. L. Cheung, and Ying Wang. Effects of climate change and fishing on the Pearl River Estuary ecosystem and fisheries. Reviews in Fish Biology and Fisheries, 29(4): 861–875, December 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-019-09574-y.

Zhang:2019:LSC

[ZDJ19] Chao Zhang, Chengzhi Ding, and Xiaoming Jiang. Large-scale cascaded dam constructions drive taxonomic and phylogenetic differentiation of fish fauna in the Lancang River, China. Reviews in Fish Biology and Fisheries, 29(4):895–916, December 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-019-09580-0.

Zischke:2013:RBW

[ZFT13] Mitchell T. Zischke, Jessica H. Farley, and Ian R. Tibbetts. Reproductive biology of wahoo, Acanthocybium solandri, off eastern Australia. Reviews in Fish Biology and Fisheries, 23(4):

491–506, December 2013. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-013-9304-z.

Zumholz:2007:EDC

[ZHP07] Karsten Zumholz, Thor Hansteen, and Uwe Piatkowski. Elemental distribution in cephalopod statoliths: NanoSIMS provides new insights into nano-scale structure. Reviews in Fish Biology and Fisheries, 17(2–3):487–491, August 2007. CO-DEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11160-006-9036-4.

Zimmerman:2005:BRJ

[Zim05] Christian E. Zimmerman. Book review: James S. Diana. Biology and Ecology of Fishes, 2nd edition. Biological Sciences Press, Cooper Publishing Group, Traverse City, Michigan, USA, 498 pp. 2004. Reviews in Fish Biology and Fisheries, 13(4):458, January 2005. CODEN RFBFEA. ISSN 0960-3166 (print),

1573-5184 (electronic). URL https://link.springer.com/
article/10.1007/s11160-004-1885-0.

Zakes:2010:EDD

[ZJDZ10] Zdzisław Zakȩś, Barbara Jankowska, and Krystyna Demska-Zakȩś. Effects of different dietary fatty acids profiles on the growth performance and body composition of juvenile tench (Tinca tinca (L.)). Reviews in Fish Biology and Fisheries, 20(3): 389–401, September 2010. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-009-9146-x.

Zhou:2019:BHC

[ZKvZ19] Shijie Zhou, Jeppe Kolding, and Paul A. M. van Zwieten. Balanced harvest: concept, policies, evidence, and management implications. Reviews in Fish Biology and Fisheries, 29(3):711– 733, September 2019. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-019-09568-w.

Zhao:2023:SDF

[ZL23] Youzhu Zhao and Yangfan Li. Spatial disparity of fishing activities overlapping the abrupt shifts for marine net primary production. Reviews in Fish Biology and Fisheries, 33(4):1409– 1421, December 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-023-09778-3.

Zillig:2021:OSD

[ZLF21] Kenneth W. Zillig, Robert A. Lusardi, and Nann A. Fangue. One size does not fit all: variation in thermal eco-physiology among Pacific salmonids. Reviews in Fish Biology and Fisheries, 31 (1):95–114, March 2021. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-020-09632-w.

Zhang:2023:AEC

[ZLH^{+}23] Zonghang Zhang, Wuhan Lin, Dongjian He, Quanming Wu, Canrui Cai, Huaxuan Chen, Yangke Shang, and Xiumei Zhang. Aquaculture environment changes fish behavioral adaptability directly or indirectly through personality traits: a case study. Reviews in Fish Biology and Fisheries, 33(4):1423–1441, December 2023. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-023-09779-2.

Zhang:2017:EOT

[ZMG17] Xiufeng Zhang, Xueying Mei, and Ramesh D. Gulati. Effects of omnivorous tilapia on water turbidity and primary production dynamics in shallow lakes: implications for ecosystem management. Reviews in Fish Biology and Fisheries, 27(1): 245–254, March 2017. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-016-9458-6.

Zimmerman:2004:I

[ZN04] Christian E. Zimmerman and Jennifer L. Nielsen. Introduction. Reviews in Fish Biology and Fisheries, 14(3):301–303, September 2004. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184

(electronic). URL https://link.springer.com/article/10.
1007/s11160-005-3618-4.

Zhang:2020:RPG

[ZRZ20] Hongling Zhang, Chao Ran, and Zhigang Zhou. Research progress on gut health of farmers teleost fish: a viewpoint concerning the intestinal mucosal barrier and the impact of its damage. Reviews in Fish Biology and Fisheries, 30(4):569– 586, December 2020. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-020-09614-y.

Zhang:2020:IFD

[ZWW20] Hui Zhang, Jinming Wu, and Qiwei Wei. Inland fisheries development versus aquatic biodiversity conservation in China and its global implications. Reviews in Fish Biology and Fisheries, 30(4):637–655, December 2020. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.

springer.com/article/10.1007/s11160-020-09622-y.

Zaporozhets:2004:IBH

[ZZ04] O. M. Zaporozhets and G. V. Zaporozhets. Interaction between hatchery and wild Pacific salmon in the Far East of Russia: a review. Reviews in Fish Biology and Fisheries, 14(3): 305–319, September 2004. CODEN RFBFEA. ISSN 0960-3166 (print), 1573-5184 (electronic). URL https://link.springer.

com/article/10.1007/s11160-005-3583-y.